### v1

# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

Abstract.
We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by explicitly breaking the symmetry. This yields a new algebraic notion of Gibbs polystability, conjecturally equivalent to K-polystability. The definition involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides with an analytic reduced stability threshold encoding the coercivity of the K-energy functional modulo automorphisms. These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves and derive a strengthened logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, which implies Aubin’s refinement of the Moser–Trudinger inequality. In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
1. Introduction
The quest for canonical metrics on complex manifolds is a classical theme in geometry, tracing back to the uniformization theorem. According to the seminal Yau–Tian–Donaldson conjecture compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e. a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold - this is now a theorem [42, 90]. The Fano condition requires that the anti-canonical line bundle $K_{X}$ of $X$ is ample, while K-polystability is a finer algebro-geometric condition that amounts to a positivity property of all polarized $\mathbb{C}^{*}$–equivariant deformations of $(X,-K_{X}),$ called test configurations. This algebro-geometric characterization has also found striking connections to the AdS/CFT correspondence in theoretical physics, where Kähler–Einstein metrics on Fano manifolds induce Sasaki–Einstein metrics on links that are dual to superconformal gauge theories [82, 44, 43]. More precisely, Fano manifolds yield regular links, while Fano orbifolds give rise to quasi-regular links (and, upon introducing weights, Sasaki–Einstein metrics on general links [6, 76]).
Even though the existence problem for Kähler–Einstein metrics is now settled, it should be emphasized that the problem of explicitly describing the Kähler–Einstein metric in terms of algebro-geometric data of $X$ is, in general, intractable, and only very few explicit examples are known. This problem is one of the motivations for the probabilistic approach to the construction of Kähler–Einstein metrics on $X$ introduced in [12, 13], that was further applied to the AdS/CFT correspondence in [28]. In the case of positive Ricci curvature, it requires, however, that the Fano manifold $X$ does not admit any holomorphic vector fields, i.e. that the group $\text{Aut}_{0}(X)$ of automorphism of $X$ homotopic to the identity, be trivial. In the present work the framework is extended to Fano manifolds with nontrivial $\text{Aut}_{0}(X)$. In companion papers we present applications to some toric quiver gauge theories appearing in the the AdS/CFT correspondence [5] and to Onsager’s statistical mechanical point vortex model on the two-sphere [4].
1.1. Recap of the case when $\text{Aut}_{0}(X)$ is trivial
Before stating our main results, let us recall the probabilistic approach on a Fano manifold $X$ with trivial group $\text{Aut}_{0}(X),$ introduced in [13]. The starting point is the basic fact that a Kähler–Einstein (KE) metric $\omega_{\text{KE}}$ with positive Ricci curvature on $X$ can be readily be recovered from its (normalized) volume form $dV_{\text{KE}}.$ The strategy of the probabilistic approach is thus to construct the normalized volume form $dV_{KE}$ by sampling a large number of $N$ points on $X.$ More precisely, the goal is to construct a canonical symmetric probability measure on the $N$-fold products $X^{N}$ – solely defined in terms of explicit algebro-geometric data of $X$ – and show that $dV_{\text{KE}}$ emerges as $N\rightarrow\infty,$ when sampling $N$ points. The construction of the probability measure on $X^{N}$ is based on the observation that $N$-fold products $X^{N}$ come with an algebraic measure $\nu^{(N)},$ uniquely determined up to scaling. More precisely, $N$ is the subsequence defined by
(1.1) $N=N_{k}\coloneqq\dim H^{0}(X,-kK_{X}),$
where the level $k$ ranges over all positive rational numbers such that $kK_{X}$ is a well-defined line bundle, using additive notation for tensor products of line bundles. To simplify the notation the subscript $k$ of $N_{k}$ will be suppressed. The measure $\nu^{(N)}$ on $X^{N}$ may be expressed as
(1.2) $\nu^{(N)}\coloneqq i^{N^{2}}\big(\mathrm{det}S^{(N)}\big)^{-1/k}\wedge\overline{\left(\det S^{(N)}\right)^{-1/k},}$
where $\det S^{(N)}$ is the holomorphic section of the line bundle $kK_{X^{N}}\rightarrow X^{N},$ defined as the Slater determinant
(1.3) $\left(\det S^{(N)}\right)(x_{1},x_{2},\dots,x_{N})\coloneqq\det\left(s_{i}(x_{j})\right),$
in terms of a given basis $s_{1},\dots,s_{N}$ in $H^{0}(X,-kK_{X})$ (changing the basis only has the effect of scaling $\nu^{(N)}).$ However, the total mass of $\nu^{(N)}$ may be infinite, since the density of $\nu^{(N)}$ blows up along the zero-locus of $\det S^{(N)}.$ Accordingly, $X$ is dubbed Gibbs stable [13], if $\nu^{(N)}$ has finite total mass, for $N$ sufficiently large (this is equivalent to an algebro-geometric criterion, recalled in the following section). Then
(1.4) $\mu^{(N)}\coloneqq\nu^{(N)}\left/\int_{X^{N}}\nu^{(N)}\right$.
defines a canonical probability measure on $X^{N},$ which is symmetric (i.e. invariant under permutations of the factors of $X^{N}).$ It is conjectured in [13] that, when sampling $N$ points on $X$ according to the probability measure $\mu^{(N)},$ the corresponding empirical measure $\delta_{N}$ on $X,$ i.e. the random discrete probability measure
(1.5) $\delta_{N}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}},$
on $X,$ converges in probability as $N\rightarrow\infty,$ to the volume form $dV_{\text{KE}}$ of a Kähler–Einstein metric $\omega_{\text{KE}}.$
However, a necessary condition for $X$ to be Gibbs stable is that $\text{Aut}_{0}(X)$ be trivial. In fact, when $\text{Aut}_{0}(X)$ is non-trivial there is no probability measure on $X^{N}$ that is invariant under the diagonal action of $\text{Aut}_{0}(X)$ (under a mild regularity assumption) [20]. The probabilistic approach in [13] is thus too canonical, in a certain sense, and we will need to break the $\text{Aut}_{0}(X)$–symmetry. This is in line with the fact the group $\text{Aut}_{0}(X)$ acts non-trivially on any given Kähler–Einstein metric. Hence, there is no canonical way of singling out a unique Kähler–Einstein metric.
1.2. Gibbs polystability and microscopic stability thresholds
We will consider the general framework of log Fano manifolds $(X,\Delta),$ consisting of a complex projective manifold $X$ and an effective $\mathbb{Q}$–divisor $\Delta$ on $X,$ whose log anti-canonical line bundle $K_{(X,\Delta)}$ is an ample $\mathbb{Q}$–line bundle [24]. This setting encompasses, in particular, Fano orbifolds [61]. Kähler–Einstein metrics on $(X,\Delta)$ are, in general, understood as positive currents $\omega_{\text{KE}}$ in $c_{1}\left(-K_{(X,\Delta)}\right)$ with locally bounded potentials, that restrict to genuine Kähler–Einstein metrics on the complement of the support of $\Delta.$ In particular, $\omega_{\text{KE }}$ satisfies the following equation in sense of currents:
$\text{Ric $\omega$}=\omega+[\Delta],$
where $\Delta]$ denotes the current of integration along the divisor $\Delta.$ Given a log Fano manifold $(X,\Delta)$ and a positive integer $k$ such that $kK_{(X,\Delta)}$ defines a line bundle we set
(1.6) $N=N_{k}\coloneqq\dim H^{0}(X,-kK_{(X,\Delta)}),\,\,\,\,\mathcal{G}\coloneqq\text{Aut}_{0}(X,\Delta)$
and we will assume that the complex Lie group $\mathcal{G}$ is reductive (this is a necessary condition for the existence of a Kähler–Einstein metric for $(X,\Delta)$ [42]).
We start by introducing a notion of Gibbs polystability, taking the group $\mathcal{G}$ into account. It is formulated in terms of the Zariski open subvariety $(X^{N})_{\text{ss}}$ of $X^{N},$ defined by the semistable locus of $X^{N},$ with respect to the ample $\mathbb{Q}$–line bundle $K_{(X,\Delta)},$ in the sense of Geometric Invariant Theory (GIT) [64]. We will denote by $\Delta_{N}$ the divisor on $X^{N}$ induced by $\Delta$ that is invariant under permutation of the factors of $X^{N}.$ Thus $(X^{N},\Delta_{N})$ is a log Fano manifold. Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [13], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined
$\gamma^{(N)}(X,\Delta)\coloneqq\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),\,\,\,\,\gamma(X,\Delta)\coloneqq\liminf_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta),$
where $\text{lct}(Y,\Delta;D)$ denotes the log canonical threshold of a divisor $D$ on a quasi-projective log pair $(Y,\Delta)$ – this number is a standard algebro-geometric invariant of a divisor – whose valuative definition is recalled in Section 2.2.1. A log Fano manifold $(X,\Delta)$ is called Gibbs stable at level $k$ if $\gamma^{(N)}(X,\Delta)>1$ and Gibbs stable iff $\gamma^{(N)}(X,\Delta)>1,$ when $N$ is sufficiently large (which is equivalent to the integrability criterion recalled in the previous section). Furthermore, $(X,\Delta)$ is called Gibbs semistable if $\gamma(X,\Delta)\geq 1$ and uniformly Gibbs stable if $\gamma(X,\Delta)>1$ [13, 57]. In order to define a notion Gibbs polystability we first introduce the reduced microscopic stability threshold $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$:
$\gamma^{(N)}(X,\Delta)^{\mathcal{G}}\coloneqq\text{lct}\left((X^{N})_{\text{ss}},\Delta_{N};\mathcal{D}^{(N)}\right),\,\,\,\gamma(X,\Delta)\mathcal{{}^{G}}\coloneqq\liminf_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)^{\mathcal{G}}.$
We propose to call $(X,\Delta)$ Gibbs polystable if it is Gibbs semistable and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ when $N$ is sufficiently large and uniformly Gibbs polystable if the latter condition is replaced by $\gamma(X,\Delta)\mathcal{{}^{G}}>1.$
Conjecture 1.1.
Let $(X,\Delta)$ be a log Fano manifold. Then $(X,\Delta)$ is Gibbs polystable iff it is uniformly Gibbs polystable iff $(X,\Delta)$ is K-polystable.
The first equivalence may, perhaps, be too optimistic, in general. But we present arguments motivating the second one. They are based on conjectural Large Deviation Principle that we shall come back to below, linking the invariant $\gamma(X,\Delta)\mathcal{{}^{G}}$ to the reduced analytic stability threshold defined by
(1.7) $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon(\mathcal{I}-\mathcal{J})^{\mathcal{G}}-C_{\epsilon}\right\},$
where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right)$ – whose minimizers are Kähler–Einstein metrics for $(X,\Delta)$ – and the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry (see Remark 3.1):
(1.8) $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}=\text{inf}_{g\in\mathcal{G}}g^{*}(\mathcal{I}-\mathcal{J}).$
In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the Dirichlet norm. On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [89, 23, 48]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [95, 48]. In the present context, however, it will be important to work with the functional $\mathcal{I}-\mathcal{J}$ as this choice leads to a thermodynamical interpretation of the corresponding reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (see Prop 3.6).
When $\mathcal{G}$ is trivial the invariant $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ equals the well-known analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta),$ which by [94], coincides with the stability threshold $\delta(X,\Delta),$ introduced in [57] and further developed in [30]. This is an algebro-geometric (valuative) invariant that has come to play a pivotal role in recent developments around the Yau–Tian–Donaldson conjecture [79]. It was conjectured in [13] that $\delta^{\mathrm{A}}(X,\Delta)=\gamma(X,\Delta).$ Here we propose the following
Conjecture 1.2.
On any log Fano manifold $(X,\Delta)$ with vanishing Futaki character the invariants $\gamma(X,\Delta)\mathcal{{}^{G}}$ and $\delta^{\mathrm{A}}(X,\Delta)\mathcal{{}^{G}}$ coincide.
The validity of this conjecture would imply the second equivalence in Conjecture 1.1. Indeed, by [48], $\delta^{\mathrm{A}}(X,L)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a Kähler–Einstein metric, which, by the solution of the YTD conjecture is equivalent to K-polystability (see Theorem 3.7).
Our first main result concerns the case of log Fano curves $(X,\Delta),$ i.e. the case when $n=1.$ In this case $X$ is the Riemann sphere, $X=\mathbb{P}^{1}$ and $\Delta$ is any divisor of $r$ points on $X$ with weights $w_{i}\in]0,1[$ satisfying
$2-\sum_{i=1}^{r}w_{i}>0.$
When $\mathcal{G}$ is non-trivial it is well-known [74] that $(X,\Delta)$ is K-polystable iff $\Delta$ is the divisor $\Delta_{w}$ on $\mathbb{P}^{1}$ supported on two points with the same weight $w\in[0,1)$ (thus $\Delta_{w}=0$ when $w=0).$
Theorem 1.3.
The previous two conjectures hold for any log Fano curve $(X,\Delta).$ More precisely,
$\gamma^{(N)}(X,\Delta)=\min\left\{\frac{2(1-1/N)}{V},\min_{l}\frac{2(1-w_{l})}{V}\right\}$
and when the group $\mathcal{G}$ is non-trivial $(X,\Delta)$ is Gibbs polystable iff $\Delta$ is of the form $\Delta_{w}$ for some $w\in[0,1[.$ Moreover, for $N>3$,
$\gamma^{(N)}(X,0)^{\mathcal{G}}=\frac{N-1}{\lfloor N/2\rfloor},\,\,\,\,\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{2(1-1/N)}{V},\frac{N-1}{\lfloor N/2\rfloor-1}\right\}\,\text{if \,$w\neq 0.$}$
For $N=2,3$, $\gamma^{(N)}(X,0)^{\mathcal{G}}=\infty$, and $\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}=\frac{2(1-1/N)}{V}$. Additionally,
$\gamma(X,0)^{\mathcal{G}}=\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2,\,\text{and\,}\,\gamma(X,\Delta_{w})^{\mathcal{G}}=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}\,\text{if}\,w\neq 0.$
The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ [37, 38] (see Section 4). We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1). This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant implicitly appearing in [63, 77], obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its centralizer) – see Section 4.2.
Beyond complex curves, Gibbs stability has turned out to be notoriously difficult to establish, and at present no examples are known apart from the case of exceptional Fano orbifolds, where it holds automatically. But in the companion paper [5] we exhibit non-exceptional log Fano manifolds $(\mathbb{P}^{2},\Delta)$ that are Gibbs polystable already at the minimal level (i.e. for the smallest $k\in]0,\infty[$ such that $kK_{(X,\Delta)}$ defines a line bundle):
Theorem 1.4.
[5] Let $\Delta_{w}$ be the torus-invariant divisor on $\mathbb{P}^{2}$ assigning the same weight $w\in]0,1[$ to each of the three torus-invariant hyperplanes. Then the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ is Gibbs polystable at the minimal level iff $w>3/4.$
The log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ may, when w=11/mw=1-1/m for a positive integer $m,$ be identified with the Fano orbifold defined as the quotient of $\mathbb{P}^{2}$ under the standard action of $(\mathbb{Z}/m\mathbb{Z})^{2}.$ In this case the previous theorem applies to the AdS/CFT correspondence, where the Fano orbifold in question encodes a particular toric quiver gauge theory.
1.3. Explicit symmetry breaking and the emergence of a unique Kähler–Einstein metric
Next, we break the $\mathcal{G}$–symmetry explicitly by fixing a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and a $\mathcal{K}$–invariant metric on $K_{(X,\Delta)}$ with positive curvature. This metric defines a $\mathcal{K}$–invariant Kähler form $\omega_{0}$ in $c_{1}(-K_{(X,\Delta)})$ and thus a symplectic form on $X.$ The natural action of $\mathcal{K}$ on $K_{(X,\Delta)}$ gives rise to a moment map
(1.9) $\boldsymbol{m}\colon X\rightarrow\mathfrak{k}^{*}$
from $X$ into the dual of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}.$ We will show that if $(X,\Delta)$ admits some Kähler–Einstein metric, then there exists a unique such metric whose normalized volume form $\mu_{\text{KE}}$ has vanishing moment, in the sense that
(1.10) $\mu_{\text{KE}}\in\mathcal{P}(X)_{0}\coloneqq\left\{\mu\in\mathcal{P}(X)\colon\int_{X}\boldsymbol{m}\mu=0\right\},$
where $\mathcal{P}(X)$ denotes the space of all probability measures $\mu$ on $X$ (see Theorem 3.7). The volume form $\mu_{\text{KE}}$ is $\mathcal{K}$-invariant, and we will propose an essentially explicit probabilistic construction of $\mu_{\text{KE}}.$ To this end denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ normalized so that
$\boldsymbol{m}_{N}(x_{1},\dots,x_{N})=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{m}(x_{i})\in\mathfrak{k}^{*}.$
Given a positive integer $N$ such that 0 is a regular value of $\boldsymbol{m}_{N}$ we denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 5.1 for the general definition. The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 5.11), i.e. that
(1.11) $\mu_{\boldsymbol{m}_{N}}^{(N)}\coloneqq\frac{1}{Z_{\boldsymbol{m}_{N}}}\nu_{\boldsymbol{m}_{N}}^{(N)},\,\,\,\,Z_{\boldsymbol{m}_{N}}\coloneqq\int\nu_{\boldsymbol{m}_{N}}^{(N)}$
defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ is symmetric, i.e. invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant. The original $\mathcal{G}$–symmetry has thus been reduced to a $\mathcal{K}$–symmetry. Accordingly, the probability measure 1.11 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e. $\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.11.
Before stating our two main conjectures we note that in contrast to the $\mathcal{K}$–reduced probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)},$ the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ depends on a choice of bases in $H^{0}(X,-kK_{(X,\Delta)}).$ As a consequence, so does its total mass $Z_{\boldsymbol{m}_{N}}.$ In what follows we will take the basis to be orthonormal with respect to the $L^{2}$–norm on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the fixed metric on $K_{(X,\Delta)}.$ The corresponding constant $Z_{\boldsymbol{m}_{N}}$ shall then be called the $\mathcal{K}$-reduced partition function, in view of its statistical mechanical interpretation described in Section 5.1. The fixed metric on $K_{(X,\Delta)}$ also determines the Mabuchi functional $\mathcal{M}$ on the space of Kähler metrics in $c_{1}(-K_{(X,\Delta)})$ (which, a priori, is only defined up to an additive constant).
Conjecture 1.5.
Let $(X,\Delta)$ be a log Fano manifold and assume that $(X,\Delta)$ admits a Kähler–Einstein metric. Then
(1.12) $\text{(i)}-\lim_{N\rightarrow\infty}N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}.$
As a consequence, $Z_{\boldsymbol{m}_{N}}<\infty$ for $N$ sufficiently large and viewing the empirical measure $\delta_{N}$ as a random variable on $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right),$
(1.13) $\text{(ii) }\lim_{N\rightarrow\infty}\delta_{N}=\mu_{\text{KE}},$
in probability, where $\mu_{\text{KE}}$ is the normalized volume form of the unique KE metric for $(X,\Delta)$ with vanishing moment, i.e. $\mu_{\text{KE}}\in\mathcal{P}(X)_{0}.$
This conjecture is motivated by a heuristic mean field type argument, outlined in Section 5.2, suggesting the following Large Deviation Principle (LDP) on the space $\mathcal{P}(X)_{0},$ as $N\rightarrow\infty$:
(1.14) $(\delta_{N})_{*}\nu_{\boldsymbol{m}_{N}}^{(N)}\sim e^{-NF(\mu)},$
where $F$ is the free energy functional, introduced in [10], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$. We show that the validity of such a LDP would imply both the asymptotics 1.12 and the convergence 1.13, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$. Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
Our next main result establishes the first part of the previous conjecture when $n=1$ with an explicit rate of convergence:
Theorem 1.6.
The convergence 1.12 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1),$ when $N$ is taken as the sequence of odd positive integers. More precisely,
$N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}+O\left(\frac{\log N}{N}\right).$
As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.13) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 5.2).
1.4. Outline of the proof of Theorem 1.6 and connections to arithmetic geometry
The starting point of the proof of Theorem 1.6 is the observation that the top form on $X^{N}$ defined by $\nu^{(N)}$ is $\mathcal{G}$–invariant (since the quantized Futaki character of $(X,\Delta)$ vanishes). Moreover, the corresponding integral over $\left\{\boldsymbol{m}_{N}=0\right\}$ is shown to be comparable to the following integral over the GIT quotient of $X^{N}$ under $\mathcal{G}$:
(1.15) $\int_{(X^{N})_{\mathrm{ss}}/\mathcal{G}}\nu^{(N),\mathrm{cc}},$
where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.15 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60]. Such integrals can be computed explicitly in terms of Gamma values, as first shown in [53] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.6, by exploiting some connections to K-stable arithmetic log Fano curves in [3].
It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.15 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 7, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86]. The proof of Theorem 1.6 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.
1.5. Outlook on valuative aspects
The present probabilistic framework also suggests the following definition of a divisorial polystability threshold $\delta(X,\Delta)_{\text{ps}\text{ }}$ :
$\delta(X,\Delta)_{\text{ps}\text{ }}\coloneqq\inf_{\mu\in\mathcal{P}(X^{\text{div}})_{\text{ss}}}\frac{\int_{X^{\text{div}}}A_{(X,\Delta)}\mu}{\left\|\mu\right\|},$
where $X^{\text{div}}$ denotes the space of all divisorial valuations $v$ over $X,$ $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ (formula 2.6), $\left\|\mu\right\|$ denotes the non-Archimedean pluricomplex energy of a probability measure $\mu$ on $X^{\text{div}}$ with respect to $K_{(X,\Delta)}$ [35] and $\mathcal{P}(X^{\text{div}})_{\text{ss}}$ denotes the space of all probability measures $\mu$ on $X^{\text{div}}$ of the following form (where $N$ is the sequence defined by formula 1.6):
$\mu=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\delta_{v_{j}^{(N)}},$
for a sequence of $N$-tuples of divisorial valuations $(v_{1}^{(N)},\dots,v_{N}^{(N)})$ over $X,$ for which the set $\{v_{j}^{(N)}\}$ is finite (as $j$ and $N$ both vary) – so that $\mu$ has finite support – and the center of the associated decomposable divisorial valuation over $X^{N}$ intersects the semistable locus $(X^{N})_{\text{ss }}$ of $X^{N}.$ When $\mathcal{G}$ is trivial, $\delta(X,\Delta)_{\text{ps}\text{ }}$ coincides with the standard stability threshold $\delta(X,\Delta)$ [57, 30], since one may then restrict to measures $\mu$ supported on a single divisorial valuation $v$. We propose the following conjecture, whose validity would imply Conjecture 1.2:
Conjecture 1.7.
Let $(X,\Delta)$ be a log Fano manifold with vanishing Futaki character. Then
  • $(X,\Delta)$ is $K$-polystable iff $\delta(X,\Delta)_{\text{$\mathrm{ps}$}\text{ }}>1.$
  • $\lim_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)^{\mathcal{G}}=\delta(X,\Delta)_{\text{ps }}$
  • $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}=\delta(X,\Delta)_{\text{ps }}$
The convergence in the second item above is motivated by the approach for proving the convergence of $\gamma^{(N)}(X,\Delta)$ towards $\delta(X,\Delta)_{\text{ }}$ outlined in [17, Section 5], using non-Archimedean pluripotential theory. The validity of the third item would imply the first item, by [48]. It should be emphasized that, in general, the invariant $\delta(X,\Delta)_{\text{ps }}$ differs from the reduced $\delta$-invariant of $(X,\Delta),$ introduced in [93], which is defined with respect to a fixed maximal complex torus in $\mathcal{G}.$ Our results imply Conjecture 1.7 for log Fano curves, but we leave the proof and further investigations of Conjecture 1.7 for the future.
1.6. Acknowledgements
This work was supported by a Wallenberg Scholar grant from the Knut and Alice Wallenberg foundation.
1.7. Organization of the paper
  • Section 2 is concerned with algebro-geometric aspects of the $N$-fold product $X^{N},$ leveraging the Fulton–MacPherson compactification. The algebro-geometric part of Theorem 1.3 is established, giving explicit formulas for the microscopic stability thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ of log Fano curves. As a consequence, Conjecture 1.1 (linking Gibbs polystability to K-polystability) holds on log Fano curves.
  • In Section 3 the role of $X^{N}$ is played by the space $\mathcal{P}(X)$ of all probability measures on $X.$ The remaining analytic part of Theorem 1.3 is established, giving a formula for the reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ on any log Fano curve with vanishing Futaki invariant in Theorem 3.13. Combined with the results in the previous section this shows that the large $N$-limit of $\gamma^{(N)}(X,\Delta)\mathcal{{}^{G}}$ coincides with $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}},$ thus confirming Conjecture 1.2 on log Fano curves. The starting point of the proof of Theorem 3.13 is a thermodynamical formula for $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}.$ This formula is established on any log Fano manifold in Section 3.3, where the structure of the space $\mathcal{P}(X)_{0}$ of all probability measures on $X$ with vanishing moments is studied and new analytic characterizations of log Fano manifolds admitting Kähler–Einstein metrics are given in Theorem 3.7.
  • In Section 4 applications of Theorem 3.13 to logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
  • In Section 5 we study the $\mathcal{K}$–reduced probability measures $\mu_{\boldsymbol{m}_{N}}^{(N)}$ on $X^{N}$ and the corresponding partition function $Z_{\boldsymbol{m}_{N}}$ appearing in Conjecture 1.5. The asymptotics as $N\rightarrow\infty$ are described in a overarching conjectural Large Deviation Principle, formulated in Conjecture 5.4, that we show implies both Conjecture 1.1 and Conjecture 1.2.
  • In Section 6 Theorem 1.6 is established, showing that $N^{-1}\log Z_{\boldsymbol{m}_{N}}$ converges towards the inf of the Mabuchi functional on any log Fano curve with vanishing Futaki invariant.
  • In Section 7 relations to arithmetic geometry are described
  • The appendix provides a formula for the free energy of a toric log Fano curve
2. Gibbs polystability and microscopic stability thresholds of log Fano manifolds
In this section we will calculate the microscopic stability thresholds of log Fano curves appearing in Theorem 1.3. This will prove Conjecture 1.1 on log Fano curves, linking Gibbs polystability to K-polystability.
2.1. Setup
We start by describing in more detail the general setup, introduced in Section 2. Let $(X,\Delta)$ be a log Fano manifold, i.e. $X$ is a complex projective manifold and $\Delta$ is an effective $\mathbb{Q}$–divisor on $X$ such that the log anticanonical line bundle $K_{(X,\Delta)}$ of $(X,\Delta)$ is ample. Recall that $K_{(X,\Delta)}$ is the $\mathbb{Q}$–line bundle defined as the tensor power of the canonical line bundle $K_{X}$ of $X$ and the $\mathbb{Q}$–line bundle defined by $\Delta,$ i.e. $K_{(X,\Delta)}\coloneqq K_{X}+\Delta,$ using the standard additive notation for tensor powers of line bundles that we shall adopt. Fix a positive integer $k$ such that $k\Delta$ is a divisor – rather than a mere $\mathbb{Q}$–divisor – so that $kK_{(X,\Delta)}$ defines a line bundle and set
$N\coloneqq\dim H^{0}(X,-kK_{(X,\Delta}).$
Denote by $s_{\Delta}$ a (multi-valued) section cutting out the $\mathbb{Q}$–divisor $\Delta$ and by $\Delta_{N}$ the $\mathbb{Q}$–divisor on the $N$-fold product $X^{N}$ cut out by $\bigotimes_{i=1}^{N}\pi_{i}^{*}s_{\Delta}$, where $\pi_{i}\colon X^{N}\rightarrow X$ is the projection to the $i^{\mathrm{th}}$ factor. Thus $(X^{N},\Delta_{N})$ defines a log Fano manifold. We will be concerned with the effective $\mathbb{Q}$–divisor $\mathcal{D}^{(N)}$ on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ whose support consists of all configurations $(x_{1},\dots,x_{N})$ of $N$ points on $X$ which are in “bad position” with respect to $H^{0}(X,-kK_{(X,\Delta)})$:
$\text{Supp}\left(\mathcal{D}^{(N)}\right)\coloneqq\left\{(x_{1},\dots,x_{N})\colon\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})\right\}$
Concretely, $k\mathcal{D}^{(N)}$ is the divisor on $X^{N}$ cut out by the section $\det S^{(N)}$ of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ defined by
(2.1) $(\det S^{(N)})(x_{1},x_{2},\dots,x_{N})\coloneqq\det(s_{i}(x_{j})),$
in terms of a fixed basis $s_{1},\dots,s_{N}$ in $H^{0}(X,-kK_{(X,\Delta)}).$ We will denote by $\mathcal{G}$ the complex Lie group defined by
$\mathcal{G}\coloneqq\text{Aut}_{0}(X,\Delta),$
consisting of automorphisms of $X,$ homotopic to the identity, preserving $\Delta.$ The group $\mathcal{G}$ naturally acts (diagonally) on $X^{N}$ and it also acts on $H^{0}(X,-kK_{(X,\Delta)}),$ using that the action of $\mathcal{G}$ naturally lifts to $kK_{(X,\Delta)}.$
2.1.1. Microscopic stability thresholds and Gibbs polystability
In [13], the following microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ was introduced:
(2.2) $\gamma^{(N)}(X,\Delta)\coloneqq\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),\,\,\,\,\gamma(X,\Delta)\coloneqq\liminf_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)$
expressed as the log canonical threshold of $\mathcal{D}^{(N)}$ on $(X^{N},\mathcal{D}^{(N)})$ (whose definition is recalled in Section 2.2.1 below). Following [13], $(X,\Delta)$ is called Gibbs stable at level $k$ if $\gamma^{(N)}(X,\Delta)>1$ and Gibbs stable iff $\gamma^{(N)}(X,\Delta)>1,$ when $N$ is sufficiently large. Furthermore, $(X,\Delta)$ is called Gibbs semistable if $\gamma(X,\Delta)\geq 1$ and uniformly Gibbs stable if $\gamma(X,\Delta)>1$ [13, 57]. If $(X,\Delta)$ is Gibbs stable at some level $k,$ then the group $\mathcal{G}$ is necessarily trivial [20]. Assuming that $\mathcal{G}$ is reductive, we define the reduced microscopic stability threshold $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ by
(2.3) $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}\coloneqq\text{lct}\left((X^{N})_{\text{ss}},\Delta_{N};\mathcal{D}^{(N)}\right),\,\,\,\,\gamma(X,\Delta)^{\mathcal{G}}\coloneqq\liminf_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$
where $(X^{N})_{\text{ss}}$ denotes the Zariski open subset of $X^{N}$ defined as the semistable locus of $X^{N}$ with respect to the natural action of the complex reductive Lie group $\mathcal{G}$ on the polarized manifold $(X^{N},-K_{(X^{N},\Delta_{N})})$ (whose definition is recalled in Section 2.2.2 below). We propose the following
Definition 2.1.
A log Fano manifold $(X,\Delta)$ is called (uniformly) Gibbs polystable if it is Gibbs semistable and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ for $N$ sufficiently large ($\gamma(X,\Delta)^{\mathcal{G}}>1).$
The reductivity of the group $\mathcal{G}$ is thus part of the definition of (uniform) Gibbs polystability. We expect that, in the previous definition, the condition of Gibbs semistability can be replaced by the condition that the Futaki character $\text{Fut}_{(X,\Delta)}$ of $(X,\Delta)$ vanishes (the definition of $\text{Fut}_{(X,\Delta)}$ and its quantizations $\text{Fut}_{(X,\Delta),k}$ is recalled in Section 3.2.6). This would follow from the validity of the Large Deviation Principle appearing in Conjecture 5.4 (and we will show that it is the case on log Fano curves). Accordingly, it seems natural to also introduce the following definition, involving the quantized Futaki character $\text{Fut}_{(X,\Delta),k}$:
Definition 2.2.
A log Fano manifold $(X,\Delta)$ is called Gibbs polystable at level $k$ if $\gamma^{(N_{k})}(X,\Delta)^{\mathcal{G}}>1$ and the quantized Futaki character $\text{Fut}_{(X,\Delta),k}$ vanishes.
This definition also applies to some positive rational values for $k,$ as long as $kK_{(X,\Delta)}$ is defined as a line bundle. For example, when $X=\mathbb{P}^{n}$, $kK_{X}$ is defined as a line bundle iff $k(n+1)$ is an integer, since $K_{X}=\mathcal{O}(n+1).$
Remark 2.3.
When $\mathcal{G}$ is non-trivial we do not expect that Gibbs polystability is equivalent to Gibbs polystability at level $k$ for all sufficiently large positive integers $k$ (or even suitable rational $k).$ Indeed, there are examples of toric Fano manifolds, where the Futaki character $\text{Fut}_{(X,\Delta)}$ vanishes, while its quantizations do not vanish. However, we do expect that in the definitions of both $\gamma(X,\Delta)$ and $\gamma(X,\Delta)^{\mathcal{G}}$ the liminf can be taken either over all positive integers $k$ such that $k\Delta$ is a divisor, or over all rational $k$ for which $kK_{(X,\Delta)}$ is defined as a line bundle (for example, we will show that this is the case on log Fano curves). Anyhow, in the definitions we will, for simplicity, demand that $N$ only ranges over all $N_{k}$ for which $k$ is a positive integer and $k\Delta$ is a divisor.
2.2. Preliminaries
To compute the microscopic stability thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ on log Fano curves $(X,\Delta)$ we first recall some preliminary notions and results.
2.2.1. Log canonical thresholds
Let $(X,\Delta)$ be a log pair, defined here as a normal projective variety $X,$ that we for simplicity assume is non-singular, together with an effective $\mathbb{Q}$–divisor $\Delta$ on $X$ such that $K_{X}+\Delta$ defines a $\mathbb{Q}$–line bundle. Moreover, we will always fix a (multivalued) holomorphic section $s_{\Delta}$ cutting out $\Delta$. The log canonical threshold of a $\mathbb{Q}$–divisor $D$ on $X$ with respect to $(X,\Delta)$ may, in analytic terms, be defined as the following integrability threshold:
[…]
where $s_{D}$ is a (multivalued) holomorphic section cutting out $D$ and $||\cdot||_{D}$, $||\cdot||_{\Delta}$ denotes any fixed smooth metrics on the line bundles corresponding to $D$ and $\Delta$, respectively. In order to calculate the log canonical thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ we will instead use the well-known valuative definition of $\mathrm{lct}(X,\Delta;D)$ (see, for example, [25, Thm B.5]):
(2.5) $\mathrm{lct}(X,\Delta;D)=\inf_{v}\mathrm{lct}(X,\Delta;D)_{v},\,\,\,\,\mathrm{lct}_{v}(X,\Delta;D)\coloneqq\frac{A_{(X,\Delta)}(v)}{v(D)},$
where the inf ranges over all divisorial valuations $v$ over $X$ and $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ with respect to $(X,\Delta).$ We recall that a divisorial valuation $v$ over $X$ is an integer-valued function on the space $\mathbb{C}(X)$ of all rational functions $f$ on $X$ of the form
$v(f)=\mathrm{ord}_{E}(\pi^{*}f)$
where $\pi\colon Y\rightarrow X$ is a birational morphism from a non-singular variety $Y,$ $E$ is a prime divisor on $Y$ and $\mathrm{ord}_{E}g$ denotes the order of vanishing of a given rational function $g$ along $E.$ The subvariety $\pi(E)$ is called the center of the divisorial valuation. The action of a valuation on $\mathbb{C}(X)$ naturally extends to sections of line bundles (since the action is invariant under scaling by a non-zero holomorphic function) and thus to divisors $D.$ Moreover, by imposing homogeneity it also extends to $\mathbb{Q}$–divisors, so that $v$ takes rational values. [^1] These divisorial valuations are usually called normalized divisorial valuations, with a general divisorial valuation being a positive multiple of a normalized divisorial valuation, but here we will stick to the above notation. The log discrepancy $A_{(X,\Delta)}(v)$ of a divisorial valuation $v$ is defined as
(2.6) $A_{(X,\Delta)}(v)=1+\mathrm{ord}_{E}(\pi^{*}K_{(X,D)}-K_{Y}),$
where $\pi^{*}K_{(X,D)}-K_{Y}$ defines a divisor, supported on the exceptional divisors of $\pi.$
Example 2.4.
Any smooth subvariety $Z$ of $X$ determines a divisorial valuation that we will refer to as $v_{Z}$, defined by $v=\mathrm{ord}_{E}(\pi^{*}(\cdot))$, where $E$ is the exceptional divisor of the blowup $\pi\colon\mathrm{Bl}(X)\rightarrow X$. The center of $v$ is simply $Z$ and $A_{(X,\Delta)}(v)$ is the codimension of $Z.$
Let $\pi\colon Y\rightarrow X$ be a log resolution of $(X,D+\Delta),$ i.e. a birational morphism that is an isomorphism over the complement of $\text{supp}(D+\Delta)$ such that $\pi^{*}(D+\Delta)$ is a divisor on $Y$ with simple normal crossings. Denoting by $E_{1},\dots,E_{m}$ its components, the inf in 2.5 can be restricted to the divisorial valuations defined by $E_{1},\dots,E_{m}$. Using the existence of log resolutions it then follows readily that the two definitions 2.4 and 2.5 of the log canonical threshold agree.
We will use the following basic product property: if $(X_{1},\Delta_{1})$ and $(X_{2},\Delta_{2})$ are log pairs, and $D_{1},D_{2}$ are effective $\mathbb{Q}$–divisors on $X_{1}$ and $X_{2}$ respectively, then
(2.7) $\mathrm{lct}(X_{1}\times X_{2},\mathrm{pr_{1}^{*}(\Delta_{1})+\mathrm{pr}_{2}^{*}(\Delta_{2});\mathrm{pr_{1}^{*}(D_{1})+\mathrm{pr}_{2}^{*}(D_{2}))=\min_{i=1,2}\mathrm{lct}(X_{i},\Delta_{i};D_{i})}.}$
[…]
Next, if $U$ is an open subset in the Euclidean topology of $X$, then $\mathrm{lct}(U,\Delta|_{U},D|_{U})$ designates the inf 2.5 restricted to valuations whose center have non-zero intersection with $U$. With this definition, analogous versions of the above properties hold, and moreover, if $X=\cup_{i}U_{i}$, then clearly,
$\mathrm{lct}(X,\Delta;D)=\inf_{i}\mathrm{lct}(U_{i},\Delta|_{U_{i}},D|_{U_{i}}).$
A divisorial valuation $v$ over $X$ is said to compute $\mathrm{lct}(X,\Delta;D)$ if the inf 2.5 is attained at $v.$ More generally, when considering a sequence $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ a sequence $v_{N}$ of valuations over $X_{N}$ will be said to asymptotically compute $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ if $\text{lct}_{v_{N}}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right)$ converges towards $\liminf_{N\rightarrow\infty}\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),$ as $N\rightarrow\infty.$ Finally, note that any $N$-tuple of divisorial valuations $v_{1},\dots,v_{N}$ over $X$ induces a valuation over $X^{N}$ by blowing up the variety $E_{1}\times\cdots\times E_{N}$ over $X^{N},$ where $E_{i}$ is the divisor over $X$ corresponding to $v_{i}.$ Such a divisorial valuation over $X^{N}$ will be called decomposable.
2.2.2. Semistability in GIT
Let $L\rightarrow Y$ be an ample line bundle over a complex projective manifold $Y$ and $G$ a complex reductive group acting on $(Y,L).$ The corresponding semistable locus (in sense of Geometric Invariant Theory) is the Zariski open subset $Y_{\text{ss}}$ of $Y$ consisting of all $y\in Y$ such that there exists a GG–invariant section $s_{k}\in H^{0}(Y,kL)$ for some positive integer $k$ such that $s_{k}(y)\neq 0$ [64]. When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathrm{PGL}(2,\mathbb{C}),$ acting diagonally, $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the the fraction $I/N$ of points $x_{i}$ coinciding satisfies $I/N\leq 1/2$, i.e.
(2.8) $\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}\right)\{x\}\leq 1/2$
for all $x\in\mathbb{P}^{1}$ (see [85, Section 3]). When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathbb{C}^{*}$ the semistable locus $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the inequality 2.8 holds for the two fixed points $x$ of the action of $\mathbb{C}^{*}$ on $\mathbb{P}^{1}.$ Just like the case $G=\mathrm{PGL}(2,\mathbb{C}$) this follow readily from the Hilbert–Mumford criterion for one-parameter subgroups of $\mathbb{C}^{*}\subset G.$
2.2.3. The Fulton–MacPherson compactification of configuration spaces
In [58] a canonical algebraic compactification $X^{[N]}$ of the configuration space $X^{N}\setminus\cup_{i\neq j}\{x_{i}=x_{j}\}$ of a non-singular algebraic variety $X$ was constructed. It has the following useful properties:
  1. (1)
    $X^{[N]}$ is a non-singular variety, projective if $X$ is [58, Theorems 1 and 2].
  2. (2)
    There is a canonical morphism $\pi_{\mathrm{FM}}\colon X^{[N]}\rightarrow X^{N}$, which restricts to an isomorphism $\pi_{\mathrm{FM}}\colon X^{[N]}\setminus Z\rightarrow X^{N}\setminus\cup_{i\neq j}\{x_{i}=x_{j}\},$ where $Z$ defines a simple normal crossings divisor [58, Theorem 3].
  3. (3)
    There is a bijection between the set $\mathcal{W}$ of all submanifolds $W$ of $X^{N}$ of the form
    $W=\left\{x_{i_{1}}=x_{i_{2}}=\cdots=x_{i_{r}}=0\right\}\subset X^{N}$
    and a certain set of prime divisors $E_{W}^{\mathrm{}}$ on $X^{[N]}$, such that for every $W\in\mathcal{W}$,
    $\pi_{FM}^{*}(\mathcal{I}(W))=\mathcal{I}\left(\sum_{W^{\prime}\subset W}E_{W^{\prime}}\right)$
    where $\mathcal{I}(\cdot)$ denotes the corresponding ideal sheaves and the sum ranges over all $W^{\prime}$ in $\mathcal{W}$ contained in $E$ [58, Theorem 3].
Note that when $\mathrm{dim(X)=1}$, the first two properties imply that the map $\pi_{FM}$ is a log resolution of the log pair $(X^{N},\sum_{i<j}\{x_{i}=x_{j}\})$.
Given $W\in\mathcal{W},$ we denote by $v_{W}^{\mathrm{FM}}$ the divisorial valuation corresponding to $E_{W}$. We can also consider the valuation $v_{\mathrm{W}}$ as in Example 2.4. The following Lemma shows that these two valuations associated to a $W\in\mathcal{\mathcal{W}}$ are the same.
Lemma 2.5.
For any $W\in\mathcal{W},$ $v_{W}^{\mathrm{FM}}=v_{W}$.
Proof.
[…]
$\begin{array}[]{rcl}X^{[N]}&\xrightarrow{f}&\mathrm{{Bl}}_{W}(X^{N})\\ \pi_{FM}\searrow&&\swarrow\pi_{W}\\ &X^{N}\end{array}$
where $\pi_{W}\colon\mathrm{Bl}_{W}(X^{N})\rightarrow X^{N}$ is the blow-up morphism, whose exceptional divisor we denote by $E_{\pi_{W}}.$ It will thus be enough to show that $f_{*}^{-1}(E_{\pi_{W}})=E_{W^{\prime}},$ where $f_{*}^{-1}(E_{\pi_{W}})$ denotes the strict transform of $E_{\pi_{W}}$ under $f.$ To this end let us first observe that there exists $W^{\prime}$ in $\mathcal{W}$ such that
(2.9) $f_{*}^{-1}(E_{\pi_{W}})=E_{W^{\prime}},\,\,\,W^{\prime}\subset W$
Indeed, $\pi_{W}^{*}(\mathcal{I}(W))=\mathcal{I}(E_{\pi_{W}})$ and thus by the commutativity of the diagram, $f^{*}(\mathcal{I}(E_{\pi_{W}}))=f^{*}\circ\pi_{W}^{*}(\mathcal{I}(W))=\pi_{FM}(\mathcal{I}(E_{\pi_{W}})),$ which, by property 3 of the Fulton–MacPherson compactification above, coincides with the ideal defined by the divisor $\sum_{W^{\prime}\subset W}E_{W^{\prime}}.$ Since the $E_{W}$’s are prime this proves the identity 2.9. It follows from 2.9 that
$\pi_{FM}(E_{W^{\prime}})=W$
(since $\pi_{FM}(E_{W^{\prime}})=\pi_{W}\circ f(E_{W^{\prime}})=\pi_{W}(W)).$ But this implies that $W^{\prime}=W.$ Indeed, seeking a contradiction, assume that $W^{\prime}\neq W$. Then, by property 3 of the Fulton–MacPherson compactification above, $\pi_{FM}(E_{W^{\prime}})\subset\pi_{FM}(\bigcup_{W^{\prime\prime}\subset W^{\prime}}E_{W^{\prime\prime}})=W^{\prime}\subsetneq W$, which is a contradiction. ∎
When $\mathrm{dim}(X)=1$, the subset $\cup_{i<j}\{x_{i}=x_{j}\}$ defines a divisor on $X^{N}$ that we denote by $D_{N}$. The following lemma describes the log discrepancy and action of the valuations $v_{W}$ on $D_{N}$.
Lemma 2.6.
Let $\pi_{FM}\colon X^{[N]}\rightarrow X^{N}$ be the log resolution described above. For any $W\in\mathcal{W}$ we have that
$A_{(\mathbb{P}^{1})^{N}}(v_{W})=\mathrm{codim}W.$
Additionally,
$v_{W}(D_{N})=\frac{1}{k}\frac{\mathrm{codim}(W)(\mathrm{codim}W+1)}{2}.$
Proof.
[…]
2.3. The microscopic stability thresholds of log Fano curves
We now turn to the case when $(X,\Delta)$ is a log Fano curve. This means that $X=\mathbb{P}^{1}$ and $\Delta$ is any divisor of $r$ points $p_{1},\dots,p_{m}$ on $X$ with weights $w_{i}\in]0,1[$ satisfying
$V\coloneqq\deg(K_{X}+\Delta)=2-\sum_{i=1}^{m}w_{i}>0.$
We can thus identify $kK_{(\mathbb{P}^{1},\Delta)}$ with $kV\mathcal{O}(1),$ which shows that $N=kV+1.$ Now identifying $\mathbb{P}^{1}$ with $\mathbb{C}\cup\{\infty\}$ and taking the basis $e_{1},\dots,e_{N}$ appearing in the definition 2.1 of $\det S^{(N)}$ to consist of monomials $1,z,z^{2},\dots,z^{kV}$ we may, in the the affine chart $\mathbb{C},$ express
(2.10) $\det S^{(N)}=\prod_{i<j}^{N}(z_{i}-z_{j}),\,\,\,\,\,N=Vk+1,$
using the classical Vandermonde identity (and in general any point in $(\mathbb{P}^{1})^{N}$ has a Zariski–neighbourhood of this type, by symmetry).
2.3.1. The case $\Delta=0$
As a warm-up, we start by computing $\gamma^{(N)}(\mathbb{P}^{1}),$ thus reproving the Gibbs semistability of $\mathbb{P}^{1}$, first shown in [56], by algebraic means, and then in [18], by analytic means.
Proposition 2.7.
$\mathbb{P}^{1}$ is Gibbs semistable. More precisely, for any integer $N\geq 2$
$\gamma^{(N)}(\mathbb{P}^{1})\coloneqq\mathrm{lct}((\mathbb{P}^{1})^{N};\mathcal{D}^{(N)})=1-1/N$
and $\gamma^{(N)}(\mathbb{P}^{1})$ is computed by the divisorial valuation induced by the diagonal of $X^{N}$ (viewed as a one-dimensional submanifold). Moreover, for any given point $p$ on $\mathbb{P}^{1}$ the sequence of decomposable valuations $v_{N}$ over $(\mathbb{P}^{1})^{N}$ induced by $(p,p,\dots,p)\in(\mathbb{P}^{1})^{N}$ asymptotically computes $\mathrm{lct}((\mathbb{P}^{1})^{N};\mathcal{D}^{(N)}).$
Proof.
We have,
$\mathrm{lct}((\mathbb{P}^{1})^{N};\mathcal{D}^{(N)})=\min_{v}\frac{A_{(\mathbb{P}^{1})^{N}}(v)}{v(\mathcal{D}^{(N)})}=\min_{v=v_{W}\colon W\in\mathcal{W}}\frac{A_{(\mathbb{P}^{1})^{N}}(v)}{v(\mathcal{D}^{(N)})}$
where the first minimum ranges over divisorial valuations on $X$, which we can restrict to the ones of the form $v_{W},W\in\mathcal{W}$ since $\pi_{FM}\colon(\mathbb{P}^{1})^{[N]}\rightarrow(\mathbb{P}^{1})^{N}$ is a log resolution of $((\mathbb{P}^{1})^{N},\mathcal{D}^{(N)})$, and due to Lemma 2.5. Lemma 2.6 now yields
$\mathrm{lct}((\mathbb{P}^{1})^{N},\mathcal{D}^{(N)})$ $=k\min_{W\in\mathcal{W}}\frac{\mathrm{codim}W}{\mathrm{codim}W(\mathrm{codim}W+1)/2}=\min_{W\in\mathcal{W}}\frac{2k}{\mathrm{codim}W+1}$
$=\frac{2k}{2k+1}=\frac{N-1}{N}.$
For the last claim, note that the minimum is computed by any $W^{\prime}$ such that $\mathrm{codim}(W^{\prime})=N-1$ implying that in fact, $W^{\prime}=(p_{1}=\cdots=p_{N})$. Finally, for the valuation $v_{N}$ induced by $(p,p,\dots,p)\in(\mathbb{P}^{1})^{N}$ we have, since by Example 2.4 $A(v_{N})=N$ and $v_{N}(z_{i}-z_{j})=1,$
$\frac{A(v_{N})}{v_{N}(\mathcal{D}^{(N)})}=\frac{N}{N(N-1)/2k}=1,\,\,\,\,\,\,(2k=N-1).$
In a similar way we next prove the uniform Gibbs polystability of $\mathbb{P}^{1}$:
Proposition 2.8.
$\mathbb{P}^{1}$ is uniformly Gibbs polystable. More precisely, for an integer $N>3$,
[…]
For $N=2,3$, $\gamma^{(N)}(\mathbb{P}^{1}\mathrm{)^{\mathcal{G}}}=\infty$.
  • In case $N>3$, $\gamma^{(N)}(\mathbb{P}\mathrm{{}^{1})^{\mathcal{G}}}$ is computed by the the divisorial valuations over $(X^{N})_{\mathrm{ss}}$ induced by the submanifolds $\{x_{i_{1}}=\cdots=x_{i_{\lfloor N/2\rfloor}}\}\subset X^{N}$ for any given index $I$ of length $\lfloor N/2\rfloor.$
  • A decomposable sequence $v_{N}$ of divisorial valuations over $(X^{N})_{\mathrm{ss}}$ asymptotically computing $\gamma^{(N)}(\mathbb{P}\mathrm{{}^{1})^{\mathcal{G}}}$ is obtained by fixing three distinct points $p_{1},p_{2}$ and $p$ on $\mathbb{P}^{1}$ and letting $v_{N}$ be the valuation induced by $\{p_{1}\}^{\lfloor N/2\rfloor}\times\{p\}\times\{p_{2}\}^{\lfloor N/2\rfloor}\in(\mathbb{P}^{1})^{N}$ when $N$ is odd and by $\{p_{1}\}^{N/2}\times\{p_{2}\}^{N/2}\in(\mathbb{P}^{1})^{N}$ when $N$ is even.
Proof.
By Proposition 2.7, $\mathbb{P}^{1}$ is Gibbs semistable. The map $\pi_{FM}\colon Y\rightarrow(\mathbb{P}^{1})_{\mathrm{ss}}^{N}$ given by restricting $\pi_{FM}$ to $Y\coloneqq\pi_{FM}^{-1}((\mathbb{P}^{1})_{\mathrm{ss}}^{N})$ is a log resolution of $((\mathbb{P}^{1})_{\mathrm{ss}}^{N},\mathcal{D}^{(N)})$. Denote by $\mathcal{W}^{\mathrm{ss}}$ the subvarieties in $\mathcal{W}$ that have non-empty intersection with the semistable locus. By the explicit description in Section 2.2.2, $\mathcal{W}^{\mathrm{ss}}$ are the elements $W$ of $\mathcal{W}$ such that $\mathrm{codim}W\leq\lfloor N/2\rfloor-1$, which is non-empty iff $N>3$. Consequently $\mathrm{lct}((\mathbb{P}^{1})_{\mathrm{ss}}^{N},\mathcal{D}^{(N)})=\infty$ if $N=2,3$ and otherwise
$\mathrm{lct}((\mathbb{P}^{1})_{\mathrm{ss}}^{N},\mathcal{D}^{(N)})=k\min_{W\in\mathcal{W}^{\mathrm{ss}}}\frac{\mathrm{codim}W}{\mathrm{codim}W(\mathrm{codim}W+1)/2}=\min_{W\in\mathcal{W}^{\mathrm{ss}}}\frac{2k}{\mathrm{codim}W+1}=\frac{N-1}{\lfloor N/2\rfloor}$
since the center of $v_{W}$ is precisely $W$. Finally, consider the valuation $v_{N}$ induced by $\{p_{1}\}^{N/2}\times\{p_{2}\}^{N/2}\in(\mathbb{P}^{1})^{N}$ when $N$ is even. By Example 2.4 $A(v_{N})=N$ and decomposing $\prod_{i<j}^{N}(z_{i}-z_{j})$ into the three groups of indices $(i,j)$ where $i,j\leq N/2,$ $i,j>N/2$ and the rest reveals that $kv_{N}(\mathcal{D}^{(N)})=2\left((N/2)\left((N/2)-1\right)/2\right).$ Hence,
$\frac{A(v_{N})}{v_{N}(\mathcal{D}^{(N)})}=\frac{N}{(N/2)\left((N/2)-1\right)/k}=\frac{N(N-1)/2}{(N/2)\left((N/2)-1\right)}=2\frac{N-1}{N-2}$
converging to $2$ as $N\rightarrow\infty.$ The computation in the case when $N$ is odd is similar. ∎
2.3.2. The case $\Delta\neq 0$
We now turn to the case of $\Delta\neq 0$. Instead of constructing an explicit log resolution of $((\mathbb{P}^{1})^{N},\mathcal{D}^{(N)}+\Delta_{N})$, generalizing the Fulton–MacPherson compactification of $(\mathbb{P}^{1})^{N}$ (which, presumable, can be done using the results in [49] or [78]) we will pursue a different route that reduces the calculations to an application of the Fulton–MacPherson compactification of $(\mathbb{P}^{1})^{N+1}.$ We begin by computing the invariant $\gamma^{(N)}(X,\Delta).$
Proposition 2.9.
For any log Fano curve $(\mathbb{P}^{1},\Delta)$ and integer $N\geq 2$,
(2.11) $\gamma^{(N)}(X,\Delta)=\mathrm{lct}(X^{N},\Delta_{N};\mathcal{D}^{(N)})=\min\left\{\frac{2(1-1/N)}{V},\min_{l}\frac{2(1-w_{l})}{V}\right\}$
Moreover, if $\frac{2(N-1)}{VN}\leq\min_{l}\frac{2(1-w_{l})}{V},$ then $v_{\{x_{1}=\cdots=x_{N}\}}$ computes $\mathrm{lct}(X^{N},\Delta_{N};\mathcal{D}^{(N)})$ and if $\frac{2(N-1)}{VN}\geq\min_{l}\frac{2(1-w_{l})}{V}$, then $\mathrm{lct}(X^{N},\Delta_{N};\mathcal{D}^{(N)})$ is computed by any divisorial valuation of the form $v_{\{x_{1}=\cdots=x_{N}=p_{l}\}}$, where $l$ realizes the minimum in $\min_{l}\frac{2(1-w_{l})}{V}.$
Proof.
Given $w\in]0,1[$ and $p\in\mathbb{P}^{1}$ denote by $\Delta_{N}^{p,w}$ the divisor on $(\mathbb{P}^{1})^{N},$ invariant under permutations, that projects to the divisor $wp$ under the $N$ projections to $X.$
Step 1: The following formula holds:
(2.12) $\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N};\mathcal{D}^{(N)})=\min_{i}\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p_{i},w_{i}},\mathcal{D}^{(N)}).$
To prove this, consider the following covering of $(\mathbb{P}^{1})^{N}$ by open subsets,
(2.13) $(\mathbb{P}^{1})^{N}=U\cup\bigcup_{q\in\{p_{1},\dots,p_{m}\}^{\tilde{N}},\tilde{N}=1,\dots,N}U_{q},$
where $U=\{(x_{1},\dots,x_{N}\in(\mathbb{P}^{1})^{N}\colon x_{i}\neq p_{j}\forall i,j\}$ and $U_{q}=\{(x_{1},\dots,x_{N})\in(\mathbb{P}^{1})^{N}\colon d(x_{i},q_{i})<d_{0}\forall i\}$ where $d$ is the round metric on $\mathbb{P}^{1}$ and $d_{0}<\min_{i,j}d(p_{i},p_{j})/2$. Then
(2.14) $\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N};\mathcal{D}^{(N)})=\min(\mathrm{lct}(U,\Delta_{N}|_{U};\mathcal{D}^{(N)}|_{U}),\min_{q\in\{p_{1},\dots,p_{m}\}^{\tilde{N}},\tilde{N}=1,\dots,N}\mathrm{lct}(U_{q},\Delta_{N}|_{U_{q}};\mathcal{D}^{(N)}|_{U_{q}})).$
Since $\Delta_{N}|_{U}$ is trivial, we can use the Fulton–MacPherson compactification as a log resolution there, and we turn to the other cases. Without loss of generality we may assume that $q=(p_{i_{1}},\dots,p_{i_{1}},p_{i_{2}},\dots,p_{i_{2}},\dots,p_{i_{l}},\dots,p_{i_{l}})$ for $1\leq l\leq m$ and some $p_{i_{1}},\dots,p_{i_{l}}$ distinct and where $p_{i_{1}}$ is repeated $k_{1}$ times and $p_{i_{2}}$ is repeated $k_{2}$ times, and so on. Then $U_{q}=V_{1}\times\cdots\times V_{l}\times(\mathbb{P}^{1})^{r}$, for some $0\leq r\leq N-1$, where $V_{j}=\{(x_{1},\dots,x_{k_{j}})\in(\mathbb{P}^{1})^{k_{j}}\colon d(x_{j},p_{i_{j}})<d_{0}\forall j\}$. Observe that $\Delta_{N}|_{U_{q}}=\Delta_{k_{1}}^{p_{i_{1}},w_{i_{1}}}\boxplus\cdots\boxplus\Delta_{k_{l}}^{p_{i_{l}},w_{i_{l}}}\boxplus 0$. Furthermore, $\mathcal{D}^{(N)}|_{U_{q}}\eqqcolon\mathcal{D}_{k_{1}}\boxplus\cdots\boxplus\mathcal{D}_{k_{l}}\times\mathcal{D}_{r}$. By 2.7
$\mathrm{lct}(U_{q},\Delta_{N}|_{U_{q}};\mathcal{D}^{(N)}|_{U_{q}})=\min\left(\min_{1\leq j\leq l}\left(\mathrm{lct}\left(\left.V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}}\right|_{V_{j}};\left.\mathcal{D}_{k_{j}}\right|_{V_{j}}\right)\right),\mathrm{lct}\left((\mathbb{P}^{1})^{r};\mathcal{D}_{r}\right)\right).$
But for any $1\leq j\leq l$, $\pi^{*}(\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}})\leq\Delta_{N}^{p_{i_{j}},w_{i_{j}}}$, and $\pi^{*}(\mathcal{D}_{k_{j}})\leq\mathcal{D}^{(N)}$ where $\pi\colon(\mathbb{P}^{1})^{N}\rightarrow(\mathbb{P}^{1})^{k_{j}}$ is the projection and similarly $\pi^{*}(\mathcal{D}_{r})\leq\mathcal{D}^{(N)}$. In view of the definition 2.5 and 2.7, $\mathrm{lct}(V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}_{k_{j}})\geq\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ with $q^{\prime}=(p_{i_{j}},\dots,p_{i_{j}})$ but the numbers $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ are contenders in the inf 2.13. Moreover $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})=\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}^{(N)})$ and any such term is lesser or equal to $\mathrm{lct}(U,\Delta_{N}|_{U};\mathcal{D}^{(N)}|_{U})$ in 2.14, and also $\mathrm{lct}((\mathbb{P}^{1})^{r};\mathcal{D}_{r})$. This proves formula 2.12.
Step 2: The following formula holds
$\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p,w};\mathcal{D}^{(N)})=\frac{2(1-w)}{V}.$
To prove this, consider the affine chart $\mathbb{C}\subset\mathbb{P}^{1}$ centered at $p$ and the restrictions to $\mathbb{C}^{N}\subset(\mathbb{P}^{1})^{N}$ of the (multivalued) sections cutting out $\mathcal{D}^{(N)}$ and $\Delta_{N}^{p,w},$ respectively:
(2.15) $\left(\prod_{i<j}(z_{i}-z_{j})\right)^{-1/k}\text{and\,\,}\left(\prod_{i}z_{i}\right)^{w}.$
Now regard these as functions of $N+1$ variables $z_{1},\dots,z_{N}$ and $z_{N+1}$, which are independent of $z_{N+1}.$ In other words, we consider their pullbacks to $\mathbb{C}^{N+1}$ under the natural projection from $\mathbb{C}^{N+1}$ to $\mathbb{C}^{N}.$ This does not affect the computation of the corresponding log canonical threshold, by 2.7. After making the affine change of variables $z_{i}=u_{i}-u_{N+1}$ for $i=1,\dots,N$ and $z_{N+1}=u_{N+1}$ the two expressions 2.15 become
(2.16) $\left(\prod_{1\leq i<j\leq N}(u_{i}-u_{j})\right)^{-1/k}\text{and\,\,}\left(\prod_{1\leq i\leq N}(u_{i}-u_{N+1})\right)^{w}.$
We can thus use the Fulton–MacPherson compactification in the $N+1$ variables $u_{1},\dots,u_{N+1},$ $\pi_{FM}\colon\mathbb{P}^{[N+1]}\rightarrow\mathbb{P}^{N+1},$ as a log resolution of the pullback of $\mathcal{D}^{(N)}+\Delta_{N}^{p,w}.$ However, now there are two types of subvarieties $W$ in $\mathcal{W}$ that we need to treat differently when computing $\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p,w};\mathcal{D}^{(N)})$: those not involving $u_{N+1},$ that we denote by $\mathcal{W}^{\prime},$ and those involving $u_{N+1},$ that we denote by $\mathcal{W}_{p}.$ Denoting by $v_{W}$ the corresponding divisorial valuations over $\mathbb{P}^{N+1}$ we have
$v_{W}(\Delta^{p,w})=0,\,\text{if $W\in\mathcal{W}^{\prime}$},\,\,\,\,\,\,v_{W}(\Delta^{p,w})=w\mathrm{codim}W\,\text{if \,W$\in\mathcal{W}_{p}$,}$
as follows exactly as in Lemma 2.6, using Lemma 2.5. From the definition it follows that
$A_{((\mathbb{P}^{1})^{N},\Delta^{p,w})}(v_{W})=A_{(\mathbb{P}^{1})^{N}}(v_{W})-v_{W}(\Delta^{p,w})$
and consequently,
$\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p,w};\mathcal{D}^{(N)})$ $=\min\left(\min_{W\in\mathcal{W}^{\prime}}\frac{A_{((\mathbb{P}^{1})^{N+1},\Delta_{N}^{p,w})}(v_{W})}{v_{W}(\mathcal{D}^{(N)})},\min_{W\in\mathcal{W}_{p}}\frac{A_{((\mathbb{P}^{1})^{N+1},\Delta_{N}^{p,w})}(v_{W})}{v_{W}(\mathcal{D}^{(N)})}\right)$
$=\min\left(\min_{W\in\mathcal{W}^{\prime}}\frac{2k}{\mathrm{codim}W+1},\min_{W\in\mathcal{W}^{p}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}\right).$
Since
$\min_{W\in\mathcal{W}^{\prime}}\frac{2k}{\mathrm{codim}W+1}=\frac{2k}{N}=\frac{2(N-1)}{VN},$
we thus get
$\min_{W\in\mathcal{W}^{p}}\frac{\mathrm{codim}W-w_{l}\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}=\min_{W\in\mathcal{W}^{p}}\frac{2k(1-w)}{\mathrm{codim}W-1}=\frac{2k(1-w)}{N-1}=\frac{2(1-w)}{V},$
which concludes the proof of the formula in Step 2.
Finally, letting $\pi$ be the blowup along the subvariety $W^{\prime}\coloneqq\{x_{1}=\cdots=x_{N}\}$ and $\pi_{W^{\prime}}$ the corresponding divisorial valuation, we find that
$\frac{A_{((\mathbb{P}^{1})^{N},\Delta_{N})}(v_{W^{\prime}})}{v_{W^{\prime}}(\mathcal{D}^{(N)})}=\frac{2k}{N}=\frac{2(N-1)}{VN}$
which proves the claim when $\frac{2(N-1)}{VN}\leq\min_{l}\frac{2(1-w_{l})}{V}.$ The other case follows similarly by blowing up along a subvariety of the form $\{x_{1}=\dots=x_{N}=p_{l}\}$. ∎
The formula for $\gamma^{(N)}(X,\Delta)$ in the previous proposition corrects the formula given in [18, Thm 4.5], which is only correct modulo an error term of order $O(1/N)$ (that comes from the discrepancy between the upper and the lower bound on $\gamma^{(N)}(X,\Delta)$ established by analytic means in [18]).
2.3.3. The case of non-trivial $\mathcal{G}$
To investigate which log Fano curves are Gibbs polystable, but not Gibbs stable, we need to restrict to the cases with non-trivial $\mathcal{G}.$ This means the number $m$ of components of $\Delta$ satisfies $m\leq 2.$ It is also required that $(X,\Delta)$ be Gibbs semistable. This implies that there is some point $p_{i}$ in the support of $\Delta$ such that $(1-w_{i})=V/2,$ by the formula for $\gamma(X,\Delta)$ resulting from Proposition 2.9. If $m=1,$ then this equality implies that $V=0$ in which case $(X,\Delta)$ is not log Fano. We can thus restrict to $m=2$, since $m=0$ is the content of Proposition 2.8. In this case $w_{i}=w$ for some $w\in]0,1[,$ by Proposition 2.9. Such a divisor $\Delta$ on $\mathbb{P}^{1}$ will be denoted by $\Delta_{w}.$
Proposition 2.10.
Let $\Delta_{w}$ be the divisor on $\mathbb{P}^{1}$ supported on two distinct points $p_{1}$ and $p_{2}$ with the same weight $w\in]0,1[.$ $(\mathbb{P}^{1},\Delta_{w})$ is uniformly Gibbs polystable. More precisely, for $N=2,3$, $\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}=\frac{2(1-1/N)}{V}$ and for any $N\geq 3$,
$\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}\coloneqq\mathrm{lct}\left((\mathbb{P}^{1})_{\mathrm{ss}}^{N},\Delta_{N};\mathcal{D}^{(N)}\right)=\min\left(\frac{2(1-1/N)}{V},\frac{N-1}{\lfloor N/2\rfloor-1}\right)$
  • Now assume $N>3$. When $V\geq 2(\lfloor N/2\rfloor-1)/N$, $\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}$ is computed by the valuation induced by the diagonal in $X^{N}$ and when $V\leq 2(\lfloor N/2\rfloor-1)/N$, $\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}$ is computed by the divisorial valuations defined by the submanifolds in $X^{N}$ given by $\{x_{i_{1}}=\cdots=x_{i_{\lfloor N/2\rfloor}}=p_{1}\}$ and $\{x_{i_{1}}=\cdots=x_{i_{\lfloor N/2\rfloor}}=p_{2}\}$ for any given index $I$ of length $\lfloor N/2\rfloor$.
  • When $V\geq 1$, $\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}$ is asymptotically computed by the decomposable valuation $v_{N}$ induced by $(p,\dots,p)\in(\mathbb{P}^{1})^{N}$ for any given point $p$ not in the support of $\Delta_{w}$ and when $V\leq 1$, $\gamma^{(N)}(X,\Delta_{w})^{\mathcal{G}}$ is asymptotically computed by the sequence $v_{N}$ appearing in Prop 2.10, if $p_{1}$ and $p_{2}$ are the points in the support of $\Delta.$
Proof.
In this case the group $\mathcal{G}$ is isomorphic to $\mathbb{C}^{*}.$ Denote by $\mathcal{W}_{p}^{\mathrm{ss}}$ the elements of $\mathcal{W}_{p}$ from Lemma 2.9 that have a non-empty intersection with the corresponding semistable locus on $X^{N}$ (explicitly described in Section 2.2.2). Following closely the proof of Proposition 2.9, but taking into account that we restrict to the semistable locus, as in 2.8, we find that
$\mathrm{lct}((\left.(\mathbb{P}^{1})_{\mathrm{ss}}^{N})\right|_{U_{p}},\Delta_{N};\mathcal{D}^{(N)})=\min\left(\min_{W\in\mathcal{W}^{\prime}}\frac{2k}{\mathrm{codim}W+1},\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}\right)$
when $N>3$. We have, as in Proposition 2.9,
$\min_{W\in\mathcal{W}^{\prime}}\frac{2k}{\mathrm{codim}W+1}=\frac{2(N-1)}{VN},$
while
$\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}$ $=\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{2k(1-w)}{\mathrm{codim}W-1}$
$=\frac{2(N-1)(1-w)}{V(\lfloor N/2\rfloor-1)}=\frac{N-1}{(\lfloor N/2\rfloor-1)}.$
The cases $N=2,3$ is similar except that $\mathcal{W}_{p}^{\mathrm{ss}}$ is empty. The last part, concerning divisorial valuations that compute the log canonical threshold, follows similarly as in Proposition 2.9. ∎
2.4. Gibbs polystability of log Fano curves
We recall [74] that a log Fano curve $(X,\Delta)$ is K-(semi)stable iff $V/2>(\geq)1-w_{l}$ for all $l$ (and K-stable iff it is uniformly K-stable). Moreover, when $\mathcal{G}$ is non-trivial $(X,\Delta)$ is a K-polystable iff the Futaki character of $(X,\Delta)$ vanishes iff $\Delta=0$ or $\Delta=\Delta_{w}$ for $w\in]0,1[$ (as in Prop 2.10). Comparing with Propositions 2.7, 2.8, 2.9 and 2.10 we thus find:
Theorem 2.11.
Let $(X,\Delta)$ be a log Fano curve. Then $(X,\Delta)$ is Gibbs semistable if and only if $(X,\Delta)$ K-semistable. Moreover, $(X,\Delta)$ is (uniformly) Gibbs [poly]stable if and only if $(X,\Delta)$ is (uniformly) K-[poly]stable.
This proves Conjecture 1.1 on log Fano curves. It was previously known that Gibbs semistability was equivalent to K-semistability and that (uniform) Gibbs stability was equivalent to (uniform) K-stability (see [18]).
3. Analytic stability thresholds and thermodynamics
In this section we will, in particular, compute the reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (defined by formula 1.7) when $(X,\Delta)$ is a log Fano curve, thus completing the proof of Theorem 1.3.
3.1. Setup
We continue with the notation introduced in Section 2.1. In particular, the complex Lie group $\text{Aut}_{0}(X,\Delta)$ is denoted by $\mathcal{G}$ and we will assume that it is reductive, i.e. $\mathcal{G}$ is the complexification of any maximally compact subgroup of $\mathcal{G}.$ We will break the $\mathcal{G}$–symmetry by fixing a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G};$ the corresponding Lie algebras will be denoted by $\mathfrak{k}$ and $\mathfrak{g}.$ The elements of the complex Lie algebra $\mathfrak{g}$ may be identified with holomorphic vector fields $V^{1,0}$ on $X$ of type $(1,0)$ (whose flow preserves the divisor $\Delta).$ Identifying a holomorphic vector field $V^{1,0}$ with the real vector field $\text{Re}V^{1,0}$ defined by its real part, the reductivity of $\mathcal{G}$ translates into the following decomposition of the Lie algebra $\mathfrak{g},$ viewed as a real vector space:
$\mathfrak{g}=\mathfrak{k}\oplus J\mathfrak{k},$
where $J$ denotes the endomorphism of the real tangent bundle $TX$ defined by the complex structure on $X.$
Furthermore, we fix a $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ with positive curvature form $\omega_{0}(=dd^{c}\phi_{0})$ (using additive notation for metrics on line bundles, recalled below). The fixed metric $\phi_{0}$ on $K_{(X,\Delta)}$ corresponds to a volume form on $X$ with divisorial singularities that we denote by $\mu_{0}$ (see Section 3.2.1). We will denote by $\mathcal{P}(X)$ the space of all probability measures on $X,$ endowed with its weak topology and by $\mathcal{P}(X)_{0}$ the subspace of all measures $\mu$ with vanishing moment:
(3.1) $\int_{X}\boldsymbol{m}\mu=0,$
where $\boldsymbol{m}$ denotes the moment map induced by the action of $\mathcal{K}$ on $(X,-K_{(X,\Delta})$ and the symplectic form defined by $\omega_{0}$ (see Section 3.2.7). Functionals whose arguments are measures $\mu$ will be denoted by capital letters (e.g. $E(\mu)),$ while those whose arguments are functions $u$ or metrics $\phi$ will be denoted by calligraphic letters, such as $\mathcal{E}(u).$
3.2. Preliminaries
3.2.1. Metrics on $K_{(X,\Delta)}$ versus volume forms with divisorial singularities
We will use additive notation for tensor powers of line bundles $L\rightarrow X$ and their Hermitian metrics. This means that we identify a Hermitian metric $\left\|\cdot\right\|$ on $L$ with a collection of smooth local functions $\phi_{U}\coloneqq-\log(\left\|e_{U}\right\|^{2})$ associated to a given covering of $X$ by open subsets $U$ and trivializing holomorphic sections $e_{U}$ of $L\rightarrow U.$ The normalized curvature form on $X$ of the metric may then, locally, be expressed as
$dd^{c}\phi_{U}\coloneqq\frac{i}{2\pi}\partial\bar{\partial}\phi_{U}$
Accordingly, as is customary, we will symbolically denote by $\phi$ a given Hermitian metric on $L$ and by $dd^{c}\phi$ its curvature form, which represent the first Chern class $c_{1}(L)$ of $L.$ We will denote by $\mathcal{H}(L)$ the space of all smooth metrics $\phi$ on $L$ such that $dd^{c}\phi>0$ and by $\text{PSH}(L)$ the space of all psh metrics on $L,$ i.e. $\phi_{U}$ is a plurisubharmonic (psh) function on $X.$ Then $dd^{c}\phi$ defines a positive current on $L.$ The subspace of $\text{PSH}(L)$ consisting of locally bounded metrics, i.e. $\phi_{U}\in L_{\text{loc}}^{\infty},$ will be denoted by $\text{PSH}(L)_{\mathrm{b}}.$ More generally, when $L$ is merely defined as $\mathbb{Q}$–line bundle a metric $\phi$ on $L$ may be defined by the collection $\phi_{U}\coloneqq k^{-1}\phi_{U}^{(k)},$ where $\phi^{(k)}$ is a metric on $kL$ for a positive integer such that $kL$ is a bona fide line bundle.
Now assume that $(X,\Delta$) is a log Fano manifold and take $L$ to be the $\mathbb{Q}$–line bundle $K_{(X,\Delta)}$. Then, following [24], any metric $\phi$ in $\mathrm{PSH}(L)_{\mathrm{b}}$ induces a measure $\mu_{\phi}$ on $X$, absolutely continuous with respect to Lebesgue measure. Recall that we have fixed a (multivalued) section $s_{\Delta}$ cutting out $\Delta$. Then, locally in a holomorphic coordinate chart $U$,
(3.2) $\mu_{\phi}\coloneqq e^{-\phi_{U}}|s_{U}|^{-2}\frac{i^{n^{2}}}{2^{n}}dz\wedge d\bar{z},\,\,\,dz\coloneqq dz_{1}\wedge\cdots\wedge dz_{n}$
by taking $e_{U}=\partial/\partial z_{1}\wedge\cdots\wedge\partial/\partial z_{n}\otimes e_{U}^{*}$ and $s_{\Delta}=s_{U}e_{U}$. A measure $\mu$ on $X$ will be called a volume form with divisorial singularities (along $\Delta$) if $\mu=\mu_{\phi}$ for some $\phi$ in $\text{PSH}(L)_{\mathrm{b}}$ with the property that $\phi$ is smooth on $X-\Delta.$
3.2.2. Kähler–Einstein metrics for $(X,\Delta)$
[…]
3.2.3. Pluripotential theory and energy
In this section we recall the notion of pluricomplex energy of a probability measure $\mu$ introduced in [23], but using the notation in [12, 13]. Given a Kähler form $\omega_{0}$ on a compact complex manifold consider the space of all Kähler potentials:
$\mathcal{H}(X,\omega_{0})\coloneqq\left\{u\in C^{\infty}(X)\colon\,\,\omega_{u}\coloneqq\omega_{0}+dd^{c}u>0\right\}$
and denote by $\text{PSH}(X,\omega)$ its closure in $L^{1}(X).$ Given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$, expressing $u=\phi-\phi_{0}$ for a metric $\phi$ on $L,$ we may identify the spaces $\mathcal{H}(X,\omega_{0})$ and $\text{PSH}(X,\omega_{0})$ with $\mathcal{H}(L)$ and $\text{PSH}(\mathcal{L}),$ respectively. We also set
$\mathcal{H}_{1,1}(X,\omega)\coloneqq\left\{u\colon\omega_{u}\in L_{\text{loc}}^{\infty},\,\,\omega_{u}\geq 0\right\}.$
Thus, $\mathcal{H}(X,\omega_{0})\subset\mathcal{H}_{1,1}(X,\omega)\subset\text{PSH}(X,\omega_{0})$. Given a log Fano manifold $(X,\Delta)$ we will take $L=-K_{(X,\Delta)}$ and $\phi_{0}$ as in the setup introduced above. Consider the following functional on $C^{\infty}(X)$:
(3.3) $\mathcal{E}(u)\coloneqq\frac{1}{(n+1)V}\int_{X}\sum_{j=0}^{n}u\omega_{u}^{n-j}\wedge\omega_{u}^{j},\,\,\,\,V=\int_{X}\omega_{0}^{n}$
whose differential on $\mathcal{H}(X,\omega_{0})$ is given by
(3.4) $d\mathcal{E}|_{u}=\omega_{u}^{n}/V.$
As a consequence, the restriction of $\mathcal{E}$ to $\mathcal{H}(X,\omega_{0})$ is increasing and thus admits a canonical extension to $\text{PSH}(X,\omega_{0}),$ namely the smallest extension which is upper semi-continuous with respect to the $L^{1}$–topology [23]. The subspace of $\text{PSH}(X,\omega_{0})$ consisting of all $u$ such that $\mathcal{E}(u)>$-$\infty$ is denoted by $\mathcal{E}^{1}(X,\omega_{0}).$ Formula3.4 still applies on $\mathcal{E}^{1}(X,\omega_{0}),$ when $\omega_{u}^{n}$ is the Monge–Ampère measure defined in terms of non-pluripolar products [23].
The pluricomplex energy $E(\mu)$ of a probability measure $\mu$ on $X$ (with respect to $\omega_{0})$ is defined by following lower semi-continuous (lsc) convex function on $\mathcal{P}(X)$:
(3.5) $E(\mu)\coloneqq\sup_{u\in C^{0}(X)\cap\text{PSH}(X,\omega_{0})}\mathcal{E}(u)-\left\langle u,\mu\right\rangle.$
By [23], $E(\mu)<\infty$ iff there exists $u_{\mu}\in\text{PSH}(X,\omega_{0})$ realizing the sup above such that $\mathcal{E}(u_{\mu})>$-$\infty$ (but $u_{\mu}$ need not be in $C^{0}(X)).$ Moreover, $u_{\mu}$ is uniquely determined, up to an additive constant, by
(3.6) $\omega_{u_{\mu}}^{n}/V=\mu,$
using the non-pluripolar product introduced in [32], to define $\omega_{u_{\mu}}^{n}$. We will fix the additive constant by demanding that $\int u_{\mu}\omega_{0}^{n}=0$ and call $u_{\mu}$ the potential (function) of $\mu.$ Equivalently, the corresponding potential metric $\phi_{\mu}$ satisfies
(3.7) $(dd^{c}\phi_{\mu})_{\,}^{n}/V=\mu,\,\,\,\,\int(\phi_{\mu}-\phi_{0})(dd^{c}\phi_{0})^{n}=0.$
By the Calabi–Yau theorem $\mu$ is a smooth volume form iff $\phi_{\mu}\in\mathcal{H}(L).$ The functional $E$ on $\mathcal{P}(X)$ satisfies
(3.8) $E(\omega_{0}^{n}/V)=0$
and on the the convex subspace $\{E<\infty\}\subset\mathcal{P}(X)$ its differential may be represented as follows [10]:
(3.9) $dE_{|\mu}=-u_{\mu}\,\,\,\left(=\phi_{0}-\phi_{\mu}\right),$
For example, any measure $\mu$ on $X$ with divisorial singularities (or more, generally with an $L^{p}$–density for some $p>1)$ has finite energy, $E(\mu)<\infty.$ Thus the properties 3.8 and 3.9 determine $E$ on $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ for any $p>1$ (and the functional $E$ on $\mathcal{P}(X)$ can then be characterized as the greatest lower semi-continuous extension of $E$ from $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ to $\mathcal{P}(X)$).
Remark 3.1.
In terms of standard functionals in Kähler geometry, when $\mu$ is smooth,
(3.10) $E(\mu)=(\mathcal{I}-\mathcal{J})(u_{\mu})=\mathcal{E}(u_{\mu})-\left\langle u_{\mu},\mu\right\rangle,$
where $\mathcal{I}$ and $\mathcal{J}$ are the standard energy functionals in Kähler geometry, defined by
$\mathcal{J}(u)\coloneqq-\mathcal{E}(u)+\frac{1}{V}\int_{X}u\frac{\omega_{0}^{n}}{V},\,\,\,\,\,\mathcal{I}(u)\coloneqq\int_{X}u\frac{(\omega_{0}^{n}-\omega_{u}^{n})}{V}.$
Example 3.2.
When $n=1$ integration by parts gives
$V\mathcal{E}(u)=-\frac{1}{2}\int_{X}du\wedge d^{c}u+\int_{X}u\omega_{0},\,\,2VE(\mu)=-\int_{X}u_{\mu}\mu=\int_{X}du_{\mu}\wedge d^{c}u_{\mu},$
if $E(\mu)<\infty$ or equivalently $u_{\mu}$ is in the space $H^{1}(X)\subset L^{2}(X)$ of all $u$ such that $du\in L^{2}(X).$ Moreover, in this case the sup in formula 3.5 can be taken over all of $C^{0}(X)$ or over $H^{1}(X).$ Indeed, by classical Hilbert space theory the latter sup is attained by $u_{\mu}\in H^{1}(X)$ solving the (Laplace) equation 3.6 and hence $u_{\mu}\in\text{PSH}(X,\omega_{0})$.
3.2.4. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
The space $\mathcal{H}(X,\omega_{0})$ can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to $\mathcal{H}_{1,1}(X,\omega_{0})$ in [40]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
3.2.5. The Mabuchi functional $\mathcal{M}$ versus the free energy functional $F$
Given an ample line bundle $L$ the Mabuchi functional $\mathcal{M}$ was originally defined when $\Delta$ is trivial, by declaring that its differential on $\mathcal{H}(L)$ at $\phi$ is identified with $(R_{\phi}-\overline{R})(dd^{c}\phi)^{n},$ where $R_{\phi}$ denotes the scalar curvature of the Kähler metric $dd^{c}\phi$ and $\overline{R}$ its average [80]. Here we will instead follow the thermodynamical formalism in [10, 24], which applies in the general setup of log Fano varieties $(X,\Delta).$ We thus define the Mabuchi functional $\mathcal{M}$ on $\text{PSH}(-K_{(X,\Delta)})_{\mathrm{b}}$ with respect to the log pair $(X,\Delta)$ and the fixed reference metric $\phi_{0}\in\text{PSH}(-K_{(X,\Delta)})_{\mathrm{b}}$, by
(3.11) $\mathcal{M}(\phi)\coloneqq F((dd^{c}\phi_{\mu})_{\,}^{n}/V)\coloneqq F_{-1}((dd^{c}\phi_{\mu})_{\,}^{n}/V)$
where $F_{\beta}(\mu)$ is the free energy functional on $\mathcal{P}(X)$ at inverse temperature $\beta\in\mathbb{R},$ defined as $F_{\beta}(\mu)=\infty$ if $E(\mu)=\infty$ and when $E(\mu)<\infty,$
$F_{\beta}(\mu)\coloneqq\beta E(\mu)+D(\mu|\mu_{0})\in]0,\infty],$
where $D(\mu|\nu)$ denotes the entropy of $\mu$ relative to $\nu$ for given measures $\mu$ and $\nu$ on a topological space $Y,$ also known as the Kullback–Leibler divergence:
(3.12) $D(\mu|\nu)=\int_{X}\log\left(\frac{\mu}{\nu}\right)\mu$
if $\mu$ is absolutely continuous with respect to $\nu$ and otherwise $D(\mu|\nu)\coloneqq\infty.$ Set
(3.13) $F(\beta)\coloneqq\inf_{\mu\in\mathcal{P}(X)}F_{\beta}(\mu),\,\,\,\,F(\beta)_{0}\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}F_{\beta}(\mu)$
We recall that the function $\mu\mapsto D(\mu|\nu)$ is lower semicontinuous and convex on $\mathcal{P}(X)$ as follows, for example, from the following well-known Legendre duality formula:
(3.14) $D(\mu|\nu)=\sup_{u\in C^{0}(Y)}\left\{\int u\mu-\log\int e^{u}\nu\right\}.$
3.2.6. The Futaki character of $(X,\Delta)$ and its quantization
Let $V^{1,0}$ be a holomorphic vector field on $X$ of type $(1,0)$ corresponding to an element in $\mathfrak{g}.$ Its Futaki invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})$ may be defined by
(3.15) $\text{Fut}_{(X,\Delta)}(V^{1,0})\coloneqq-\int_{X}\mathcal{L}_{V^{1,0}}\phi_{0}\frac{(dd^{c}\phi_{0})^{n}}{V}\in\mathbb{C},$
where $\mathcal{L}_{V^{1,0}}\phi_{0}$ denotes the Lie derivative of $\phi_{0}$ along $V^{1,0},$ defined by decomposing $V^{1,0}=\text{Re}V^{1,0}+i\text{Im}V^{1,0}$ and setting
(3.16) $\mathcal{L}_{W}\phi\coloneqq\lim_{t\rightarrow 0}\left(\exp(tW)^{*}\phi_{0}-\phi_{0}\right)$
for any real vector field $W$ on $X$ that lifts to $K_{(X,\Delta)}$, and extended $\mathbb{C}$–linearly to any complex vector field on $X$ that lifts to $K_{(X,\Delta)}$. The corresponding complex-valued function $\text{Fut}_{(X,\Delta)}$ on the Lie algebra $\mathfrak{g},$ defining an element in $\mathfrak{g}^{*},$ is called the Futaki character and is independent of the choice of $\phi_{0}$:
Lemma 3.3.
For any given $\psi,\phi\in\text{PSH}(X,-K_{(X,\Delta)})_{\mathrm{b}}$ and holomorphic vector field $V^{1,0}$ on $X$ corresponding to an element in $\mathfrak{g}$
(3.17) $\text{$\mathrm{Fut}$}_{(X,\Delta)}(V^{1,0})=-\int_{X}\mathcal{L}_{V^{1,0}}\psi\frac{(dd^{c}\psi)^{n}}{V}=-\int_{X}\mathcal{L}_{V^{1,0}}\phi\frac{(dd^{c}\phi)^{n}}{V}$
As a consequence, the definition 3.15 of $\text{Fut}_{(X,\Delta)}(V^{1,0})$ is independent of the choice of $\phi_{0}$ and $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Moreover,
(3.18) $\frac{d}{dt}\mathcal{M}\left(\exp(t\mathrm{Re}V^{1,0})^{*}\psi\right)_{|t=0}=\mathrm{Re}\text{$\mathrm{Fut}$}_{(X,\Delta)}(V^{1,0})$
More generally, 3.17 holds if $\psi$ and $\phi$ have finite energy and 3.18 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy. In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
Proof.
This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.17 first note that, for any given any $g\in\mathcal{G},$
$\mathcal{E}\left(g^{*}\psi\right)-\mathcal{E}\left(g^{*}\phi\right)=\mathcal{E}(\psi)-\mathcal{E}(\phi).$
This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.4 defined by formula 3.4, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.17. Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$. Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e. iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.11, we can decompose
$\mathcal{M}(\psi)=-\mathcal{E}(\psi)+D(dd^{c}\psi)^{n}/V|\mu_{\psi}),$
expressed in terms of the functional $\mathcal{E}$ on $\text{PSH}(L)_{\mathrm{b}}$ and the entropy of the measure $(dd^{c}\psi)^{n}/V$ relative to the measure $\mu_{\psi},$ defined by formula 3.2. For any $g\in\mathcal{G}$ we have $g^{*}\mu_{\psi}=\mu_{g^{*}\psi}.$ Hence, setting $g_{t}\coloneqq\exp(tV),$
$\mathcal{M}(g_{t}^{*}\psi)=-\mathcal{E}(g_{t}^{*}\psi)+D(dd^{c}\psi)^{n}/V|\mu_{\psi}).$
Differentiating with respect to $t$ and using 3.3 thus proves 3.18. More generally, the proof of 3.17 only used that $\phi$ and $\psi$ have finite energy. Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [23] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [24], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy. Hence, $\frac{d}{dt}\mathcal{M}\left(\exp(t\text{Re}V^{1,0})^{*}\psi\right)_{|t=0}$, showing that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
The independence of the choice of $\phi_{0}$ is also a consequence of the fact that $\text{Fut}_{(X,\Delta)}$ arises as the leading term of the large-$k$ asymptotics of the quantized Futaki character $\text{Fut}_{(X,\Delta),k}(\cdot)$ at level $k,$ defined by
(3.19) $\text{Fut}_{(X,\Delta),k}(V^{1,0})\coloneqq-\frac{1}{kN}\text{Tr}(\mathcal{L}_{V^{1,0}}),$
where $\mathcal{L}_{V^{1,0}}$ now denotes the endomorphism of $H^{0}(X,-kK_{(X,\Delta)})$ induced by the infinitesimal action of $V^{1,0}$ on $H^{0}(X,-kK_{(X,\Delta)})$ (see the proof of [26, Prop 4.7]).
3.2.7. The moment maps $\boldsymbol{m}$ and $\boldsymbol{m}_{N}$
The compact group $\mathcal{K}$ fixing the symplectic form $\omega_{0}$ naturally acts on $(X,-K_{(X,\Delta)})$ and thus induces a moment map $\boldsymbol{m}$ from $X$ into the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}.$ This map is $\mathcal{K}$–invariant (with respect to the coadjoint action on $\mathfrak{k}^{*}$) [64]. Fixing a basis $e_{1},\dots,e_{r}$ in $\mathfrak{k}$ we can identify $\mathfrak{k}^{*}$ with $\mathbb{R}^{r}$ and represent
$\boldsymbol{m}\colon\,\,X\rightarrow\mathbb{R}^{r},\,\,\,x\mapsto(h_{e_{1}}(x),\dots,h_{e_{r}}(x)),$
where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.16). We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
$\omega_{N}\coloneqq\frac{1}{N}\sum_{i=1}^{N}\omega(x_{i}),$
where $\omega(x_{i})$ denotes the pullback of $\omega$ on $X$ to $X^{N}$ under the projection to the $i^{\mathrm{th}}$ factor of $X^{N}$. Concretely,
$\boldsymbol{m}_{N}\colon X^{N}\rightarrow\mathbb{R}^{r},\,\,\,\boldsymbol{m}_{N}(x_{1},\dots,x_{N})=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{m}(x_{i})\in\mathbb{R}^{r}$
By the Kempf–Ness theorem [64], the $\mathcal{G}$–orbit of $\left\{\boldsymbol{m}_{N}=0\right\}$ coincides with the polystable locus $(X^{N})_{\text{ps }}$ of $X^{N}$ (the subset of points $x$ in the semistable locus $(X)_{\text{ss}}^{N}$ such that $\mathcal{G}x$ is closed as a subset $(X)_{\mathrm{ss}}^{N}$):
(3.20) $\mathcal{G}\left\{\boldsymbol{m}_{N}=0\right\}=(X^{N})_{\text{ps }}\subset X_{\text{ss}}$
Moreover, 0 is a regular value of $\boldsymbol{m}_{N}$ iff all the stabilizers in $\mathcal{K}$ of the points in $\left\{\boldsymbol{m}_{N}=0\right\}$ are finite iff the semistable locus of $X^{N}$ coincides with the polystable locus [64, Cor 3.10].
3.3. The structure of the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment
With the preliminary results in place we turn to the proof of the following result, which will play a key role in what follows.
Proposition 3.4.
Consider the infimum of the pluricomplex energy $E$ on the orbit $\mathcal{G}\nu$ of any volume form in $\mathcal{P}(X)$ with divisorial singularities along $\Delta.$ It is attained at a unique $\mu$ in $(\mathcal{G}\nu)/\mathcal{K},$ which is the unique critical point of $E$ on $(\mathcal{G}\nu)/\mathcal{K}.$ Moreover, if the Futaki character of $(X,\Delta)$ vanishes, then $\mu$ is the unique element in $\left(\mathcal{G}\nu\cap\mathcal{P}(X)_{0}\right)/\mathcal{K}$ and, more generally, any $\mu\in\mathcal{P}(X)_{0}$ minimizes $E$ on $\mathcal{G}\mu.$
Proof.
First note that $E$ is $\mathcal{K}$–invariant (since $\omega_{0}$ is $\mathcal{K}$–invariant). We thus obtain a function
$\mathcal{G}/\mathcal{K}\rightarrow\mathbb{R},\,\,\,g\mapsto E(g^{*}\nu),$
The first part of the proposition to be proved says that this function has a unique critical point. To prove this first recall that $\mathcal{G}=\mathcal{K}\exp(J\mathfrak{k}),$ by polar decomposition of complex reductive Lie groups. Hence, by basic convex analysis it enough to show that $E(g^{*}\nu)$ is proper and strictly convex along curves $g_{t}$ of the form $g_{t}=\exp(tJW),$ where $W\in\mathfrak{k}-\{0\}.$ Expressing $\nu=(dd^{c}\phi_{\nu})^{n}/V,$ for some $\phi_{\nu}\in\text{PSH}(-K_{(X,\Delta)})_{b},$ gives $g_{t}^{*}\nu=dd^{c}(g_{t}^{*}\phi_{\nu}))^{n}/V.$ However, the real one-parameter curve $g_{t}^{*}\phi_{\nu}$ need not define a geodesic in the sense of [40], even when $\phi_{\nu}$ is smooth, unless $\phi_{\nu}$ is $\mathcal{K}$-invariant. To bypass this difficulty we will use a complexification argument, by setting $\tau=s+it$ for $s,t\in\mathbb{R}$ and $g_{\tau}\coloneqq\exp(sW+tJW).$ Since $E$ is $\mathcal{K}$-invariant we have $E(g_{{}_{\tau}}^{*}\nu)=E(g_{t}^{*}\nu).$ The corresponding properness on $\mathcal{G}/\mathcal{K}$ is well-known (see, for example, [21, Remark 9.2]). What remains is thus to show that
(3.21) $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)>0.$
[…]
$\phi^{\tau}\coloneqq\exp(g_{\tau})^{*}\phi_{\nu},$
we can express $g_{\tau}^{*}\nu=(dd^{c}\phi^{\tau})^{n}/V.$ We claim that there exists a positive constant $c$ such that
(3.22) $\frac{\partial}{\partial\tau}\frac{\partial}{\partial\overline{\tau}}E\left(\frac{(dd^{c}\phi^{\tau})^{n}}{V}\right)\geq c\int_{X}i\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)\wedge\overline{\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)}\wedge dd^{c}\phi_{0}\wedge(dd^{c}\phi^{\tau})^{n-2}$
(when $n=2$ and when $n=1$ the last factor above should be omitted). Accepting this for the moment, it follows that $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0.$ Moreover, if $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)=0,$ then $\partial(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}})=0$ on $X-\Delta,$ since both $dd^{c}\phi_{0}$ and $dd^{c}\phi^{\tau}$ are smooth and strictly positive on $X-\Delta.$ But this means that
$\partial\bar{\partial}\phi^{\tau}(\cdot,\overline{W^{1,0}})=0\,\,\,\text{on $X-\Delta$}$
and since $i\partial\bar{\partial}\phi^{\tau}$ is smooth and strictly positive on $X-\Delta$ it follows that $W=0$ on $X-\Delta.$ Thus, by continuity, $W=0$ on $X,$ which is a contradiction. Hence, 3.21 holds. Finally, the inequality 3.22 is shown in essentially the same way as in the proof of [21, Lemma 9.1]. More precisely, in [21, Lemma 9.1] it is assumed that the corresponding metric $\phi^{\tau}$ is independent of the real part of $\tau$ (in our notation) and that $\phi^{\tau}$ satisfies the complex geodesic equation, that is the local function $\Phi(z,\tau)\coloneqq\phi^{\tau}(z)$ satisfies
(3.23) $(dd^{c}\Phi)^{n+1}=0\,\,\,\text{on $X\times\mathbb{C}$}$
But the proof only uses the latter equation and this equation is satisfied for our $\phi^{\tau}$ (as is well-known and can be seen by locally making the change of variable, $w\coloneqq g_{\tau}(z)).$
Next, let us show that $\mu$ is a critical point of $E$ on the orbit $\mathcal{G}\nu$ iff $\mu\in\mathcal{P}(X)_{0}.$ Since $E$ is $\mathcal{K}$–invariant it will be enough to show that for any real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g},$
(3.24) $\frac{d}{dt}E(\mu_{t})_{|t=0}=-\int h_{W}\mu,\,\,\,\>\mu_{t}\coloneqq\exp(tJW)^{*}\mu$
where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ First recall that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0},$ (see Section 3.2.7). To prove formula 3.24 note that, by 3.9, denoting by $\phi$ the potential of $\mu,$
$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int(\phi-\phi_{0})\frac{d\mu_{t}}{dt}|{}_{t=0}=-\int(\phi-\phi_{0})\mathcal{L}\mu,$
Integrating by parts gives
$\frac{d}{dt}E(\mu_{t})|_{t=0}=\int\mathcal{L}(\phi-\phi_{0})\mu=\int\mathcal{L}\phi\frac{(dd^{c}\phi)^{n}}{V}-\int h_{W}\mu,$
which proves 3.24, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
Finally, take $\mu\in\mathcal{P}(X)_{0}.$ If $E(\mu)=\infty,$ then $E(g^{*}\mu)=\infty$ for all $g\in\mathcal{G}.$ Hence, we may assume that $E(\mu)<\infty,$ which means that $E(g^{*}\mu)<\infty$ for all $g\in\mathcal{G}.$ Consider, as before, the function $g\mapsto E(g^{*}\mu)$ on $\mathcal{G}/\mathcal{K}.$ It follows from the inequality $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0,$ shown above when $\nu$ is a volume form, and a standard approximation argument, that $E(g^{*}\mu)$ is convex on the symmetric space $\mathcal{G}/\mathcal{K}.$ Moreover, as before, $\frac{d}{dt}E(\mu_{t})_{|t=0}=0$ using the previous lemma, that applies since $E(\mu)<\infty$ (which equivalently means that $\mathcal{E}(\phi_{\mu})>-\infty).$ Since any critical point of a convex function $f$ on the symmetric space $\mathcal{G}/\mathcal{K}$ minimizes $f$ this shows that $\mu$ minimizes $E$ on $\mathcal{G}\mu,$ as desired. ∎
By the previous proposition, the following map is well-defined:
$\pi_{0}\colon\,\mathcal{V}(X)\rightarrow\mathcal{V}(X)_{0}/\mathcal{K},\,\,\,E\left(\pi_{0}(\nu)\right)=\inf_{\mathcal{G}\nu}E$
where $\mathcal{V}(X)$ denotes the subspace of $\mathcal{P}(X)$ consisting of volume forms, i.e. measures with smooth and strictly positive density and $\mathcal{V}(X)_{0}\coloneqq\mathcal{P}(X)_{0}\cap\mathcal{V}(X).$ More generally, $\pi_{0}$ is well-defined for any volume form $\mu$ in $\mathcal{P}(X)$ with divisorial singularities.
Lemma 3.5.
Assume that the Futaki character of $(X,\Delta)$ vanishes. Then $F_{-\gamma}\left(\pi_{0}(\nu)\right)\leq F_{-\gamma}\left(\nu\right)$ when $\gamma\leq 1$ and $F_{-\gamma}\left(\pi_{0}(\nu)\right)\geq F_{-\gamma}\left(\nu\right)$ when $\gamma\geq 1.$ As a consequence, for $\gamma\leq 1,$
(3.25) $\inf_{\mathcal{V}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{\mathcal{V}}(X)}F_{-\gamma}=\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)}F_{-\gamma}$
Proof.
First note that $F_{-\gamma}$ is $\mathcal{K}$–invariant, since $\phi_{0}$ is $\mathcal{K}$–invariant. Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.25. Thus, to prove the rest of the identities 3.25, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively. ∎
The following result relates the analytic threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}},$ defined by formula 1.7, to the free energy functional (Section 3.2.5) restricted to the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment:
Proposition 3.6.
The invariant $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ is independent of the choice of reference $\phi_{0}$ and if the Futaki character of $(X,\Delta)$ vanishes, then
$\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}=\sup_{\gamma>0}\left\{\inf_{\mu\in\mathcal{V}(X)_{0}}F_{-\gamma}(\mu)>-\infty\right\}.$
Proof.
To prove the first statement of the proposition it is enough to show that when changing the metric $\phi_{0}$ the corresponding pluricomplex energy $E_{\phi_{0}}$ only changes by an additive bounded term. To this end note that, by formula 3.9, given $\phi_{0}$ and $\psi_{0}$ in $\mathcal{H}(L)$ the differential $d(E_{\phi_{0}}-E_{\phi_{1}})|_{\mu}$ is represented by the function $u\coloneqq\phi_{0}-\phi_{1},$ for any $\mu$ in $\mathcal{P}(X)$ satisfying $E(\mu)<\infty.$ Hence, there exists a constant $C$ such that
$E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)=C+\left\langle u,\mu\right\rangle,$
showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired. Finally, the second statement follows from combining formulae 3.11 and 3.10 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.6). ∎
We are now ready for the proof of the following result, where the equivalence $1\iff 2$ follows directly from [24, 48]. The equivalence $1\iff 4$ provides partial evidence for the Large Deviation Principle (LDP) in Conjecture A, stated in Section 5.2, as it establishes the lower semicontinuity of the rate functional – a property built into the definition of an LDP.
Theorem 3.7.
On a log Fano manifold $(X,\Delta)$ the following are equivalent:
  1. (1)
    $(X,\Delta)$ admits a Kähler–Einstein metric.
  2. (2)
    $\delta^{\mathrm{A}}(X,L)^{\mathcal{G}}>1$.
  3. (3)
    The Futaki character of $(X,\Delta)$ vanishes, $\mathcal{G}$ is reductive and there exists $\epsilon>0$ such that $F_{-(1+\epsilon)}$ is bounded from below on $\mathcal{P}(X)_{0},$ i.e. $F(-\gamma)_{0}>$-$\infty$ for some $\gamma>1.$
  4. (4)
    The Futaki character of $(X,\Delta)$ vanishes, $\mathcal{G}$ is reductive and the functional $F_{-1}$ is lsc on $\mathcal{P}(X)_{0}$.
Additionally, if $(X,\Delta)$ admits a Kähler–Einstein metric then there is a unique one whose normalized volume form $\mu_{\text{KE}}$ is in $\mathcal{P}(X)_{0}$. Moreover, the measure $\mu_{\text{KE}}$ is the unique minimizer of $F_{-1}$ on $\mathcal{P}(X)_{0}$ and
(3.26) $\mu_{\text{KE }}=\lim_{\beta\searrow-1}\mu_{\beta}\,\,\text{in $\mathcal{P}(X)$},$
where $\mu_{\beta}$ is the unique minimizer of $F_{\beta}$ on $\mathcal{P}(X),$ or equivalently, on $\mathcal{P}(X)_{0}.$ Moreover, $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant.
Proof.
In the proof we will identify a metric $\phi\in\mathcal{H}(-K_{(X,\Delta)})$ with the corresponding function $u(:=\phi-\phi_{0})$ in $\mathcal{H}(X,\omega_{0})$ and denote by $u\cdot g$ the function corresponding to $g^{*}\phi.$ As shown in [48], $(X,\Delta)$ admits a Kähler–Einstein metric iff there exist $\epsilon,C>0$ such that
(3.27) $\mathcal{M}(u)\geq\epsilon\inf_{g\in\mathcal{G}}d_{1}(0,u\cdot g)-C\,\,\,\text{on $\mathcal{H}(X,\omega_{0}$})\cap\{\mathcal{E}(u)=0\},$
where $d_{1}$ is the $L^{1}$–Finsler metric on $\mathcal{H}(X,\omega_{0})$, which, on $\{\mathcal{E}(u)=0\}$ satisfies
(3.28) $A^{-1}E\left(\frac{\omega_{u}^{n}}{V}\right)-B\leq d_{1}(0,u)\leq AE\left(\frac{\omega_{u}^{n}}{V}\right)+B$
for some positive constants $A$ and $B.$ Indeed, by 3.10, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.28 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
Next, note that, $3\implies 2$ follows from the previous proposition. To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.27 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.28 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
$F(\mu)\geq\epsilon\inf_{\mathcal{G}}E(\mu)-C\text{\,\,on $\mathcal{P}(X)\cap\{E<\infty\}.$ }$
In particular, restricting this inequality to $\mathcal{P}(X)_{0},$ and using the last statement of Prop 3.4, yields $F(\mu)\geq\epsilon E(\mu)-C$ on $\mathcal{P}(X)\cap\{E<\infty\},$ i.e. $F_{-(1+\epsilon)}\geq-C$ on $\mathcal{P}(X)\cap\{E<\infty\}$ and hence, by definition, on $\mathcal{P}(X),$ as desired. To prove that $1\iff 4,$ first assume that $F_{-1}$ is lsc on $\mathcal{P}(X)_{0}.$ Since $\mathcal{P}(X)_{0}$ is compact it follows that $F_{-1}$ admits a minimizer in $\mathcal{P}(X)_{0}.$ Such a minimizer is KE, since the Futaki character vanishes (see below). Conversely, assume that $(X,\Delta)$ admits a KE metric. Take a sequence $\mu_{j}$ converging towards $\mu_{\infty}$ in $\mathcal{P}(X)_{0}.$ To prove that $F_{-1}$ is lsc we may as well assume that $F_{-1}(\mu_{j})\leq C.$ Using $1\iff 3,$ it follows that $D(\mu_{j}|\mu_{\phi_{0}})\leq C.$ Thus, by [24, Thm 2.17], $F_{-1}(\mu_{\infty})\leq\liminf F_{-1}(\mu_{j}),$ as desired.
To prove the last statement of the theorem, recall that any KE volume form $\mu$ in $\mathcal{P}(X)$ minimizes $F_{-1}$ on $\mathcal{P}(X)$ [24]. Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [21, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e. on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.26. In the present case, where $\phi_{0}$ is assumed $\mathcal{K}$–invariant, it thus follows from Prop 3.4 that $\mu_{-1}\in\mathcal{P}(X)_{0}$ and conversely, any KE volume form in $\mathcal{P}(X)_{0}$ coincides with $\mu_{-1}.$ It also follows from the aforementioned uniqueness that $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant. ∎
Remark 3.8.
When $X$ admits a KE metric and $\mathcal{G}$ is nontrivial it is well known that $F(\beta)>$-$\infty$ iff $\beta\geq-1.$ But, $F(\beta)_{0}$ provides a finite extension of $F(\beta)$ from $1,\infty[$ to some neighbourhood of $1,\infty[$ in $\mathbb{R},$ as follows from the third point in the previous theorem, combined with Lemma 3.5.
3.3.1. Comparison with Tian’s orthogonality condition
Consider the space $\mathcal{H}(X,\omega_{0})/\mathbb{R}$ of all Kähler potentials, modulo additive constants. The group $\mathcal{G}$ naturally acts from the right on $\mathcal{H}(X,\omega_{0})/\mathbb{R},$ by demanding that $g^{*}\omega_{u}=\omega_{0}+dd^{c}(u\cdot g).$ Now assume that $\omega_{0}$ is Kähler–Einstein and consider the following orthogonality condition on $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$:
(3.29) $\int_{X}u\psi\omega_{0}^{n}=0\,\,\,\forall\psi\in\text{$\ker$(}\Delta-\lambda_{1})$
where $\Delta$ denotes the Laplacian with respect to $\omega$ and $\lambda_{1}$ denotes its smallest strictly positive eigenvalue. This condition appears in Tian’s conjecture [89, Conj 5.5], settled in [48], saying that if $\omega_{0}$ is Kähler–Einstein, then there exists positive constants $\epsilon$ and $C$ such that
(3.30) $\mathcal{M}(u)\geq\epsilon(\mathcal{I}-\mathcal{J})(u)-C$
for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.29.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48]. It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [23, Section 2]). The inequality 3.30 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.30 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.29 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
Lemma 3.9.
Let $X$ be a Fano manifold and assume that the Futaki character of $X$ vanishes and that the group $\mathcal{G}$ is reductive. Fix a Kähler form $\omega_{0}$ that is invariant under a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}.$ Given $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ the following two conditions are equivalent for $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}:$
$(i)\,\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)=\mathcal{J}(v),\,\,\,\,(ii)\int_{X}v\Delta\boldsymbol{m}\omega_{0}^{n}=0,$
where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, there exists a unique $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions. When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.29 for $v.$
Proof.
First observe that $\mathcal{J}(u)$ is $\mathcal{K}$–invariant, since $\omega_{0}$ is assumed $\mathcal{K}$–invariant. It thus follows, precisely as in the proof of Prop 3.4, that the inf over $\mathcal{G}$ in $(i)$ may be replaced by the inf over $J\mathfrak{k}.$ The pullback of $\mathcal{J}$ to $J\mathfrak{k}$ is proper and strictly convex on $J\mathfrak{k}.$ Indeed, for any complex one-parameter curve $\phi^{\tau}$ in $\mathcal{H}(-K_{X})$ satisfying the complex geodesic equation 3.23 a direct calculation gives
$V\frac{\partial^{2}\mathcal{J}(\phi^{\tau}-\phi_{0})}{\partial\tau\partial\overline{\tau}}=\int_{X}\frac{\partial^{2}\phi^{\tau}}{\partial\tau\partial\overline{\tau}}\omega_{0}^{n}=\int_{X}\left|\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)\right|_{dd^{c}\phi^{\tau}}^{2}\omega_{0}^{n}$
and we can thus proceed precisely as in the proof of Prop 3.4. Hence, the inf in $(i)$ is attained for a unique $v\in u\cdot\mathcal{G}\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ and this $v$ is the unique critical point of the the pullback of $\mathcal{J}$ to $J\mathfrak{k}.$ Now, fix $W\in\mathfrak{k}$ and $v\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ and set $v_{t}\coloneqq v\cdot\exp(tJW).$ Then
(3.31) $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=\int_{X}\left.\frac{dv_{t}}{dt}\right|_{t=0}\omega_{0}^{n}=\int_{X}\left\langle dv,JW\right\rangle\omega_{0}^{n}=\int_{X}\nabla v\cdot JW\omega_{0}^{n}$
using, in the first equality, that the Futaki character vanishes (formula 3.15). Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$. Using that $JW$ is proportional to $\nabla h_{W},$ where $h_{W}$ is the Hamiltonian function corresponding to $W,$ we deduce that
$V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=-c_{n}\int_{X}\nabla v\cdot\nabla h_{W}\omega_{0}^{n}=c_{n}\int_{X}v\Delta h_{W}\omega_{0}^{n},$
Hence, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ holds, as desired. Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [8, Thm 2.2]. Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.29, when $\omega_{0}$ is Kähler–Einstein. ∎
In the case when $n=1,$ so that $X=\mathbb{P}^{1}$, the functionals $\mathcal{I}-\mathcal{J}$ and $\mathcal{J}$ coincide and thus the moment condition appearing in Theorem 3.7 then coincides with condition $(ii)$ in the previous lemma. The formula $\delta^{A}(\mathbb{P}^{1},0)^{\mathcal{G}}=2$ that we shall establish in the next section thus yields a quasi-sharp inequality of the form 3.30 on $\mathbb{P}^{1},$ saying that the inequality holds for any $\epsilon\in]0,1[$ - but not for $\epsilon>1$ - when $u$ satisfies Tian’s orthogonality condition 3.29.
3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
Now assume that $(X,\Delta)$ is a K-polystable log Fano curve and that $\mathcal{G}$ is non-trivial. This means that $X=\mathbb{P}^{1}$ and $\Delta$ is of the following form
$\Delta_{w}\coloneqq w[x_{0}]+w[x_{1}],\,\,\,\,(\implies V=2-2w)$
for two distinct points $x_{0}$ and $x_{1}$ and $w\in[0,1[$ (in contrast to Prop 2.10 we now allow $w=0$ so that $\Delta_{0}=0$). Note that when $\Delta_{w}=0$, $\mathcal{G}=\mathrm{PGL}(2,\mathbb{C})$ and $\mathcal{K}\cong\mathrm{SU(2)/{\pm\mathrm{Id}}\cong\mathrm{SO}(3)}$, and when $\Delta_{w}\neq 0$, $\mathcal{G}=\mathbb{C}^{*}$ and $\mathcal{K}=S^{1}$. Without loss of generality we may as well assume that $x_{0}=0$ and $x_{1}=\infty$ in $\mathbb{P}^{1}.$ We will identify $X$ with the two-sphere, i.e. the unit-sphere in $\mathbb{R}^{3}$, via stereographic projection.
Example 3.10.
Consider the case when the fixed $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ is induced by the Fubini–Study metric on $\mathcal{O}(1)\rightarrow\mathbb{P}^{1}$ and thus $\mathrm{SU}(2)$–invariant. Then, identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\omega_{0}=V\omega$ for the unique $\mathrm{SO}(3)$–invariant probability measure $\omega$ and
$E(\mu)=-\frac{V}{2}\int_{X\times X}\log\left(|x-y|^{2}\right)\mu(x)\otimes\mu(y)+C$
for some constant $C$ (by Example 3.2, using the explicit Green function for $(X,\omega_{0})$ ). Moreover, when $\Delta=0$ the measure $\mu_{0}$ is proportional to $\omega_{0}$ and $\boldsymbol{m}=\boldsymbol{x}.$ When $w\neq 0,$ taking $x_{0}$ and $x_{1}$ as the south and north pole on the unit sphere, $\mu_{0}$ is proportional to $e^{-|\boldsymbol{x}-\boldsymbol{x}_{0}|^{2}-|\boldsymbol{x}-\boldsymbol{x}_{1}|^{2}}\omega_{0}$ and $\boldsymbol{m}=2x_{3}/V.$
Lemma 3.11.
Given the log Fano curve $(X,\Delta_{w})$ consider any $\mu\in\mathcal{P}(X)_{0}.$ When $w=0$ $\mu(\left\{x\right\})\leq 1/2$ for all $x\in X$ and when $w\neq 0,$ $\mu(\left\{x\right\})\leq 1/2$ for $x=0$ and $x=\infty.$
Proof.
First consider the case when $w=0,$ $\mathcal{K}=\mathrm{SO}(3).$ Given $x\in X$ there exists a subgroup $\mathcal{K}_{x}$ of $\mathcal{K}$ fixing $x$ that is isomorphic to the group $S^{1}.$ Denote by $h$ the corresponding Hamiltonian: $h=\left\langle\boldsymbol{m},\xi\right\rangle$, where $\xi$ generates $\mathcal{K}_{x}.$ The assumption $\mu\in\mathcal{P}(X)_{0}$ implies that $\int h\mu=0.$ As is well-known for any (holomorphic) $S^{1}$–action on $(\mathbb{P}^{1},-K_{X})$ and fixed $S^{1}$–invariant $\phi_{0}\in\mathcal{H}(-K_{X})$ we have h(X)=[1,1]h(X)=[-1,1] where we may assume that $h^{-1}\{x_{0}\}=-1$ and $h^{-1}\{x_{1}\}=1$ (for example, this is manifestly the case when $\phi_{0}$ is induced by the Fubini–Study metric and since $\mathcal{K}_{x}$ is abelian, $h(X)$ is independent of $\phi_{0}).$ Hence, $\nu\coloneqq h_{*}\mu$ is a probability measure on $\mathbb{R}$ satisfying $\int_{[-1,1]}t\nu=0.$ But this implies that $\nu(\left\{1\right\})\leq 1/2$ and $\nu(\left\{-1\right\})\leq 1/2,$ as desired. Next, consider the case when $w\neq 0.$ Then we can identify $\mathcal{K}$ with the group $S^{1}.$ Denote by $h$ the corresponding Hamiltonian. As in the previous case it is well-known that h(X)=[V/2,V/2],h(X)=[-V/2,V/2], where V=12w.V=1-2w. Since [V/2,V/2][-V/2,V/2] is symmetric it follows precisely as before that $\mu(\left\{x_{i}\right\})\leq 1/2,$ using that the $x_{i}$ are fixed points for the $S^{1}$–action. ∎
Given a log Fano curve $(X,\Delta)$ and a fixed $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ we introduce the following “restricted” version of Tian’s $\alpha$-invariant:
$\alpha(X,\Delta)_{0}\coloneqq\sup\left\{\gamma\in]0,\infty[\colon\mathcal{I}_{\gamma,\mu_{0}}(\phi_{\mu})\coloneqq\int_{X}e^{-\gamma(\phi_{\mu}-\phi_{0})}\mu_{0}<\infty\,\,\forall\mu\in\mathcal{P}(X)_{0}\right\},$
where $\phi_{\mu}$ is the potential of $\mu$ (formula 3.7).
Lemma 3.12.
Let $(X,\Delta)$ be a log Fano curve. Given $\phi_{0},$
$\alpha(X,\Delta)_{0}=\sup\left\{t\in]0,\infty[\colon\exists C_{t}\in\mathbb{R}\,\,\,\mathcal{I}_{t,\mu_{0}}(\phi_{\mu})\leq C_{t}\,\,\,\forall\mu\in\mathcal{P}(X)_{0}\right\}.$
Moreover, $\alpha(X,0)_{0}\geq 1$ and $\alpha(X,\Delta_{w})_{0}\geq\min\left\{\frac{1}{2-2w},1\right\},$ when $w\neq 0.$
Proof.
Given $t<\alpha(X,\Delta)_{0}$ it will be enough to show that the functional $\phi\mapsto\mathcal{I}_{t,\mu_{0}}(\phi_{\mu})$ is bounded from above on the subspace $K$ of $\text{PSH(}X,-K_{(X,\Delta)})$ consisting of all potentials $\phi$ of $\mu\in\mathcal{P}(X)_{0}.$ To this end first note that $K$ is a compact subspace of $\text{PSH(}X,-K_{(X,\Delta)}),$ endowed with the $L^{1}$–topology. Indeed, $\mathcal{P}(X)_{0}$ is a closed compact subset of the compact space $\mathcal{P}(X),$ hence compact and, as is well-known, the map $\mu\mapsto\phi_{\mu}$ is a homeomorphism. It will thus be enough to show that $\mathcal{I}_{t,\mu_{0}}$ is continuous on $K.$ But since $\mathcal{I}_{t+\delta,\mu_{0}}$ is pointwise finite on $K$ for $\delta$ a sufficiently small positive number, this is shown precisely as in the proof of [10, Cor A.2], using the corresponding local result of Demailly–Kollar.
We next prove the lower bounds on $\alpha(X,\Delta_{w})_{0}.$ First recall that for any function $\psi$ in $L_{\text{loc}}^{1}(\mathbb{C})$ we have
(3.32) $(dd^{c}\psi)(\{0\})<1\implies\int_{|z|\leq r}e^{-\psi}idz\wedge d\bar{z}<\infty$
for $r$ sufficiently small (as follows, for example, from Green’s formula in $\mathbb{C}$ or Skoda’s integrability theorem in $\mathbb{C}^{n}$). Now, first consider the case $w=0$ and take $\mu\in\mathcal{P}(X)_{0}.$ By the previous lemma, $\mu(\left\{x\right\})\leq 1/2$ for any $x\in X.$ Since $V=2$ this means that $dd^{c}\phi_{\mu}(\left\{x\right\})\leq 1$ for any $x\in X.$ By 3.32, this implies that $e^{-t\phi_{\mu}}\in L_{\mathrm{loc}}^{1}$ for any $t<1.$ Hence, $\alpha(\mathbb{P}^{1},0)_{0}\geq 1.$ Finally, in the case $w\neq 0$ we may assume that $\Delta_{w}=w([0]+[\infty]).$ On a disc $D_{r}$ of radius $r$ centered at $x=0,$
$\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}\leq C\int_{|z|\leq r}e^{-\left(t\phi+w\log|z|^{2}\right)}idz\wedge d\bar{z}.$
Setting $\psi\coloneqq t\phi+w\log|z|^{2}$ we have,
$(dd^{c}\psi)(\{0\})\leq tV/2+w=t(1-w)+w\,\,\,(<1\iff t<1).$
Hence, by 3.32, $\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ if $t<1$ and the same argument applies when 0 is replaced by $\infty.$ Finally, for $x$ different from 0 and $\infty$ we can use that $(dd^{c}\psi)(\{x\})\leq tV$ (since $\int dd^{c}\phi=V$) to deduce that $\int_{D_{r}(x)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ when t<1/V=1/(22w),t<1/V=1/(2-2w), as desired. ∎
The lower bounds on $\alpha(X,\Delta_{w})_{0}$ in the previous lemma are, in fact, equalities. Indeed, take $(x_{0},x_{1})\in X^{2}$ such that $\boldsymbol{m}_{2}((x_{0},x_{1}))=0$ and $x\in X$ such that $\boldsymbol{m}(x)=0$ and set
(3.33) $\mu\coloneqq\frac{1}{2}\delta_{x_{0}}+\frac{1}{2}\delta_{x_{1}},\,\text{when}\,w\in\{0\}\sqcup[\frac{1}{2},1[\,\mathrm{\ and\ }\,\mu\coloneqq\delta_{x}\,\mathrm{when}\,w\in]0,\frac{1}{2}[.$
Then $\mu\in\mathcal{P}(X)_{0}$ and a direct calculation reveals that $\mathcal{I}_{t,\mu_{0}}(\phi_{\mu})=\infty$ for $t\coloneqq\min\{\frac{1}{2-2w},\ 1\}.$ At any event, the upper bound on $\alpha(X,\Delta_{w})_{0}$ is also a consequence of the next theorem, which contains the analytic part of Theorem 1.3. The proof of the lower bounds exploit the lower bound on $\alpha(X,\Delta_{w})_{0}$ in the previous lemma and the proof of the upper bounds leverages geodesic rays in $\mathcal{H}_{1,1}(X,\omega_{0}).$
Theorem 3.13.
For the log Fano curves $(X,\Delta_{w})$,
$\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\sup\left\{\gamma\in]0,\infty[\colon F(-\gamma)_{0}>-\infty\right\}=2\alpha(X,\Delta_{w})_{0}$
where $F(-\gamma)_{0}$ is defined by formula 3.13. Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
Proof.
Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [10] (concerning the case $\mathcal{G}$ trivial) to the present setup. Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.14 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
$D(\mu)\geq t\int_{X}u_{\mu}\mu-\log\mathcal{I}_{t,\mu_{0}}(u_{\mu})\geq t\int_{X}u_{\mu}\mu-C_{t},$
using Lemma 3.12. Next, setting $u_{\mu}\coloneqq\phi_{\mu}-\phi_{0}$ (formula 3.7) we deduce, using $E(\mu)=-\int_{X}u_{\mu}\mu/V2$ (Example 3.2),
$F_{\gamma}(\mu)\geq(-\gamma+2t)E(\mu)-C_{t},$
which is bounded from below if $\gamma\leq 2t,$ proving $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$
Next, when $w\neq 0$ we may as well assume $\Delta=w[0]+w[\infty].$ Recall that w=1V/2.w=1-V/2. Hence, the desired lower bound on $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ follows from the lower bound on $\alpha(X,\Delta)_{0}$ in Lemma 3.12. To prove the upper bound let us first show that
(3.34) $2\geq\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$
To this end recall that, $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ is independent of the reference $\phi_{0}$ (by Prop 3.6, using that the Futaki character of $(X,\Delta_{w})$ vanishes). Hence, we can take $\phi_{0}$ to be induced by the Fubini–Study metric. It will be enough to show that there exists a curve $\mu^{t}$ in $\mathcal{P}(X)_{0}$ such that for any $\gamma>2,$ $F_{\gamma}(\mu^{t})\rightarrow$-$\infty$ as $t\rightarrow\infty.$ To this end first note that if $\mu$ is $S^{1}$–invariant under the standard $S^{1}$–action of $\mathbb{P}^{1}$ fixing 0 and $\infty$ and $\mathbb{Z}_{2}$–invariant under the $\mathbb{Z}_{2}$–action $(x_{1},x_{2},x_{3})\mapsto(-x_{1},-x_{2},-x_{3})$ then $\boldsymbol{m}(x_{1},x_{2},x_{3})=-\boldsymbol{m}(-x_{1},-x_{2},-x_{3})$ and it follows that $\mu\in\mathcal{P}(X)_{0}.$ We will construct $\mu^{t}$ as $\mu^{t}=dd^{c}\phi_{t}/V,$ where $\phi_{t}$ is an $S^{1}$–invariant geodesic in $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}},$ using the notation in [16] (recalled in the appendix), where such a curve $\phi_{t}$ is realized as the Legendre transform of a function $v_{t}(y)$ on the moment polytope P=[V/2,V/2]P=[-V/2,V/2] of $(X,-K_{(X,\Delta)})$ of the form
(3.35) $\phi_{t}=v_{t}^{*},\,\,\,\,v_{t}\coloneqq v_{0}+tv,$
where $v$ is a piecewise affine convex function on [1,1][-1,1] and $v_{0}$ is the Legendre transform of $\phi_{0}.$ It follows from well-known general results that such a curve $\phi_{t}$ is a geodesic ray in $\mathcal{H}(-K_{X})_{1,1}^{S^{1}}$ (see [41, Section 7]). Here we will take
$v(y)=|y|.$
Then $v_{t}$ is invariant under $y\mapsto-y$ and as a consequence, $\mu^{t}$ is $\mathbb{Z}_{2}$-invariant. Moreover, as shown in the appendix, for any $\phi\in\mathcal{H}(-K_{X})_{1,1}^{S^{1}},$ setting $\mu=dd^{c}\phi/V$,
$F_{\gamma}(\mu)=\gamma\frac{\mathcal{L}(v)}{V}+(1-\gamma)\frac{V}{2}\int_{P}\left|v^{\prime}\right|dy-\int_{P}v^{\prime\prime}dy+O(1),\,\,\,\frac{\mathcal{L}(v)}{V}\coloneqq\int_{\partial P}\frac{v}{2}-\int_{P}v\frac{dy}{V},$
where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 8.1, combined with 8.3). In particular,
$\frac{\mathcal{L}(v_{t})-\mathcal{L}(v_{0})}{V}=t\left(\frac{V}{2}-\int_{-V/2}^{V/2}|y|\frac{dy}{V}\right)=Vt\left(\frac{1}{2}-(\frac{1}{2})^{2}\right)=t\frac{V}{4}$
and
$\int_{-V/2}^{V/2}\left|v_{t}^{\prime}(y)\right|\frac{dy}{V}=2\int_{0}^{V/2}\left|u^{\prime}_{0}(y)+t\right|\frac{dy}{V}=t+O(1),$
as $t\rightarrow\infty.$ Since $v_{t}^{\prime\prime}$ is independent of $t,$ it follows that
$F_{\gamma}(\mu^{t})=\gamma t\frac{V}{4}+(1-\gamma)t\frac{V}{2}+O(1)=tV\left(\gamma\left(\frac{1}{4}-\frac{1}{2}\right)+\frac{1}{2}\right)+O(1),$
which tends to -$\infty$ if $\gamma>2,$ proving the inequality 3.34. It remains to show that when $w\in]0,1[$
$\frac{2}{V}\geq\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}.$
To this end, let $A_{t}$ be the automorphism of $X\subset\mathbb{R}^{3}$ obtained as the flow at time $t$ of a holomorphic vector field with zeros at $(1,0,0)$ and (1,0,0),(-1,0,0), so that $(1,0,0)$ is the attractive fixed point of $A_{t}.$ More precisely, in the affine holomorphic coordinate $z$ centered at $(1,0,0)$ (and equal to $\infty$ at (1,0,0))(-1,0,0)) we have
(3.36) $A_{t}(z)=e^{-t/2}z.$
Set $\mu^{t}\coloneqq(A_{t})_{*}\omega,$ where $\omega$ denotes the $\mathrm{SO}(3)$–invariant probability measure on $X.$ This is a curve in $\mathcal{P}(X)_{0}$. Indeed, $\mu_{t}$ is invariant under the $S^{1}$–action defined by rotation around the $x_{1}$–axis. As a consequence, $\mu_{t}$ is invariant under the $\mathbb{Z}_{2}$–action $(x_{1},x_{2},x_{3})\mapsto(x_{1},x_{2},-x_{3})$ (since the $\mathbb{Z}_{2}$–orbits are on the same $S^{1}$–orbit). Since $\boldsymbol{m}(x_{1},x_{2},-x_{3})=-\boldsymbol{m}(x_{1},x_{2},-x_{3})$ it follows that $\mu_{t}$ is in $\mathcal{P}(X)_{0}$. It will thus be enough to show that, as $t\rightarrow\infty,$
$E(\mu^{t})=\frac{V}{2}t+O(1),\,\,\,\,D(\mu^{t}|\mu_{0})=t+O(1).$
In the corresponding toric coordinates, corresponding to the $S^{1}$–action fixing the $x_{1}$–axis, the map $A_{t}$ corresponds to $x\mapsto x-t$ and $\mu^{t}=dd^{c}\phi_{t}/V,$ where
(3.37) $\phi_{t}(x)=A_{-t}^{*}\phi_{0}=\phi_{0}(x+t).$
The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 8.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
$E(\mu^{t})=0+\frac{V}{2}t+O(1).$
Next, expressing $\mu_{0}=e^{-f}\omega,$ where $f$ is smooth away from north and south pole, gives
$D(\mu^{t}|\mu_{0})=D(\mu^{t}|\omega)+\int_{X}f\mu^{t}=D(\mu^{t}|\omega)+O(1),$
using that $\mu^{t}$ converges weakly, as $t\rightarrow\infty$ to a Dirac mass at the point $(1,0,0),$ where $f$ is continuous. Finally,
$D(\mu^{t}|\omega)=D((A_{t})_{*}\omega|\omega)=D(\omega_{0}|(A_{-t})_{*}\omega)\coloneqq\int_{X}\log\frac{\omega}{(A_{-t})_{*}\omega}\omega=t+O(1),$
where the last equality results from writing
$\log\frac{\omega}{(A_{-t})_{*}\omega}=\log\frac{1+|z|^{2}}{1+e^{-t}|z|^{2}}+t$
and using the dominated convergence theorem. ∎
The proof of the previous theorem shows that when $\gamma>\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ the functional $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is destabilized by a geodesic ray, i.e. $F_{\gamma}(\mu_{t})\rightarrow-\infty,$ as $t\rightarrow\infty,$ for a curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ of the form $\mu_{t}=\omega_{u_{t}}/V,$ where $u_{t}$ is a geodesic ray in $\mathcal{H}_{1,1}(X,\omega_{0}).$ This is reminiscent of the general existence/properness principle in [48, Section 3]. However, in the present setup the function $F_{\gamma}(\omega_{u_{t}}/V)$ need not be convex in $t$ (since $\gamma>1)$ which is required in [48]. Moreover, our geodesics are required to satisfy the constraint $\int_{X}\boldsymbol{m}\omega_{u_{t}}=0.$
3.5. Stability thresholds on the space of geodesic rays vs divisorial valuations
Denote by $\mathcal{R}_{1,1}(X,\omega_{0})_{0}$ the subspace of geodesics rays $\boldsymbol{u}\coloneqq\{u_{t}\}$ such that $\int\boldsymbol{m}\omega_{u_{t}}^{n}=0$ for all $t.$ Given a convex functional $f$ on $\mathcal{H}_{1,1}(X,\omega_{0})$ we will denote by $\boldsymbol{f}$ the functional on $\mathcal{R}_{1,1}(X,\omega_{0})_{0}$ defined as the slope at infinity of $f,$ i.e. the limit of $f(u_{t})/t,$ as $t\rightarrow\infty,$ in $\infty$$\infty$]. Consider now the invariant
$\overrightarrow{\delta}_{0}(X,\Delta)\coloneqq\inf_{\mathcal{R}_{1,1}(X,\omega_{0})_{0}}\frac{\boldsymbol{D}}{\boldsymbol{E}},$
where $E(u_{t})\coloneqq E(\omega_{u_{t}}^{n}/V)\coloneqq(\mathcal{I}-\mathcal{J})(u_{t}),$ $D(u_{t})\coloneqq D(\omega_{u_{t}}^{n}/V|\mu_{0})$ and $(\boldsymbol{D}/\boldsymbol{E})(\boldsymbol{u})\coloneqq\infty$ if $\boldsymbol{E}(\boldsymbol{u})=0.$ We recall that, $E(u_{t})$ and $D(u_{t})$ are convex with respect to $t$ (by the proof of [21, Lemma 9.1] and [27, Thm 4.10], respectively) and thus $\boldsymbol{D}$ and $\boldsymbol{E}$ are well-defined. The following proposition shows, in particular, that $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}.$
Proposition 3.14.
For the log Fano curves $(X,\Delta_{w})$ we have $\overrightarrow{\delta}_{0}(X,0)=2$ and $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\min\{\frac{1}{1-w},2\}$ when $w\neq 0.$ Moreover, when $\phi_{0}$ is induced by the Fubini–Study metric, there is a geodesic ray $u_{t}$ in $\boldsymbol{H}_{1,1}(X,\omega_{0})_{0},$ attaining the inf defining $\overrightarrow{\delta}_{0}(X,\Delta)$, with the property that the limit of $\omega_{u_{t}}/V$ in $\mathcal{P}(X)_{0}$ as $t\rightarrow\infty$ is the measure $\mu$ defined by formula 3.33.
Proof.
[…]
Combining the previous proposition with the results in Section 2.3 shows that $\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right)$ converges towards $\overrightarrow{\delta}_{0}(X,\Delta),$ i.e.
$\lim_{N\rightarrow\infty}\inf_{v}\frac{A_{(X^{N},\Delta_{N})}(\boldsymbol{v})}{\boldsymbol{v}(\mathcal{D}^{(N)})}=\inf_{\mathcal{R}_{1,1}(X,\omega_{0})_{0}}\frac{\boldsymbol{D}}{\boldsymbol{E}},$
where $\boldsymbol{v}$ ranges over all divisorial valuations over $(X^{N})_{\text{ss}}.$ Moreover, there exists a sequence of decomposable divisorial valuations $\boldsymbol{v}_{N}$ over $X^{N}$ asymptotically minimizing the lhs above, which are the valuations over $(X^{N})_{\text{ss}}$ induced by some $(x_{1},\dots,x_{N})\in(X^{N})_{\text{ss}},$ such that for $N$ even
(3.38) $\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}=\lim_{t\rightarrow\infty}\frac{\omega_{u_{t}}}{V}\,\,\,\,\,\,(=\mu)$
where $u_{t}$ is the geodesic ray minimizing $\boldsymbol{D}/\boldsymbol{E}$ over in $\mathcal{R}_{1,1}(X,\omega_{0})_{0}.$ More generally, for a general positive integer $N(\geq 2)$ the measure in the lhs above has to be replaced by its weak large-$N$ limit.
Remark 3.15.
The geodesic ray $u_{t}$ appearing in the previous proposition is attached to a test configuration $(\mathcal{X},\mathcal{L})$ for $(X,-K_{(X,\Delta)})$ and thus defines a psh metric on the Berkovich analytification of $(X,-K_{(X,\Delta)}),$ with respect to the trivial valuation on $\mathbb{C},$ in the sense of [33, 35]. Moreover, the Monge–Ampère measure $\text{MA(}(\mathcal{X},\mathcal{L}))$ of $(\mathcal{X},\mathcal{L})$ on the Berkovich analytification $X^{\text{an}}$ of $X$ coincides with the measure $\mu$ defined by formula 3.33, under the standard embedding of a complex curve $X$ into $X^{\text{an}}$ (where $x\in X$ is mapped to the divisorial valuation defined by $x).$ This follows from the fact that, by construction, the test configuration $(\mathcal{X},\mathcal{L})$ is toric for some $\mathbb{C}^{*}$–action on $\mathbb{P}^{1}$. Thus formula 3.38 translates into the following weak convergence of measures on the Berkovich analytification of $\mathbb{P}^{1}$:
$\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\delta_{v_{i}}=\text{MA(}(\mathcal{X},\mathcal{L}))$
for the valuations $v_{i}$ on $\mathbb{P}^{1}$ defined by the points $x_{i}.$
4. Applications to logarithmic Hardy–Littlewood–Sobolev inequalities, Moser–Trudinger inequalities and spontaneous symmetry breaking
We continue with the setup in the previous section applied to the case when $X$ is the Riemann sphere and $\Delta=\Delta_{w}$ for $w\in[0,1].$ The lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ resulting from Theorem 3.13 is equivalent to the inequality
$E(\mu)\leq\frac{1}{\gamma}D(\mu|\mu_{0})+C_{\gamma}$
for any $\mu\in\mathcal{P}(X).$ When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38]. In this case it holds without the moment constraint (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant). Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ when $\mu$ has vanishing moment (formula 3.1). In contrast to [37, 38], it does not seem possible to use symmetric decreasing rearrangements when $\gamma>1$ (see Section 4.1.1).
We next show that Theorem 3.13 also implies Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2]. To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
(4.1) $\mathcal{D}_{\gamma}(u)\coloneqq-\mathcal{E}(u)-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0}.$
It was introduced in [51] (when $\Delta=0)$ [^3] In the notation of [51], $\mathcal{D}_{\gamma}(u)=F_{\gamma}(u),$ but here the notation $F_{\gamma}$ is reserved for the free energy functional.. Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (in particular, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2. We will restrict $\mathcal{D}_{\gamma}(u)$ to the subspace $C^{1}(X)_{0}$ of all $u\in C^{1}(X)$ satisfying
(4.2) $\int_{X}\boldsymbol{m}e^{-\gamma u}\mu_{0}=0$
and say that $\mathcal{D}_{\gamma}$ is destabilized by a curve $u_{t}$ in $C^{1}(X)_{0}$ if $\mathcal{D}_{\gamma}(u_{t})\rightarrow$-$\infty$ as $t\rightarrow\infty.$
Corollary 4.1.
When $w=0$ there exists for any $\gamma\in]0,2[$ a constant $C_{\gamma}$ such that for any $u\in C^{1}(X)_{0}$
$\mathcal{D}_{\gamma}(u)\geq-C_{\gamma}.$
Moreover, when $\phi_{0}$ is $\mathrm{SU}(2)$–invariant there exists a geodesic ray $u_{t}$ in $C^{1}(X)_{0}\cap\mathcal{H}(X,\omega_{0}),$ destabilizing $\mathcal{D}_{\gamma}$ for any $\gamma>2$ and when $w\neq 0$ the corresponding results holds when the threshold $2$ is replaced by $\min\{\frac{1}{1-w},2\}.$
Proof.
First observe that for any $u\in C^{1}(X)$
(4.3) $\frac{1}{\gamma}F\left(\frac{e^{-\gamma u}\mu_{0}}{\int_{X}e^{-\gamma u}\mu_{0}}\right)\leq\mathcal{D}_{\gamma}(u).$
This is essentially shown in [10], but for completeness we recall the proof. The starting point is that when $n=1$, the sup in the definition 3.5 of $E(\mu)$ may be taken over all of $C^{0}(X),$ as pointed out in Example 3.2. Hence, $E(\mu)\leq\left\langle v,\mu\right\rangle-\mathcal{E}(v)$ for any $v\in C^{0}(X).$ Now take $\mu=e^{-\gamma u}/\int_{X}e^{-\gamma u}\mu_{0}$ and apply the previous inequality for $v=u.$ The inequality 4.3 then follows from the identity
$\frac{1}{\gamma}D\left(\left.\frac{e^{-\gamma u}}{\int_{X}e^{-\gamma u}\mu_{0}}\right|\mu_{0}\right)=-\left\langle u,\frac{e^{-\gamma u}}{\int_{X}e^{-\gamma u}\mu_{0}}\right\rangle-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0},$
which is a direct consequence of the definition of $D(\mu|\mu_{0}).$ Thanks to the inequality 4.3 the lower bound on $\mathcal{D}_{\gamma}$ in the corollary thus follows directly from the lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ in Theorem 3.13. Next, to destabilize $\mathcal{D}_{\gamma}(u),$ first recall [10] that for any $u\in\mathcal{H}(X,\omega_{0})$
(4.4) $\mathcal{D}_{\gamma}(u)\leq\frac{1}{\gamma}F_{\gamma}\left(\frac{\omega_{u}}{V}\right).$
In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed, $F_{\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$ using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e. $\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.37, has the desired properties. Indeed, $\mu_{t}$ destabilizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ and, just as before, the $\mathbb{Z}_{2}$–invariance of $u_{t}$ (resulting from the $\mathbb{Z}_{2}$–invariance of $\phi_{t}$ and $\phi_{0})$ implies that $e^{-\gamma u_{t}}\mu_{0}$ has vanishing moment. Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.35, has the desired properties. Again, this follows from the symmetry argument employed in the proof of Theorem 3.13, since $u_{t}$ is $\mathbb{Z}_{2}$–invariant. ∎
Remark 4.2.
As the proof above reveals, the assumed $\mathrm{SU}(2)$–invariance of $\phi_{0}$ in the corollary, may be replaced by $\mathbb{Z}_{2}$–invariance for the $\mathbb{Z}_{2}$–actions appearing in the proof (that depend on wether $w=0$ or $w\neq 0).$
The connection to Moser–Trudinger inequalities stems from the following simple
Lemma 4.3.
Given a log Fano curve $(X,\Delta),$ $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ and $\gamma>0$ the constrained inequality
[…]
is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint:
[…]
Proof.
Note that both the constrained inequalities appearing in the lemma are invariant under the additive action of $\mathbb{R}$ on $C^{1}(X).$ Hence, we may as well assume that $\int u\omega_{0}=0$ and set $v\coloneqq\gamma u.$ Then $\mathcal{E}(u)=-\mathcal{J}(u)$ and the equivalence thus follows directly from $\mathcal{J}(\gamma^{-1}v)=\gamma^{-2}\mathcal{J}(v).$
Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10). In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.5 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the Corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
Building on Aubin’s result the following sharp form of the the Moser–Trudinger inequality with Aubin’s constraint was established in [72], that settles a conjecture in [39]:
Theorem 4.4.
[72]Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.5 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
We propose the following strengthening of the previous theorem:
Conjecture 4.5.
Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant and $\gamma\in]0,2].$ Then
$F_{\gamma}(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0}.$
In other words, the conjecture says that $F_{\gamma}(\mu)$ is minimized on $\mathcal{P}(X)_{0}$ by the $\mathrm{SO}(3)$–invariant probability measure $\omega_{0}/V.$ When $\gamma\leq 1$ this follows from the general fact that the Mabuchi functional is minimized on the Kähler potentials of Kähler–Einstein metrics [8]. We will show below that Conjecture 4.5 holds when $\gamma\leq 3/2,$ using Theorem 4.4. Note that the conjecture is equivalent to the inequality 3.30 for the Mabuchi functional $\mathcal{M}(u)$ on $S^{2}$ subject to Tian’s orthogonality condition for $u\in\mathcal{H}(X,\omega_{0})$ for $\epsilon=1$ and $C=0.$
4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
In [89], Tian conjectured that for any Fano manifold $X$ and Kähler–Einstein metric $\omega_{0}$ on $X$ there exist constants $\epsilon$ and $C$ such that
(4.6) $\mathcal{D}_{1}(u)\geq\epsilon(\mathcal{I}-\mathcal{J})(u)-C$
for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.29. The conjecture was settled in [48] and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.5 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.29 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87]. Here we show, building on Theorem 4.4, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.4 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.6):
Proposition 4.6.
Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X)$ (i.e. $\int\omega_{0}=2).$ For any $v$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.29
$\frac{1}{2}\mathcal{J}(v)-\int_{X}v\omega_{0}-\log\int_{X}e^{-v}\frac{\omega_{0}}{2}\geq 0,$
i.e. the inequality is saturated when $v$ is constant.
Proof.
First note that it is is equivalent to show that for any $v$ in $\mathcal{H}(X,\omega_{0})$
(4.7) $\mathcal{D}_{1}(v)\geq\frac{1}{2}\mathcal{J}(v)$
if $v$ satisfies Tian’s orthogonality condition (which, in this case, is equivalent to $\int\boldsymbol{x}\omega_{v}=0;$ see Lemma 3.9 and its proof). Indeed, since both $\mathcal{D}_{1}$ and $\mathcal{J}$ are invariant under $v\mapsto v+C,$ for any $C\in\mathbb{R},$ we may as well assume that $\int v\omega_{0}=0.$ Then $\mathcal{E}(v)=\mathcal{J}(v)$ which yields the equivalence in question. Now take $u$ in $\mathcal{H}(X,\omega_{0})$ and consider the right action of $\mathcal{G}$ on $\mathcal{H}(X,\omega_{0})$ induced by the identification between $\phi\in\text{PSH}(L)$ and $u\coloneqq\phi-\phi_{0},$ where $\phi_{0}$ is a fixed metric on $K_{X}$ with (normalized) curvature form $\omega_{0}.$ In other words, $g^{*}\phi=u\cdot g+\phi_{0}$ and it follows that
$g^{*}\left(e^{-u}\omega_{0}\right)=e^{-u\cdot g}\omega_{0}$
Indeed, since $\omega_{0}$ is KE we may suppose that $\omega_{0}=\mu_{\phi_{0}},$ i.e. $e^{-u}\omega_{0}=\mu_{\phi_{0}+u}.$ Using $g^{*}\mu_{\psi}=\mu_{g^{*}\psi}$ thus gives
$g^{*}\left(e^{-u}\omega_{0}\right)=\mu_{g^{*}(\phi_{0}+u)}\eqqcolon\mu_{\phi_{0}+u\cdot g}=e^{-u\cdot g}\mu_{\phi_{0}}=e^{-u\cdot g}\omega_{0}.$
Next, by Prop 3.4, there exists $g_{0}\in\mathcal{G}$ such that $\int_{X}\boldsymbol{x}g_{0}^{*}\left(e^{-u}\omega_{0}\right)=0,$ i.e. $\int_{X}\boldsymbol{x}e^{-u\cdot g_{0}}\omega_{0}=0.$ Since $\mathcal{D}$ is $\mathcal{G}$–invariant we conclude that
$\mathcal{D}_{1}(u)=\mathcal{D}_{1}(u\cdot g_{0})\geq\frac{1}{2}\mathcal{J}(u\cdot g_{0})\geq\frac{1}{2}\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)$
by applying the inequality 4.7 to $v\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.7 indeed holds if $u$ satisfies Tian’s constraint. ∎
As a consequence, we deduce the inequality in Conjecture 4.5 for $\gamma\leq 3/2$ (which corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.6):
Proposition 4.7.
Assume $\gamma\leq 3/2.$ Then, for all measures $\mu$ in $\mathcal{P}(X)_{0}$
$F_{\gamma}(\mu)\geq F_{\gamma}\left(\frac{\omega_{0}}{V}\right)(=0)$
If it is moreover assumed that $\mu\geq(1-1/\gamma)\omega_{0}/V,$ then the previous inequality holds when $\gamma\in[0,2].$
Proof.
Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.4 with the inequality 4.7 gives
$F_{1}\left(\frac{\omega_{u}}{V}\right)\geq\frac{1}{2}\mathcal{J}(u)\,\,\,\,\text{if $\int\boldsymbol{x}\omega_{u}=0$. }$
Since $\mathcal{J}(u)=E(\omega_{u}/V)$ this equivalently means that $F_{1}(\omega_{u}/V)-E(\omega_{u}/V)/2\geq 0.$ In other words, $F_{3/2}(\omega_{u}/V)\geq 0.$ Using that $F_{\gamma}(\mu)$ is decreasing with respect to $\gamma$ (since $E\geq 0)$ this proves the inequality in the corollary. To prove the last statement first note that it follows from the previous proposition, combined with the scaling argument in the proof of Lemma 4.3, that
$\mathcal{D}_{\gamma}(v)\geq 0,\,\,\text{for}\,\gamma\in[0,2]$
if $\gamma v$ satisfies Tian’s orthogonality condition and $\gamma v\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{v}=0$ (using that, by homogeneity, $v$ satisfies Tian’s orthogonality condition, i.e. $0=\int\boldsymbol{x}v\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{v},$ using integration by parts). Furthermore, the condition that $\gamma v\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{v}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption. We can thus apply the inequality 4.4 to $v$ to conclude that $F_{\gamma}(\frac{\omega_{v}}{V})\geq 0.$
4.1. Spontaneous symmetry breaking
In this section we take the corresponding fixed element $\phi_{0}\in\mathcal{H}(-K_{(\mathbb{P}^{1},\Delta_{w})})$ to be induced by the Fubini–Study metric, i.e. to be $\mathrm{SU}(2)$–invariant (Example 3.10) and the points $x_{0}$ and $x_{1}$ as the north and south pole of $X,$ identified with the unit sphere in $\mathbb{R}^{3}.$ Given $(X,\Delta_{w}),$ denote by $\mathcal{P}(X)_{0}^{S^{1}}$ the subspace of $\mathcal{P}(X)_{0}$ consisting of all $S^{1}$–invariant measures $\mu,$ with respect to the standard $S^{1}$–action fixing $x_{0}$ and $x_{1},$ i.e. rotation around the $x_{3}$–axis. Likewise, we denote by $C^{1}(X)_{0}^{S^{1}}$ the space of all $S^{1}$–invariant $u\in C^{1}(X)$ satisfying the moment condition 4.2.
Proposition 4.8.
For any $\gamma<2,$ the functional $F_{-\gamma}(\mu)$ is bounded from below on $\mathcal{P}(X)_{0}^{S^{1}}$ and, as a consequence, $\mathcal{D}_{\gamma}(u)$ is bounded from below on $C^{1}(X)_{0}^{S^{1}}.$
Proof.
By the argument in the beginning of the proof of Theorem 3.13 it is enough to show that
$\alpha(X,\Delta)_{0}^{S^{1}}\coloneqq\sup\left\{\gamma\in]0,\infty[\colon\mathcal{I}_{\gamma,\mu_{0}}(\phi_{\mu}-\phi_{0})<\infty\,\,\forall\mu\in\mathcal{P}(X)_{0}^{S^{1}}\right\}\geq 1.$
But this follows from the proof of Lemma 3.12. Indeed, for $x=0,\infty$, the argument is the same, while for $x\neq 0,\infty$, we use that if $\mu\in\mathcal{P}(X)_{0}^{S^{1}},$ then $\phi_{\mu}$ is locally bounded on $\mathbb{C}^{*},$ identified with $\mathbb{P}^{1}-\{0,\infty\}.$ Indeed, in the standard trivialization over $\mathbb{C}^{*}$ the function $\phi_{\mu}$ is $S^{1}$–invariant and psh on $\mathbb{C}^{*}$ and can thus be expressed as $\phi_{\mu}(z)=\Phi(x),$ for $x=\log(|z|^{2}),$ where $\Phi$ is a convex continuous function $\Phi$ on $\mathbb{R}.$ Finally, the lower bound on $\mathcal{D}_{\gamma}(u)$ follows from the inequality 3.12. ∎
The previous proposition reveals that a phenomenon of spontaneous symmetry breaking arises when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[.$ Indeed, in this case, by Thm 3.13, Cor 4.1 and the previous proposition,
(4.8) $\infty=\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}<\inf_{\mathcal{P}(X)_{0}^{S^{1}}}F_{-\gamma},\,\,\,-\infty=\inf_{C^{1}(X)_{0}}\mathcal{D}_{\gamma}(u)<\inf_{C^{1}(X)_{0}^{S^{1}}}\mathcal{D}_{\gamma}(u)$
although both $F_{\gamma}$ and $\mathcal{P}(X)_{0}$ are $S^{1}$–invariant (and likewise for $\mathcal{D}_{\gamma}$ and $C^{1}(X)_{0})$. The point is that the destabilizing curve $\mu^{t}$, exhibited in the proof of Theorem 3.13, is not $S^{1}$–invariant (with respect to the $S^{1}$–action fixing 0 and $\infty).$ This is in contrast to the case when $\gamma=1,$ where the two infima above do coincide. Indeed, when $\gamma=-1$, the inf of $F_{\gamma}$ is attained at the unique Kähler–Einstein volume form in $\mathcal{P}(X)_{0}$ and this volume form is $S^{1}$–invariant (and likewise for $\mathcal{D}_{1}(u)).$
The previous phenomenon of spontaneous symmetry breaking is reinforced on the space $\mathcal{R}_{1,1}(X,\omega_{0})_{0}$ of geodesic rays with vanishing moment (defined in the beginning of Section 3.5). Indeed, combining Prop 3.14 and Prop 4.8 reveals that, when $w\in]0,1/2[,$
$\frac{1}{1-w}=\inf_{\mathcal{R}_{1,1}(X,\omega_{0})_{0}}\frac{\boldsymbol{D}}{\boldsymbol{E}}<\inf_{\mathcal{R}_{1,1}(X,\omega_{0})_{0}^{S^{1}}}\frac{\boldsymbol{D}}{\boldsymbol{E}}=2$
where $\mathcal{R}_{1,1}(X,\omega_{0})_{0}^{S^{1}}$ denotes the $S^{1}$–invariant subspace of $\mathcal{R}_{1,1}(X,\omega_{0})_{0},$ with respect to the $S^{1}$–action on $\mathcal{R}_{1,1}(X,\omega_{0})_{0}$ induced by the $S^{1}$–action on $X.$ More precisely, by the proof of Theorem 3.13, the inf in the lhs above is attained for a geodesic ray that is not $S^{1}$–invariant. The $S^{1}$–invariant function $\boldsymbol{D}/\boldsymbol{E}$ on the space $\mathcal{R}_{1,1}(X,\omega_{0})_{0}$ thus resembles the classical “Mexican hat potential”, exemplifying the prototypical phenomenon of spontaneous symmetry breaking in the physics literature – an $S^{1}$–invariant function whose minima are not $S^{1}$–invariant.
4.1.1. Failure of symmetric rearrangements
In [37, 38] and [9], the sharp logarithmic HLS inequality on the two-sphere $X$ (corresponding to the case $w=0$ and $\gamma=1)$ was shown using symmetric decreasing rearrangements. The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [37, 38] (using stereographic projection), while in [9] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere. Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.8 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$
4.2. Restricting to $\mathcal{K}$-invariant metrics
Denote by $\mathcal{H}(X,\omega_{0})^{\mathcal{K}}$ the $\mathcal{K}$–invariant subspace of $\mathcal{H}(X,\omega_{0})$ and by the $C(\mathcal{G})$ the center of $\mathcal{G}.$ The following analytic invariant appears implicitely in [48, 63, 77]:
$\delta_{C(\mathcal{G})}^{A}(X,\Delta)\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon\inf_{g\in C(g)}(\mathcal{I}-\mathcal{J})-C_{\epsilon}\,\,\text{on\,$\mathcal{H}(X,\omega_{0})^{\mathcal{K}}$}\right\}$
By [77, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $(X,\Delta)$ admits a KE metric. Hence, by [48, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $\delta_{\mathcal{G}}^{A}(X,\Delta)>1.$ However, in general these two analytic invariants may differ:
Proposition 4.9.
For the log Fano curves $(X,\Delta_{w})$ we have $\delta_{C(\mathcal{G})}^{A}(X,0)=\infty$ and $\delta_{C(\mathcal{G})}^{A}(X,\Delta_{w})=2$ when $w\neq 0.$ As a consequence, $\delta_{C(\mathcal{G})}^{A}(X,\Delta_{w})=\delta_{\mathcal{G}}^{A}(X,\Delta_{w})$ iff $w\in[1/2,1].$
Proof.
Consider first the case $w=0.$ As in Prop 3.6 we may assume that $\phi_{0}$ is the Fubini–Study metric. Then $\mathcal{K}=\mathrm{SO}(3)$ acts transitively on $X$ and fixes $\omega_{0}.$ Hence, $\mathcal{H}(X,\omega_{0})^{\mathcal{K}}$ consists of the constant functions. As a consequence, $\delta_{C(\mathcal{G})}^{A}(X,0)=\infty.$ Next, when $w\neq 0,$ $\mathcal{G}=C(\mathcal{G})=\mathbb{C}^{*}$ and $\mathcal{K}=S^{1}.$ Hence, $\delta_{C(\mathcal{G})}^{A}(X,\Delta_{w})$ coincides with the $S^{1}$–invariant version of $\delta_{\mathcal{G}}^{A}(X,\Delta_{w}),$ which equals $2,$ by Prop 4.8. ∎
5. The $\mathcal{K}$–reduced probability measures on $X^{N}$
We continue with the setup introduced in Section 3. We thus fix a compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and, given a log Fano manifold $(X,\Delta),$ we fix a $\mathcal{K}$–invariant positively curved metric on $K_{(X,\Delta)}.$ Recall that $\boldsymbol{m}_{N}$ denotes the corresponding moment map for the diagonal action of $\mathcal{K}$ on $X^{N}$, defined $X^{N}$ with values in the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}$ (Section 3.2.7).
Before introducing the probabilistic setup we start with some preparations. First assume that 0 is a regular value of the moment map $\boldsymbol{m}_{N}.$ Given a measure $\nu$ on $X^{N}$ with a smooth density and a basis $e_{1},\dots,e_{r}$ of $\mathfrak{k}$, we will denote by $\nu_{\boldsymbol{m}_{N}}$ the corresponding Gelfand–Leray form on the submanifold $\left\{\boldsymbol{m}_{N}=0\right\}$ of $X^{N},$ uniquely determined by the following property:
(5.1) $\nu_{\boldsymbol{m}_{N}}\wedge dh_{1}\wedge\cdots\wedge dh_{r}=\nu$
when evaluated at any given point in $\left\{\boldsymbol{m}_{N}=0\right\}$. Here $h_{1},\dots,h_{r}$ are Hamiltonians for the vector fields on $X$ corresponding to $e_{1},\dots,e_{r}$ with respect to the action of $\mathcal{K}$. Note that $\mathbf{m}_{N}=\sum_{i=1}^{r}h_{i}e_{i}^{*}$. In particular, $\nu_{\boldsymbol{m}_{N}}$ defines a measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}$, that we shall call the Gelfand–Leray measure.
Another measure supported on $\left\{\boldsymbol{m}_{N}=0\right\}$ may be obtained as follows, by exploiting the complex structure $J$ on $X.$ Denote by $\nu^{\mathrm{c}}$ the (Nr,Nr)(N-r,N-r)-form on $X^{N}$ obtained by contracting the given measure $\nu$ on $X^{N}$ with $Je_{1},..,Je_{r}$ on $U$:
$\nu^{\mathrm{c}}=\nu(Je_{1},..,Je_{r},\cdot).$
Its restriction to $X^{N}$ thus defines a measure on $\left\{\boldsymbol{m}_{N}=0\right\}$, that we shall also denote by $\nu^{\mathrm{c}}.$ More generally, when $\nu$ is absolutely continuous with respect to Lebesgue measure and smooth on an open subset $U$ of $X^{N}$ we replace $X^{N}$ by $U$ and proceed as before to get measures $\nu_{\boldsymbol{m}_{N}}$ and $\nu^{\mathrm{c}}$ supported on on $U\cap\left\{\boldsymbol{m}_{N}=0\right\}.$
5.1. Probabilistic and statistical mechanical setup
5.1.1. The Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ and the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ and partition function $Z_{\boldsymbol{m}_{N}}.$
Given a log Fano manifold $(X,\Delta),$ denote by $\nu^{(N)}$ the measure on $X^{N}$ defined by
$\nu^{(N)}=i^{N^{2}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
where $\det S^{(N)}$ denotes the section of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ obtained by taking $s_{1},\dots,s_{N}$ in formula 1.3 to be a basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$ When 0 is a regular value of $\boldsymbol{m}_{N}$ we will denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the corresponding Gelfand–Leray measures, supported on $X^{N}.$ While this measure depends on the choice of basis in $\mathfrak{k},$ changing the basis only has the effect of multiplying $\nu_{\boldsymbol{m}_{N}}$ with a positive constant $C$ (the determinant of the corresponding change of bases matrix). Since $C$ is independent of $N$ this does not effect the large-$N$ limits in Conjecture 1.5 and Theorem 1.6.
When 0 is not a regular value of $\boldsymbol{m}_{N}$ we propose, tentatively, to define $\nu_{\boldsymbol{m}_{N}}$ as
[…]
Remark 5.1.
When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 5.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed). It would be interesting to know if, in general, this limit exists and if it coincides with the measure supported on the regular locus, defined by the Gelfand–Leray type measure on the regular locus.
Set
(5.3) $\mu_{\boldsymbol{m}_{N}}^{(N)}\coloneqq\nu_{\boldsymbol{m}_{N}}^{(N)}/Z_{\boldsymbol{m}_{N}},\,\,\,Z_{\boldsymbol{m}_{N}}\coloneqq\int\nu_{\boldsymbol{m}_{N}}^{(N)}.$
Assuming that $Z_{\boldsymbol{m}_{N}}<\infty$ the measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ defines a probability measure on $X^{N},$ that we shall call the $\mathcal{K}$–reduced canonical probability measure and the corresponding normalizing constant $Z_{\boldsymbol{m}_{N}}$ will be called the $\mathcal{K}$–reduced partition function.
5.1.2. The contracted setup
One could also define an alternative probability measures $(\mu^{(N)})^{\mathrm{c}}$ by replacing the Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with the contracted measure $(\nu^{(N)})^{\mathrm{c}}$ and the $\mathcal{K}$–reduced partition function $Z_{\boldsymbol{m}_{N}}$ with the contracted partition function, defined by
(5.4) $Z_{N}^{\mathrm{c}}\coloneqq\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}$
From the complex–geometric perspective this has some advantages, illustrated by Proposition 5.14 below. But, in view of the relations to the statistical mechanics literature discussed in the following section, we will mainly stick to the definition 5.3. In any event, we expect that the large-$N$ asymptotics obtained from the two constructions coincide (for example, this is the case when $n=1,$ as shown in the course of the proof of Theorem 1.6).
5.1.3. The $\mathcal{K}$–reduced Gibbs measures at inverse temperature $\beta$
Given a number $\beta\in\mathbb{R},$ set
$\nu_{\beta}^{(N)}\coloneqq e^{-\beta NE^{(N)}}\mu_{0}^{\otimes N},\,\,\,\,\,E^{(N)}(x_{1},\dots,x_{N})\coloneqq-(kN)^{-1}\log\left(\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}\right).$
When $\beta=-1$ it coincides with the measure $\nu^{(N)}$ on $X^{N}$ (but $\nu_{\beta}^{(N)}$ depends on the fixed metric $\phi_{0}$ when $\beta\neq-1).$ The measures $\nu_{\beta}^{(N)}$ on $X^{N}$ induce a family of Gelfand–Leray measures $\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}.$ Assuming that the $\mathcal{K}$–reduced partition function $Z_{\boldsymbol{m}_{N}}(\beta)$ at inverse temperature $\beta$ is finite,
$Z_{\boldsymbol{m}_{N}}(\beta)\coloneqq\int\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}<\infty,$
we obtain a measure on $X^{N}$ that we shall call the $\mathcal{K}$–reduced Gibbs measures at inverse temperature $\beta$ by setting
(5.5) $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}\coloneqq\frac{\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}}{Z_{\boldsymbol{m}_{N}}(\beta)}=\frac{e^{-\beta NE^{(N)}}\left(\mu_{0}^{\otimes N}\right)_{\boldsymbol{m}_{N}}}{Z_{\boldsymbol{m}_{N}}(\beta)}.$
When $\beta=-1$ this measure coincides with the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$, defined in the previous section, and when $\beta=0$ it coincides with the classical microcanonical measure associated to the Hamiltonians $h_{1},\dots,h_{r},$ that is ubiquitous in the statistical mechanics literature.
Remark 5.2.
From a statistical mechanics perspective, the function $E^{(N)}$ plays the role of the $N$-particle energy (or Hamiltonian) and the corresponding Gibbs measure $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ describes the corresponding equilibrium state of $N$ interacting particles confined to $\left\{\boldsymbol{m}_{N}=0\right\},$ at inverse temperature $\beta.$ The corresponding probability space $(X^{N},\mu_{\boldsymbol{m}_{N},\beta}^{(N)})$ may be viewed as a singular incarnation of the notion of mixed ensemble introduced in [54]. In particular, when $X=\mathbb{P}^{1},$ identified with the two-sphere and $\omega_{0}$ is the standard $\mathrm{SO}(3)$–invariant symplectic form on $X,$ $E^{(N)}$ coincides with the energy particle in Onsager’s statistical mechanical description of the 2D incompressible Euler equations, adapted to the two-sphere [67, 68]. The corresponding Gibbs measures $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ are invariant under the Hamiltonian flow of $E^{(N)}$ on $X^{N}$ and, building on Theorem 1.6, we show in the companion paper [4] that the empirical measure $\delta_{N},$ viewed as a random variable on the probability space $(X^{N},\mu_{\boldsymbol{m}_{N},\beta}^{(N)})$ converges in law towards the $\mathrm{SO}(3)$–invariant probability measure on the two-sphere, for any $\beta>-1.$
Proposition 5.3.
For any $\beta\in\mathbb{R}$ the probability measure $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ is $\mathcal{K}$–invariant, when it is well-defined.
Proof.
First observe that for any $g\in\mathcal{G},$ $g^{*}\det S^{(N)}=\chi_{N}(g)\det S^{(N)}$ for a homomorphism $\chi_{N}\colon\mathcal{G}\rightarrow\mathbb{C}^{*}$ (see Lemma 5.9 below). It follows that the restriction of $\chi_{N}$ to $\mathcal{K}$ takes values in $S^{1}\subset\mathbb{C}^{*}.$ Indeed, since $\mathcal{K}$ is compact the image $\chi_{N}(\mathcal{G})$ is a compact subgroup of $\mathbb{C}^{*}$ and thus contained in the unique maximally compact subgroup $S^{1}$ of $\mathbb{C}^{*}.$ Since $\phi_{0}$ is $\mathcal{K}$–invariant, it follows that $\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}$ is $\mathcal{K}$–invariant on $X^{N}.$ Finally, since the measure $\mu_{0}$ and $|\boldsymbol{m}|$ are also $\mathcal{K}$–invariant we deduce that $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ is $\mathcal{K}$–invariant. ∎
5.2. Motivation: A conjectural Large Deviation Principle
We begin by formulating a conjecture, which serves as motivation for the conjectures stated in the introduction of the paper. Recall that a sequence of measures $\Gamma_{N}$ on a topological space $\mathcal{M}$ satisfies a Large Deviation Principle (LDP) with speed $N$ and rate function $F\colon\mathcal{M}\rightarrow$$\infty$$\infty$] if $F$ is lsc and for any (Borel) measurable subset $\mathcal{B}$ of $\mathcal{M}$
[…]
Conjecture 5.4.
The following conjectures hold:
  • A(i) If $F(\beta-\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ then the measures $(\delta_{N})_{*}\left((\nu_{\beta}^{(N)})_{\boldsymbol{m}_{N}}\right)$ on $\mathcal{P}(X)_{0}$ satisfy an LDP with speed $N$ and rate functional $F_{\beta}(\mu)$. (ii) If $\mu_{\beta}^{(N)}$ are well-defined probability measures for $N\gg 1,$ then the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ on $\mathcal{P}(X)_{0}$ satisfy an LDP with speed $N$ and rate functional $F_{\beta}(\mu)-C$ for some $C\in\mathbb{R}.$
  • B(i) If $F(\beta-\epsilon)>-\infty,$ for some $\epsilon>0,$ then the measures $(\delta_{N})_{*}\left(\nu_{\beta}^{(N)}\right)$ on $\mathcal{P}(X)$ satisfy an LDP with speed $N$ and rate functional $F_{\beta}(\mu)$. (ii) if $\nu_{\beta}^{(N)}/\int\nu_{\beta}^{(N)}$ are well-defined probability measures for $N\gg 1,$ then the probability measures $(\delta_{N})_{*}\text{$\left(\nu_{\beta}^{(N)}/\int\nu_{\beta}^{(N)}\right)$}$ on $\mathcal{P}(X)$ satisfy an LDP with speed $N$ and rate functional $F_{\beta}(\mu)-C$ for some $C\in\mathbb{R}.$
Conjecture B follows from results in [18], when $n=1$ and when $\beta>0$ it is established in [12, 13] for any $n.$ In the general case, Conjecture A is motivated by the following heuristic argument. First, as shown in [12], the following mean field type approximation holds:
$E^{(N)}(x_{1},\dots,x_{N})\rightarrow E(\mu),\,\,\text{as $\delta_{N}(x_{1},\dots,x_{N})\rightarrow\mu$}$
in the sense of $\Gamma$-convergence. This suggests the formal approximation
$(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}\sim e^{-N\left(\beta E+\chi\right)}(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right),$
as probability measures on $\mathcal{P}(X),$ where $\chi=0$ on $\mathcal{P}(X)_{0}$ and $\chi=\infty$ away from $\mathcal{P}(X)_{0}.$ Next, according to Sanov’s classical LDP [50]
$(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right)\sim e^{-ND(\mu|\mu_{0})}.$
Hence, formally, $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ satisfies a LDP with rate functional $\beta F+\chi.$ Finally, since $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ is supported on $\mathcal{P}(X)_{0}$ such an LDP is equivalent to the LDP for $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ in Conjecture 5.4. The argument for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ is similar.
We next show that Conjecture 5.4 implies the conjectures stated in the introduction of the paper.
Proposition 5.5.
If Conjecture 5.4 A holds for all $\beta<0,$ then
$\gamma(X,\Delta)^{\mathcal{G}}=\sup_{\gamma>0}\left\{F(-\gamma)_{0}>-\infty\right\},\,\,\,\forall\gamma\in]0,\infty[.$
As a consequence, if Conjecture 5.4 A holds, then $\gamma(X,\Delta)^{\mathcal{G}}>1$ and the Futaki character of $(X,\Delta)$ vanishes iff $(X,\Delta)$ is K-polystable. Moreover, Conjecture 1.5 also follows. Assuming also the validity of Conjecture 5.4 B(i), it follows that $(X,\Delta)$ is uniformly Gibbs polystable iff $(X,\Delta)$ if K-polystable. Finally, assuming only the validity of Conjecture B,
$\gamma(X,\Delta)=\sup_{\gamma>0}\left\{F(-\gamma)>-\infty\right\},\,\,\,\forall\gamma\in]0,\infty[.+$
Proof.
Assume that $F(\beta-\epsilon)_{0}>-\infty,$ for some $\epsilon>0.$ Then it follows directly from the definition of the assumed LDP for $(\delta_{N})_{*}\left((\nu_{\beta}^{(N)})_{\boldsymbol{m}_{N}}\right)$ that
(5.6) $\lim_{N\rightarrow\infty}N^{-1}\log\int(\nu_{\beta}^{(N)})_{\boldsymbol{m}_{N}}=\inf_{\mathcal{P}(X)_{0}}F_{\beta}(\mu)\eqqcolon F(\beta)_{0}$
As a consequence, $\int(\nu_{\beta}^{(N)})_{\boldsymbol{m}_{N}}<\infty,$ for $N\gg 1,$ showing that
(5.7) $\gamma(X,\Delta)^{\mathcal{G}}\geq\sup_{\gamma>0}\left\{F(-\gamma)_{0}>-\infty\right\}.$
Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 5.11. Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
$0=-\lim_{N\rightarrow\infty}N^{-1}\log\int_{X^{N}}\mu_{\beta}^{(N)}=\inf_{\mathcal{P}(X)_{0}}F_{\beta}(\mu)-C,$
showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 5.7 is an equality. As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [42] for the case $\Delta=0$ with $X$ smooth and [77, 79] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
Next, to prove that Conjecture 5.4 A implies Conjecture 1.5, assume that $(X,\Delta)$ admits a KE metric. It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 5.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.12 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric). As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large. This means that $\mu_{-1}^{(N)}(=\mu_{\boldsymbol{m}_{N}}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional. But this is shown in Theorem 3.7.
Now assume that $(X,\Delta)$ is uniformly Gibbs polystable, i.e. $\gamma(X,\Delta)^{\mathcal{G}}>1$ and $(X,\Delta)$ is Gibbs semistable. Then, by [57], $(X,\Delta)$ is K-semistable, which implies that the Futaki character of $(X,\Delta)$ vanishes. Hence, as shown above, $(X,\Delta)$ is K-polystable. Conversely, if $(X,\Delta)$ is K-polystable, then it admits a KE metric and, as a consequence, the inf of $F_{-1}(\mu)$ over $\mathcal{P}(X)$ is finite. Assuming that Conjecture 5.4 B holds it follows, as before, that, for $\beta>-1,$
(5.8) $\lim_{N\rightarrow\infty}N^{-1}\log\int_{X^{N}}\nu_{\beta}^{(N)}=\inf_{\mathcal{P}(X)}F_{\beta}(\mu),\,\,\,\gamma(X,\Delta)\geq\sup_{\gamma>0}\left\{F(-\gamma)>-\infty\right\}$
Hence, for any $\beta>-1,$ $\int_{X^{N}}\nu_{\beta}^{(N)}<\infty$ for $N$ sufficiently large, which means that $(X,\Delta)$ is Gibbs semistable. Since, as shown above, Conjecture 5.4 A implies that $\gamma(X,\Delta)^{\mathcal{G}}>1,$ it follows that $(X,\Delta)$ is uniformly Gibbs polystable. ∎
Remark 5.6.
If Conjecture 5.4 holds, then $(X,\Delta)$ is Gibbs polystable iff its Futaki character vanish and $\gamma(X,\Delta)^{\mathcal{G}}>1,$ by the previous proposition.
Before continuing we recall that, by [94],
(5.9) $\sup_{\gamma>0}\left\{F(-\gamma)>-\infty\right\}=\delta(X,\Delta),$
where $\delta(X)$ is the algebro-geometric invariant, introduced in [57] and realized as a valuative stability threshold in [30]. As shown in [57] (see also [19] for an analytic proof),
$\gamma(X,\Delta)\coloneqq\liminf_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)\leq\delta(X,\Delta)$
5.2.1. A digression on the need for $\epsilon>0$ in Conjecture 5.4
It should be stressed that, in general, it is not enough to assume that Conjecture B holds with $\epsilon=0$ (and, most likely, not in Conjecture A either). Indeed, if Conjecture B holds with $\epsilon=0,$ then it would imply the implication
$F_{-1}\,\,\text{bounded below on $\mathcal{P}(X)\implies F_{-1}\,\,\text{is lsc on $\mathcal{P}(X)$}$}$
But this implication fails, in general. Indeed, by [15, Thm 5.11], $F_{-1}$ is lsc on $\mathcal{P}(X)$ iff $X$ admits a KE metric and $\mathcal{G}$ is trivial (which, in turn, is equivalent to $X$ being K-stable [42]). Hence, already $X=\mathbb{P}^{1}$ is a counterexample. But there are also counterexamples when $\mathcal{G}$ is trivial. Indeed, $F_{-1}$ is bounded from below on $\mathcal{P}(X)$ iff $(X,\Delta)$ is K-semistable (see [75] for the case when $X$ is smooth and $\Delta$ is trivial and the proof of [1, Thm 2.5] for the general case). However, there are examples of K-semistable log Fano manifolds $(X,\Delta)$ with $\mathcal{G}$ is trivial, such that $(X,\Delta)$ is not K-stable (for example, this can be achieved when $X=\mathbb{P}^{1}$ and $\Delta$ consists of three points).
It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)). If inequality 5.9 can be strengthened as follows, then the second assumption implies the first:
Lemma 5.7.
Assume that there exists $N$ such that $\delta(X,\Delta)\geq\gamma^{(N)}(X,\Delta).$ Then $\int\nu_{-\gamma}^{(N)}<\infty$ implies that there exists $\epsilon>0$ such that $F(-\gamma-\epsilon)>-\infty.$ Likewise, if $\sup_{\gamma>0}\left\{F(-\gamma)_{0}>-\infty\right\}\geq\gamma^{(N)}(X,\Delta)^{\mathcal{G}},$ then $\int(\nu_{\beta}^{(N)})_{\boldsymbol{m}_{N}}<\infty$ implies that there exists $\epsilon>0$ such that $F(-\gamma-\epsilon)_{0}>-\infty.$
Proof.
The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds). The second statement is shown in a similar way, using Prop 5.11 below. ∎
For example, when $n=1$, it follows from Theorem 1.3 that $\gamma^{(N)}(X,\Delta)\leq\gamma(X,\Delta)$ for all $N$ and, as a consequence, $\delta(X,\Delta)\geq\gamma^{(N)}(X,\Delta).$ Moreover, in the toric case we have the following result, obtained by combining [57] with [65]:
Proposition 5.8.
Let $X$ be a toric Fano variety which is Gorenstein. If $\delta(X)<1$ then $\delta(X)>\gamma^{(N)}(X)$ for $N$ sufficiently large. If, additionally, $n=2,$ then $\delta(X)\leq 1$ implies that $\delta(X)\geq\gamma^{(N)}(X)$ and if $\delta(X)<1,$ then $\delta(X)>\gamma^{(N)}(X)$ for all $N.$
Proof.
As shown in [57] (see also [19] for an analytic proof) $\delta_{k}(X)\geq\gamma^{(N_{k})}(X),$ where $\delta_{k}(X)$ is the invariant introduced in [57], defined in terms of log canonical thresholds of basis divisors on $X.$ Next, in the case of a toric Fano variety which is Gorenstein (i.e. the corresponding moment polytope is reflexive) it follows from [65, Cor 2.13] that $\delta_{k}(X)<\delta(X)$ for $k$ sufficiently large, when $\delta(X)<1.$ Finally, when $n=2$ it is shown in [65, Cor 2.13] that
$\delta_{k}(X)^{-1}=1+c_{k}\left(\delta(X)^{-1}-1\right),$
where $c_{k}>1.$ Hence, $\delta(X)=1\iff\delta_{k}(X)=1$ for all $k$ and if $\delta(X)<1,$ then $\delta_{k}(X)^{-1}>1+\delta(X)^{-1}-1,$ i.e. $\delta_{k}(X)<\delta(X),$ proving the first statement of the proposition. ∎
5.3. Gibbs polystability vs finiteness of the restricted partition function $Z_{\boldsymbol{m}_{N}}$
Lemma 5.9.
Given a log Fano manifold $(X,\Delta),$ there exists a real character $\chi_{N}$ on $\mathcal{G}$ such that $g^{*}\nu_{N}=\chi_{N}(g)\nu_{N}$ for any $g\in\mathcal{G}.$ More precisely, given $W$ in the Lie algebra of $\mathcal{G}$, $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)},$ for $t\in\mathbb{C},$ where $\text{Fut}_{(X,\Delta),k}(W)$ denotes the quantized Futaki invariant at level $k.$
Proof.
By basic linear algebra, $g^{*}\nu_{N}=\left|\det A(g)\right|^{-2/k}\nu_{N}$ where $A(\cdot)$ denotes the representation of $\mathcal{G}$ on $H^{0}(X,-kK_{(X,\Delta)}).$ Hence, $\chi_{N}(g)\coloneqq\left|\det A(g)\right|^{-2/k}$ is a real character on $\mathcal{G}$ and $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)}$ by the very definition of $\text{Fut}_{(X,\Delta),k}(W).$
Lemma 5.10.
$\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right)$ for any $\epsilon_{N}$ sufficiently small to ensure that $\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\subset(X^{N})_{\mathrm{ss}}.$
Proof.
First, by the analytic definition of the lct (formula 2.4), $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}).$ Next, note that there exists $\epsilon_{N}\in]0,\infty[$ such that the closure of $\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ is contained in $(X^{N})_{\text{ss}}$ (since $\left\{\boldsymbol{m}_{N}=0\right\}$ is compact and contained in the open set $(X^{N})_{\text{ss}}).$ Hence, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ implies that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right).$ To prove the converse, assume that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right)$ and consider the integral of $\nu^{(N)}$ over a neighbourhood $U$ of a given point $(x_{1},\dots,x_{N})$ in $(X^{N})_{\text{ss }}.$ By the Kempf–Ness theorem [64, Thm 3.3], a point $(x_{1},\dots,x_{N})\in(X^{N})_{\text{ss}}$ iff $\overline{\mathcal{G}(x_{1},\dots,x_{N})}$ intersects $\left\{\left\|\boldsymbol{m}_{N}\right\|=0\right\}.$ This means that there exists $g\in\mathcal{G}$ such that $g(x_{1},\dots,x_{N})\in\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ and, as a consequence, there exists a neighborhood $U$ of $(x_{1},\dots,x_{N})$ such that $gU\subset\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}.$ Thus, by assumption, $\int_{gU}\nu^{(N)}<\infty$ and since the integrability of $\nu^{(N)}$ is $\mathcal{G}$–invariant (by the previous lemma) we deduce that $\int_{U}\nu^{(N)}<\infty,$ showing that $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}),$ as desired. ∎
Proposition 5.11.
$\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{\boldsymbol{m}_{N}}<\infty.$ In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{\boldsymbol{m}_{N}}<\infty$ for all sufficiently large $N.$
Proof.
When 0 is not a regular value of $\boldsymbol{m}_{N}$ this follows directly from the previous lemma. Now assume that 0 is a regular value of $\boldsymbol{m}_{N}$ and consider the following smooth map
(5.10) $\Phi\colon\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}\rightarrow X^{N},\,(\boldsymbol{y},\boldsymbol{t})\mapsto g_{\boldsymbol{t}}(\boldsymbol{y})\coloneqq\exp(\sum_{i\leq r}t_{i}Je_{i})\boldsymbol{y}$
This map induces a diffeomorphism between $\left\{\boldsymbol{m}_{N}=0\right\}\times[-\epsilon,\epsilon]^{r}$ and a neighbourhood $U_{\epsilon}$ of $\left\{\boldsymbol{m}_{N}=0\right\},$ for $\epsilon$ a sufficiently small positive number. This follows from the observation that, along $\left\{\boldsymbol{m}_{N}=0\right\},$ the tangent bundle of $X$ is the direct sum of the tangent bundle of $\left\{\boldsymbol{m}_{N}=0\right\}$ and the bundle spanned by the vector fields induced by $Je_{1},\dots,Je_{r}.$ Indeed, the vector fields induced by $e_{1},\dots,e_{r}$ are tangent to $\left\{\boldsymbol{m}_{N}=0\right\}$ (since $\mathcal{K}$ preserves $\left\{\boldsymbol{m}_{N}=0\right\})$ and the vector fields corresponding to $Je_{1},\dots,Je_{r}$ are linearly independent (since $Je_{i}$ is the gradient of $\left\langle\text{$\boldsymbol{m}_{N}$},e_{i}\right\rangle$ with respect to the Kähler metric $\omega_{0}$ and 0 is a regular value of $\boldsymbol{m}_{N}).$ Hence, by the previous lemma, it will be enough to show that there exists a constant $C$ (depending on $N)$ such that
(5.11) $C^{-1}\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}$
But this follows from combining the inequalities 5.12 below with the first identity in Proposition 5.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
5.4. Relating Gibbs polystability and the restricted partition function $Z_{\boldsymbol{m}_{N}}$ to integrals over $(X^{N})_{\mathrm{s}}/\mathcal{G}$
The main virtue of the following proposition, as compared with Prop 5.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
Proposition 5.12.
If $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then
$Z_{N}^{\mathrm{c}}\coloneqq\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}<\infty$
Proof.
Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 5.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small. Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 5.10. As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite. It will thus be enough to show that
(5.12) $\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}C^{-1}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}$
for some positive constant $C,$ depending on $\epsilon.$ In turn, to prove this it will be enough to show that
(5.13) $\Phi^{*}(\nu^{(N)})=e^{\left\langle\boldsymbol{t},\boldsymbol{b}_{N}\right\rangle}\Phi^{*}\left(\left.\left(\nu^{(N)}\right)^{\mathrm{c}}\right|_{\{\mathbf{m}_{N}=0\}}\right)\wedge d\boldsymbol{t}$
for some vector $\boldsymbol{b}_{N}$ (representing the quantized Futaki character at level $k$). To this end, cover $\left\{\boldsymbol{m}_{N}=0\right\}$ with a finite number of coordinate charts $V_{i}.$ In a given such chart we can locally identify $\boldsymbol{y}$ with a coordinate vector. Thus
$\Phi^{*}(\nu^{(N)})=\rho_{N}(\boldsymbol{y},\boldsymbol{t})d\boldsymbol{y}\wedge d\boldsymbol{t}$
for a local non-negative function $\rho_{N},$ which is smooth on an open subset. Next, consider the additive action of $\mathbb{R}^{r}$ on $\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}$ defined by the maps
$\widetilde{g}_{\boldsymbol{s}}(\boldsymbol{y},\boldsymbol{t})\coloneqq(\boldsymbol{y},\boldsymbol{t}+\boldsymbol{s})$
and note that
$(\widetilde{g}_{\boldsymbol{s}}^{-1})^{*}\left(\rho_{N}(\boldsymbol{y},\boldsymbol{t})d\boldsymbol{y}\wedge d\boldsymbol{t}\right)=\rho_{N}(\boldsymbol{y},\boldsymbol{t}-\boldsymbol{s}))d\boldsymbol{y}\wedge d\boldsymbol{t}.$
Next, by the very definition of the map $\Phi,$
$\Phi\left(\widetilde{g}_{\boldsymbol{t}}(\boldsymbol{y},0)\right)=g_{\boldsymbol{t}}(\boldsymbol{y})\coloneqq\exp(\sum_{i\leq r}t_{i}Je_{i})\boldsymbol{y}$
Hence, at the point $(\boldsymbol{y},\boldsymbol{t})$ in $\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r},$
$(\widetilde{g}_{\boldsymbol{t}}^{-1})^{*}\Phi^{*}(\nu^{(N)})=\Phi^{*}\left(g_{\boldsymbol{t}}^{-1}\right)^{*}\nu^{(N)}.$
But, by Lemma 5.9,
$\left(g_{\boldsymbol{t}}^{-1}\right)^{*}\nu^{(N)}=e^{-\left\langle\boldsymbol{t},\boldsymbol{b}_{N}\right\rangle}\nu^{(N)}.$
All in all, this means that
(5.14) $\rho_{N}(\boldsymbol{y},0)d\boldsymbol{y}\wedge d\boldsymbol{t}=e^{\left\langle\boldsymbol{t},\boldsymbol{b}_{N}\right\rangle}\rho_{N}(\boldsymbol{y},\boldsymbol{t})d\boldsymbol{y}\wedge d\boldsymbol{t}.$
Finally, since the derivative of $\Phi$ at $(\boldsymbol{y},0)$ maps $\frac{\partial}{\partial t_{i}}$ to the tangent vector at $\boldsymbol{y}\in X^{N}$ induced by $Je_{i},$ we have
$\rho_{N}(\boldsymbol{y},0)d\boldsymbol{y}=\left.\Phi^{*}((\nu^{(N)})^{\mathrm{c}})\right|_{(\mathbf{y},0)}.$
Hence, the desired formula 5.13 follows from formula 5.14. ∎
Remark 5.13.
Formula 5.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
Consider now the following function on $X^{N},$
(5.15) $f_{N}\coloneqq\omega_{N}(e_{1},Je_{1})\cdots\omega_{N}(e_{r},Je_{r}),$
which vanishes at some points of $\left\{\boldsymbol{m}_{N}=0\right\}$ iff 0 is not a regular value of $\boldsymbol{m}_{N}.$
Proposition 5.14.
Assume that the quantized Futaki character at level $k$ vanishes. Then the contracted form $(\nu^{(N)})^{c}$ on $X^{N}$ is $\mathcal{G}$–invariant and if 0 is a regular value of $\boldsymbol{m}_{N},$ then
$\int_{\left\{\boldsymbol{m}_{N}=0\right\}}f_{N}\nu_{\boldsymbol{m}_{N}}^{(N)}=\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}$
for a positive constant $C,$ depending only on the choice of $e_{1},\dots,e_{r}$ (and thus independent of $N)$. Here $\delta_{V}$ $V$ denotes contraction with a given vector field $V$. Moreover, in the general case when 0 is potentially not a regular value of $\mathbf{m}_{N}$, these identities hold if $\left\{\boldsymbol{m}_{N}=0\right\}$ is replaced by its intersection with the stable locus $(X^{N})_{\text{s}}.$
We recall that 0 is a regular value of $\boldsymbol{m}_{N}$ iff the stable locus $(X^{N})_{\text{s}}$ coincides with both the semistable and the polystable locus [64, Cor 3.10]. In general, $(X^{N})_{\text{s}}$ is a Zariski open subset of the semistable locus and $(X^{N})_{\text{s}}/\mathcal{G}$ is a Zariski open subset (which as an orbifold) of the projective variety defined by the GIT quotient of $X^{N}$ under the action of $\mathcal{G}.$ Before turning to the proof of this proposition we note that, combining the previous two propositions gives the following
Proposition 5.15.
Assume that the quantized Futaki character at level $k$ vanishes and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1.$ Then
$\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}<\infty$
5.4.1. Proof of Prop 5.14
We will consider a general setup of a complex reductive Lie group $G$ acting on a polarized complex projective manifold $(X,L)$ (in the previous proposition the role of $X$ is played by $X^{N})$. Fix a maximally compact subgroup $K$ of $G$ and denote by $m$ the corresponding moment map. We start with the case when 0 is a regular value of $m.$ Fix a basis $e_{1},\dots,e_{r}$ in the Lie algebra of $K.$ Let $\nu$ be a GG–invariant measure on $X$ (possibly with infinity mass) which has a smooth density on an open subset. We can then define a positive (nr,nr)(n-r,n-r)–current on $X$ by defining where $\nu$ has smooth density:
$\nu^{\mathrm{cc}}\coloneqq\delta_{e_{1}}\cdots\delta_{e_{m}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu,$
and extending it by zero to all of $X$.
Lemma 5.16.
The current $\nu^{\mathrm{cc}}$ is GG–invariant
Proof.
It suffices to show $\nu^{\mathrm{cc}}$ is GG–invariant wherever $\nu^{\mathrm{cc}}$ defines a smooth form. At such a point, we compute for arbitrary $e_{i}$ in our given basis,
$\mathcal{L}_{e_{i}}\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu=\delta_{e_{1}}\mathcal{L}_{e_{i}}\delta_{e_{2}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu$$\delta_{[e_{i},e_{1}]}\delta_{e_{2}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu$.
Now if $G$ is abelian, we can continue in a similar manner and find that
$\mathcal{L}_{e_{i}}\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu=\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\mathcal{L}_{e_{i}}\nu=0$
since $\nu$ is assumed invariant. If instead $G$ is simple, we may assume without loss of generality that $e_{1},\dots,e_{r}$ is a Chevalley basis of the Lie algebra of $K$. We proceed as before and find that $\mathcal{L}_{e_{i}}\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu$ can be written as a sum of terms that are either given by $\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\mathcal{L}_{e_{i}}\nu$, or are on the form $\delta_{e_{1}}\delta_{e_{2}}\cdots\delta_{[e_{i},e_{j}]}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu$, or $\delta_{e_{1}}\delta_{e_{2}}\dots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{J[e_{i},e_{j}]}\cdots\delta_{Je_{r}}\nu$, for some $j$. In the first case, the term is zero because $\nu$ is GG–invariant. For the second and third cases, recall that in a Chevalley basis, the Lie bracket of two basis elements are either proportional to one of the basis elements or vanishes. Thus terms of this type are non-vanishing iff $e_{i},e_{j}]$ happens to be proportional to $v_{j}$. This in turn happens iff $v_{i}=H$ where $H$ is an element of the Cartan subalgebra and $e_{j}=E_{\alpha}$ where $E_{\alpha}$ is the basis element corresponding to the root of $\alpha$. In this case
$e_{i},e_{j}]=[H,E_{\alpha}]=\alpha_{H}E_{\alpha}$
where $\alpha_{H}$ is the component of $\alpha$ along the basis vector in the root lattice corresponding to $H$. But since the roots come in pairs $\alpha$ and $\alpha$, two such terms will cancel so that in the end $\mathcal{L}_{e_{i}}\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu=0$. In a similar manner $\mathcal{L}_{Je_{i}}\delta_{e_{1}}\cdots\delta_{e_{r}}\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu=0$ for arbitrary $e_{i}$ in case $G$ is simple or abelian, and thus $\nu^{cc}$ is $G$-invariant in these cases. The general case of reductive $G$ is easily reduced to the abelian or simple cases. ∎
It follows that $\nu^{\mathrm{cc}}$ induces a measure $\alpha$ on the complex quotient $X_{\text{ps }}/G.$ By the Kempf–Ness theorem [64] there is a natural homeomorphism
[…]
Consider the embedding defined by inclusion:
$j:\,\{m=0\}\hookrightarrow X$
Lemma 5.17.
The form $F^{*}\alpha$ coincides with the form $\beta$ on $\{m=0\}/K$ induced by the KK–invariant form $j^{*}\nu^{\mathrm{cc}}$ on $\{m=0\}.$
Proof.
Consider the following commutative diagram (by the Kempf–Ness theorem the map $F$ is a diffeomorphism in the sense of orbifolds);
(5.16) $\begin{array}[]{rlcrl}&&j\\ &X_{\text{s}}&\leftarrow&\left\{m=0\right\}\\ \pi&\downarrow&&\downarrow&p\\ &X_{\text{s}}/G&\leftarrow&\left\{m=0\right\}/K\\ &&F\end{array}$
By the commutativity, $j^{*}\pi^{*}\alpha=p^{*}F^{*}\alpha.$ Since $\pi^{*}\alpha=\nu^{\mathrm{cc}}$ this means that $j^{*}\nu^{\mathrm{cc}}=p^{*}F^{*}\alpha$ which equivalently means that $j^{*}\nu^{\mathrm{cc}}$ descends to $F^{*}\alpha$ on $\left\{m=0\right\}/K,$ as desired. ∎
Now consider the contracted form $\nu^{\mathrm{c}}\coloneqq\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu,$ which, by restriction, induces a form on $\left\{m=0\right\}.$
Proposition 5.18.
The following identity holds
$\int_{X_{\text{s}}/G}\nu^{\mathrm{cc}}=C\int_{\left\{m=0\right\}}\nu^{\mathrm{c}},$
where $C$ is the constant $\int_{K}dk,$ where $dk$ denotes the invariant measure on $K$ satisfying $dk(e_{1},\dots,e_{m})=1,$ where we have identified $e_{i}$ with invariant vector fields on $K.$
Proof.
By the previous lemma, we need to show that
$\int_{\left\{m=0\right\}/K}\beta=C\int\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu.$
To this end it is enough to show that the projection $p$ in 5.16 satisfies
$p_{*}\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu=C\beta,$
i.e. that
$p^{*}(p_{*}\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu)=C\delta_{e_{1}}\cdots\delta_{e_{m}}\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu,$
We will use that $p$ defines a principal KK–bundle over $\left\{m=0\right\}/K$. Consider thus a general principle KK–bundle
$p\colon M\rightarrow B$
and let $\eta$ be a KK–invariant top form on $M.$ It will be enough to show that
$p^{*}(p_{*}\eta)=C\delta_{e_{1}}\cdots\delta_{e_{m}}\eta.$
The idea of the proof is to exploit that, since $\eta$ is KK–invariant, integration along the fibers of $p$ simply amounts to contracting with $\delta_{e_{1}}\cdots\delta_{e_{m}}.$ Since the problem is local over the base $B$ it is enough to consider the case when $M=B\times K.$ Since $\eta$ is KK–invariant we can express
$\eta=p^{*}\beta\wedge dk,$
where $\beta$ is a form on $B.$ Hence,
$p_{*}\eta=p^{*}\beta\wedge\int_{K}dk=C\beta,\,\,\,\,C\coloneqq\int_{K}dk$
giving and $p^{*}p_{*}\eta=p^{*}\beta=C\delta_{e_{1}}\cdots\delta_{e_{m}}\eta,$ as desired. ∎
Denote by $h_{i}$ the Hamiltonian functions corresponding to $e_{i}$:
$dh_{i}=\delta_{e_{i}}\omega.$
Recall that $\nu_{m}$ denote the Gelfand–Leray type measure, defined in Section 5.1.1.
Lemma 5.19.
The following identity of forms on $\{m=0\}$ holds
$f\nu_{m}=\left.(\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu)\right|_{\{m=0\}}$
where $f$ is the function on $\{m=0\}$ defined by
$f\coloneqq\omega(e_{1},Je_{1})\cdots\omega(e_{r},Je_{r})$
Proof.
By the defining property of $\nu_{m}$ and Leibniz rule for contractions we have
$\delta_{Je_{1}}\cdots\delta_{Je_{r}}\nu=\delta_{Je_{1}}\cdots\delta_{Je_{r}}(\nu_{m}\wedge dh_{1}\wedge\cdots\wedge dh_{r})=\nu_{m}dh_{1}(Je_{1})\cdots dh_{r}(Je_{r})+\cdots$
where the dots indicates a sum of forms that contain some factor of $dh_{i}$ for some $i.$ In particular, such forms vanish when restricted to $\{m=0\}.$
Finally, Prop 5.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
5.5. Examples where the quantized Futaki character vanishes and reductivity holds
We next provide examples from [29] where the quantized Futaki character vanishes (and thus $\nu^{(N)}$ is $\mathcal{G}$–invariant, by Lemma 5.9) and $\mathcal{G}$ is reductive, without assuming that $X$ admits a Kähler-Einstein metric (in the toric case):
Proposition 5.20.
[29] On a homogenous Fano manifold $X$ the group $\mathcal{G}$ is reductive and the quantized Futaki character $\text{Fut}_{(X,\Delta),k}$ on $\mathfrak{g}$ vanishes for any $k.$ Moreover, for an $n$-dimensional toric log Fano variety $(X,\Delta)$, $\text{Fut}_{(X,\Delta),k}$ vanishes on $\mathcal{G}$ iff the quantized barycenter $\boldsymbol{b}_{k}$ of the moment polytope $P$ of $(X,-K_{(X,\Delta)})$ vanishes:
$\boldsymbol{b}_{k}\coloneqq\frac{1}{kN_{k}}\sum_{\boldsymbol{p}_{i}\in kP\cap\mathbb{Z}^{n}}\boldsymbol{p}_{i}=0$
As consequence, $\mathcal{G}$ is then reductive, if $\Delta$ is assumed to have simple normal crossings.
For example, $\boldsymbol{b}_{k}=0$ for all $k$ when $X$ is $\mathbb{P}^{m}\times\mathbb{P}^{l}$ (since $X$ is homogenous) and when $X$ is $\mathbb{P}^{2}$ blown up in three non-collinear points (as follows from [65, Thm 6.1]). In general, given a toric Fano manifold $X,$ $\boldsymbol{b}_{k}=0$ for all $k$ iff $\boldsymbol{b}_{k}=0$ for at least $n+1$ distinct values of $k,$ as a shown in shown in [65]; see also the survey [62].
6. Asymptotics of the $\mathcal{K}$-reduced partition function on log Fano curves
In this section we prove Theorem 1.6. To this end it will be convenient to replace the basis of $H^{0}(X,-kK_{(X,\Delta)}),$ appearing in the definition of $\nu^{(N)},$ by the basis of monomials $1,z,z^{2},\dots$ induced by standard identification of $kK_{(X,\Delta)}$ with $kV\mathcal{O}(1)\rightarrow\mathbb{P}^{1}.$ This amounts to adding an appropriate constant to the corresponding Mabuchi functional $\mathcal{M}$ of $(X,\Delta)$ to get a new functional that we denote by $\mathcal{M}_{\mathbb{Z}}$ (the subscript indicates the arithmetic interpretation of this setup, discussed in Section 7). The inf of $\mathcal{M}_{\mathbb{Z}}$ was explicitly computed in [3, Thm 6.1], when $\Delta$ is any divisor on $\mathbb{P}^{1}$ with support $\{0,1,\infty\},$ assuming that $(X,\Delta)$ is K-stable (in particular, $\mathcal{G}$ is then trivial). In a nutshell, we will reduce the proof of Theorem 1.6, where $\mathcal{G}$ is non-trivial to the computation in [3, Thm 6.1], by showing that breaking the $\mathcal{G}$-symmetry corresponds, effectively, to perturbing the given K-polystable log Fano curve $(X,\Delta)$ to a K-stable log Fano curve $(X,\Delta_{N})$.
Lemma 6.1.
On any log Fano curve $(X,\Delta)$ with vanishing Futaki character, $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$
Proof.
We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11). Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 5.20, which implies the lemma. ∎
6.1. The case when when $\Delta=0.$
Recall that in the the case of $(\mathbb{P}^{1},0)$ we have $\mathcal{G}=\mathrm{PGL}(2,\mathbb{C})$ and $\mathcal{K}=\mathrm{SO}(3).$
6.1.1. Expressing the contracted partition function $Z_{N}^{\mathrm{c}}$ as a complex Selberg integral $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
By Prop 5.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 5.4) may be expressed as
$Z_{N}^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}(\nu^{(N)})^{\mathrm{cc}}$
for some positive constant $C.$ Next, given a positive integer $N$ and $\boldsymbol{w}=(w_{1},w_{2},w_{3})\in]0,1[^{3}$ set, when $N\geq 2$
(6.1) $\mathcal{Z}_{N}^{\boldsymbol{w}}=\int_{\mathbb{C}^{N}}\rho_{N}^{\boldsymbol{w}}\,\,\,,\rho_{N}^{\boldsymbol{w}}\coloneqq\left(\prod_{i\neq j\leq N}\left|z_{i}-z_{j}\right|\right)^{\frac{2-w_{1}-w_{2}-w_{3}}{N-1}}\prod_{i\leq N}\left|z_{i}\right|^{-2w_{1}}\left|z_{i}-1\right|^{-2w_{2}}\left(\frac{i}{2}dz\wedge d\bar{z}\right)^{\otimes N},$
which is a complex Selberg integral [60] (for $N\coloneqq 1$ the first factor in the right hand side above is set to be $1).$ We will show that, for a particular choice of $e_{1},e_{2}$ and $e_{3}$ in the Lie algebra of $\mathcal{G}$ (that has an arithmetic interpretation, as discussed in Section 7):
(6.2) $\int_{(X^{N})_{\text{s}}/\mathcal{G}}\left(\nu^{(N)}\right)^{\mathrm{cc}}=\mathcal{Z}_{N-3}^{\boldsymbol{w}},\,\,\,\text{$\boldsymbol{w}\coloneqq k^{-1}(1,1,1)$, }$
where $k^{-1}=2/(N-1).$ When computing the integral in the lhs above we may restrict $\nu^{(N)}$ to the Zariski open subset $\mathbb{C}^{N}$ of $X^{N},$ where $\mathbb{C}\coloneqq\mathbb{P}^{1}-\{\infty\}.$ Given $(x_{1},\dots,x_{N})\in X^{N}$ denote by $g\in\mathcal{G}$ the unique automorphism of $X(=\mathbb{P}^{1})$ mapping $(x_{1,}x_{2},x_{3})$ to $(0,1,\infty).$ Set
$y_{i}\coloneqq g^{*}z_{i}\coloneqq\frac{(z_{1}-z_{2})(z_{i}-z_{3})}{(z_{1}-z_{3})(z_{i}-z_{2})},\,\,\,i=4,5,\dots,N,$
viewed as a function of $(z_{1},\dots,z_{N})$ on a Zariski open subset of $\mathbb{C}^{N}.$ A direct computation reveals that $y_{i}$ is $\mathcal{G}$–invariant (using that $\mathcal{G}$ is generated by translations, scalings and inversion). Moreover, as shown below,
(6.3) $\nu^{(N)}=\eta\wedge\rho_{N-3}^{\boldsymbol{w}}(y_{4},\dots,y_{N}),\,\,\,\eta\coloneqq\frac{(\frac{i}{2})^{3}dz_{1}\wedge d\overline{z}_{1}\wedge\cdots\wedge dz_{3}\wedge d\overline{z}_{3}}{\left|(z_{1}-z_{2})(z_{1}-z_{3})(z_{2}-z_{3})\right|^{2}}$
Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 6.1, applied to $N=3).$ Thus formula 6.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 5.9. Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant. In fact, taking $e_{1},e_{2}$ and $e_{3}$ as the standard generators of the Lie algebra $\mathfrak{sl}(2,\mathbb{C})$ corresponding to the holomorphic vector fields $\partial/\partial z,$ $z\partial/\partial z$ and $z^{2}\partial/\partial z$ on $\mathbb{C}$ a direct computation gives
$\eta^{\mathrm{cc}}=1.$
Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 5.4.1. Thus, accepting formula 6.3, we conclude that
$(\nu^{(N)})^{\mathrm{cc}}=\rho_{N-3}^{\boldsymbol{w}}(y_{4},\dots,y_{N}),$
which implies the desired formula 6.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
Finally, as next described, the deferred proof of formula 6.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 6.2). First note that since $V=2$ and $N=2k+1$ [18, Lemma 4.3] yields
$\nu^{(N)}=\left(\prod_{i\neq j}\left|z_{i}-z_{j}\right|\right)^{-\frac{1}{k}}\left(\frac{i}{2}dz\wedge d\overline{z}\right)^{\otimes N}.$
Next, a direct computation gives
$\prod_{4\leq i<j\leq N}|y_{i}-y_{j}|^{-2/k}$ $=\bigg|\frac{(z_{1}-z_{2})(z_{2}-z_{3})}{z_{1}-z_{3}}\bigg|^{-2\binom{N-3}{2}/k}\prod_{4\leq i<j\leq N}\bigg|\frac{z_{i}-z_{j}}{(z_{i}-z_{2})(z_{j}-z_{2})}\bigg|^{-2/k}=$
$=\bigg|\frac{(z_{1}-z_{2})(z_{2}-z_{3})}{z_{1}-z_{3}}\bigg|^{-2\binom{N-3}{2}/k}\prod_{4\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\prod_{\ell=4}^{N}|z_{\ell}-z_{2}|^{2(N-4)/k}$
$\prod_{\ell=4}^{N}|y_{\ell}|^{-2/k}$ $=\bigg|\frac{z_{1}-z_{2}}{z_{1}-z_{3}}\bigg|^{-2(N-3)/k}\prod_{\ell=4}^{N}\bigg|\frac{z_{\ell}-z_{3}}{z_{\ell}-z_{2}}\bigg|^{-2/k}$
$\prod_{\ell=4}^{N}|1-y_{\ell}|^{-2/k}$ $=\bigg|\frac{z_{2}-z_{3}}{z_{1}-z_{3}}\bigg|^{-2(N-3)/k}\prod_{\ell=4}^{N}\bigg|\frac{z_{\ell}-z_{1}}{z_{\ell}-z_{2}}\bigg|^{-2/k}$
Moreover, for $4\leq i\leq N$,
$\prod_{i=4}^{N}\bigg|\frac{\partial y_{i}}{\partial z_{i}}\bigg|^{2}=\prod_{i=4}^{N}\bigg|\frac{\partial}{\partial z_{i}}\bigg(\frac{(z_{1}-z_{2})(z_{i}-z_{3})}{(z_{1}-z_{3})(z_{i}-z_{2})}\bigg)\bigg|^{2}=\bigg|\frac{(z_{1}-z_{2})(z_{2}-z_{3})}{z_{1}-z_{3}}\bigg|^{2(N-3)}\prod_{i=4}^{N}|z_{i}-z_{2}|^{-4}.$
Combining the above, we find that
$\eta_{3}\wedge\rho_{N-3}$ $=\eta_{3}\wedge|z_{1}-z_{2}|^{2M_{1}}|z_{2}-z_{3}|^{2M_{1}}|z_{1}-z_{3}|^{2M_{2}}\times\prod_{4\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\times$
$\qquad\qquad\qquad\times\bigwedge_{\ell=4}^{N}|(z_{\ell}-z_{1})(z_{\ell}-z_{3})|^{-2/k}|z_{\ell}-z_{2}|^{2M_{3}}\mathrm{d}z_{\ell}\wedge\mathrm{d}\bar{z}_{\ell},$
where $M_{1}=-\binom{N-3}{2}\frac{1}{k}-\frac{N-3}{k}+N-3,$ $M_{2}=\binom{N-3}{2}\frac{1}{k}+\frac{2(N-3)}{k}-(N-3)$ and $M_{3}=\frac{2}{k}-2+\frac{N-4}{k}.$ Using that k=(N1)/2k=(N-1)/2, gives $M_{1}=M_{2}=1-1/k$ and $M_{3}=-1/k$. Thus,
$\eta_{3}\wedge\rho_{N-3}$ $=\eta_{3}\wedge\prod_{1\leq i<j\leq 3}|z_{i}-z_{j}|^{2(1-1/k)}\prod_{4\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\prod_{\begin{subarray}{c}1\leq i\leq 3\\ 4\leq j\leq N\end{subarray}}|z_{i}-z_{j}|^{-2/k}\bigwedge_{\ell=4}^{N}\mathrm{d}z_{\ell}\wedge\mathrm{d}\bar{z}_{\ell}$
$=\prod_{1\leq i<j\leq 3}|z_{i}-z_{j}|^{-2/k}\prod_{4\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\prod_{\begin{subarray}{c}1\leq i\leq 3\\ 4\leq j\leq N\end{subarray}}|z_{i}-z_{j}|^{-2/k}\bigwedge_{\ell=1}^{N}\mathrm{d}z_{\ell}\wedge\mathrm{d}\bar{z}_{\ell}$
$=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\bigwedge_{\ell=1}^{N}\mathrm{d}z_{\ell}\wedge\mathrm{d}\bar{z}_{\ell}=\nu^{(N)}(z),$
which concludes the proof of formula 6.3.
6.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 5.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
(6.4) $N^{-1}C\leq\omega_{N}(e_{3},Je_{3})\leq C\,\,\text{on $\{\left\langle\boldsymbol{m}_{N},e_{3}\right\rangle=0\}\subset X^{N}$.}$
The upper bound follows directly from the compactness of $X.$ To prove the lower one, set $p\coloneqq\left\langle\boldsymbol{m},e_{3}\right\rangle,$ which is the Hamiltonian for the $S^{1}$–action generated by $e_{3}.$ Using that h(X)=[1,1]h(X)=[-1,1] (see the beginning of the proof of Lemma 3.11) and that the $S^{1}$–action has precisely two zeros $x_{0}$ and $x_{1}$ – and that these are simple – yields $\omega_{N}(e_{3},Je_{3})\geq C(1-p^{2})$ (since $p(x_{i})=\pm 1).$ Hence,
$\omega_{N}(e_{3},Je_{3})/C\leq\frac{1}{N}\sum_{i=1}^{N}(p_{i}^{2}-1),\,\,\,\,\sum_{i=1}^{N}p_{i}=0,\,|p_{i}|\leq 1.$
Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above. Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 6.4. Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
$\Delta_{N/2}\coloneqq\left\{(t_{1},\dots,t_{N})\in[0,1]^{N}\colon\sum_{i=1}^{N}t_{i}=N/2\right\}$
We need to show that its extremal points satisfy $t_{i}\in\{0,1/2,1\}$ when $N$ is assumed odd. To do this we will argue by contradiction. Assume that $v=(v_{1},\dots,v_{N})$ is a vertex of $\Delta_{N/2}$ with not all coordinates in $\{0,1/2,1\}$. Claim: at least two coordinates are not in $\{0,1/2,1\}$. To see this, we will argue by contradiction. Without loss of generality, $v_{1}\notin\{0,1/2,1\}$ and $v_{i}\in\{0,1/2,1\}\forall i\neq 1$. Then $v_{1}=\sum_{i=1}^{N}v_{i}-\sum_{i=2}^{N}v_{i}=N/2-\sum_{i=2}^{N}v_{i}$ showing that $v_{1}$ is a multiple of $1/2$, proving the claim.
Letting $v_{1},v_{2}\notin\{0,1/2,1\}$, consider for $a>0$ the line segment $L\colon[-a,a]\rightarrow\mathbb{R}^{N},\ t\mapsto v+te_{1}-te_{2}$. For $a<\min_{i=1,2}(\mathrm{dist}(v_{i},\{0,1\}))$, the line segment L[a,a]L[-a,a] lies in $\Delta_{N/2}$, while $L(0)=v$, proving that $v$ is not extremal.
6.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 6.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial). By [3, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
(6.5) $\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathcal{Z}_{N}^{\boldsymbol{w}}=\inf\mathcal{M}_{\mathbb{Z}}$
where $\mathcal{M}_{\mathbb{Z}}$ denotes the arithmetic Mabuchi functional attached to $(X,\Delta).$ The challenge in the present case is that we are considering a joint limit where $\boldsymbol{w}\in]0,1[^{3}$ depends on $N$ and converges towards $(0,0,0).$ Luckily, as next explained, this situation can be handled by exploiting the explicit formula for the integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ first established in [53] when $\Delta$ is supported on at most three points. By replacing N3N-3 with $N$ it will be equivalent to show that
(6.6) $\boldsymbol{w}_{N}\coloneqq\frac{2}{N+2}(1,1,1)\implies-\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathcal{Z}_{N}^{\boldsymbol{w}_{N}}=\inf\mathcal{M}_{\mathbb{Z}},$
where $\mathcal{M}_{\mathbb{Z}}$ denote the arithmetic Mabuchi functional attached to $(\mathbb{P}^{1},0).$ As pointed out in [3, Section 6.2], it follows from [53, Formula B.9] (and [60, formula 3.1] and [92, Section 2.3]) that for any weight vector $\boldsymbol{w}\in]0,1[^{3}$ such that $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ is a log Fano curve that is K-stable, i.e. such that
(6.7) $\left|V\right|\coloneqq 2-w_{1}-w_{2}-w_{3}>0,\,\,\,w_{i}<\sum_{i\neq j}w_{j},\,\,\,\forall i,$
the following explicit formula holds when $N\geq 2$:
(6.8) $\mathcal{Z}_{N}^{\boldsymbol{w}}/\pi^{N}=\left(\frac{N!}{-l\left(-\frac{1}{2}\frac{\left|V\right|}{N-1}\right)^{N}}\right)\prod_{j=0}^{N-1}\frac{-l\left(-(j+1)\frac{\left|V\right|/2}{N-1}\right)}{l\left(w_{1}+j\frac{\left|V\right|/2}{N-1}\right)l\left(w_{2}+j\frac{V/2}{N-1}\right)l\left(w_{3}+j\frac{\left|V\right|/2}{N-1}\right)},$
where $l(x)\coloneqq\Gamma(x)/\Gamma(1-x),$ in terms of the classical Gamma function $\Gamma(x).$ When $N=1$, $\mathcal{Z}_{N}^{\boldsymbol{w}}$ is a complex version of the classical beta integral:
(6.9) $\mathcal{Z}_{1}^{\boldsymbol{w}}\coloneqq\int_{\mathbb{C}}\left|z\right|^{-2w_{1}}\left|z-1\right|^{-2w_{2}}\frac{i}{2}dz\wedge d\bar{z}=\pi\frac{\Gamma(1-w_{1})}{\Gamma(w_{1})}\frac{\Gamma(1-w_{1})}{\Gamma(w_{1})}\frac{\Gamma(-1+w_{1}+w_{2})}{\Gamma(2-w_{1}-w_{2})}$
(as follows from formula 6.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
$\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right)$
Combining this convergence with the convergence 6.5 gives the following explicit formula
(6.10) $\inf\mathcal{M}_{\mathbb{Z}}=\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right),$
which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 6.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 6.6. Note that
$w_{N,i}=O(N^{-1}),\,\,\,\left|V_{N}\right|=\left|V\right|-O(N^{-1}),\,\,V=2,$
where the error terms are positive. Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 6.8:
$\frac{1}{N}\left(\log N!-N\log\left(-l\left(-\frac{1}{2}\frac{\left|V_{N}\right|}{N-1}\right)\right)\right)=\log\frac{|V|}{2}-1+\frac{1}{2}\frac{\log N}{N}+O\left(N^{-1}\right).$
Next, let us show that
(6.11) $\frac{1}{N}\sum_{j=0}^{N-1}\log\Gamma\left(w_{N}+j\frac{\left|V_{N}\right|/2}{N-1}\right)=\int_{0}^{1}\log\Gamma\left(w_{N}+t\frac{V_{N}}{2}\right)dt+O(N^{-1}\log N)$
To this end first observe that, for any decreasing function $g$ on the real line and $N$ fixed,
$\int_{N^{-1}}^{1}gdt+g(1)\frac{1}{N}\geq\sum_{j=1}^{N}g\left(\frac{j}{N}\right)\frac{1}{N}\geq\int_{0}^{1}gdt,$
using that $g(j/N)\geq g(t)$ for $t\in[j/N,(j+1)/N]$ and $g(t)\geq g(j/N)$ for $t\in[(j-1)/N,j/N].$ We apply these inequalities to
$g(t)\coloneqq-\log\Gamma(w+\epsilon_{N}+t\frac{\left|V\right|}{2}),\,\,\,\,\epsilon_{N}=O(N^{-1}),$
where $\epsilon_{N}$ is a sequence of numbers such that $g(j/N)=-\log\Gamma(w_{N}+j\frac{\left|V_{N}\right|/2}{N-1});$ the function $g(t)$ is decreasing, since $x\mapsto\Gamma(x)$ is increasing. Next, note that
$\int_{N^{-1}}^{1}\log\Gamma\left(w_{N}+t\frac{V_{N}}{2}\right)dt=\int_{0}^{1}\log\Gamma\left(w+t\frac{V}{2}\right)dt+O(N^{-1}\log N),\,\,\,\,g(1)=O(\log N)$
and
$\int_{0}^{1}\log\Gamma\left(w_{N}+t\frac{V_{N}}{2}\right)dt=\int_{0}^{1}\log\Gamma\left(w+t\frac{V}{2}\right)dt+O(N^{-1}\log N),$
using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 6.11. The nominator in the second factor in formula 6.8 is handled in essentially the same manner, giving
$N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}_{N}}=\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right)+O(N^{-1}\log N)$
for $w_{i}=0$ and $V=2.$ The proof of formula 6.6 is thus concluded by invoking formula 6.10.
6.2. Proof of Theorem 1.6 when $\Delta$ is nontrivial
In this case $\mathcal{G}=\mathbb{C}^{*}$ and $\mathcal{K}=S^{1}.$ In the light of the previous proof, concerning the case when $\Delta$ is trivial, it will be enough to show that
(6.12) $\int_{(X^{N})_{\text{s}}/\mathcal{G}}\left(\nu^{(N)}\right)^{\mathrm{cc}}=\mathcal{Z}_{N-1}^{\boldsymbol{w}},\,\,\,\boldsymbol{w}\coloneqq\left(w,\frac{1-w}{N-1},w\right),$
when the generator $e$ of $\mathcal{G}$ is taken to be the one defined by the vector field $z\frac{\partial}{\partial z}$ on $\mathbb{P}^{1}$ (the rest of the proof then proceeds essentially as before). To this end, consider $\mathbb{C}^{N}\subset X^{N}$ with standard coordinates $z_{i}$ and set
$y_{i}\coloneqq z_{i}/z_{1},\,\,i=2,3,\dots,N$
viewed as a function of $(z_{1},\dots,z_{N})$ on $\mathbb{C}^{N}.$ These functions are manifestly $\mathcal{G}$–invariant and it will thus be enough to show that
(6.13) $\nu^{(N)}=\eta\wedge\rho_{N-1}^{\boldsymbol{w}}(y_{2},\dots,y_{N-2})\left(\frac{i}{2}dy\wedge d\overline{y}\right)^{\otimes N-1},\,\,\,\eta=\frac{i}{2}\frac{dz_{1}}{z_{1}}\wedge\overline{(\frac{dz_{1}}{z_{1}})}$
To prove this formula recall that, by [18, Lemma 4.3], we may express
$\nu^{(N)}=\left(\prod_{i\neq j}\left|z_{i}-z_{j}\right|\right)^{-\frac{V}{N-1}}\prod_{i\leq N}\left|z_{i}\right|^{-2w}\left(\frac{i}{2}dz\wedge d\overline{z}\right)^{\otimes N},\,\,\,V\coloneqq 2-2w.$
It thus follows directly from the homogeneity of the density of $\nu^{(N)}$ with respect the standard $\mathbb{C}^{*}$–action on $\mathbb{C}^{N}$ that
$\nu^{(N)}=|z_{1}|^{2r}\left(\left(\prod_{i\neq j,i,j\geq 2}\left|y_{i}-y_{j}\right|\right)^{\frac{2-2w}{N-1}}\prod_{2\leq i\leq N}\left|y_{i}\right|^{-2w}\prod_{2\leq i\leq N}\left|1-y_{i}\right|^{-2w}\frac{i}{2}\frac{dz_{1}}{z_{1}}\wedge\overline{(\frac{dz_{1}}{z_{1}})}\wedge\left(\frac{i}{2}dy\wedge d\overline{y}\right)^{\otimes N-1}\right)$
for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e. $\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 6.13.
7. Relations to arithmetic geometry
In this section we delve into the connections to arithmetic geometry, alluded to in the beginning of Section 6.
[…]
$\mathcal{M}_{\mathbb{Z}}(\phi)=-2\hat{h}_{\text{$\phi$}}(-\mathcal{K}_{(\mathcal{X},\mathcal{D})})+D\left(\left.\frac{(dd^{c}\phi)^{n}}{V}\right|\mu_{\phi}\right),\,\,\,\,\hat{h}_{\text{$\phi$}}(\mathcal{L})=(\mathcal{L},\phi)^{n+1}/(L^{n}(n+1)),$
[…]
$\inf\mathcal{M}_{\mathbb{Z}}=-2\hat{h}_{\text{can}}(\mathcal{X},\mathcal{D}),$
where $h_{\text{can}}(\mathcal{X},\mathcal{D})$ is the normalized canonical height of $(\mathcal{X},\mathcal{D}),$ introduced in [2], i.e. the normalized height of $(\mathcal{X},\mathcal{D})$ with respect to any KE metric $\phi_{\text{KE}}$ on $K_{(X,\Delta)},$ normalized so that $\int_{X}\mu_{\phi_{\text{KE }}}=1.$
As next explained, the arithmetic structure also induces a sequence of canonical measures $\nu_{\mathbb{Z}}^{(N)}$ on $X^{N}.$ Consider a positive integer $k$ such that $k\mathcal{K}_{(\mathcal{X},\mathcal{D})}$ is defined as a line bundle. Then $H^{0}(\mathcal{X},-k\mathcal{K}_{(\mathcal{X},\mathcal{D})})$ is a free $\mathbb{Z}$-module whose complexification is $H^{0}(X,-kK_{(X,\Delta)})$:
$H^{0}(X,-kK_{(X,\Delta)})=H^{0}\left(\mathcal{X},-k\mathcal{K}_{(\mathcal{X},\mathcal{D})}\right)\otimes_{\mathbb{Z}}\mathbb{C}.$
The measure $\nu_{\mathbb{Z}}^{(N)}$ is defined by replacing the orthonormal basis in $H^{0}(X,-kK_{(X,\Delta)}),$ appearing in the definition of $\nu^{(N)}$ (Section 5.1), with any basis of the $\mathbb{Z}$-module $H^{0}(\mathcal{X},-k\mathcal{K}_{(\mathcal{X},\mathcal{D})})$ (the independence of the choice of basis follows from the fact that the absolute value of the corresponding change of basis matrices equals one).
Now fix a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and a $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ with positive curvature. The canonical measure $\nu_{\mathbb{Z}}^{(N)}$ on $X^{N}$ then induces a Gelfand–Leray type measure $(\nu_{\mathbb{Z}}^{(N)})_{\boldsymbol{m}_{N}}$ (defined in Section 5.1).
Conjecture 7.1.
Let $(\mathcal{X},\mathcal{D})$ be an arithmetic log pair over $\mathbb{Z}$ such that $\mathcal{K}_{(\mathcal{X},\mathcal{D})}$ is a relatively ample $\mathbb{Q}$-line bundle over $\mathcal{X}$ and assume that its complexification $(X,\Delta)$ admits a KE metric (or equivalently, that $(X,\Delta)$ is K-polystable). Then
(7.1) $\lim_{N\rightarrow\infty}N^{-1}\log\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}=\inf\mathcal{M}_{\mathbb{Z}}.$
In fact, this conjecture is equivalent to the convergence in Conjecture 1.5, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [3, Thm 5.3]). One drawback of Conjecture 1.5 – from the arithmetic point of view – is that left hand side in formula 7.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).). This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 5.5) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
(7.2) $\mathfrak{g}=\mathfrak{g_{\mathbb{Z}}}\otimes_{\mathbb{Z}}\mathbb{C}.$
Indeed, then $\nu^{(N)}$ is $\mathcal{G}$–invariant (by Lemma 5.9) and taking any basis $e_{1},\dots,e_{r}$ for the $\mathbb{Z}$–module $\mathfrak{g_{\mathbb{Z}}}$ induces a contracted measure $(\nu^{(N)})^{\mathrm{cc}}$ (independent of the choice of bases), which descends to the stable locus $(X^{N})_{\text{s}}/\mathcal{G}$ of the GIT quotient (by Lemma 5.16). We thus obtain a sequence of arithmetic invariants
$\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})\coloneqq\int_{(X^{N})_{\text{s}}/\mathcal{G}}(\nu^{(N)})^{\mathrm{cc}}$
realized as periods. The following conjecture generalizes [18, Conjecture 5.2] (see also [3, Section 5.2]) to the case when $\mathcal{G}$ is nontrivial:
Conjecture 7.2.
Let $(\mathcal{X},\mathcal{D})$ be an arithmetic log pair over $\mathbb{Z}$ such that $\mathcal{K}_{(\mathcal{X},\mathcal{D})}$ is a relatively ample $\mathbb{Q}$–line bundle over $\mathcal{X}$ and assume that $(X,\Delta)$ is K-polystable and that the quantized Futaki character vanishes for $k$ sufficiently large. Then
$\lim_{N\rightarrow\infty}N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})=\inf\mathcal{M}_{\mathbb{Z}}$
By Prop 5.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 5.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.6).
7.1. Examples
The condition 7.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e. $\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [66, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [81, Section 2]. By Prop 5.20, the assumptions of Conjecture 7.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish. It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial. One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases. For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 6.1 and formula 6.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
$Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1})=\pi^{N-3}\frac{(N-3)!}{-l(-\frac{1}{N-4})^{N}}\prod_{j=0}^{N-4}\frac{-l(-(j+1)\frac{1}{N-4})}{l(\frac{2}{N-1}+j\frac{1}{N-4})^{3}},\,\,\,\,\,l(x)\coloneqq\frac{\Gamma(x)}{\Gamma(1-x)}$
(and a similar formula holds when $N=4,$ by formula 6.9). Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 6.12 and formula 6.8 gives, when $N\geq 3,$
$Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{w})=\pi^{N-1}\left(\frac{(N-1)!}{-l(-\frac{1-w}{N-2})^{N}}\right)\prod_{j=0}^{N-2}\frac{-l(-(j+1)\frac{1-w}{N-2})}{l(w+j\frac{1-w}{N-2})^{2}l(\frac{1-w}{N-1}+j\frac{1-w}{N-2})}.$
Note that the left hand side in this equality extends to a holomorphic function of $w\in\mathbb{C}$ in the strip $\Re w<1\}$ and the right hand side defines a meromorphic function on all of $\mathbb{C},$ which is thus the unique meromorphic extension of $Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{w}).$
8. Appendix
8.1. The free energy of a toric log Fano curve
Let $(X,\Delta)$ be a toric log Fano curve and denote by $P$ the corresponding moment polytope, which is an interval in $\mathbb{R}.$ Following [16, 22], we will identify the metric $\phi$ on $K_{(X,\Delta)}$ with a continuous convex function on $\mathbb{R}.$ More precisely, on $\mathbb{C}^{*}\hookrightarrow X$, let $x=\log(|z|^{2})$. Trivializing $K_{(X,\Delta)}$ with $\frac{\mathrm{d}z}{z}\otimes s_{U}e_{\Delta}$ over $U=\mathbb{C}^{*}$ we let $\phi(x)\coloneqq\phi_{U}(z)$ in the chosen trivialization over $U=\mathbb{C}^{*}$ (abusing notation slightly). Then
(8.1) $\text{(Log)}_{*}dd^{c}\phi=\partial^{2}\phi dx,$
where $\partial^{2}\phi$ denotes the second order distributional derivatives of $\phi.$ Denote by $v$ the convex function on $P$ defined by the Legendre transform of $\phi,$ $v(y)=\sup_{x\in\mathbb{R}}xy-\phi(x).$
Proposition 8.1.
Denote by $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}}$ the space of all $S^{1}$–invariant psh metrics $\phi$ on $K_{(X,\Delta)}$ such that $\phi(x)$ is in $C^{1,1}(\mathbb{R}).$ Given $\phi_{0}\in\text{PSH(}-K_{(X,\Delta})_{\mathrm{b}}$ which is assumed $S^{1}$–invariant, the corresponding free energy $F_{\gamma}(\mu)$ satisfies the following formula, when $\mu=dd^{c}\phi/V$ for $\phi\in\mathcal{H}(-K_{(X,\Delta)})_{1,1}^{S^{1}}$:
$VF_{-\gamma}(\mu)=\gamma\mathcal{L}(v)+(1-\gamma)\int_{P}\phi_{0}(v^{\prime})dy-\int_{P}v^{\prime\prime}dy,\,\,\,\mathcal{L}(v)\coloneqq\frac{V}{2}\int_{\partial P}v-\int_{P}vdy$
where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v.$
Proof.
When $\gamma=1$ it is shown in the proof of [16, Lemma 5, step 1] that
$V\mathcal{M}(\phi)=\mathcal{L}(v)+\int_{\mathbb{R}}(\log\phi^{\prime\prime})\phi^{\prime\prime}dx$
(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.11, this means that
$F_{-1}(\mu)=\gamma E(\mu)+D(\mu|\mu_{\phi})=\mathcal{L}(v)-\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx+\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$
Note that, by formula 8.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
(8.2) $E(\mu)=-\mathcal{L}(v)+\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx,$
giving
$F_{-\gamma}(\mu)=\gamma\mathcal{L}(v)+(1-\gamma)\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx+\int_{\mathbb{R}}(\log\phi^{\prime\prime})\phi^{\prime\prime}dx.$
As shown in proof of [16, Lemma 5, step 2], the last term above equals $\int_{P}v^{\prime\prime}dy$ (as shown by making the change of variables $y=\phi^{\prime}(x)).$ Likewise,
$\int_{\mathbb{R}}\phi_{0}(x)\phi^{\prime\prime}(x)dx=\int_{P}\phi_{0}(v^{\prime})dy,$
using that $x=v^{\prime}(y).$
Recall [22] that then $P$ is symmetric, P=[V/2,V/2],P=[-V/2,V/2], any $\phi_{0}\in\text{PSH (}-K_{(X,\Delta})_{\mathrm{b}}$ which is $S^{1}$–invariant satisfies
(8.3) $\phi_{0}(x)=\frac{V}{2}|x|+O(1),$
where $O(1)$ denotes a bounded term.
8.2. The twisted Ding functional of a toric log Fano manifold
The functional $\mathcal{D}_{\gamma}$ may in the toric case be expressed as follows, using the notation above, with $u\coloneqq\phi-\phi_{0}$:
$\mathcal{D}_{\gamma}(u)=-\frac{\mathcal{E}(u)}{V}--\frac{1}{\gamma}\log\int e^{-\left(\gamma\phi+(1-\gamma)\phi_{0}\right)}dx,\,\,\,\,-\frac{\mathcal{E}(u)}{V}=\int_{P}v\frac{dy}{V}$
(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [22, Prop 2.9] in the second equality). For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.1 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.35 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
$\phi_{t}(x)=0,\,\text{when }|x|\leq t,\,\,\,\,\phi_{t}(x)=|x-t|,\,\text{when }t\leq|x|$
A simple direct calculation gives $\mathcal{D}_{\gamma}(u_{t})=(1/2-\frac{\gamma-1}{\gamma})t+O(1),$ tending to -$\infty$ as $t\rightarrow\infty$ if $\gamma>2.$ This shows that $u_{t}$ is a destabilizing curve for $\mathcal{D}_{\gamma}$ (which provides an alternative direct proof of the destabilizing result in the proof of Cor 4.1). A similar argument applies when $w\in]0,1/2],$ by multiplying the formula for $\phi_{t}(x)$ above by $V/2.$

### v2

@@
 The definition involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides with an analytic reduced stability threshold encoding the coercivity of the K-energy functional modulo automorphisms.
-These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves and derive a strengthened logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, which implies Aubin’s refinement of the Moser–Trudinger inequality.
+These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves.
+By imposing a moment constraint, we derive a strengthened form of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, that implies the sharp form of Aubin’s refinement of the Moser–Trudinger inequality.
 In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
@@
 The quest for canonical metrics on complex manifolds is a classical theme in geometry, tracing back to the uniformization theorem.
-According to the seminal Yau–Tian–Donaldson conjecture compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e.
+According to the seminal Yau–Tian–Donaldson conjecture a compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e.
 a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold - this is now a theorem [42, 90].
@@
 The measure $\nu^{(N)}$ on $X^{N}$ may be expressed as
-(1.2) $\nu^{(N)}\coloneqq i^{N^{2}}\big(\mathrm{det}S^{(N)}\big)^{-1/k}\wedge\overline{\left(\det S^{(N)}\right)^{-1/k},}$
-where $\det S^{(N)}$ is the holomorphic section of the line bundle $kK_{X^{N}}\rightarrow X^{N},$ defined as the Slater determinant
+(1.2) $\nu^{(N)}\coloneqq\frac{i^{(Nn)^{2}}}{2^{Nn}}\big(\mathrm{det}S^{(N)}\big)^{-1/k}\wedge\overline{\left(\det S^{(N)}\right)^{-1/k},}$
+where $n$ is the dimension of $X$ and $\det S^{(N)}$ is the holomorphic section of the line bundle $kK_{X^{N}}\rightarrow X^{N},$ defined as the Slater determinant
 (1.3) $\left(\det S^{(N)}\right)(x_{1},x_{2},\dots,x_{N})\coloneqq\det\left(s_{i}(x_{j})\right),$
@@
 We will denote by $\Delta_{N}$ the divisor on $X^{N}$ induced by $\Delta$ that is invariant under permutation of the factors of $X^{N}.$ Thus $(X^{N},\Delta_{N})$ is a log Fano manifold.
-Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [13], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined
+Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [13], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined by
 $\gamma^{(N)}(X,\Delta)\coloneqq\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),\,\,\,\,\gamma(X,\Delta)\coloneqq\liminf_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta),$
@@
 They are based on conjectural Large Deviation Principle that we shall come back to below, linking the invariant $\gamma(X,\Delta)\mathcal{{}^{G}}$ to the reduced analytic stability threshold defined by
-(1.7) $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon(\mathcal{I}-\mathcal{J})^{\mathcal{G}}-C_{\epsilon}\right\},$
-where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right)$ – whose minimizers are Kähler–Einstein metrics for $(X,\Delta)$ – and the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry (see Remark 3.1):
-(1.8) $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}=\text{inf}_{g\in\mathcal{G}}g^{*}(\mathcal{I}-\mathcal{J}).$
-In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the Dirichlet norm.
+(1.7) $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon(\mathcal{I}-\mathcal{J})^{\mathcal{G}}+C_{\epsilon}\right\},$
+where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right),$ whose minimizers are Kähler–Einstein metrics for $(X,\Delta),$ and $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry
+(1.8) $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}=\text{inf}_{g\in\mathcal{G}}g^{*}(\mathcal{I}-\mathcal{J})$
+(the definition of $\mathcal{I}$ and $\mathcal{J}$ are recalled in Remark 3.1).
+In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the square of the Dirichlet norm of the Kähler potential.
 On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [89, 23, 48]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [95, 48].
@@
 The validity of this conjecture would imply the second equivalence in Conjecture 1.1.
-Indeed, by [48], $\delta^{\mathrm{A}}(X,L)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a Kähler–Einstein metric, which, by the solution of the YTD conjecture is equivalent to K-polystability (see Theorem 3.7).
+Indeed, as shown in [48], confirming a conjecture of Tian, $\delta^{\mathrm{A}}(X,L)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a Kähler–Einstein metric, which, by the solution of the YTD conjecture is equivalent to K-polystability (see Theorem 3.7).
 Our first main result concerns the case of log Fano curves $(X,\Delta),$ i.e.
@@
 $\gamma(X,0)^{\mathcal{G}}=\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2,\,\text{and\,}\,\gamma(X,\Delta_{w})^{\mathcal{G}}=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}\,\text{if}\,w\neq 0.$
-The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ [37, 38] (see Section 4).
-We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1).
-This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant implicitly appearing in [63, 77], obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its centralizer) – see Section 4.2.
+The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ in [37, 38, 9] under a moment constraint, i.e.
+a strengthening of the following inequality for all probability measures $\mu$ on $S^{2}:$
+(1.9) $E(\mu)\leq D(\mu|\mu_{0}),$
+where $D(\mu|\mu_{0})$ denotes the entropy of $\mu,$ relative to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2},$ and $E(\mu)$ denotes the logarithmic energy of $\mu$ (explicitely expressed in formula 4.3).
+In fact, combining our formula for $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}$ with the uniqueness result for the Liouville equation on $S^{2}$ established in [72] yields:
+Theorem 1.4.
+Let $X=\mathbb{P}^{1},$ identified with the unit-sphere $S^{2}$ in $\mathbb{R}^{3},$ and denote by $\mathcal{M}$ the Mabuchi functional on the space of all Kähler forms $\omega$ in $c_{1}(X),$ normalized so that $\mathcal{M}(\omega)=0$ when $\omega$ is proportional to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2}.$ For any $\omega$ in $c_{1}(X),$
+$\mathcal{M}(\omega)\geq(\mathcal{I}-\mathcal{J})^{\mathcal{G}}(\omega)$
+with equality iff $\omega$ is proportional to $\mu_{0}$ modulo the action of $\mathcal{G}$ (i.e.
+modulo conformal diffeomorphisms of $S^{2}).$ Moreover, the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ may be replaced by $\mathcal{I}-\mathcal{J}$ if the Kähler potential of $\omega$ satisfies Tian’s orthogonality condition 3.31.
+More generally, for any probability measure $\mu$ on $S^{2},$ satisfying the moment constraint $\int_{S^{2}\subset\mathbb{R}^{3}}\boldsymbol{x}\mu=0,$
+(1.10) $E(\mu)\leq\frac{1}{2}D(\mu|\mu_{0})$
+with equality iff $\mu=\mu_{0}.$
+The inequality for $\mathcal{M}$ in formula 1.7, for some (non-effective) positive constant $\epsilon$ and $C_{\epsilon}\in\mathbb{R}$ was first established in [48].
+Since $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ (by Theorem 1.3) the inequality for $\mathcal{M}$ fails on $\mathbb{P}^{1}$ when $\epsilon>1$ (for any $C_{\epsilon})$ and thus the multiplicative constants in Theorem 1.4 are sharp.
+More generally, the formula for $\delta^{\mathrm{A}}(X,w)^{\mathcal{G}}$ in Theorem 1.3 yields a quasi-sharp generalization of Aubins’ improvement of the Moser-Trudinger inequality to $w\neq 0.$ We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1).
+This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its center), appearing implicitly in [48, 63, 77] – see Section 4.2.
 Beyond complex curves, Gibbs stability has turned out to be notoriously difficult to establish, and at present no examples are known apart from the case of exceptional Fano orbifolds, where it holds automatically.
@@
 Given a positive integer $N$ such that 0 is a regular value of $\boldsymbol{m}_{N}$ we denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 6.1 for the general definition.
-The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 5.11), i.e.
+The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 6.11), i.e.
 that
@@
 The original $\mathcal{G}$–symmetry has thus been reduced to a $\mathcal{K}$–symmetry.
-Accordingly, the probability measure 1.11 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
+Accordingly, the probability measure 1.13 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
 We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e.
-$\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.11.
+$\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.13.
 Before stating our two main conjectures we note that in contrast to the $\mathcal{K}$–reduced probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)},$ the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ depends on a choice of bases in $H^{0}(X,-kK_{(X,\Delta)}).$ As a consequence, so does its total mass $Z_{\boldsymbol{m}_{N}}.$ In what follows we will take the basis to be orthonormal with respect to the $L^{2}$–norm on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the fixed metric on $K_{(X,\Delta)}.$ The corresponding constant $Z_{\boldsymbol{m}_{N}}$ shall then be called the $\mathcal{K}$-reduced partition function, in view of its statistical mechanical interpretation described in Section 6.1.
@@
 where $F$ is the free energy functional, introduced in [10], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$.
-We show that the validity of such a LDP would imply both the asymptotics 1.12 and the convergence 1.13, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$.
+We show that the validity of such a LDP would imply both the asymptotics 1.14 and the convergence 1.15, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$.
 Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
@@
 Theorem 1.7.
-The convergence 1.12 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1),$ when $N$ is taken as the sequence of odd positive integers. More precisely,
+The convergence 1.14 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1$), when $N$ is taken as the sequence of odd positive integers. More precisely,
 $N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}+O\left(\frac{\log N}{N}\right).$
-As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.13) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 5.2).
+As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.15) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 6.2).
 1.4. Outline of the proof of Theorem 1.7 and connections to arithmetic geometry
@@
 Moreover, the corresponding integral over $\left\{\boldsymbol{m}_{N}=0\right\}$ is shown to be comparable to the following integral over the GIT quotient of $X^{N}$ under $\mathcal{G}$:
-(1.15) $\int_{(X^{N})_{\mathrm{ss}}/\mathcal{G}}\nu^{(N),\mathrm{cc}},$
-where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.15 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60].
+(1.17) $\int_{(X^{N})_{\mathrm{ss}}/\mathcal{G}}\nu_{\mathbf{m}_{N}}^{(N),\mathrm{cc}},$
+where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.17 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60].
 Such integrals can be computed explicitly in terms of Gamma values, as first shown in [53] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.7, by exploiting some connections to K-stable arithmetic log Fano curves in [3].
-It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.15 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 7, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86].
+It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.17 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 8, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86].
 The proof of Theorem 1.7 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.
@@
 The starting point of the proof of Theorem 3.13 is a thermodynamical formula for $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}.$ This formula is established on any log Fano manifold in Section 3.3, where the structure of the space $\mathcal{P}(X)_{0}$ of all probability measures on $X$ with vanishing moments is studied and new analytic characterizations of log Fano manifolds admitting Kähler–Einstein metrics are given in Theorem 3.7.
-  • In Section 4 applications of Theorem 3.13 to logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
+  • In Section 4 applications of Theorem 3.13 to constrained logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
+  • In Section 5 we establish a sharp version of the constrained logarithmic Hardy–Littlewood–Sobolev inequalities on the two-sphere, by combining Theorem 3.13 with the uniqueness result for the Liouville equation on $S^{2}$ established in [72].
   • In Section 6 we study the $\mathcal{K}$–reduced probability measures $\mu_{\boldsymbol{m}_{N}}^{(N)}$ on $X^{N}$ and the corresponding partition function $Z_{\boldsymbol{m}_{N}}$ appearing in Conjecture 1.6.
@@
 $\begin{array}[]{rcl}X^{[N]}&\xrightarrow{f}&\mathrm{{Bl}}_{W}(X^{N})\\ \pi_{FM}\searrow&&\swarrow\pi_{W}\\ &X^{N}\end{array}$
-where $\pi_{W}\colon\mathrm{Bl}_{W}(X^{N})\rightarrow X^{N}$ is the blow-up morphism, whose exceptional divisor we denote by $E_{\pi_{W}}.$ It will thus be enough to show that $f_{*}^{-1}(E_{\pi_{W}})=E_{W^{\prime}},$ where $f_{*}^{-1}(E_{\pi_{W}})$ denotes the strict transform of $E_{\pi_{W}}$ under $f.$ To this end let us first observe that there exists $W^{\prime}$ in $\mathcal{W}$ such that
+where $\pi_{W}\colon\mathrm{Bl}_{W}(X^{N})\rightarrow X^{N}$ is the blow-up morphism, whose exceptional divisor we denote by $E_{\pi_{W}}.$ It will thus be enough to show that $f_{*}^{-1}(E_{\pi_{W}})=E_{W},$ where $f_{*}^{-1}(E_{\pi_{W}})$ denotes the strict transform of $E_{\pi_{W}}$ under $f.$ To this end let us first observe that there exists $W^{\prime}$ in $\mathcal{W}$ such that
 (2.9) $f_{*}^{-1}(E_{\pi_{W}})=E_{W^{\prime}},\,\,\,W^{\prime}\subset W$
@@
 while
-$\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}$ $=\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{2k(1-w)}{\mathrm{codim}W-1}$
-$=\frac{2(N-1)(1-w)}{V(\lfloor N/2\rfloor-1)}=\frac{N-1}{(\lfloor N/2\rfloor-1)}.$
+$\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{\mathrm{codim}W-w\mathrm{codim}W}{\frac{1}{k}\mathrm{codim}W(\mathrm{codim}W-1)/2}=\min_{W\in\mathcal{W}_{p}^{\mathrm{ss}}}\frac{2k(1-w)}{\mathrm{codim}W-1}=\frac{2(N-1)(1-w)}{(\lfloor N/2\rfloor-1)V}=\frac{N-1}{(\lfloor N/2\rfloor-1)}$
 The cases $N=2,3$ is similar except that $\mathcal{W}_{p}^{\mathrm{ss}}$ is empty.
@@
 Furthermore, we fix a $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ with positive curvature form $\omega_{0}(=dd^{c}\phi_{0})$ (using additive notation for metrics on line bundles, recalled below).
-The fixed metric $\phi_{0}$ on $K_{(X,\Delta)}$ corresponds to a volume form on $X$ with divisorial singularities that we denote by $\mu_{0}$ (see Section 3.2.1).
+The fixed metric $\phi_{0}$ on $K_{(X,\Delta)}$ corresponds to a volume form on $X$ with divisorial singularities that we denote by $\mu_{0},$ which (up to scaling the metric on $K_{(X,\Delta)}$)) we may assume is a probability measure; see Section 3.2.1.
 We will denote by $\mathcal{P}(X)$ the space of all probability measures on $X,$ endowed with its weak topology and by $\mathcal{P}(X)_{0}$ the subspace of all measures $\mu$ with vanishing moment:
@@
 […]
+3.2.3. Kähler metrics vs Kähler potentials and metrics on line bundles
+Given a Kähler form $\omega_{0}$ on a compact complex manifold, consider the space of all Kähler potentials:
+$\mathcal{H}(X,\omega_{0})\coloneqq\left\{u\in C^{\infty}(X)\colon\,\,\omega_{u}\coloneqq\omega_{0}+dd^{c}u>0\right\}$
+Using the map $u\mapsto\omega_{u}$ we will identify the space of all Kähler metrics $\omega$ in the cohomology class $\omega_{0}]\in H^{2}(X,\mathbb{R})$ with $\mathcal{H}(X,\omega_{0})/\mathbb{R},$ under the additive action of $\mathbb{R}$ on $\mathcal{H}(X,\omega_{0})$ (acting as $u\rightarrow u+C).$ Moreover, given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$, expressing
+(3.3) $u=\phi-\phi_{0}$
+for a metric $\phi$ on $L,$ we may identify the space $\mathcal{H}(X,\omega_{0})$ with the space with $\mathcal{H}(L)$ of smooth metrics on $L$ with positive curvature form.
 3.2.4. Pluripotential theory and energy
 In this section we recall the notion of pluricomplex energy of a probability measure $\mu$ introduced in [23], but using the notation in [12, 13].
-Given a Kähler form $\omega_{0}$ on a compact complex manifold consider the space of all Kähler potentials:
-$\mathcal{H}(X,\omega_{0})\coloneqq\left\{u\in C^{\infty}(X)\colon\,\,\omega_{u}\coloneqq\omega_{0}+dd^{c}u>0\right\}$
-and denote by $\text{PSH}(X,\omega)$ its closure in $L^{1}(X).$ Given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$, expressing $u=\phi-\phi_{0}$ for a metric $\phi$ on $L,$ we may identify the spaces $\mathcal{H}(X,\omega_{0})$ and $\text{PSH}(X,\omega_{0})$ with $\mathcal{H}(L)$ and $\text{PSH}(\mathcal{L}),$ respectively.
-We also set
-$\mathcal{H}_{1,1}(X,\omega)\coloneqq\left\{u\colon\omega_{u}\in L_{\text{loc}}^{\infty},\,\,\omega_{u}\geq 0\right\}.$
-Thus, $\mathcal{H}(X,\omega_{0})\subset\mathcal{H}_{1,1}(X,\omega)\subset\text{PSH}(X,\omega_{0})$.
+Given a Kähler form $\omega_{0}$ on a compact complex manifold $X,$ denote by $\text{PSH}(X,\omega)$ the closure of $\mathcal{H}(X,\omega_{0})$ in $L^{1}(X),$ called the space of all $\omega_{0}$-plurisubharmonic functions.
+Given a line bundle $L$ over $X$ and a metric $\phi_{0}$ on $L$ with curvature form $\omega_{0}$ we can identify $\text{PSH}(X,\omega_{0})$ with $\mathcal{H}(L)$ and $\text{PSH}(\mathcal{L}),$ using formula 3.3.
 Given a log Fano manifold $(X,\Delta)$ we will take $L=-K_{(X,\Delta)}$ and $\phi_{0}$ as in the setup introduced above.
@@
 As a consequence, the restriction of $\mathcal{E}$ to $\mathcal{H}(X,\omega_{0})$ is increasing and thus admits a canonical extension to $\text{PSH}(X,\omega_{0}),$ namely the smallest extension which is upper semi-continuous with respect to the $L^{1}$–topology [23].
-The subspace of $\text{PSH}(X,\omega_{0})$ consisting of all $u$ such that $\mathcal{E}(u)>$-$\infty$ is denoted by $\mathcal{E}^{1}(X,\omega_{0}).$ Formula3.4 still applies on $\mathcal{E}^{1}(X,\omega_{0}),$ when $\omega_{u}^{n}$ is the Monge–Ampère measure defined in terms of non-pluripolar products [23].
+The subspace of $\text{PSH}(X,\omega_{0})$ consisting of all $u$ such that $\mathcal{E}(u)>$-$\infty$ is denoted by $\mathcal{E}^{1}(X,\omega_{0}).$ Formula3.5 still applies on $\mathcal{E}^{1}(X,\omega_{0}),$ when $\omega_{u}^{n}$ is the Monge–Ampère measure defined in terms of non-pluripolar products [23].
 The pluricomplex energy $E(\mu)$ of a probability measure $\mu$ on $X$ (with respect to $\omega_{0})$ is defined by following lower semi-continuous (lsc) convex function on $\mathcal{P}(X)$:
@@
 (3.10) $dE_{|\mu}=-u_{\mu}\,\,\,\left(=\phi_{0}-\phi_{\mu}\right),$
-For example, any measure $\mu$ on $X$ with divisorial singularities (or more, generally with an $L^{p}$–density for some $p>1)$ has finite energy, $E(\mu)<\infty.$ Thus the properties 3.8 and 3.9 determine $E$ on $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ for any $p>1$ (and the functional $E$ on $\mathcal{P}(X)$ can then be characterized as the greatest lower semi-continuous extension of $E$ from $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ to $\mathcal{P}(X)$).
+For example, any measure $\mu$ on $X$ with divisorial singularities (or more, generally with an $L^{p}$–density for some $p>1)$ has finite energy, $E(\mu)<\infty.$ Thus the properties 3.9 and 3.10 determine $E$ on $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ for any $p>1$ (and the functional $E$ on $\mathcal{P}(X)$ can then be characterized as the greatest lower semi-continuous extension of $E$ from $\mathcal{P}(X)\cap L_{\text{loc}}^{p}$ to $\mathcal{P}(X)$).
 Remark 3.1.
@@
 When $n=1$ integration by parts gives
-$V\mathcal{E}(u)=-\frac{1}{2}\int_{X}du\wedge d^{c}u+\int_{X}u\omega_{0},\,\,2VE(\mu)=-\int_{X}u_{\mu}\mu=\int_{X}du_{\mu}\wedge d^{c}u_{\mu},$
-if $E(\mu)<\infty$ or equivalently $u_{\mu}$ is in the space $H^{1}(X)\subset L^{2}(X)$ of all $u$ such that $du\in L^{2}(X).$ Moreover, in this case the sup in formula 3.5 can be taken over all of $C^{0}(X)$ or over $H^{1}(X).$ Indeed, by classical Hilbert space theory the latter sup is attained by $u_{\mu}\in H^{1}(X)$ solving the (Laplace) equation 3.6 and hence $u_{\mu}\in\text{PSH}(X,\omega_{0})$.
+(3.12) $\mathcal{E}(u)=-\frac{1}{2V}\int_{X}du\wedge d^{c}u+\int_{X}u\frac{\omega_{0}}{V},\,\,E(\mu)=\frac{1}{2V}\int_{X}du_{\mu}\wedge d^{c}u_{\mu}=-\frac{1}{2}\int_{X}u_{\mu}\mu$
+if $E(\mu)<\infty$ or equivalently $u_{\mu}$ is in the Sobolev space $W^{2,1}(X)$ of all $u\in L^{2}(X)$ such that $du\in L^{2}(X).$ Moreover, when $n=1$ the sup in formula 3.6 can be taken over all of $C^{0}(X)$ or over $W^{2,1}(X).$ Indeed, by classical Hilbert space theory the latter sup is attained by $u_{\mu}\in W^{2,1}(X)$ solving the (Laplace) equation 3.7 and hence $u_{\mu}\in\text{PSH}(X,\omega_{0})$.
 3.2.5. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
-The space $\mathcal{H}(X,\omega_{0})$ can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to $\mathcal{H}_{1,1}(X,\omega_{0})$ in [40].
-We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
-3.2.5. The Mabuchi functional $\mathcal{M}$ versus the free energy functional $F$
+The space $\mathcal{H}(X,\omega_{0})$ of Kähler potentials can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to the space
+$\mathcal{H}_{1,1}(X,\omega)\coloneqq\left\{u\colon\omega_{u}\in L_{\text{loc}}^{\infty},\,\,\omega_{u}\geq 0\right\}.$
+in [40]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
+3.2.6. The Mabuchi functional $\mathcal{M}$ versus the free energy functional $F_{\beta}$
 Given an ample line bundle $L$ the Mabuchi functional $\mathcal{M}$ was originally defined when $\Delta$ is trivial, by declaring that its differential on $\mathcal{H}(L)$ at $\phi$ is identified with $(R_{\phi}-\overline{R})(dd^{c}\phi)^{n},$ where $R_{\phi}$ denotes the scalar curvature of the Kähler metric $dd^{c}\phi$ and $\overline{R}$ its average [80].
@@
 (3.15) $F(\beta)\coloneqq\inf_{\mu\in\mathcal{P}(X)}F_{\beta}(\mu),\,\,\,\,F(\beta)_{0}\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}F_{\beta}(\mu)$
-We recall that the function $\mu\mapsto D(\mu|\nu)$ is lower semicontinuous and convex on $\mathcal{P}(X)$ as follows, for example, from the following well-known Legendre duality formula:
+We recall that the function $(\mu,\nu)\mapsto D(\mu|\nu)$ is non-negative and lower semicontinuous (lsc) on $\mathcal{P}(X)^{2}$ and $\mu\mapsto D(\mu|\nu)$ is convex on $\mathcal{P}(X).$ These properties follow directly from the following well-known Legendre duality formula [50]:
 (3.16) $D(\mu|\nu)=\sup_{u\in C^{0}(Y)}\left\{\int u\mu-\log\int e^{u}\nu\right\}.$
@@
 (3.20) $\frac{d}{dt}\mathcal{M}\left(\exp(t\mathrm{Re}V^{1,0})^{*}\psi\right)_{|t=0}=\mathrm{Re}\text{$\mathrm{Fut}$}_{(X,\Delta)}(V^{1,0})$
-More generally, 3.17 holds if $\psi$ and $\phi$ have finite energy and 3.18 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy.
+More generally, 3.19 holds if $\psi$ and $\phi$ have finite energy and 3.20 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy.
 In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
 Proof.
-This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.17 first note that, for any given any $g\in\mathcal{G},$
+This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.19 first note that, for any given any $g\in\mathcal{G},$
 $\mathcal{E}\left(g^{*}\psi\right)-\mathcal{E}\left(g^{*}\phi\right)=\mathcal{E}(\psi)-\mathcal{E}(\phi).$
-This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.4 defined by formula 3.4, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.17.
+This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.5 defined by formula 3.5, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.19.
 Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$.
 Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e.
-iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.11, we can decompose
+iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.13, we can decompose
 $\mathcal{M}(\psi)=-\mathcal{E}(\psi)+D(dd^{c}\psi)^{n}/V|\mu_{\psi}),$
@@
 $\mathcal{M}(g_{t}^{*}\psi)=-\mathcal{E}(g_{t}^{*}\psi)+D(dd^{c}\psi)^{n}/V|\mu_{\psi}).$
-Differentiating with respect to $t$ and using 3.3 thus proves 3.18.
-More generally, the proof of 3.17 only used that $\phi$ and $\psi$ have finite energy.
+Differentiating with respect to $t$ and using 3.4 thus proves 3.20.
+More generally, the proof of 3.19 only used that $\phi$ and $\psi$ have finite energy.
 Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [23] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [24], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy.
@@
 $\boldsymbol{m}\colon\,\,X\rightarrow\mathbb{R}^{r},\,\,\,x\mapsto(h_{e_{1}}(x),\dots,h_{e_{r}}(x)),$
-where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.16).
+where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.18).
 We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
@@
 and since $i\partial\bar{\partial}\phi^{\tau}$ is smooth and strictly positive on $X-\Delta$ it follows that $W=0$ on $X-\Delta.$ Thus, by continuity, $W=0$ on $X,$ which is a contradiction.
-Hence, 3.21 holds.
-Finally, the inequality 3.22 is shown in essentially the same way as in the proof of [21, Lemma 9.1].
+Hence, 3.23 holds.
+Finally, the inequality 3.24 is shown in essentially the same way as in the proof of [21, Lemma 9.1].
 More precisely, in [21, Lemma 9.1] it is assumed that the corresponding metric $\phi^{\tau}$ is independent of the real part of $\tau$ (in our notation) and that $\phi^{\tau}$ satisfies the complex geodesic equation, that is the local function $\Phi(z,\tau)\coloneqq\phi^{\tau}(z)$ satisfies
@@
 where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ First recall that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0},$ (see Section 3.2.8).
-To prove formula 3.24 note that, by 3.9, denoting by $\phi$ the potential of $\mu,$
+To prove formula 3.26 note that, by 3.10, denoting by $\phi$ the potential of $\mu,$
 $\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int(\phi-\phi_{0})\frac{d\mu_{t}}{dt}|{}_{t=0}=-\int(\phi-\phi_{0})\mathcal{L}\mu,$
@@
 $\frac{d}{dt}E(\mu_{t})|_{t=0}=\int\mathcal{L}(\phi-\phi_{0})\mu=\int\mathcal{L}\phi\frac{(dd^{c}\phi)^{n}}{V}-\int h_{W}\mu,$
-which proves 3.24, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
+which proves 3.26, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
 Finally, take $\mu\in\mathcal{P}(X)_{0}.$ If $E(\mu)=\infty,$ then $E(g^{*}\mu)=\infty$ for all $g\in\mathcal{G}.$ Hence, we may assume that $E(\mu)<\infty,$ which means that $E(g^{*}\mu)<\infty$ for all $g\in\mathcal{G}.$ Consider, as before, the function $g\mapsto E(g^{*}\mu)$ on $\mathcal{G}/\mathcal{K}.$ It follows from the inequality $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0,$ shown above when $\nu$ is a volume form, and a standard approximation argument, that $E(g^{*}\mu)$ is convex on the symmetric space $\mathcal{G}/\mathcal{K}.$ Moreover, as before, $\frac{d}{dt}E(\mu_{t})_{|t=0}=0$ using the previous lemma, that applies since $E(\mu)<\infty$ (which equivalently means that $\mathcal{E}(\phi_{\mu})>-\infty).$ Since any critical point of a convex function $f$ on the symmetric space $\mathcal{G}/\mathcal{K}$ minimizes $f$ this shows that $\mu$ minimizes $E$ on $\mathcal{G}\mu,$ as desired.
@@
 First note that $F_{-\gamma}$ is $\mathcal{K}$–invariant, since $\phi_{0}$ is $\mathcal{K}$–invariant.
-Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.25.
-Thus, to prove the rest of the identities 3.25, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively.
+Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.27.
+Thus, to prove the rest of the identities 3.27, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively.
 ∎
@@
 To prove the first statement of the proposition it is enough to show that when changing the metric $\phi_{0}$ the corresponding pluricomplex energy $E_{\phi_{0}}$ only changes by an additive bounded term.
-To this end note that, by formula 3.9, given $\phi_{0}$ and $\psi_{0}$ in $\mathcal{H}(L)$ the differential $d(E_{\phi_{0}}-E_{\phi_{1}})|_{\mu}$ is represented by the function $u\coloneqq\phi_{0}-\phi_{1},$ for any $\mu$ in $\mathcal{P}(X)$ satisfying $E(\mu)<\infty.$ Hence, there exists a constant $C$ such that
+To this end note that, by formula 3.10, given $\phi_{0}$ and $\psi_{0}$ in $\mathcal{H}(L)$ the differential $d(E_{\phi_{0}}-E_{\phi_{1}})|_{\mu}$ is represented by the function $u\coloneqq\phi_{0}-\phi_{1},$ for any $\mu$ in $\mathcal{P}(X)$ satisfying $E(\mu)<\infty.$ Hence, there exists a constant $C$ such that
 $E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)=C+\left\langle u,\mu\right\rangle,$
 showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired.
-Finally, the second statement follows from combining formulae 3.11 and 3.10 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.6).
+Finally, the second statement follows from combining formulae 3.13 and 3.11 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.7).
 ∎
@@
 (3.30) $A^{-1}E\left(\frac{\omega_{u}^{n}}{V}\right)-B\leq d_{1}(0,u)\leq AE\left(\frac{\omega_{u}^{n}}{V}\right)+B$
-for some positive constants $A$ and $B.$ Indeed, by 3.10, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.28 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
+for some positive constants $A$ and $B.$ Indeed, by 3.11, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.30 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
 Next, note that, $3\implies 2$ follows from the previous proposition.
-To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.27 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.28 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
+To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.29 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.30 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
 $F(\mu)\geq\epsilon\inf_{\mathcal{G}}E(\mu)-C\text{\,\,on $\mathcal{P}(X)\cap\{E<\infty\}.$ }$
@@
 Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [21, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e.
-on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.26.
+on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.28.
 In the present case, where $\phi_{0}$ is assumed $\mathcal{K}$–invariant, it thus follows from Prop 3.4 that $\mu_{-1}\in\mathcal{P}(X)_{0}$ and conversely, any KE volume form in $\mathcal{P}(X)_{0}$ coincides with $\mu_{-1}.$ It also follows from the aforementioned uniqueness that $\mu_{\text{KE}}$ is $\mathcal{K}$–invariant.
@@
 (3.32) $\mathcal{M}(u)\geq\epsilon(\mathcal{I}-\mathcal{J})(u)-C$
-for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.29.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48].
+for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.31.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48].
 It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [23, Section 2]).
-The inequality 3.30 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.30 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.29 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
+The inequality 3.32 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.32 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.31 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
 Lemma 3.9.
@@
 where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, there exists a unique $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions.
-When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.29 for $v.$
+When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.31 for $v.$
 Proof.
@@
 (3.33) $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=\int_{X}\left.\frac{dv_{t}}{dt}\right|_{t=0}\omega_{0}^{n}=\int_{X}\left\langle dv,JW\right\rangle\omega_{0}^{n}=\int_{X}\nabla v\cdot JW\omega_{0}^{n}$
-using, in the first equality, that the Futaki character vanishes (formula 3.15).
+using, in the first equality, that the Futaki character vanishes (formula 3.17).
 Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$.
@@
 Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [8, Thm 2.2].
-Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.29, when $\omega_{0}$ is Kähler–Einstein.
+Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.31, when $\omega_{0}$ is Kähler–Einstein.
 ∎
 In the case when $n=1,$ so that $X=\mathbb{P}^{1}$, the functionals $\mathcal{I}-\mathcal{J}$ and $\mathcal{J}$ coincide and thus the moment condition appearing in Theorem 3.7 then coincides with condition $(ii)$ in the previous lemma.
-The formula $\delta^{A}(\mathbb{P}^{1},0)^{\mathcal{G}}=2$ that we shall establish in the next section thus yields a quasi-sharp inequality of the form 3.30 on $\mathbb{P}^{1},$ saying that the inequality holds for any $\epsilon\in]0,1[$ - but not for $\epsilon>1$ - when $u$ satisfies Tian’s orthogonality condition 3.29.
 3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
@@
 Consider the case when the fixed $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ is induced by the Fubini–Study metric on $\mathcal{O}(1)\rightarrow\mathbb{P}^{1}$ and thus $\mathrm{SU}(2)$–invariant.
-Then, identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\omega_{0}=V\omega$ for the unique $\mathrm{SO}(3)$–invariant probability measure $\omega$ and
-$E(\mu)=-\frac{V}{2}\int_{X\times X}\log\left(|x-y|^{2}\right)\mu(x)\otimes\mu(y)+C$
-for some constant $C$ (by Example 3.2, using the explicit Green function for $(X,\omega_{0})$ ).
-Moreover, when $\Delta=0$ the measure $\mu_{0}$ is proportional to $\omega_{0}$ and $\boldsymbol{m}=\boldsymbol{x}.$ When $w\neq 0,$ taking $x_{0}$ and $x_{1}$ as the south and north pole on the unit sphere, $\mu_{0}$ is proportional to $e^{-|\boldsymbol{x}-\boldsymbol{x}_{0}|^{2}-|\boldsymbol{x}-\boldsymbol{x}_{1}|^{2}}\omega_{0}$ and $\boldsymbol{m}=2x_{3}/V.$
+Then, identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\omega_{0}=V\omega$ for the unique $\mathrm{SO}(3)$–invariant probability measure $\omega$ on $S^{2}.$ Moreover, when $\Delta=0$ the measure $\mu_{0}$ is proportional to $\omega_{0}$ and $\boldsymbol{m}=\boldsymbol{x}.$ When $w\neq 0,$ taking $x_{0}$ and $x_{1}$ as the south and north pole on the unit sphere, $\mu_{0}$ is proportional to $|\boldsymbol{x}-\boldsymbol{x}_{0}|^{-2w}|\boldsymbol{x}-\boldsymbol{x}_{1}|^{-2w}\omega_{0}$ and $\boldsymbol{m}=2x_{3}/V.$
 Lemma 3.11.
@@
 $\alpha(X,\Delta)_{0}\coloneqq\sup\left\{\gamma\in]0,\infty[\colon\mathcal{I}_{\gamma,\mu_{0}}(\phi_{\mu})\coloneqq\int_{X}e^{-\gamma(\phi_{\mu}-\phi_{0})}\mu_{0}<\infty\,\,\forall\mu\in\mathcal{P}(X)_{0}\right\},$
-where $\phi_{\mu}$ is the potential of $\mu$ (formula 3.7).
+where $\phi_{\mu}$ is the potential of $\mu$ (formula 3.8).
 Lemma 3.12.
@@
 for $r$ sufficiently small (as follows, for example, from Green’s formula in $\mathbb{C}$ or Skoda’s integrability theorem in $\mathbb{C}^{n}$).
-Now, first consider the case $w=0$ and take $\mu\in\mathcal{P}(X)_{0}.$ By the previous lemma, $\mu(\left\{x\right\})\leq 1/2$ for any $x\in X.$ Since $V=2$ this means that $dd^{c}\phi_{\mu}(\left\{x\right\})\leq 1$ for any $x\in X.$ By 3.32, this implies that $e^{-t\phi_{\mu}}\in L_{\mathrm{loc}}^{1}$ for any $t<1.$ Hence, $\alpha(\mathbb{P}^{1},0)_{0}\geq 1.$ Finally, in the case $w\neq 0$ we may assume that $\Delta_{w}=w([0]+[\infty]).$ On a disc $D_{r}$ of radius $r$ centered at $x=0,$
+Now, first consider the case $w=0$ and take $\mu\in\mathcal{P}(X)_{0}.$ By the previous lemma, $\mu(\left\{x\right\})\leq 1/2$ for any $x\in X.$ Since $V=2$ this means that $dd^{c}\phi_{\mu}(\left\{x\right\})\leq 1$ for any $x\in X.$ By 3.34, this implies that $e^{-t\phi_{\mu}}\in L_{\mathrm{loc}}^{1}$ for any $t<1.$ Hence, $\alpha(\mathbb{P}^{1},0)_{0}\geq 1.$ Finally, in the case $w\neq 0$ we may assume that $\Delta_{w}=w([0]+[\infty]).$ On a disc $D_{r}$ of radius $r$ centered at $x=0,$
 $\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}\leq C\int_{|z|\leq r}e^{-\left(t\phi+w\log|z|^{2}\right)}idz\wedge d\bar{z}.$
@@
 $(dd^{c}\psi)(\{0\})\leq tV/2+w=t(1-w)+w\,\,\,(<1\iff t<1).$
-Hence, by 3.32, $\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ if $t<1$ and the same argument applies when 0 is replaced by $\infty.$ Finally, for $x$ different from 0 and $\infty$ we can use that $(dd^{c}\psi)(\{x\})\leq tV$ (since $\int dd^{c}\phi=V$) to deduce that $\int_{D_{r}(x)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ when t<1/V=1/(22w),t<1/V=1/(2-2w), as desired.
+Hence, by 3.34, $\int_{D_{r}(0)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ if $t<1$ and the same argument applies when 0 is replaced by $\infty.$ Finally, for $x$ different from 0 and $\infty$ we can use that $(dd^{c}\psi)(\{x\})\leq tV$ (since $\int dd^{c}\phi=V$) to deduce that $\int_{D_{r}(x)}e^{-t(\phi-\phi_{0})}\mu_{0}<\infty$ when t<1/V=1/(22w),t<1/V=1/(2-2w), as desired.
 ∎
@@
 $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\sup\left\{\gamma\in]0,\infty[\colon F(-\gamma)_{0}>-\infty\right\}=2\alpha(X,\Delta_{w})_{0}$
-where $F(-\gamma)_{0}$ is defined by formula 3.13.
+where $F(-\gamma)_{0}$ is defined by formula 3.15.
 Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
@@
 Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [10] (concerning the case $\mathcal{G}$ trivial) to the present setup.
-Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.14 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
+Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.16 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
 $D(\mu)\geq t\int_{X}u_{\mu}\mu-\log\mathcal{I}_{t,\mu_{0}}(u_{\mu})\geq t\int_{X}u_{\mu}\mu-C_{t},$
-using Lemma 3.12. Next, setting $u_{\mu}\coloneqq\phi_{\mu}-\phi_{0}$ (formula 3.7) we deduce, using $E(\mu)=-\int_{X}u_{\mu}\mu/V2$ (Example 3.2),
+using Lemma 3.12. Next, setting $u_{\mu}\coloneqq\phi_{\mu}-\phi_{0}$ (formula 3.8) we deduce, using $E(\mu)=-\int_{X}u_{\mu}\mu/V2$ (Example 3.2),
 $F_{\gamma}(\mu)\geq(-\gamma+2t)E(\mu)-C_{t},$
@@
 $F_{\gamma}(\mu)=\gamma\frac{\mathcal{L}(v)}{V}+(1-\gamma)\frac{V}{2}\int_{P}\left|v^{\prime}\right|dy-\int_{P}v^{\prime\prime}dy+O(1),\,\,\,\frac{\mathcal{L}(v)}{V}\coloneqq\int_{\partial P}\frac{v}{2}-\int_{P}v\frac{dy}{V},$
-where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 8.1, combined with 8.3).
+where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 9.1, combined with 9.3).
 In particular,
@@
 $F_{\gamma}(\mu^{t})=\gamma t\frac{V}{4}+(1-\gamma)t\frac{V}{2}+O(1)=tV\left(\gamma\left(\frac{1}{4}-\frac{1}{2}\right)+\frac{1}{2}\right)+O(1),$
-which tends to -$\infty$ if $\gamma>2,$ proving the inequality 3.34. It remains to show that when $w\in]0,1[$
+which tends to -$\infty$ if $\gamma>2,$ proving the inequality 3.36. It remains to show that when $w\in]0,1[$
 $\frac{2}{V}\geq\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}.$
@@
 (3.39) $\phi_{t}(x)=A_{-t}^{*}\phi_{0}=\phi_{0}(x+t).$
-The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 8.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
+The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 9.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
 $E(\mu^{t})=0+\frac{V}{2}t+O(1).$
@@
 Proposition 3.14.
-For the log Fano curves $(X,\Delta_{w})$ we have $\overrightarrow{\delta}_{0}(X,0)=2$ and $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\min\{\frac{1}{1-w},2\}$ when $w\neq 0.$ Moreover, when $\phi_{0}$ is induced by the Fubini–Study metric, there is a geodesic ray $u_{t}$ in $\boldsymbol{H}_{1,1}(X,\omega_{0})_{0},$ attaining the inf defining $\overrightarrow{\delta}_{0}(X,\Delta)$, with the property that the limit of $\omega_{u_{t}}/V$ in $\mathcal{P}(X)_{0}$ as $t\rightarrow\infty$ is the measure $\mu$ defined by formula 3.33.
+For the log Fano curves $(X,\Delta_{w})$ we have $\overrightarrow{\delta}_{0}(X,0)=2$ and $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\min\{\frac{1}{1-w},2\}$ when $w\neq 0.$ Moreover, when $\phi_{0}$ is induced by the Fubini–Study metric, there is a geodesic ray $u_{t}$ in $\boldsymbol{H}_{1,1}(X,\omega_{0})_{0},$ attaining the inf defining $\overrightarrow{\delta}_{0}(X,\Delta)$, with the property that the limit of $\omega_{u_{t}}/V$ in $\mathcal{P}(X)_{0}$ as $t\rightarrow\infty$ is the measure $\mu$ defined by formula 3.35.
 Proof.
@@
 The geodesic ray $u_{t}$ appearing in the previous proposition is attached to a test configuration $(\mathcal{X},\mathcal{L})$ for $(X,-K_{(X,\Delta)})$ and thus defines a psh metric on the Berkovich analytification of $(X,-K_{(X,\Delta)}),$ with respect to the trivial valuation on $\mathbb{C},$ in the sense of [33, 35].
-Moreover, the Monge–Ampère measure $\text{MA(}(\mathcal{X},\mathcal{L}))$ of $(\mathcal{X},\mathcal{L})$ on the Berkovich analytification $X^{\text{an}}$ of $X$ coincides with the measure $\mu$ defined by formula 3.33, under the standard embedding of a complex curve $X$ into $X^{\text{an}}$ (where $x\in X$ is mapped to the divisorial valuation defined by $x).$ This follows from the fact that, by construction, the test configuration $(\mathcal{X},\mathcal{L})$ is toric for some $\mathbb{C}^{*}$–action on $\mathbb{P}^{1}$.
-Thus formula 3.38 translates into the following weak convergence of measures on the Berkovich analytification of $\mathbb{P}^{1}$:
+Moreover, the Monge–Ampère measure $\text{MA(}(\mathcal{X},\mathcal{L}))$ of $(\mathcal{X},\mathcal{L})$ on the Berkovich analytification $X^{\text{an}}$ of $X$ coincides with the measure $\mu$ defined by formula 3.35, under the standard embedding of a complex curve $X$ into $X^{\text{an}}$ (where $x\in X$ is mapped to the divisorial valuation defined by $x).$ This follows from the fact that, by construction, the test configuration $(\mathcal{X},\mathcal{L})$ is toric for some $\mathbb{C}^{*}$–action on $\mathbb{P}^{1}$.
+Thus formula 3.40 translates into the following weak convergence of measures on the Berkovich analytification of $\mathbb{P}^{1}$:
 $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\delta_{v_{i}}=\text{MA(}(\mathcal{X},\mathcal{L}))$
 for the valuations $v_{i}$ on $\mathbb{P}^{1}$ defined by the points $x_{i}.$
-4. Applications to logarithmic Hardy–Littlewood–Sobolev inequalities, Moser–Trudinger inequalities and spontaneous symmetry breaking
+4. Applications to logarithmic HLS inequalities, Moser–Trudinger inequalities and spontaneous symmetry breaking
 We continue with the setup in the previous section applied to the case when $X$ is the Riemann sphere and $\Delta=\Delta_{w}$ for $w\in[0,1].$ The lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ resulting from Theorem 3.13 is equivalent to the inequality
-$E(\mu)\leq\frac{1}{\gamma}D(\mu|\mu_{0})+C_{\gamma}$
-for any $\mu\in\mathcal{P}(X).$ When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38].
+(4.1) $E(\mu)\leq\frac{1}{\gamma}D(\mu|\mu_{0})+C_{\gamma}\,\,\forall\mu\in\mathcal{P}(X):\int_{X}\boldsymbol{m}\mu=0.$
+When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38, 9].
 In this case it holds without the moment constraint (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant).
-Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ when $\mu$ has vanishing moment (formula 3.1).
-In contrast to [37, 38], it does not seem possible to use symmetric decreasing rearrangements when $\gamma>1$ (see Section 4.1.1).
-We next show that Theorem 3.13 also implies Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2].
-To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
+Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ under the constraint that $\mu$ has vanishing moment.
+In contrast to [37, 38, 9], in this case it does not seem possible to use symmetric decreasing rearrangements, due to a phenomenon of spontaneous symmetry breaking (see Section 4.1.1).
+To facilitate the comparison between our complex-geometric notation and the PDE-literature [37, 39, 38, 9, 72] we provide the following basic lemma, proved in the last section of the appendix:
+Lemma 4.1.
+(normalizations) Let $\omega_{\text{$FS$}}$ denote the Fubini-Study Kähler form on $\mathbb{P}^{1}$ (which is $SU(2)$-invariant and thus coincides with the $SO(3)$-invariant probability measure on $S^{2})$ and denote by $\Delta_{\text{FS}}$ the Laplacian on $\mathbb{P}^{1}$ defined by
+(4.2) $dd^{c}u=\Delta_{\text{FS}}u\omega_{\text{FS}},$
+Then $\Delta_{\text{FS}}$ coincides with the Laplacian $\Delta_{g_{S^{2}}}$ wrt the metric $g_{S^{2}}$ on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}.$ Moreover, when $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta_{w})})$ is $SU(2)$-invariant, i.e.
+when $\omega_{0}=V\omega_{\text{FS}}$ for V=(22w),V=(2-2w),
+(4.3) $E(\mu)=(1-w)\int_{S^{2}\times S^{2}}G\mu\otimes\mu,$
+where $G(\boldsymbol{x},\boldsymbol{y})$ is the Green function for $\Delta_{g_{S^{2}}},$ i.e. $G(\boldsymbol{x},\boldsymbol{y})=\log\left(\frac{|\boldsymbol{x}-\boldsymbol{y}|^{2}}{4}\right)+1.$
+In Section 5 we will establish a sharp version of the constrained logarithmic HLS inequality 4.1 when $w=0.$ But first, in this section we show that Theorem 3.13 implies a generalization of Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2] to $w\neq 0.$ To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
 (4.4) $\mathcal{D}_{\gamma}(u)\coloneqq-\mathcal{E}(u)-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0}.$
 It was introduced in [51] (when $\Delta=0)$ [^3] In the notation of [51], $\mathcal{D}_{\gamma}(u)=F_{\gamma}(u),$ but here the notation $F_{\gamma}$ is reserved for the free energy functional..
-Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (in particular, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2.
+Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (thus, importantely, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2.
 We will restrict $\mathcal{D}_{\gamma}(u)$ to the subspace $C^{1}(X)_{0}$ of all $u\in C^{1}(X)$ satisfying
@@
 The starting point is that when $n=1$, the sup in the definition 3.6 of $E(\mu)$ may be taken over all of $C^{0}(X),$ as pointed out in Example 3.2.
-Hence, $E(\mu)\leq\left\langle v,\mu\right\rangle-\mathcal{E}(v)$ for any $v\in C^{0}(X).$ Now take $\mu=e^{-\gamma u}/\int_{X}e^{-\gamma u}\mu_{0}$ and apply the previous inequality for $v=u.$ The inequality 4.3 then follows from the identity
+Hence, $E(\mu)\leq\left\langle v,\mu\right\rangle-\mathcal{E}(v)$ for any $v\in C^{0}(X).$ Now take $\mu=e^{-\gamma u}/\int_{X}e^{-\gamma u}\mu_{0}$ and apply the previous inequality for $v=u.$ The inequality 4.6 then follows from the identity
 $\frac{1}{\gamma}D\left(\left.\frac{e^{-\gamma u}}{\int_{X}e^{-\gamma u}\mu_{0}}\right|\mu_{0}\right)=-\left\langle u,\frac{e^{-\gamma u}}{\int_{X}e^{-\gamma u}\mu_{0}}\right\rangle-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0},$
-which is a direct consequence of the definition of $D(\mu|\mu_{0}).$ Thanks to the inequality 4.3 the lower bound on $\mathcal{D}_{\gamma}$ in the corollary thus follows directly from the lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ in Theorem 3.13.
+which is a direct consequence of the definition of $D(\mu|\mu_{0}).$ Thanks to the inequality 4.6 the lower bound on $\mathcal{D}_{\gamma}$ in the corollary thus follows directly from the lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ in Theorem 3.13.
 Next, to destabilize $\mathcal{D}_{\gamma}(u),$ first recall [10] that for any $u\in\mathcal{H}(X,\omega_{0})$
@@
 In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed, $F_{\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$ using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e.
-$\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.37, has the desired properties.
+$\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.39, has the desired properties.
 Indeed, $\mu_{t}$ destabilizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ and, just as before, the $\mathbb{Z}_{2}$–invariance of $u_{t}$ (resulting from the $\mathbb{Z}_{2}$–invariance of $\phi_{t}$ and $\phi_{0})$ implies that $e^{-\gamma u_{t}}\mu_{0}$ has vanishing moment.
-Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.35, has the desired properties.
+Finally, when $w\geq 1/2$ the curve $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is the Legendre transform of $v_{t},$ defined by formula 3.37, has the desired properties.
 Again, this follows from the symmetry argument employed in the proof of Theorem 3.13, since $u_{t}$ is $\mathbb{Z}_{2}$–invariant.
@@
 […]
-is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint:
+is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint [^4] Strictly speaking, Aubin’s inequalities [7, Cor 2] are obtained by replacing $\varphi$ with $\varphi$:
 […]
 Proof.
-Note that both the constrained inequalities appearing in the lemma are invariant under the additive action of $\mathbb{R}$ on $C^{1}(X).$ Hence, we may as well assume that $\int u\omega_{0}=0$ and set $v\coloneqq\gamma u.$ Then $\mathcal{E}(u)=-\mathcal{J}(u)$ and the equivalence thus follows directly from $\mathcal{J}(\gamma^{-1}v)=\gamma^{-2}\mathcal{J}(v).$
+Note that both the constrained inequalities appearing in the lemma are invariant under the additive action of $\mathbb{R}$ on $C^{1}(X).$ Hence, we may as well assume that $\int u\omega_{0}=0$ and set $\varphi\coloneqq\gamma u.$ Then $\mathcal{E}(u)=-\mathcal{J}(u)$ and the equivalence thus follows directly from $\mathcal{J}(\gamma^{-1}\varphi)=\gamma^{-2}\mathcal{J}(\varphi).$
 Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10).
-In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.5 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the Corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
+In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.8 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
 Building on Aubin’s result the following sharp form of the the Moser–Trudinger inequality with Aubin’s constraint was established in [72], that settles a conjecture in [39]:
 Theorem 4.5.
-[72]Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.5 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
-We propose the following strengthening of the previous theorem:
-Conjecture 4.5.
-Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant and $\gamma\in]0,2].$ Then
-$F_{\gamma}(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0}.$
-In other words, the conjecture says that $F_{\gamma}(\mu)$ is minimized on $\mathcal{P}(X)_{0}$ by the $\mathrm{SO}(3)$–invariant probability measure $\omega_{0}/V.$ When $\gamma\leq 1$ this follows from the general fact that the Mabuchi functional is minimized on the Kähler potentials of Kähler–Einstein metrics [8].
-We will show below that Conjecture 4.5 holds when $\gamma\leq 3/2,$ using Theorem 4.4.
-Note that the conjecture is equivalent to the inequality 3.30 for the Mabuchi functional $\mathcal{M}(u)$ on $S^{2}$ subject to Tian’s orthogonality condition for $u\in\mathcal{H}(X,\omega_{0})$ for $\epsilon=1$ and $C=0.$
+[72] Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.8 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
+In Theorem 5.1 we will prove strengthening of this result, saying that the minimum of the corresponding free energy $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is attained at the $SO(3)$-invariant probability measure for any $\gamma\in]0,2]$ (by the inequality 4.6 this indeed implies Theorem 4.5).
+But we first give some applications of Theorem 4.5 in Kähler geometry and indicate why the constrained Moser-Trudinger inequality in Theorem 4.5 is, a priori, weaker than the sharp lower bound on $F_{\gamma}$ that will be established in Theorem 5.1.
 4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
@@
 (4.9) $\mathcal{D}_{1}(u)\geq\epsilon(\mathcal{I}-\mathcal{J})(u)-C$
-for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.29.
-The conjecture was settled in [48] and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.5 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.29 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87].
-Here we show, building on Theorem 4.4, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.4 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.6):
+for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.31.
+The conjecture was settled in [48] (with non-effective constants $\epsilon$ and $C)$ and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.8 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.31 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87].
+Here we show, using Theorem 4.5, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.5 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.9):
 Proposition 4.6.
-Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X)$ (i.e.
-$\int\omega_{0}=2).$ For any $v$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.29
-$\frac{1}{2}\mathcal{J}(v)-\int_{X}v\omega_{0}-\log\int_{X}e^{-v}\frac{\omega_{0}}{2}\geq 0,$
-i.e. the inequality is saturated when $v$ is constant.
+Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X).$ For any $\varphi$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.31
+$\frac{1}{2}\mathcal{J}(\varphi)-\int_{X}\varphi\frac{\omega_{0}}{2}-\log\int_{X}e^{-\varphi}\frac{\omega_{0}}{2}\geq 0,$
+i.e. the inequality is saturated when $\varphi$ is constant. Equivalently,
+(4.10) $\mathcal{D}_{1}(\varphi)\geq\frac{1}{2}\mathcal{J}(\varphi).$
 Proof.
-First note that it is is equivalent to show that for any $v$ in $\mathcal{H}(X,\omega_{0})$
-(4.7) $\mathcal{D}_{1}(v)\geq\frac{1}{2}\mathcal{J}(v)$
-if $v$ satisfies Tian’s orthogonality condition (which, in this case, is equivalent to $\int\boldsymbol{x}\omega_{v}=0;$ see Lemma 3.9 and its proof).
-Indeed, since both $\mathcal{D}_{1}$ and $\mathcal{J}$ are invariant under $v\mapsto v+C,$ for any $C\in\mathbb{R},$ we may as well assume that $\int v\omega_{0}=0.$ Then $\mathcal{E}(v)=\mathcal{J}(v)$ which yields the equivalence in question.
+First observe that the equivalence stated in the proposition follows from the fact that both $\mathcal{D}_{1}$ and $\mathcal{J}$ are invariant under $\varphi\mapsto\varphi+C,$ for any $C\in\mathbb{R}.$ Hence, we may as well assume that $\int\varphi\omega_{0}=0.$ Then $\mathcal{E}(\varphi)=\mathcal{J}(\varphi)$ which yields the equivalence in question.
+Note also that $\varphi$ satisfies Tian’s orthogonality condition iff $\int\boldsymbol{x}\omega_{\varphi}=0;$ see Lemma 3.9 and its proof.
 Now take $u$ in $\mathcal{H}(X,\omega_{0})$ and consider the right action of $\mathcal{G}$ on $\mathcal{H}(X,\omega_{0})$ induced by the identification between $\phi\in\text{PSH}(L)$ and $u\coloneqq\phi-\phi_{0},$ where $\phi_{0}$ is a fixed metric on $K_{X}$ with (normalized) curvature form $\omega_{0}.$ In other words, $g^{*}\phi=u\cdot g+\phi_{0}$ and it follows that
@@
 $\mathcal{D}_{1}(u)=\mathcal{D}_{1}(u\cdot g_{0})\geq\frac{1}{2}\mathcal{J}(u\cdot g_{0})\geq\frac{1}{2}\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)$
-by applying the inequality 4.7 to $v\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.7 indeed holds if $u$ satisfies Tian’s constraint.
+by applying the inequality 4.10 to $\varphi\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.10 indeed holds if $u$ satisfies Tian’s constraint.
 ∎
-As a consequence, we deduce the inequality in Conjecture 4.5 for $\gamma\leq 3/2$ (which corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.6):
+As a consequence, we deduce the following weaker form of Theorem 5.1:
 Proposition 4.7.
-Assume $\gamma\leq 3/2.$ Then, for all measures $\mu$ in $\mathcal{P}(X)_{0}$
+Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X)$ and assume $\gamma\leq 3/2.$ Then, for all measures $\mu$ in $\mathcal{P}(X)_{0}$
 $F_{\gamma}(\mu)\geq F_{\gamma}\left(\frac{\omega_{0}}{V}\right)(=0)$
@@
 Proof.
-Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.4 with the inequality 4.7 gives
+Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.7 with the inequality 4.10 gives
 $F_{1}\left(\frac{\omega_{u}}{V}\right)\geq\frac{1}{2}\mathcal{J}(u)\,\,\,\,\text{if $\int\boldsymbol{x}\omega_{u}=0$. }$
@@
 To prove the last statement first note that it follows from the previous proposition, combined with the scaling argument in the proof of Lemma 4.4, that
-$\mathcal{D}_{\gamma}(v)\geq 0,\,\,\text{for}\,\gamma\in[0,2]$
-if $\gamma v$ satisfies Tian’s orthogonality condition and $\gamma v\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{v}=0$ (using that, by homogeneity, $v$ satisfies Tian’s orthogonality condition, i.e.
-$0=\int\boldsymbol{x}v\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{v},$ using integration by parts).
-Furthermore, the condition that $\gamma v\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{v}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption.
-We can thus apply the inequality 4.4 to $v$ to conclude that $F_{\gamma}(\frac{\omega_{v}}{V})\geq 0.$
+$\mathcal{D}_{\gamma}(\varphi)\geq 0,\,\,\text{for}\,\gamma\in[0,2]$
+if $\gamma\varphi$ satisfies Tian’s orthogonality condition and $\gamma\varphi\in\mathcal{H}(X,\omega_{0}).$ The former condition is equivalent to $\int\boldsymbol{x}\omega_{\varphi}=0$ (using that, by homogeneity, $\varphi$ satisfies Tian’s orthogonality condition, i.e.
+$0=\int\boldsymbol{x}\varphi\omega_{0},$ which is proportional to $\int\boldsymbol{x}\omega_{\varphi},$ using integration by parts).
+Furthermore, the condition that $\gamma\varphi\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{\varphi}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption.
+We can thus apply the inequality 4.7 to $\varphi$ to conclude that $F_{\gamma}(\frac{\omega_{\varphi}}{V})\geq 0.$
 4.1. Spontaneous symmetry breaking
-In this section we take the corresponding fixed element $\phi_{0}\in\mathcal{H}(-K_{(\mathbb{P}^{1},\Delta_{w})})$ to be induced by the Fubini–Study metric, i.e.
-to be $\mathrm{SU}(2)$–invariant (Example 3.10) and the points $x_{0}$ and $x_{1}$ as the north and south pole of $X,$ identified with the unit sphere in $\mathbb{R}^{3}.$ Given $(X,\Delta_{w}),$ denote by $\mathcal{P}(X)_{0}^{S^{1}}$ the subspace of $\mathcal{P}(X)_{0}$ consisting of all $S^{1}$–invariant measures $\mu,$ with respect to the standard $S^{1}$–action fixing $x_{0}$ and $x_{1},$ i.e.
+In this section we take the corresponding fixed element $\phi_{0}\in\mathcal{H}(-K_{(\mathbb{P}^{1},\Delta_{w})})$ to be $\mathrm{SU}(2)$–invariant (Example 3.10) and the points $x_{0}$ and $x_{1}$ as the north and south pole of $X,$ identified with the unit sphere $S^{2}$ in $\mathbb{R}^{3}.$ Given $(X,\Delta_{w}),$ denote by $\mathcal{P}(X)_{0}^{S^{1}}$ the subspace of $\mathcal{P}(X)_{0}$ consisting of all $S^{1}$–invariant measures $\mu,$ with respect to the standard $S^{1}$–action fixing $x_{0}$ and $x_{1},$ i.e.
 rotation around the $x_{3}$–axis.
-Likewise, we denote by $C^{1}(X)_{0}^{S^{1}}$ the space of all $S^{1}$–invariant $u\in C^{1}(X)$ satisfying the moment condition 4.2.
+Likewise, we denote by $C^{1}(X)_{0}^{S^{1}}$ the space of all $S^{1}$–invariant $u\in C^{1}(X)$ satisfying the moment condition 4.5.
 Proposition 4.8.
@@
 ∎
-The previous proposition reveals that a phenomenon of spontaneous symmetry breaking arises when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[.$ Indeed, in this case, by Thm 3.13, Cor 4.1 and the previous proposition,
+The previous proposition reveals that a phenomenon of spontaneous symmetry breaking arises when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[.$ Indeed, in this case, by Thm 3.13, Cor 4.2 and the previous proposition,
 (4.11) $\infty=\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}<\inf_{\mathcal{P}(X)_{0}^{S^{1}}}F_{-\gamma},\,\,\,-\infty=\inf_{C^{1}(X)_{0}}\mathcal{D}_{\gamma}(u)<\inf_{C^{1}(X)_{0}^{S^{1}}}\mathcal{D}_{\gamma}(u)$
@@
 The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [37, 38] (using stereographic projection), while in [9] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere.
-Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.8 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$
+Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.11 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$
 4.2. Restricting to $\mathcal{K}$-invariant metrics
@@
 ∎
+5. A sharp constrained log HLS inequality on $S^{2}$
+In this section we will prove the following result, which is equivalent to Theorem 1.4 stated in the introduction (by formula 3.13 and Prop 3.4):
+Theorem 5.1.
+Let $X$ be the two-sphere, $\Delta=0$ and assume that $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then for any given $\gamma\in]0,2]$
+$F_{\gamma}(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0},$
+[…]
+When $\gamma=1$ the inequality for $F_{1}(\mu)$ follows from the general fact that, when a Fano manifold $X$ admits a Kähler–Einstein metric $\omega_{\text{KE}}\in C^{1}(X),$ the Mabuchi functional $\mathcal{M}$ is minimized on the Kähler potential $u$ of $\omega_{\text{KE}}$ [8] and $\mathcal{M}$ is then $\mathcal{G}$-invariant – so that the moment condition on $\mu$ can be dispensed with.
+In the case of the two-sphere (or more generally, the unit sphere in $\mathbb{R}^{n+1})$ the inequality $F_{1}(\mu)\geq 0$ on $\mathcal{P}(X)$ is shown in [37, 38, 9] using symmetrization (the inequality $F_{\gamma}(\mu)\geq 0$ for $\gamma\in[0,1]$ then follows directly).
+Before embarking on the detailed proof of Theorem 5.1 it may be illuminating to compare its structure to the proof of Theorem 4.5 in [72].
+The starting point of the latter is Aubin’s constrained Moser–Trudinger inequality which, in our notations, implies that the twisted Ding functional $\mathcal{D}_{\gamma}$ admits a minimizer $u_{\gamma}$ on the space of all functions in the Hilbert space $W^{2,1}(X)$ satisfying Aubin’s constraint.
+Since this is an infinite dimensional Hilbert manifold it follows that $u_{\gamma}$ satisfies an Euler–Lagrange equation, involving the differential of $\mathcal{D}_{\gamma}$ and three Lagrange multipliers (incorporating Aubin’s constraint).
+As shown in [39, Lemma 3.1], the Lagrange multipliers vanish and, as a result, $u_{\gamma}$ is, in fact, a critical point of $\mathcal{D}_{\gamma}$ on $W^{2,1}(X)$ and thus satisfies a Liouville equation on $S^{2},$ that is shown to have only have constant solutions in [72] (thus confirming a conjecture in [39]).
+In order to prove Theorem 5.1 we replace $\mathcal{D}_{\gamma}$ with the free energy functional $F_{\gamma}$ on $\mathcal{P}(X)_{0}.$ It follows from the formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ in Theorem 3.13 that $F_{\gamma}$ admits a minimizer $\mu_{\gamma}$ in $\mathcal{P}(X)_{0}$ for any given $\in]0,2[.$ However, since the space $\mathcal{P}(X)$ has a boundary it seems hard to directly show that $\mu_{\gamma}$ satisfies an appropriate equation.
+Instead, we leverage a duality formula, expressing the infimum of $F_{\gamma}$ over $\mathcal{P}(X)_{0}$ as the infimum over $W^{2,1}(X)$ of a perturbation of the twisted Ding functional:
+$\inf_{\mathcal{P}(X)_{0}}F_{\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),$
+where $\delta_{\gamma}$ is an entropic distance, defining a function on $W^{2,1}(X)$ taking values in $0,\infty[,$ that vanishes precisely when $\gamma u$ satisfies Aubin’s constraint 4.8 (in particular, the unboundedness from above of $\delta_{\gamma}(u)$ compensates for the unboundedness from below of $\mathcal{D}_{\gamma}(u)).$ This bypasses the boundary problem and allows us to reduce the proof of Theorem 5.1 to the uniqueness result in [72].
+In fact, the proof will show that the inf in the right hand side is attained precisely when $u$ is constant and thus vanishes.
+Remark 5.2.
+A consequence of the proof is that the infimum of $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ over $W^{2,1}(X)$ coincides with the one obtained by replacing the functional $\delta_{\gamma}(u)$ with the indicator function $\chi(u)$ attached to Aubin’s constraint (i.e.
+$\chi(u)$ vanishes when $\gamma u$ satisfies Aubin’s constraint and $\chi(u)=\infty$ otherwise).
+From a general optimization perspective this is somewhat miraculous.
+5.1. Proof of Theorem 5.1
+Recall that $W^{2,1}(X)$ (or $W^{2,1}$ for short) denotes the Hilbert space of all $u\in L^{2}(X)$ such that
+$\left\|u\right\|_{W^{2,1}}^{2}\coloneqq\int_{X}du\wedge d^{c}u+\int_{X}u^{2}\omega_{0}<\infty.$
+We will denote by $W^{2,1}(X)^{*}$ the dual Hilbert space, that we identify with a subspace of the space of all distributions $\nu$ on $X.$ We recall the following standard regularity results that hold on any compact complex curve:
+Lemma 5.3.
+$u\in W^{2,1}(X)\implies\int_{X}e^{-\gamma u}\omega_{0}<\infty$ for all $\gamma\in]0,\infty[.$ Moreover, for $\mu\in\mathcal{P}(X)$ we have that
+$E(\mu)<\infty\iff\mu\in W^{2,1}(X)^{*}\iff u_{\mu}\in W^{2,1}(X).$
+Finally, $D(\mu)<\infty\implies E(\mu)<\infty.$
+Proof.
+The first implication follows from the Moser–Trudinger inequality (which is a consequence of the inequality $\mathcal{D}_{\gamma_{0}}\geq-C_{0}$ on $C^{\infty}(X)$ for a sufficently small positive number $\gamma_{0}$) and the equivalence statement follows from the remark in Example 3.2.
+The final implication follows from the fact that the Green’s function of the Laplacian on $X$ has a logarithmic singularity along the diagonal (or, alternatively, from the inequality $F_{\gamma_{0}}\geq-C_{0}$ on $\mathcal{P}(X)$ for a sufficently small positive number $\gamma_{0}).$
+Given $\gamma>0$ and a function $u\in W^{2,1}(X)$ we set
+$\mu_{\gamma u}:=e^{-\gamma u}\mu_{0}/\int_{X}e^{-\gamma u}\mu_{0},$
+which defines a probability measure on $X,$ by the previous lemma.
+It will be enough to prove the inequality in Theorem 5.1 for a fixed $\gamma\in]1,2[.$ By Theorem 3.13 the infimum of $F_{\gamma}$ on $\mathcal{P}(X)_{0}$ is then finite.
+By the previous lemma we thus have
+$\inf_{\mathcal{P}(X)_{0}}F_{\gamma}=\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{\gamma}>-\infty.$
+5.1.1. A duality formula
+We start by proving the following formula, which applies in the general setup of a log Fano curve $(X,\Delta):$
+(5.1) $\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),\,\,\,\delta_{\gamma}(u)\coloneqq\frac{1}{\gamma}\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\mu_{\gamma u}),$
+where $\mathcal{D}_{\gamma}(u)$ is defined by formula 4.4.
+We will adopt the sign convention for Legendre–Fenchel transforms used in [10]: given a functional $\mathcal{A}$ on $W^{2,1}(X)$ and a functional $A$ on $W^{2,1}(X)^{*}$ we denote by $\mathcal{A}^{*}$ and $A^{*}$ the functionals on $W^{2,1}(X)^{*}$ and $W^{2,1}(X)$ defined by
+$\mathcal{A}^{*}(\mu):=\sup_{u\in W^{2,1}(X)}\mathcal{A}(u)-\left\langle u,\mu\right\rangle,\,\,\,A^{*}(u)=\inf_{\mu\in W^{2,1}(X)^{*}}A(\mu)+\left\langle u,\mu\right\rangle$
+Thus $\mathcal{A}^{*}$ is lsc and convex on $W^{2,1}(X)^{*},$ while $A^{*}$ is usc and concave on $W^{2,1}(X).$ Moreover, $A$ is lsc and convex and $\mathcal{A}$ is usc and concave iff
+(5.2) $(A^{*})^{*}=A,\,\,\,(\mathcal{A}^{*})^{*}=\mathcal{A}$
+using that $W^{2,1}(X)$ is a Hilbert space and, in particular, locally convex (so that the Hahn–Banach theorem applies [50]). Consider now the functional on $W^{2,1}(X)^{*}$ defined by
+$\tilde{D}(\mu)=D(\mu|\mu_{0})\,\,\text{if $\mu$$\in$}\mathcal{P}(X)_{0},\,\,\,\tilde{D}(\mu)=\infty\,\,\text{if $\mu$$\notin$}\mathcal{P}(X)_{0}$
+We can then express, using the remark concerning $E(\mu)$ in Example 3.2 that is specific for $\dim X=1,$
+[…]
+Next note that
+(5.4) $\left(\frac{1}{\gamma}\tilde{D}\right)^{*}(u)=-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0}+\delta_{\gamma}(u):=\mathcal{A}(u)$
+Indeed,
+$\left(\frac{1}{\gamma}\tilde{D}\right)^{*}(u):=\inf_{\mu\in\mathcal{P}(X)_{0}\cap W^{2,1*}}\left(\frac{1}{\gamma}D(\mu|\mu_{0})+\left\langle u,\mu\right\rangle\right)=\inf_{\mu\in\mathcal{P}(X)_{0}}\left(\frac{1}{\gamma}D(\mu|\mu_{0})+\left\langle u,\mu\right\rangle\right),$
+using Lemma 5.3 in the last equality. Rewriting
+$D(\mu|\mu_{0})+\gamma\left\langle u,\mu\right\rangle=D(\mu|\mu_{u})-\log\int_{X}e^{-\gamma u}\mu_{0}$
+then proves 5.4.
+Next, note that the functional $\tilde{D}(\mu)$ on $W^{2,1}(X)^{*}$ is convex and lsc (since $\mathcal{P}(X)_{0}$ is compact and convex in $W^{2,1}(X)^{*})$ and $D(\mu|\mu_{0})$ is convex on $\mathcal{P}(X)_{0}).$ It thus follows from combining formula 5.4 with the duality 5.2 that
+$\frac{1}{\gamma}\tilde{D}=\mathcal{A}^{*},$
+so that formula 5.3 translates into
+$\inf_{\mathcal{P}(X)_{0}\cap W^{2,1}(X)^{*}}F_{\gamma}=\inf_{W^{2,1}(X)^{*}}\left(-\mathcal{E}^{*}+\mathcal{A}^{*}\right).$
+Lemma 5.4.
+Let $\mathcal{E}$ and $\mathcal{A}$ be usc concave functionals on a Hilbert space $H$ and assume that $\mathcal{E}$ and $\mathcal{E}^{*}$ are finite on $H$ and $H^{*}$ respectively. Then
+(5.5) $\inf_{H^{*}}\left(-\mathcal{E}^{*}+\mathcal{A}^{*}\right)=\inf_{H}\left(-\mathcal{E}+\mathcal{A}\right)$
+Proof.
+Assume that $(-\mathcal{E}^{*}+\mathcal{A}^{*})(\mu)\geq C\in\mathbb{R}$ for all $\mu\in H^{*}.$ Then, for any fixed $u\in H,$
+$\mathcal{A}^{**}(u):=\inf_{\mu\in H^{*}}\left(\mathcal{A}^{*}(\mu)+\left\langle u,\mu\right\rangle\right)\geq\inf_{\mu\in H^{*}}\left(\mathcal{E}^{*}(\mu)+\left\langle u,\mu\right\rangle\right)+C:=\mathcal{E}^{**}(u)+C$
+By 5.2, this means that $\mathcal{A}(u)\geq\mathcal{E}(u)+C$ for any $u\in H.$ Conversely, if the latter inequality holds, then, by taking the sup over all $u\in H$ we directly deduce that $\mathcal{A}^{*}(\mu)\geq\mathcal{E^{*}}(\mu)+C$ for all $\mu\in H^{*}.$ Thus $(-\mathcal{E}^{*}+\mathcal{A}^{*})(\mu)\geq C$ for all $\mu\in H^{*}.$
+In the present case $\mathcal{A}$ is a Legendre–Fenchel transform (by formula 5.4) and thus usc and concave on $W^{2,1}(X).$ Furthermore, $\mathcal{E}$ is also usc and concave on $W^{2,1}(X)$ (using the formula in Example 3.2 and the fact that the Hilbert norm on a Hilbert space is convex and continuous on the Hilbert space.
+Thus applying formula 5.5 concludes the proof of the duality formula 5.1.
+5.1.2. Existence of a dual minimizer in $W^{2,1}(X)$
+We next specialize to the log Fano curve $(X,\Delta)=(X,0)$ and take $\phi_{0}$ to be SU(2)SU(2)–invariant, so that $\mu_{0}$ and $\omega_{0}/2$ both coincide with $\omega_{\text{FS }}$ and hence with the $SO(3)$-invariant probability measure on $S^{2}.$ We will show that the functional $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ admits a minimizer in $W^{2,1}(X)$ (the same argument applies to any log Fano curve).
+To prove this we make a change of variable from $u$ to $v$ on $W^{2,1}(X),$ defined by
+$\gamma u=2v,\,\,\,\text{i.e }\mu_{\gamma u}=\frac{e^{2v}\omega_{\text{FS}}}{\int e^{2v}\omega_{\text{FS}}}$
+First assume that $\int u\omega_{0}=0.$ Using that $V=2$ and $\omega_{\text{FS }}=\omega_{0}/2$ we then get
+$\gamma\mathcal{D}_{\gamma}(u)=\frac{1}{\gamma}\int_{X}dv\wedge d^{c}v+\int_{X}2v\omega_{\text{FS}}-\log\int_{X}e^{2v}\omega_{\text{FS}}:=\mathcal{F}_{\gamma}(v)$
+and
+$\gamma\delta_{\gamma}(u)=\delta(v):=\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\frac{e^{2v}\omega_{\text{FS}}}{\int e^{2v}\omega_{\text{FS}}})$
+Thus, we can express
+$\gamma\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)=\mathcal{F}_{\gamma}(v)+\delta(v).$
+Now take $\gamma^{\prime}>\gamma$ such that $\gamma^{\prime}<2.$ Combining Theorem 3.13 for $(X,0)$ with the duality formula 5.1 shows that there exists a constant $C$ such that
+$\gamma^{\prime}\left(\mathcal{D}_{\gamma^{\prime}}(u)+\delta{}_{\gamma^{\prime}}(u)\right)\geq-C$
+i.e.
+$\mathcal{F}_{\gamma^{\prime}}(v)+\delta(v)\geq-C$
+Hence, there exists $\epsilon>0$ such that the following coercivity inequality holds:
+$\mathcal{F}_{\gamma}(v)+\delta(v)\geq\epsilon\int_{X}dv\wedge d^{c}v-C$
+Now take a sequence $v_{j}$ converging towards the inf of $\mathcal{F}_{\gamma}(v)+\delta(v)$ on $W^{2,1}(X).$ Since $\mathcal{F}_{\gamma}(v)+\delta(v)$ is invariant under the additive action of $\mathbb{R}$ we may assume that $\int v_{j}\omega_{0}=0.$ The previous coercivity yields
+(5.6) $\int_{X}dv_{j}\wedge d^{c}v_{j}\leq C,\,\,\,\,\left(\int_{X}v_{j}\omega_{0}=0\right)$
+for some positive constant $C^{\prime}.$ Thus, by the weak compactness of balls in Hilbert spaces we may, after passing to a subsequence, assume that $v_{j}\rightarrow v_{\infty}\in W^{2,1}(X)$ in the weak topology of the Hilbert space $W^{2,1}(X).$ We will show that, after perhaps passing to a subsequence, $v_{\infty}$ minimizes $\mathcal{F}_{\gamma}(v)+\delta(v)$ on $W^{2,1}(X).$ First note that, since the Hilbert norm is lsc wrt the weak topology $\int_{X}dv_{j}\wedge d^{c}v_{j},$ is lower continuous along the sequence $v_{j},$ i.e.
+$\int_{X}dv_{\infty}\wedge d^{c}v_{\infty}\leq\lim\inf_{j}\int_{X}dv_{j}\wedge d^{c}v_{j}$
+By the following lemma, the integral term appearing in $\mathcal{F}_{\gamma}(v)$ is continuous along the sequence $v_{j}:$
+Lemma 5.5.
+The uniform $W^{2,1}$-bound 5.6 implies, that, after passing to a subsequence,
+(5.7) $e^{2v_{j}}\omega_{0}\rightarrow e^{2v_{\infty}}\omega_{0}\,\,\,\left(e^{pv_{\infty}}\in L^{1}(X)\right),$
+in the weak topology of measures.
+Proof.
+This is standard, but for completeness we provide a proof.
+First, by the Rellich–Kondrachov theorem, the uniform bound 5.6 implies that, after passing to a subsequence, $v_{j}\rightarrow v_{\infty}$ in $L^{1}(X).$ As a consequence, after again passing to a subsequence, we may assume that $v_{j}(x)\rightarrow v_{\infty}(x)$ for almost any $x$ in $X.$ Next, by the Moser-Trudinger inequality on $X$ the uniform bound 5.6 also implies that, for any given $\delta>0,$ there exists $A_{\delta}>0$ such that
+(5.8) $\int e^{(1+\delta)2v}\omega_{0}\leq A_{\delta}$
+Now, fix $g\in C^{0}(X).$ Given a large constant $M,$ we split
+$\int e^{2v_{j}}g\omega_{0}=\int_{\left\{v_{j}\leq M\right\}}e^{2v_{j}}g\omega_{0}+\epsilon_{j,M},\,\,\,\,\epsilon_{j,M}:=\int_{\left\{v_{j}>M\right\}}e^{2v_{j}}g\omega_{0}\leq\frac{A_{\delta}\left\|g\right\|_{L^{\infty}}}{e^{2\delta M}},$
+using the uniform bound 5.8 for some $\delta>0.$ Finally, letting first $j\rightarrow\infty$ and using the dominated convergence theorem and then letting $M\rightarrow\infty$ and using the monotone convergence theorem, shows that $\int e^{2v_{j}}g\omega_{0}$ converges towards $\int e^{2v_{\infty}}g\omega_{0},$ as desired.
+∎
+All that remains is to show that $\delta(v)$ is lsc along the sequence $v_{j}.$ To this end, take $\mu_{j}\in\mathcal{P}(X)$ such that $D(\mu_{j}|e^{2v_{j}}\omega_{\text{FS}}/\int e^{2v_{j}}\omega_{\text{FS}})\leq\delta(v_{j})+1/j.$ Since $\mathcal{P}(X)$ is compact, we may, after passing to a subsequence, assume that $\mu_{j}\rightarrow\mu_{\infty}$ in $\mathcal{P}(X)$ for some $\mu_{\infty}.$ Since $D(\cdot|\cdot)$ is lsc with respect to the product topology on $\mathcal{P}(X)^{2}$ it thus follows from the weak convergence 5.7 that
+$\delta(v_{\infty})\leq D(\mu_{\infty}|e^{2v_{\infty}}\omega_{\text{FS}}/\int e^{2v_{\infty}}\omega_{\text{FS}})\leq\liminf_{j\rightarrow\infty}\delta(v_{j}),$
+where the first inequality follows from the very definition of $\delta(v_{\infty}).$
+5.1.3. The minimizer satisfies a Liouville type PDE
+The functional $\mathcal{F}_{\gamma}(v)$ is Gateaux differentiable on $W^{2,1}(X)$ and a direct differentiation gives, for any $v\in W^{2,1}(X)$ and $\dot{v}\in C^{\infty}(X),$
+$\frac{d\mathcal{F}_{\gamma}(v+t\dot{v})}{dt}=\left\langle\left(-\frac{1}{\gamma}2dd^{c}v+2\omega_{\text{FS}}-2\frac{e^{2v}\omega_{\text{FS}}}{\int e^{2v}\omega_{\text{FS}}}\right),\dot{v}\right\rangle.$
+This means that
+$\frac{d\mathcal{F}_{\gamma}(v+t\dot{v})}{dt}=2\int_{X}\left(-\frac{1}{\gamma}\Delta_{\text{FS}}v+1-e^{2v}\right)\dot{v}\omega_{\text{FS}}$
+after, perhaps adding a constant to $v$ so that the following normalization condition holds:
+(5.9) $\int e^{2v}\omega_{\text{FS}}=1,$
+which will henceforth be assumed. Moreover, the functional $\delta(v)$ is also Gateaux differentiable and there exists a vector $\boldsymbol{\lambda}\in\mathbb{R}^{3}$ such that
+(5.10) $\frac{d\delta(v+t\dot{v})}{dt}=\left(\boldsymbol{\lambda}\cdot\boldsymbol{x}-\left\langle\boldsymbol{\lambda}\cdot\boldsymbol{x}\right\rangle\right)e^{2v}\omega_{\text{FS}},$
+where, given a function $f\in L^{1}(X),$ we will write
+$\left\langle f\right\rangle:=\int_{X}fe^{2v}\omega_{\text{FS}},$
+assuming that the normalization condition 5.9 holds. The vector $\boldsymbol{\lambda}\in\mathbb{R}^{3},$ depending on $v,$ is uniquely determined by the condition that
+(5.11) $\int_{X}\boldsymbol{x}e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}e^{2v}\omega_{\text{FS}}=0.$
+To see this, we express
+$\delta(v)\coloneqq\delta(\nu)\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\nu),\,\,\,\nu\coloneqq e^{2v}\omega_{\text{FS}}$
+It follows from classical convex analysis that $\delta(\nu)$ is differentiable along smooth curves $\nu_{t}$ in $\mathcal{P}(X).$ Indeed, $D(\cdot|\nu)$ is strictly convex and not identically $\infty$ on the non-empty convex and compact subspace $\mathcal{P}(X)_{0}$ of $\mathcal{P}(X),$ which ensures that the inf above is attained at unique minimizer $\mu\in\mathcal{P}(X)_{0}.$ Moreover, when $\nu_{0}:=\nu,$
+$\frac{d\delta(\nu_{t})}{dt}_{|t=0}=\left\langle\log(\frac{\mu}{\nu}),\frac{d\nu_{t}}{dt}\right\rangle.$
+Furthermore, the minimizer $\mu$ is given by
+(5.12) $\mu=\frac{e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu}{\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu},$
+where $\boldsymbol{\lambda}\in\mathbb{R}^{3}$ is uniquely determined by the condition that $e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu\in\mathcal{P}(X)_{0}.$ Indeed, the existence and uniqueness of $\boldsymbol{\lambda}$ follows directly from the fact that the function $\boldsymbol{\lambda}\mapsto\log\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu$ on $\mathbb{R}^{3}$ is strictly convex and proper (using that $\nu$ and $\omega_{\text{FS }}$are mutually absolutely continuous) and thus admits a critical point $\boldsymbol{\lambda}.$ Moreover, a direct computation reveals that the measure $\mu$ in $\mathcal{P}(X)_{0}$ defined by formula 5.12 is a critical point of $D(\cdot|\nu)$ in $\mathcal{P}(X)_{0}$ and thus minimizers $D(\cdot|\nu).$ Finally, taking $\nu_{t}$ to be the curve corresponding to $v+t\dot{v}$ proves formula 5.10.
+Let now $v$ be a minimizer of $\mathcal{F}_{\gamma}+\delta$ on $W^{2,1}(X).$ In particular, the Gateaux differential of $\mathcal{F}_{\gamma}+\delta$ vanishes at $v,$ forcing
+$\left(-\frac{1}{\gamma}\Delta_{\text{FS}}v+1-e^{2v}\right)+\left(\boldsymbol{\lambda}\cdot\boldsymbol{x}-\left\langle\boldsymbol{\lambda}\cdot\boldsymbol{x}\right\rangle\right)e^{2v}=0,$
+i.e. $v$ satisfies the PDE
+$\Delta_{\text{FS}}v+\gamma(1-(\boldsymbol{\lambda}\cdot\boldsymbol{x}-\left\langle\boldsymbol{\lambda}\cdot\boldsymbol{x}\right\rangle)e^{2v}=\gamma$
+5.1.4. The minimizer $v$ is constant
+To conclude the proof it will be enough to show that any minimizer $v$ of $\mathcal{F}_{\gamma}+\delta$ on $W^{2,1}(X)$ is constant.
+By the uniqueness result for the Liouville equation on $S^{2}$ for $\gamma\in]1,2]$ in [72] it is enough to show that $\boldsymbol{\lambda}=0.$ To this end first recall that, by a result of Kazdan-Warner [39, formula 2.11] (and the normalizations in Lemma 4.1) any solution $u$ of the equation
+$\Delta_{\text{FS}}u+he^{u}=c$
+satisfies the condition
+(5.13) $\int\nabla h\cdot\nabla x_{i}e^{u}\omega_{\text{FS}}=(2-c)\int hx_{i}e^{u}\omega_{\text{FS}}$
+Taking $u=2v,$ $c=2\gamma$ and $h=2(1-(\boldsymbol{\lambda}\cdot\boldsymbol{x}-\left\langle\boldsymbol{\lambda}\cdot\boldsymbol{x}\right\rangle)$ we deduce that
+$\left\langle\nabla h\cdot\nabla x_{i}\right\rangle=(2-2\gamma)\left\langle hx_{i}\right\rangle$
+Next, note that we may, without loss of generality, assume that
+$\boldsymbol{\lambda}=(0,0,\lambda)$
+Indeed, taking $F$ to be a rotation of $S^{2}$ having the property that $F^{*}\boldsymbol{\lambda}\cdot\boldsymbol{x}=\lambda x_{3}$ for some $\lambda\in\mathbb{R}$ we can replace $v$ with $F^{*}v.$ Now, applying the Kazdan-Warner condition 5.13 to $i=3$ gives, by multiplying both sides with $\lambda,$
+(5.14) $\left\langle\nabla(\lambda x_{3})\cdot\nabla(\lambda x_{3})\right\rangle=(2-2\gamma)\left\langle-\lambda x_{3}+\lambda^{2}x_{3}-\left\langle\lambda x_{3}\right\rangle\lambda x_{3}\right\rangle.$
+Using that $\left\langle\lambda^{2}x_{3}-\left\langle\lambda x_{3}\right\rangle\lambda x_{3}\right\rangle=\left\langle\lambda x_{3}-\left\langle\lambda x_{3}\right\rangle\right\rangle^{2}$ we deduce that
+$\left\langle\nabla(\lambda x_{3})\cdot\nabla(\lambda x_{3})\right\rangle=(2-2\gamma)\left(-\left\langle\lambda x_{3}\right\rangle+\left\langle\lambda x_{3}-\left\langle\lambda x_{3}\right\rangle\right\rangle^{2}\right).$
+Next, note that
+$\left\langle\lambda x_{3}\right\rangle\leq 0.$
+Indeed, setting $f:=\lambda x_{3}$ we have, by the defining property 5.11 of $\lambda,$ that $\left\langle fe^{f}\right\rangle=0.$ But, by Jensen’s inequality
+$\frac{\left\langle fe^{f}\right\rangle}{\left\langle e^{f}\right\rangle}\geq\log\left\langle e^{f}\right\rangle\geq\left\langle f\right\rangle.$
+All in all this means that the bracket in the rhs of formula 5.14 is non-negative and strictly positive unless $\lambda=0.$ Since the same thing goes for the lhs in formula 5.14 and $(2-2\gamma)<0$ we deduce that $\lambda=0,$ as desired.
+5.1.5. Any minimizer of $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ equals $\omega_{\text{FS}}$ for any $\gamma\in]1,2]$
+Given $\mu\in\mathcal{P}(X)$ such that $D(\mu)<\infty$ the following formula follows directly from the definitions:
+$\gamma^{-1}F_{\gamma}(\mu)=\mathcal{D}_{\gamma}(u_{\mu})+\gamma^{-1}D(\mu|\mu_{\gamma u_{\mu}}),$
+where $u_{\mu}$ denotes the potential of $\mu$ which, by Lemma 5.3, is in $W^{2,1}(X),$ since $D(\mu)<\infty.$ Now assume furthermore that $\mu\in\mathcal{P}(X)_{0}.$ Then it follows directly from the definition 5.1 of $\delta_{\gamma}(u)$ that
+$\gamma^{-1}F_{\gamma}(\mu)\geq\mathcal{D}_{\gamma}(u_{\mu})+\delta_{\gamma}(u_{\mu}).$
+Thus, if $\mu$ minimizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ then, by the first part of Theorem 5.1, proved above and the the duality formula 5.1, $u_{\mu}$ minimizes $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ on $W^{2,1}(X).$ But, as shown in Section 5.1.4, this implies that $u_{\mu}$ is constant, i.e.
+$\mu=\omega_{\text{FS}},$ as desired.
 6. The $\mathcal{K}$–reduced probability measures on $X^{N}$
@@
 6.1. Probabilistic and statistical mechanical setup
-5.1.1.
+6.1.1.
 The Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ and the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ and partition function $Z_{\boldsymbol{m}_{N}}.$
 Given a log Fano manifold $(X,\Delta),$ denote by $\nu^{(N)}$ the measure on $X^{N}$ defined by
-$\nu^{(N)}=i^{N^{2}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
+$\nu^{(N)}=\frac{i^{(Nn)^{2}}}{2^{Nn}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
 where $\det S^{(N)}$ denotes the section of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ obtained by taking $s_{1},\dots,s_{N}$ in formula 1.3 to be a basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$ When 0 is a regular value of $\boldsymbol{m}_{N}$ we will denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the corresponding Gelfand–Leray measures, supported on $X^{N}.$ While this measure depends on the choice of basis in $\mathfrak{k},$ changing the basis only has the effect of multiplying $\nu_{\boldsymbol{m}_{N}}$ with a positive constant $C$ (the determinant of the corresponding change of bases matrix).
@@
 Remark 6.1.
-When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 5.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed).
+When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 6.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed).
 It would be interesting to know if, in general, this limit exists and if it coincides with the measure supported on the regular locus, defined by the Gelfand–Leray type measure on the regular locus.
@@
 (6.7) $\gamma(X,\Delta)^{\mathcal{G}}\geq\sup_{\gamma>0}\left\{F(-\gamma)_{0}>-\infty\right\}.$
-Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 5.11.
+Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 6.11.
 Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
 $0=-\lim_{N\rightarrow\infty}N^{-1}\log\int_{X^{N}}\mu_{\beta}^{(N)}=\inf_{\mathcal{P}(X)_{0}}F_{\beta}(\mu)-C,$
-showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 5.7 is an equality.
+showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 6.7 is an equality.
 As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [42] for the case $\Delta=0$ with $X$ smooth and [77, 79] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
 Next, to prove that Conjecture 6.4 A implies Conjecture 1.6, assume that $(X,\Delta)$ admits a KE metric.
-It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 5.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.12 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric).
+It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 6.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.14 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric).
 As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large.
@@
 It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)).
-If inequality 5.9 can be strengthened as follows, then the second assumption implies the first:
+If inequality 6.9 can be strengthened as follows, then the second assumption implies the first:
 Lemma 6.7.
@@
 The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds).
-The second statement is shown in a similar way, using Prop 5.11 below.
+The second statement is shown in a similar way, using Prop 6.11 below.
 ∎
@@
 (6.11) $C^{-1}\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}$
-But this follows from combining the inequalities 5.12 below with the first identity in Proposition 5.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
+But this follows from combining the inequalities 6.12 below with the first identity in Proposition 6.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
 6.4. Relating Gibbs polystability and the restricted partition function $Z_{\boldsymbol{m}_{N}}$ to integrals over $(X^{N})_{\mathrm{s}}/\mathcal{G}$
-The main virtue of the following proposition, as compared with Prop 5.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
+The main virtue of the following proposition, as compared with Prop 6.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
 Proposition 6.12.
@@
 Proof.
-Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 5.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small.
-Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 5.10.
+Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 6.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small.
+Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 6.10.
 As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite.
@@
 $\rho_{N}(\boldsymbol{y},0)d\boldsymbol{y}=\left.\Phi^{*}((\nu^{(N)})^{\mathrm{c}})\right|_{(\mathbf{y},0)}.$
-Hence, the desired formula 5.13 follows from formula 5.14. ∎
+Hence, the desired formula 6.13 follows from formula 6.14. ∎
 Remark 6.13.
-Formula 5.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
+Formula 6.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
 Consider now the following function on $X^{N},$
@@
 $\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}<\infty$
-5.4.1. Proof of Prop 5.14
+6.4.1. Proof of Prop 6.14
 We will consider a general setup of a complex reductive Lie group $G$ acting on a polarized complex projective manifold $(X,L)$ (in the previous proposition the role of $X$ is played by $X^{N})$.
@@
 $\int_{\left\{m=0\right\}/K}\beta=C\int\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu.$
-To this end it is enough to show that the projection $p$ in 5.16 satisfies
+To this end it is enough to show that the projection $p$ in 6.16 satisfies
 $p_{*}\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu=C\beta,$
@@
 where the dots indicates a sum of forms that contain some factor of $dh_{i}$ for some $i.$ In particular, such forms vanish when restricted to $\{m=0\}.$
-Finally, Prop 5.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
+Finally, Prop 6.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
 6.5. Examples where the quantized Futaki character vanishes and reductivity holds
@@
 We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11).
-Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 5.20, which implies the lemma.
+Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 6.20, which implies the lemma.
 ∎
@@
 7.1.1. Expressing the contracted partition function $Z_{N}^{\mathrm{c}}$ as a complex Selberg integral $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
-By Prop 5.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 5.4) may be expressed as
+By Prop 6.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 6.4) may be expressed as
 $Z_{N}^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}(\nu^{(N)})^{\mathrm{cc}}$
@@
 (7.3) $\nu^{(N)}=\eta\wedge\rho_{N-3}^{\boldsymbol{w}}(y_{4},\dots,y_{N}),\,\,\,\eta\coloneqq\frac{(\frac{i}{2})^{3}dz_{1}\wedge d\overline{z}_{1}\wedge\cdots\wedge dz_{3}\wedge d\overline{z}_{3}}{\left|(z_{1}-z_{2})(z_{1}-z_{3})(z_{2}-z_{3})\right|^{2}}$
-Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 6.1, applied to $N=3).$ Thus formula 6.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 5.9.
+Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 7.1, applied to $N=3).$ Thus formula 7.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 6.9.
 Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant.
@@
 $\eta^{\mathrm{cc}}=1.$
-Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 5.4.1. Thus, accepting formula 6.3, we conclude that
+Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 6.4.1. Thus, accepting formula 7.3, we conclude that
 $(\nu^{(N)})^{\mathrm{cc}}=\rho_{N-3}^{\boldsymbol{w}}(y_{4},\dots,y_{N}),$
-which implies the desired formula 6.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
-Finally, as next described, the deferred proof of formula 6.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 6.2).
+which implies the desired formula 7.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
+Finally, as next described, the deferred proof of formula 7.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 7.2).
 First note that since $V=2$ and $N=2k+1$ [18, Lemma 4.3] yields
@@
 $=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\bigwedge_{\ell=1}^{N}\mathrm{d}z_{\ell}\wedge\mathrm{d}\bar{z}_{\ell}=\nu^{(N)}(z),$
-which concludes the proof of formula 6.3.
+which concludes the proof of formula 7.3.
 7.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
-We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 5.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
+We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 6.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
 (7.4) $N^{-1}C\leq\omega_{N}(e_{3},Je_{3})\leq C\,\,\text{on $\{\left\langle\boldsymbol{m}_{N},e_{3}\right\rangle=0\}\subset X^{N}$.}$
@@
 Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above.
-Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 6.4.
+Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 7.4.
 Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
@@
 7.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
-As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 6.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial).
+As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 7.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial).
 By [3, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
@@
 (7.9) $\mathcal{Z}_{1}^{\boldsymbol{w}}\coloneqq\int_{\mathbb{C}}\left|z\right|^{-2w_{1}}\left|z-1\right|^{-2w_{2}}\frac{i}{2}dz\wedge d\bar{z}=\pi\frac{\Gamma(1-w_{1})}{\Gamma(w_{1})}\frac{\Gamma(1-w_{1})}{\Gamma(w_{1})}\frac{\Gamma(-1+w_{1}+w_{2})}{\Gamma(2-w_{1}-w_{2})}$
-(as follows from formula 6.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
+(as follows from formula 7.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
 $\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right)$
-Combining this convergence with the convergence 6.5 gives the following explicit formula
+Combining this convergence with the convergence 7.5 gives the following explicit formula
 (7.10) $\inf\mathcal{M}_{\mathbb{Z}}=\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right),$
-which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 6.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 6.6.
+which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 7.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 7.6.
 Note that
@@
 where the error terms are positive.
-Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 6.8:
+Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 7.8:
 $\frac{1}{N}\left(\log N!-N\log\left(-l\left(-\frac{1}{2}\frac{\left|V_{N}\right|}{N-1}\right)\right)\right)=\log\frac{|V|}{2}-1+\frac{1}{2}\frac{\log N}{N}+O\left(N^{-1}\right).$
@@
 $\int_{0}^{1}\log\Gamma\left(w_{N}+t\frac{V_{N}}{2}\right)dt=\int_{0}^{1}\log\Gamma\left(w+t\frac{V}{2}\right)dt+O(N^{-1}\log N),$
-using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 6.11.
-The nominator in the second factor in formula 6.8 is handled in essentially the same manner, giving
+using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 7.11.
+The nominator in the second factor in formula 7.8 is handled in essentially the same manner, giving
 $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}_{N}}=\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right)+O(N^{-1}\log N)$
-for $w_{i}=0$ and $V=2.$ The proof of formula 6.6 is thus concluded by invoking formula 6.10.
+for $w_{i}=0$ and $V=2.$ The proof of formula 7.6 is thus concluded by invoking formula 7.10.
 7.2. Proof of Theorem 1.7 when $\Delta$ is nontrivial
@@
 for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e.
-$\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 6.13.
+$\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 7.13.
 8. Relations to arithmetic geometry
@@
 In fact, this conjecture is equivalent to the convergence in Conjecture 1.6, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [3, Thm 5.3]).
-One drawback of Conjecture 1.5 – from the arithmetic point of view – is that left hand side in formula 7.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).).
+One drawback of Conjecture 1.6 – from the arithmetic point of view – is that left hand side in formula 8.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).).
 This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 6.5) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
@@
 $\lim_{N\rightarrow\infty}N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})=\inf\mathcal{M}_{\mathbb{Z}}$
-By Prop 5.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 5.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.6).
+By Prop 6.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 6.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.7).
 8.1. Examples
-The condition 7.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e.
+The condition 8.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e.
 $\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [66, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [81, Section 2].
-By Prop 5.20, the assumptions of Conjecture 7.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish.
+By Prop 6.20, the assumptions of Conjecture 8.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish.
 It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial.
 One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases.
-For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 6.1 and formula 6.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
+For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 7.1 and formula 7.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
 $Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1})=\pi^{N-3}\frac{(N-3)!}{-l(-\frac{1}{N-4})^{N}}\prod_{j=0}^{N-4}\frac{-l(-(j+1)\frac{1}{N-4})}{l(\frac{2}{N-1}+j\frac{1}{N-4})^{3}},\,\,\,\,\,l(x)\coloneqq\frac{\Gamma(x)}{\Gamma(1-x)}$
-(and a similar formula holds when $N=4,$ by formula 6.9).
-Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 6.12 and formula 6.8 gives, when $N\geq 3,$
+(and a similar formula holds when $N=4,$ by formula 7.9).
+Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 7.12 and formula 7.8 gives, when $N\geq 3,$
 $Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{w})=\pi^{N-1}\left(\frac{(N-1)!}{-l(-\frac{1-w}{N-2})^{N}}\right)\prod_{j=0}^{N-2}\frac{-l(-(j+1)\frac{1-w}{N-2})}{l(w+j\frac{1-w}{N-2})^{2}l(\frac{1-w}{N-1}+j\frac{1-w}{N-2})}.$
@@
 $V\mathcal{M}(\phi)=\mathcal{L}(v)+\int_{\mathbb{R}}(\log\phi^{\prime\prime})\phi^{\prime\prime}dx$
-(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.11, this means that
+(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.13, this means that
 $F_{-1}(\mu)=\gamma E(\mu)+D(\mu|\mu_{\phi})=\mathcal{L}(v)-\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx+\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$
-Note that, by formula 8.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
+Note that, by formula 9.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
 (9.2) $E(\mu)=-\mathcal{L}(v)+\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx,$
@@
 (using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [22, Prop 2.9] in the second equality).
-For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.1 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.35 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
+For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.2 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.37 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
 $\phi_{t}(x)=0,\,\text{when }|x|\leq t,\,\,\,\,\phi_{t}(x)=|x-t|,\,\text{when }t\leq|x|$
-A simple direct calculation gives $\mathcal{D}_{\gamma}(u_{t})=(1/2-\frac{\gamma-1}{\gamma})t+O(1),$ tending to -$\infty$ as $t\rightarrow\infty$ if $\gamma>2.$ This shows that $u_{t}$ is a destabilizing curve for $\mathcal{D}_{\gamma}$ (which provides an alternative direct proof of the destabilizing result in the proof of Cor 4.1).
+A simple direct calculation gives $\mathcal{D}_{\gamma}(u_{t})=(1/2-\frac{\gamma-1}{\gamma})t+O(1),$ tending to -$\infty$ as $t\rightarrow\infty$ if $\gamma>2.$ This shows that $u_{t}$ is a destabilizing curve for $\mathcal{D}_{\gamma}$ (which provides an alternative direct proof of the destabilizing result in the proof of Cor 4.2).
 A similar argument applies when $w\in]0,1/2],$ by multiplying the formula for $\phi_{t}(x)$ above by $V/2.$
+9.3. Proof of Lemma 4.1 (normalizations)
+For any metric $g$ on $S^{2}$ compatible with the complex structure on $\mathbb{P}^{1}$
+$dd^{c}u:=\frac{1}{\pi}\frac{i}{2}\partial\bar{\partial}u=\Delta_{g}u\frac{dV_{g}}{4\pi}$
+In particular, if $g=g_{S^{2}}$ then $dV_{g}/4\pi$ has unit-mass and thus coincides with $\omega_{\text{FS}}.$ Hence, the defining formula 4.2 for $\Delta_{\text{FS}}$ forces $\Delta_{\text{FS}}=\Delta_{g_{S^{2}}}.$ Next, by formula 3.12,
+$E(\mu)=-\frac{1}{2}\int_{X}u_{\mu}\mu,\,\,\text{where}\,dd^{c}(\frac{u_{\mu}}{V})+\omega_{\text{FS}}=\mu,\,\int u_{\mu}\omega_{\text{FS}}=0.$
+Let $G(x,y)$ be the symmetric lsc function on $X^{2}$ defined by $dd^{c}G(,\cdot y)=\omega_{\text{FS}}-\delta_{y},$ where $\delta_{y}$ denotes the Dirac measure at $y$ and $\int_{X}G(\cdot,y)\omega_{\text{FS }}=0.$ Then the previous equation yields
+$E(\mu)=\frac{V}{2}\int_{X^{2}}G(x,y)\mu\otimes\mu=(1-w)\int_{X^{2}}G(x,y)\mu\otimes\mu.$
+Finally, in the holomorphic coordinate $z$ on $\mathbb{C}\subset\mathbb{P}^{1}$ we have $dd^{c}(\log(1+|z|^{2})=\omega_{\text{FS}}$ and $dd^{c}(\log|z|^{2})=\delta_{0}.$ Thus
+$G(z,w)=\log\frac{\left|z-w\right|^{2}}{\left(1+\left|z\right|^{2}\right)\left(1+\left|w\right|^{2}\right)}-C,$
+$C:=\int_{\mathbb{C}}\log\frac{\left|z\right|^{2}}{1+\left|z\right|^{2}}\omega_{\text{FS}}=\int_{\mathbb{C}}\log\frac{\left|z\right|^{2}}{1+\left|z\right|^{2}}\frac{1}{\pi}\frac{i}{2}\frac{dz\wedge d\bar{z}}{(1+\left|z\right|^{2})^{2}}=\int_{0}^{\infty}\log\frac{r^{2}}{1+r^{2}}\frac{d(r^{2})}{(1+r^{2})^{2}}=-1$
+Finally, using stereographic projection gives the explicit formula for $G(\boldsymbol{x},\boldsymbol{y}).$ Indeed, by $SO(3)$-symmetry, the argument of log above equals $(c\left|\boldsymbol{x}-\boldsymbol{y}\right|)^{2}$ for some $c>0$ and using that the diameter of $S^{2}$ equals $2$ forces $c=1/2.$

### v3

@@
 Abstract.
-We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by explicitly breaking the symmetry.
-This yields a new algebraic notion of Gibbs polystability, conjecturally equivalent to K-polystability.
-The definition involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides with an analytic reduced stability threshold encoding the coercivity of the K-energy functional modulo automorphisms.
-These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves.
-By imposing a moment constraint, we derive a strengthened form of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, that implies the sharp form of Aubin’s refinement of the Moser–Trudinger inequality.
+We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by breaking the symmetry using a moment map constraint.
+In particular, an algebraic notion of Gibbs polystability is introduced, ensuring that the corresponding point processes on $X$ are well-defined.
+We conjecture that the Gibbs polystability of $(X,\Delta)$ is equivalent to the existence of a Kähler–Einstein metric and that the unique such metric with vanishing moment emerges when sampling a large number of $N$ points on $X.$ The definition of Gibbs polystability involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides – as $N$ tends to infinity – with an analytic reduced stability threshold, encoding the coercivity of the K-energy functional modulo automorphisms.
+These conjectures follow from an overarching conjectural Large Deviation Principle for the large NN-limit.
+We prove several of our conjectures on log Fano curves and derive a strengthened form of the sharp logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere, under a moment constraint.
+It yields quantitative stability results for the sharp logarithmic HLS inequality with optimal stability constants.
+Furthermore, we show that any log Fano manifold that is strongly uniformly Gibbs polystable admits a Kähler–Einstein metric.
 In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
@@
 According to the seminal Yau–Tian–Donaldson conjecture a compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e.
-a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold - this is now a theorem [42, 90].
+a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold – this is now a theorem [53, 108].
 The Fano condition requires that the anti-canonical line bundle $K_{X}$ of $X$ is ample, while K-polystability is a finer algebro-geometric condition that amounts to a positivity property of all polarized $\mathbb{C}^{*}$–equivariant deformations of $(X,-K_{X}),$ called test configurations.
@@
 In the case of positive Ricci curvature, it requires, however, that the Fano manifold $X$ does not admit any holomorphic vector fields, i.e.
-that the group $\text{Aut}_{0}(X)$ of automorphism of $X$ homotopic to the identity, be trivial.
+that the group $\text{Aut}_{0}(X)$ of automorphisms of $X$ homotopic to the identity be trivial.
 In the present work the framework is extended to Fano manifolds with nontrivial $\text{Aut}_{0}(X)$.
@@
 The starting point is the basic fact that a Kähler–Einstein (KE) metric $\omega_{\text{KE}}$ with positive Ricci curvature on $X$ can be readily be recovered from its (normalized) volume form $dV_{\text{KE}}.$ The strategy of the probabilistic approach is thus to construct the normalized volume form $dV_{KE}$ by sampling a large number of $N$ points on $X.$ More precisely, the goal is to construct a canonical symmetric probability measure on the $N$-fold products $X^{N}$ – solely defined in terms of explicit algebro-geometric data of $X$ – and show that $dV_{\text{KE}}$ emerges as $N\rightarrow\infty,$ when sampling $N$ points.
-The construction of the probability measure on $X^{N}$ is based on the observation that $N$-fold products $X^{N}$ come with an algebraic measure $\nu^{(N)},$ uniquely determined up to scaling.
+The construction of the probability measure on $X^{N}$ is based on the observation that $N$-fold products $X^{N}$ come with an algebraic measure $\nu^{(N)},$ canonically determined up to scaling.
 More precisely, $N$ is the subsequence defined by
@@
 In fact, when $\text{Aut}_{0}(X)$ is non-trivial there is no probability measure on $X^{N}$ that is invariant under the diagonal action of $\text{Aut}_{0}(X)$ (under a mild regularity assumption) [21].
-The probabilistic approach in [13] is thus too canonical, in a certain sense, and we will need to break the $\text{Aut}_{0}(X)$–symmetry.
-This is in line with the fact the group $\text{Aut}_{0}(X)$ acts non-trivially on any given Kähler–Einstein metric.
-Hence, there is no canonical way of singling out a unique Kähler–Einstein metric.
+The probabilistic approach in [14] is thus too canonical, in a certain sense.
+This is in line with the fact the group $\text{Aut}_{0}(X)$ acts non-trivially on any given Kähler–Einstein metric – as a consequence there is no canonical way of singling out a unique Kähler–Einstein metric.
+In order to probabilistically construct a Kähler–Einstein metric we will thus need to break the $\text{Aut}_{0}(X)$–symmetry.
+This is explained in Section 1.5 below, but we first introduce an algebro-geometric notion of Gibbs polystability, taking the action of the group $\text{Aut}_{0}(X)$ into account and explore the relations to stability thresholds.
 1.2. Gibbs polystability and microscopic stability thresholds
@@
 This setting encompasses, in particular, Fano orbifolds [78].
-Kähler–Einstein metrics on $(X,\Delta)$ are, in general, understood as positive currents $\omega_{\text{KE}}$ in $c_{1}\left(-K_{(X,\Delta)}\right)$ with locally bounded potentials, that restrict to genuine Kähler–Einstein metrics on the complement of the support of $\Delta.$ In particular, $\omega_{\text{KE }}$ satisfies the following equation in sense of currents:
+Kähler–Einstein metrics on $(X,\Delta)$ are, in general, understood as positive currents $\omega_{\text{KE}}$ in $c_{1}\left(-K_{(X,\Delta)}\right)$ with locally bounded potentials, that restrict to genuine Kähler–Einstein metrics on the complement of the support of $\Delta.$ In particular, $\omega_{\text{KE }}$ satisfies the following equation in the sense of currents:
 $\text{Ric $\omega$}=\omega+[\Delta],$
@@
 and we will assume that the complex Lie group $\mathcal{G}$ is reductive (this is a necessary condition for the existence of a Kähler–Einstein metric for $(X,\Delta)$ [53]).
-We start by introducing a notion of Gibbs polystability, taking the group $\mathcal{G}$ into account.
+We start by introducing an algebro-geometric notion of Gibbs polystability, taking the group $\mathcal{G}$ into account.
 It is formulated in terms of the Zariski open subvariety $(X^{N})_{\text{ss}}$ of $X^{N},$ defined by the semistable locus of $X^{N},$ with respect to the ample $\mathbb{Q}$–line bundle $K_{(X,\Delta)},$ in the sense of Geometric Invariant Theory (GIT) [83].
@@
 But we present arguments motivating the second one.
-They are based on conjectural Large Deviation Principle that we shall come back to below, linking the invariant $\gamma(X,\Delta)\mathcal{{}^{G}}$ to the reduced analytic stability threshold defined by
+They are based on a conjectural Large Deviation Principle that we shall come back to below, linking the invariant $\gamma(X,\Delta)\mathcal{{}^{G}}$ to the reduced analytic stability threshold defined by
 (1.7) $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon(\mathcal{I}-\mathcal{J})^{\mathcal{G}}+C_{\epsilon}\right\},$
-where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right),$ whose minimizers are Kähler–Einstein metrics for $(X,\Delta),$ and $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry
+where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right),$ whose minimizers are Kähler–Einstein metrics for $(X,\Delta),$ and $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the classical non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry
 (1.8) $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}=\text{inf}_{g\in\mathcal{G}}g^{*}(\mathcal{I}-\mathcal{J})$
-(the definition of $\mathcal{I}$ and $\mathcal{J}$ are recalled in Remark 3.1).
-In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the square of the Dirichlet norm of the Kähler potential.
+(the definition of $\mathcal{I}$ and $\mathcal{J}$ are recalled in Remark 3.1.
+In the case of a complex curve, all the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ are proportional to the Dirichlet norm of the Kähler potential.
 On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [107, 25, 61]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [113, 61].
 In the present context, however, it will be important to work with the functional $\mathcal{I}-\mathcal{J}$ as this choice leads to a thermodynamical interpretation of the corresponding reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (see Prop 3.6).
+It should be emphasized that, while the definition of $\mathcal{I}-\mathcal{J}$ depends on the choice of a reference Kähler metric $\omega_{0}\in c_{1}(X)$, this does not affect the definition 1.7 of $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}},$ since it only alters the additive constant $C_{\epsilon}.$
 When $\mathcal{G}$ is trivial the invariant $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ equals the well-known analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta),$ which by [112], coincides with the stability threshold $\delta(X,\Delta),$ introduced in [73] and further developed in [35].
@@
 $\gamma(X,0)^{\mathcal{G}}=\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2,\,\text{and\,}\,\gamma(X,\Delta_{w})^{\mathcal{G}}=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}\,\text{if}\,w\neq 0.$
-The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ in [37, 38, 9] under a moment constraint, i.e.
+The formula for $\delta^{\mathrm{A}}(X,w)^{\mathcal{G}}$ in Theorem 1.3 yields a generalization of Aubins’ improvement of the Moser–Trudinger inequality to $w\neq 0$ with a quasi-sharp multiplicative constant.
+We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1).
+This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its center), appearing implicitly in [61, 81, 95] – see Section 4.2.
+Coming back to Conjecture 1.2for a general log Fano manifold, we introduce a slight modification of the log canonical threshold $\gamma^{(N)}(X,\Delta)^{\mathcal{G}},$ denoted by $\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}},$ obtained by replacing the semi-stable locus $(X^{N})_{\text{ss}}$ by a Zariski open subset $(X^{N})_{\text{ss},\epsilon_{N}}$ of $X^{N}$ containing slightly unstable configurations (see Definition 7.1).
+This leads to a notion of strong uniform Gibbs polystability, defined by replacing $(X^{N})_{\text{ss}}$ with $(X^{N})_{\text{ss},\epsilon_{N}}$ (which is expected to be equivalent to uniform Gibbs polystability, as illustrated by the case of log Fano curves; see Example 7.2).
+We establish the following quantitative lower bound on the reduced analytic stability threshold $\delta^{A}(X,\Delta)^{\mathcal{G}}:$
+Theorem 1.4.
+Let $(X,\Delta)$ be a log Fano manifold whose Futaki character vanishes. Then there exists a constant $C$ such that
+$(1-Ck^{-1})\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}}\leq\delta^{A}(X,\Delta)^{\mathcal{G}}.$
+Hence, if $(X,\Delta)$ is strongly uniformly Gibbs polystable, then $(X,\Delta)$ admits a Kähler–Einstein metric and, as consequence, $(X,\Delta)$ is K-polystable.
+The previous theorem generalizes [73, Cor 2.6], where it was shown – using purely algebro-geometric techniques – that $\gamma^{(N)}(X,\Delta)$ is bounded from above by Fujita–Odaka’s invariant $\delta_{k}(X,\Delta)$ converging to $\delta(X,\Delta)$ and, as a consequence,
+$\limsup_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)\leq\delta(X,\Delta)$
+(which can be made quantiative, using that $\delta_{k}(X,\Delta)(1-O(1/k))=\delta(X,\Delta)$, by [55]).
+However, it seems challenging to extend the algebro-geometric argument in [73] so as to incorporate the action of the group $\mathcal{G}.$ Instead, our proof of Theorem 1.4 adapts the analytic approach developed in [20], by proving an effective large deviation bound (Theorem 7.4), motivated by the conjectural Large Deviation Principle alluded to above.
+1.3. Higher dimensional examples
+Beyond complex curves, Gibbs stability has turned out to be notoriously difficult to establish, except for the case of exceptional log Fano manifolds, where it holds automatically at any level $k.$ In Section 2.5.1 we provide the first non-exceptional higher-dimensional examples of Gibbs stable log Fano varieties.
+More precisely, we show that the Fano orbifold defined as the quotient of the Fermat hypersurface $X_{n+1}$ of degree $(n+1)$ in $\mathbb{P}^{n+1},$ under the standard action of the group $(\mathbb{Z}/\mathbb{Z}_{n})^{n},$ is Gibbs stable at any level $k.$ This Fano orbifold may be identified with a log pair $(\mathbb{P}^{n},\Delta),$ where $\Delta$ is a divisor that assigns the same weight 11/(n+2)1-1/(n+2) to $n+2$ normally crossing hyperplanes in $\mathbb{P}^{n}.$ Although the group $\text{Aut}_{0}(\mathbb{P}^{n},\Delta)$ is trivial, the proof of the Gibbs stability of $(\mathbb{P}^{n},\Delta)$ exploits that the group $\text{Aut}_{0}(\mathbb{P}^{n})$ is non-trivial.
+Furthermore, in the companion paper [6] we exhibit log Fano manifolds $(\mathbb{P}^{2},\Delta)$ with non-trivial group $\text{Aut}_{0}(\mathbb{P}^{n},\Delta)$ that are Gibbs polystable at the minimal level (i.e.
+for the smallest $k\in]0,\infty[$ such that $kK_{(X,\Delta)}$ defines a line bundle):
+Theorem 1.5.
+[6] Let $\Delta_{w}$ be the torus-invariant divisor on $\mathbb{P}^{2}$ assigning the same weight $w\in[0,1[$ to each of the three torus-invariant hyperplanes.
+Then the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ is Gibbs polystable at the minimal level iff $w>3/4$ or $w=0$.
+The conclusion that $\mathbb{P}^{2}$ is Gibbs polystable at the minimal level, but $(\mathbb{P}^{2},\Delta_{w})$ is not Gibbs polystable for $w>0$ sufficiently small shows that the discontinuity phenomenon observed for $(\mathbb{P}^{1},\Delta_{w})$ in Theorem 1.3 persists in higher dimensions, at least for this example and at the minimal level.
+The case of $w=0$ in Theorem 1.5 is a special case of Proposition 2.16, stating that $\mathbb{P}^{n}$ is Gibbs polystable at the minimal level for any $n$.
+When w=11/mw=1-1/m for a positive integer $m,$ the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ may be identified with the Fano orbifold defined as the quotient of $\mathbb{P}^{2}$ under the standard action of $(\mathbb{Z}/m\mathbb{Z})^{2}.$ In this case the previous theorem applies to the AdS/CFT correspondence, where the Fano orbifold in question encodes a particular toric quiver gauge theory.
+1.4. A sharp constrained log HLS inequality on $S^{2}$ and application to quantitative stability
+Using the formula $\delta^{\mathrm{A}}(\mathbb{P}^{1},0)^{\mathcal{G}}=2$ in Theorem 1.3 we establish a strengthening of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ in [43, 44, 10] under a moment constraint, i.e.
 a strengthening of the following inequality for all probability measures $\mu$ on $S^{2}:$
 (1.9) $E(\mu)\leq D(\mu|\mu_{0}),$
-where $D(\mu|\mu_{0})$ denotes the entropy of $\mu,$ relative to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2},$ and $E(\mu)$ denotes the logarithmic energy of $\mu$ (explicitely expressed in formula 4.3).
-In fact, combining our formula for $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}$ with the uniqueness result for the Liouville equation on $S^{2}$ established in [72] yields:
+where $D(\mu|\mu_{0})$ denotes the entropy of $\mu,$ relative to the $\mathrm{SO}(3)$–invariant probability measure $\mu_{0}$ on $S^{2}$ (aka the Kullback–Leibler divergence) and $E(\mu)$ denotes the logarithmic energy of $\mu$ (explicitely expressed in formula 4.3).
+More precisely, combining our formula for $\delta^{\mathrm{A}}(\mathbb{P}^{1},0)^{\mathcal{G}}$ with the uniqueness result for the Liouville equation on $S^{2}$ established in [80] leads – via a somewhat miraculous optimization identity (formula 5.3), established in the course of the proof – to the following theorem:
 Theorem 1.6.
-Let $X=\mathbb{P}^{1},$ identified with the unit-sphere $S^{2}$ in $\mathbb{R}^{3},$ and denote by $\mathcal{M}$ the Mabuchi functional on the space of all Kähler forms $\omega$ in $c_{1}(X),$ normalized so that $\mathcal{M}(\omega)=0$ when $\omega$ is proportional to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2}.$ For any $\omega$ in $c_{1}(X),$
-$\mathcal{M}(\omega)\geq(\mathcal{I}-\mathcal{J})^{\mathcal{G}}(\omega)$
+Let $X=\mathbb{P}^{1},$ identified with the unit-sphere $S^{2}$ in $\mathbb{R}^{3},$ and denote by $\mathcal{M}$ the Mabuchi functional on the space of all Kähler forms $\omega$ in $c_{1}(X),$ normalized so that $\mathcal{M}(\omega)=0$ when $\omega$ is proportional to the $\mathrm{SO}(3)$–invariant probability measure $\mu_{0}$ on $S^{2}.$ For any $\omega$ in $c_{1}(X),$
+(1.10) $\mathcal{M}(\omega)\geq(\mathcal{I}-\mathcal{J})^{\mathcal{G}}(\omega)$
 with equality iff $\omega$ is proportional to $\mu_{0}$ modulo the action of $\mathcal{G}$ (i.e.
-modulo conformal diffeomorphisms of $S^{2}).$ Moreover, the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ may be replaced by $\mathcal{I}-\mathcal{J}$ if the Kähler potential of $\omega$ satisfies Tian’s orthogonality condition 3.31.
+modulo conformal diffeomorphisms of $S^{2}).$ Moreover, the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ may be replaced by $\mathcal{I}-\mathcal{J}$ if the Kähler potential of $\omega$ satisfies Tian’s orthogonality condition 3.30.
 More generally, for any probability measure $\mu$ on $S^{2},$ satisfying the moment constraint $\int_{S^{2}\subset\mathbb{R}^{3}}\boldsymbol{x}\mu=0,$
@@
 with equality iff $\mu=\mu_{0}.$
-The inequality for $\mathcal{M}$ in formula 1.7, for some (non-effective) positive constant $\epsilon$ and $C_{\epsilon}\in\mathbb{R}$ was first established in [48].
+The inequality for $\mathcal{M}$ in formula 1.7 on a general log Fano manifold for some (non-effective) positive constant $\epsilon$ and $C_{\epsilon}\in\mathbb{R}$ was first established in [61].
 Since $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ (by Theorem 1.3) the inequality for $\mathcal{M}$ fails on $\mathbb{P}^{1}$ when $\epsilon>1$ (for any $C_{\epsilon})$ and thus the multiplicative constants in Theorem 1.6 are sharp.
-More generally, the formula for $\delta^{\mathrm{A}}(X,w)^{\mathcal{G}}$ in Theorem 1.3 yields a quasi-sharp generalization of Aubins’ improvement of the Moser-Trudinger inequality to $w\neq 0.$ We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1).
-This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its center), appearing implicitly in [48, 63, 77] – see Section 4.2.
-Beyond complex curves, Gibbs stability has turned out to be notoriously difficult to establish, and at present no examples are known apart from the case of exceptional Fano orbifolds, where it holds automatically.
-But in the companion paper [5] we exhibit non-exceptional log Fano manifolds $(\mathbb{P}^{2},\Delta)$ that are Gibbs polystable already at the minimal level (i.e.
-for the smallest $k\in]0,\infty[$ such that $kK_{(X,\Delta)}$ defines a line bundle):
-Theorem 1.5.
-[5] Let $\Delta_{w}$ be the torus-invariant divisor on $\mathbb{P}^{2}$ assigning the same weight $w\in]0,1[$ to each of the three torus-invariant hyperplanes.
-Then the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ is Gibbs polystable at the minimal level iff $w>3/4.$
-The log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ may, when w=11/mw=1-1/m for a positive integer $m,$ be identified with the Fano orbifold defined as the quotient of $\mathbb{P}^{2}$ under the standard action of $(\mathbb{Z}/m\mathbb{Z})^{2}.$ In this case the previous theorem applies to the AdS/CFT correspondence, where the Fano orbifold in question encodes a particular toric quiver gauge theory.
+Exploiting that there is no additive constant in inequality 1.10 we deduce the following quantitative stability results for the sharp logarithmic HLS inequality 1.9 on $S^{2}$ with an optimal stability constant $1/2$ (see also Cor 5.7 for a more general formulation):
+Corollary 1.7.
+The following quantitative stability inequality for the sharp logarithmic HLS inequality on $S^{2}$ holds
+(1.12) $D(\mu|\mu_{0})-E(\mu)\geq\frac{1}{2}\inf_{\nu:\,D(\nu)=E(\nu)}D\left(\mu|\nu\right),$
+for any probability measure $\mu$ on $S^{2}$ such that $D(\mu|\mu_{0})<\infty.$ Moreover, the stability constant $1/2$ is optimal, i.e.
+the stability inequality above fails when $1/2$ is replaced by any strictly larger constant.
+The inf in the right hand side of the stability inequality 1.12 can be interpreted as an entropic “distance” (or more precisely divergence) between $\mu$ and the three-dimensional submanifold of $\mathcal{P}(S^{2})$ consisting of the optimizers of the sharp HLS inequality.
+As explained in Section 5.2 the previous corollary appears to be the first instance where the optimal stability constant has been computed explicitly, in the context of fractional Sobolev and HLS inequalities or their logarithmic versions.
+It improves, in particular, on the recent quantitative stability result in [47].
 1.5. Explicit symmetry breaking and the emergence of a unique Kähler–Einstein metric
-Next, we break the $\mathcal{G}$–symmetry explicitly by fixing a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and a $\mathcal{K}$–invariant metric on $K_{(X,\Delta)}$ with positive curvature.
-This metric defines a $\mathcal{K}$–invariant Kähler form $\omega_{0}$ in $c_{1}(-K_{(X,\Delta)})$ and thus a symplectic form on $X.$ The natural action of $\mathcal{K}$ on $K_{(X,\Delta)}$ gives rise to a moment map
+Let $(X,\Delta)$ be a log Fano manifold and fix a maximally compact subgroup $\mathcal{K}$ of the corresponding symmetry group $\mathcal{G}$ and a $\mathcal{K}$–invariant metric on $K_{(X,\Delta)}$ with positive curvature.
+The curvature defines a Kähler form $\omega_{0}$ in $c_{1}(-K_{(X,\Delta)})$ and thus a $\mathcal{K}$–invariant symplectic form on $X.$ As consequence, the natural action of $\mathcal{K}$ on $K_{(X,\Delta)}$ gives rise to a moment map
 (1.13) $\boldsymbol{m}\colon X\rightarrow\mathfrak{k}^{*}$
@@
 $\boldsymbol{m}_{N}(x_{1},\dots,x_{N})=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{m}(x_{i})\in\mathfrak{k}^{*}.$
-Given a positive integer $N$ such that 0 is a regular value of $\boldsymbol{m}_{N}$ we denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 6.1 for the general definition.
-The algebraic definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 6.11), i.e.
+Consider first the case when 0 is a regular value of $\boldsymbol{m}_{N}.$ We then denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the measure on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ induced by the measure $\nu^{(N)}$ on $X^{N}$ (defined by formula 1.2 when $\Delta$ is trivial) – see Section 6.2 for the general definition.
+The algebro-geometric definition of Gibbs polystability of $(X,\Delta)$, introduced in the previous section, is equivalent to the Gibbs semistability of $(X,\Delta)$ together with the analytic assumption that the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ has finite total mass (Prop 6.13), i.e.
 that
-(1.13) $\mu_{\boldsymbol{m}_{N}}^{(N)}\coloneqq\frac{1}{Z_{\boldsymbol{m}_{N}}}\nu_{\boldsymbol{m}_{N}}^{(N)},\,\,\,\,Z_{\boldsymbol{m}_{N}}\coloneqq\int\nu_{\boldsymbol{m}_{N}}^{(N)}$
-defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ is symmetric, i.e.
-invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant.
-The original $\mathcal{G}$–symmetry has thus been reduced to a $\mathcal{K}$–symmetry.
-Accordingly, the probability measure 1.13 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
-We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e.
-$\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.13.
-Before stating our two main conjectures we note that in contrast to the $\mathcal{K}$–reduced probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)},$ the measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ depends on a choice of bases in $H^{0}(X,-kK_{(X,\Delta)}).$ As a consequence, so does its total mass $Z_{\boldsymbol{m}_{N}}.$ In what follows we will take the basis to be orthonormal with respect to the $L^{2}$–norm on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the fixed metric on $K_{(X,\Delta)}.$ The corresponding constant $Z_{\boldsymbol{m}_{N}}$ shall then be called the $\mathcal{K}$-reduced partition function, in view of its statistical mechanical interpretation described in Section 6.1.
+(1.15) $\mu_{0}^{(N)}\coloneqq\frac{1}{Z_{N,0}}\nu_{0}^{(N)},\,\,\,\,Z_{N,0}\coloneqq\int\nu_{0}^{(N)}$
+defines a probability measure on $X^{N},$ supported on $\left\{\boldsymbol{m}_{N}=0\right\}.$ The probability measure $\mu_{0}^{(N)}$ is symmetric, i.e.
+invariant under permutations of $X^{N}$ and $\mathcal{K}$–invariant, but not $\mathcal{G}$-invariant.
+This is an instance of explicit symmetry breaking, as the original $\mathcal{G}$–symmetry has been reduced to a $\mathcal{K}$–symmetry.
+Accordingly, the probability measure 1.15 on $X^{N}$ shall be called the the $\mathcal{K}$–reduced canonical probability measure.
+We can now view the empirical measure $\delta_{N}$ (formula 1.5) as a random variable on the probability space $\left(X^{N},\mu_{0}^{(N)}\right).$ Compared to the canonical probability measure 1.4 on $X^{N}$, this means that we are conditioning the $N$ points on $X^{N}$ to have vanishing moment map $\boldsymbol{m}_{N}.$ The empirical measure $\delta_{N}$ of any such configuration of points has vanishing moment, i.e.
+$\delta_{N}$ maps $X^{N}$ into the space $\mathcal{P}(X)_{0}$ defined in formula 1.15.
+Before stating our two main conjectures we note that in contrast to the $\mathcal{K}$–reduced probability measure $\mu_{0}^{(N)},$ the measure $\nu_{0}^{(N)}$ depends on a choice of bases in $H^{0}(X,-kK_{(X,\Delta)}).$ As a consequence, so does its total mass $Z_{N,0}.$ In what follows we will take the basis to be orthonormal with respect to the $L^{2}$–norm on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the fixed metric on $K_{(X,\Delta)}.$ The corresponding constant $Z_{N,0}$ shall then be called the $\mathcal{K}$-reduced partition function, in view of its statistical mechanical interpretation described in Section 6.2.
 The fixed metric on $K_{(X,\Delta)}$ also determines the Mabuchi functional $\mathcal{M}$ on the space of Kähler metrics in $c_{1}(-K_{(X,\Delta)})$ (which, a priori, is only defined up to an additive constant).
@@
 Let $(X,\Delta)$ be a log Fano manifold and assume that $(X,\Delta)$ admits a Kähler–Einstein metric. Then
-(1.14) $\text{(i)}-\lim_{N\rightarrow\infty}N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}.$
-As a consequence, $Z_{\boldsymbol{m}_{N}}<\infty$ for $N$ sufficiently large and viewing the empirical measure $\delta_{N}$ as a random variable on $\left(X^{N},\mu_{\boldsymbol{m}_{N}}^{(N)}\right),$
+(1.16) $\text{(i)}-\lim_{N\rightarrow\infty}N^{-1}\log Z_{N,0}=\inf\mathcal{M}.$
+As a consequence, $Z_{N,0}<\infty$ for $N$ sufficiently large and viewing the empirical measure $\delta_{N}$ as a random variable on $\left(X^{N},\mu_{0}^{(N)}\right),$
 (1.17) $\text{(ii) }\lim_{N\rightarrow\infty}\delta_{N}=\mu_{\text{KE}},$
@@
 This conjecture is motivated by a heuristic mean field type argument, outlined in Section 6.3, suggesting the following Large Deviation Principle (LDP) on the space $\mathcal{P}(X)_{0},$ as $N\rightarrow\infty$:
-(1.16) $(\delta_{N})_{*}\nu_{\boldsymbol{m}_{N}}^{(N)}\sim e^{-NF(\mu)},$
+(1.18) $(\delta_{N})_{*}\nu_{0}^{(N)}\sim e^{-NF(\mu)},$
 where $F$ is the free energy functional, introduced in [11], that may be identified with the Mabuchi functional on the space of all Kähler metrics $\omega$ in $c_{1}(-K_{(X,\Delta)})$ via the relation $F(\omega^{n}/V)=\mathcal{M}(\omega)$.
-We show that the validity of such a LDP would imply both the asymptotics 1.14 and the convergence 1.15, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$.
+We show that the validity of such a LDP would imply both the asymptotics 1.16 and the convergence 1.17, using that $\mu_{\text{KE}}$ is the unique minimizer of $F(\mu)$ on $\mathcal{P}(X)_{0}$.
 Moreover, by introducing temperature into the picture the corresponding LDP would also imply Conjecture 1.2, using the observation that the infimum over $\mathcal{G}$ in formula 1.8 is attained at the unique Kähler metric in a given $\mathcal{G}$–orbit whose volume form has vanishing moment (see Prop 3.4).
-Our next main result establishes the first part of the previous conjecture when $n=1$ with an explicit rate of convergence:
+Our next main result establishes the first part of the previous conjecture on log Fano curves with an explicit rate of convergence:
 Theorem 1.9.
-The convergence 1.14 holds on any K-polystable log Fano curve $(X,\Delta)$ (i.e. when $n=1$), when $N$ is taken as the sequence of odd positive integers. More precisely,
-$N^{-1}\log Z_{\boldsymbol{m}_{N}}=\inf\mathcal{M}+O\left(\frac{\log N}{N}\right).$
-As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.15) we will show, in the companion paper [4], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 6.2).
+The convergence 1.16 holds on any K-polystable log Fano curve $(X,\Delta)$ when $N$ is taken as the sequence of odd positive integers. More precisely,
+$N^{-1}\log Z_{N,0}=\inf\mathcal{M}+O\left(\frac{\log N}{N}\right).$
+As for the conjectured convergence towards $\mu_{\text{KE}}$ (formula 1.17) we will show, in the companion paper [5], that it holds up to an arbitrarily small “temperature deformation”, naturally appearing in the context of Onsager’s point vortex model on the two-sphere (see Remark 6.5).
+The assumption that $N$ be odd in the previous theorem is equivalent to 0 being a regular value of $\boldsymbol{m}_{N}.$ In particular, this ensures that the level set $\left\{\boldsymbol{m}_{N}=0\right\}$ is a smooth submanifold of $X^{N}.$ To bypass this assumption one can, for a general log Fano manifold $(X,\Delta),$ replace $\left\{\boldsymbol{m}_{N}=0\right\}$ with a neighbourhood $\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon_{N}\right\}$ for an appropriate sequence $\epsilon_{N}$ tending to zero.
+The assumption that $(X,\Delta)$ be strongly uniformly Gibbs polystable in Theorem 1.4 implies that the corresponding normalizing constant $Z_{N,\epsilon_{N}}$ is finite, when $N\gg 1$ and $\epsilon_{N}\gg N^{-1/2}.$ Under this assumption we establish in Theorem 7.4 an effective large deviation lower bound on $Z_{N,\epsilon_{N}},$ which implies both Theorem 1.4 and the bound
+$\limsup_{N\rightarrow\infty}-N^{-1}\log Z_{N,\epsilon_{N}}\leq\inf\mathcal{M},$
+assuming that the Futaki character of $(X,\Delta)$ vanishes.
 1.6. Outline of the proof of Theorem 1.9 and connections to arithmetic geometry
@@
 (1.19) $\int_{(X^{N})_{\mathrm{ss}}/\mathcal{G}}\nu_{\mathbf{m}_{N}}^{(N),\mathrm{cc}},$
-where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.17 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [60].
+where $\nu_{\boldsymbol{m}_{N}}^{(N),\mathrm{cc}}$ denotes the $\mathcal{G}$–invariant form on $X^{N}$ of degree $2(N-\text{rank}\mathcal{G})$ obtained by contracting $\nu_{0}^{(N)}$ with a generator of the top exterior power of the complex Lie algebra $\mathfrak{g}$ of $\mathcal{G}.$ The integral 1.19 is a complex–algebraic period and we show that it can be expressed as a Selberg integral over the field $\mathbb{C}$ [76].
 Such integrals can be computed explicitly in terms of Gamma values, as first shown in [69] in the context of Conformal Field Theory, and letting $N\rightarrow\infty$ allows us to deduce Theorem 1.9, by exploiting some connections to K-stable arithmetic log Fano curves in [4].
-It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.17 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 8, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [3], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [3], is a logarithmic generalization of Odaka’s modular height [86].
+It should be stressed that, in the general setup introduced above, the sequence of forms $\nu^{(N),\mathrm{cc}}$ appearing in the integral 1.19 is not canonically attached to the given log Fano manifold $(X,\Delta),$ as it depends on the choice of a bases in $H^{0}(X,-kK_{(X,\Delta)})$ and $\mathfrak{g}.$ But, as explained in Section 9, when $(X,\Delta)$ is defined over the integers – or more precisely when $(X,\Delta)$ is the complexification of an arithmetic log Fano variety $(\mathcal{X},\mathcal{D}$) – the arithmetic structure singles out a canonical normalization of $\nu^{(N),\mathrm{cc}}$ and thus a corresponding invariant $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D}).$ Generalizing the K-stable case considered in [4], we conjecture that if $(X,\Delta)$ is K-polystable and the quantized Futaki characters of $(X,\Delta)$ vanish, then $N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})$ converges to the infimum of the arithmetic Mabuchi functional of $(\mathcal{X},\mathcal{D}).$ The latter functional, studied in [4], is a logarithmic generalization of Odaka’s modular height [104].
 The proof of Theorem 1.9 reveals that this arithmetic conjecture holds for all K-polystable log Fano curves $(X,\Delta)$ defined over the integers.
@@
 $\delta(X,\Delta)_{\text{ps}\text{ }}\coloneqq\inf_{\mu\in\mathcal{P}(X^{\text{div}})_{\text{ss}}}\frac{\int_{X^{\text{div}}}A_{(X,\Delta)}\mu}{\left\|\mu\right\|},$
-where $X^{\text{div}}$ denotes the space of all divisorial valuations $v$ over $X,$ $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ (formula 2.6), $\left\|\mu\right\|$ denotes the non-Archimedean pluricomplex energy of a probability measure $\mu$ on $X^{\text{div}}$ with respect to $K_{(X,\Delta)}$ [35] and $\mathcal{P}(X^{\text{div}})_{\text{ss}}$ denotes the space of all probability measures $\mu$ on $X^{\text{div}}$ of the following form (where $N$ is the sequence defined by formula 1.6):
+where $X^{\text{div}}$ denotes the space of all divisorial valuations $v$ over $X,$ $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ (formula 2.8), $\left\|\mu\right\|$ denotes the non-Archimedean pluricomplex energy of a probability measure $\mu$ on $X^{\text{div}}$ with respect to $K_{(X,\Delta)}$ [41] and $\mathcal{P}(X^{\text{div}})_{\text{ss}}$ denotes the space of all probability measures $\mu$ on $X^{\text{div}}$ of the following form (where $N$ is the sequence defined by formula 1.6):
 $\mu=\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{j=1}^{N}\delta_{v_{j}^{(N)}},$
@@
   • $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}=\delta(X,\Delta)_{\text{ps }}$
-The convergence in the second item above is motivated by the approach for proving the convergence of $\gamma^{(N)}(X,\Delta)$ towards $\delta(X,\Delta)_{\text{ }}$ outlined in [17, Section 5], using non-Archimedean pluripotential theory.
+The convergence in the second item above is motivated by the approach for proving the convergence of $\gamma^{(N)}(X,\Delta)$ towards $\delta(X,\Delta)_{\text{ }}$ outlined in [18, Section 5], using non-Archimedean pluripotential theory.
 The validity of the third item would imply the first item, by [61].
@@
 1.8. Acknowledgements
-This work was supported by a Wallenberg Scholar grant from the Knut and Alice Wallenberg foundation.
+Thanks to Sébastien Boucksom and Mattias Jonsson for discussions. This work was supported by a Wallenberg Scholar grant from the Knut and Alice Wallenberg foundation.
 1.9. Organization of the paper
@@
 The starting point of the proof of Theorem 3.13 is a thermodynamical formula for $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}.$ This formula is established on any log Fano manifold in Section 3.3, where the structure of the space $\mathcal{P}(X)_{0}$ of all probability measures on $X$ with vanishing moments is studied and new analytic characterizations of log Fano manifolds admitting Kähler–Einstein metrics are given in Theorem 3.7.
-  • In Section 4 applications of Theorem 3.13 to constrained logarithmic Hardy–Littlewood–Sobolev inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
-  • In Section 5 we establish a sharp version of the constrained logarithmic Hardy–Littlewood–Sobolev inequalities on the two-sphere, by combining Theorem 3.13 with the uniqueness result for the Liouville equation on $S^{2}$ established in [72].
-  • In Section 6 we study the $\mathcal{K}$–reduced probability measures $\mu_{\boldsymbol{m}_{N}}^{(N)}$ on $X^{N}$ and the corresponding partition function $Z_{\boldsymbol{m}_{N}}$ appearing in Conjecture 1.6.
+  • In Section 4 applications of Theorem 3.13 to constrained logarithmic Hardy–Littlewood–Sobolev (HLS) inequalities and Moser–Trudinger inequalities are given and a phenomenon of spontaneous symmetric breaking is exhibited.
+  • In Section 5 we establish a sharp version of the constrained logarithmic HLS inequalities on the two-sphere $S^{2},$ by combining Theorem 3.13 with the uniqueness result for the Liouville equation on $S^{2}$ established in [80].
+Applications to quantiative stability results for the logarithmic HLS inequality on $S^{2}$ are then given.
+  • In Section 6 we study the $\mathcal{K}$–reduced probability measures $\mu_{0}^{(N)}$ on $X^{N}$ and the corresponding partition function $Z_{N,0}$ appearing in Conjecture 1.8.
 The asymptotics as $N\rightarrow\infty$ are described in a overarching conjectural Large Deviation Principle, formulated in Conjecture 6.7, that we show implies both Conjecture 1.1 and Conjecture 1.2.
-  • In Section 7 Theorem 1.7 is established, showing that $N^{-1}\log Z_{\boldsymbol{m}_{N}}$ converges towards the inf of the Mabuchi functional on any log Fano curve with vanishing Futaki invariant.
+  • In Section 7 we prove an effective large deviation bound for the thickened partition functions that implies Theorem 1.4
+  • In Section 8 Theorem 1.9 is established, showing that $N^{-1}\log Z_{N,0}$ converges towards the inf of the Mabuchi functional on any log Fano curve with vanishing Futaki invariant.
   • In Section 9 relations to arithmetic geometry are described
@@
 We start by describing in more detail the general setup, introduced in Section 2.
-Let $(X,\Delta)$ be a log Fano manifold, i.e.
+Let $(X,\Delta)$ be an $n$-dimensional log Fano manifold, i.e.
 $X$ is a complex projective manifold and $\Delta$ is an effective $\mathbb{Q}$–divisor on $X$ such that the log anticanonical line bundle $K_{(X,\Delta)}$ of $(X,\Delta)$ is ample.
@@
 We will be concerned with the effective $\mathbb{Q}$–divisor $\mathcal{D}^{(N)}$ on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ whose support consists of all configurations $(x_{1},\dots,x_{N})$ of $N$ points on $X$ which are in “bad position” with respect to $H^{0}(X,-kK_{(X,\Delta)})$:
-$\text{Supp}\left(\mathcal{D}^{(N)}\right)\coloneqq\left\{(x_{1},\dots,x_{N})\colon\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})\right\}$
+(2.1) $\text{Supp}\left(\mathcal{D}^{(N)}\right)\coloneqq\left\{(x_{1},\dots,x_{N})\colon\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})\right\}$
 Concretely, $k\mathcal{D}^{(N)}$ is the divisor on $X^{N}$ cut out by the section $\det S^{(N)}$ of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ defined by
@@
 Anyhow, in the definitions we will, for simplicity, demand that $N$ only ranges over all $N_{k}$ for which $k$ is a positive integer and $k\Delta$ is a divisor.
+It follows directly from the definitions that
+(2.5) $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}\geq\gamma^{(N)}(X,\Delta)\geq\alpha(X,\Delta):=\inf_{D\sim-K_{(X,\Delta)}}\text{lct}\left(X,\Delta;D\right),$
+where $D$ ranges over all effective $\mathbb{Q}$-divisors $D$ linearly equivalent to $K_{(X,\Delta)}$.
+The invariant $\alpha(X,\Delta)$ coincides with Tian’s alpha invariant of $(X,\Delta),$ which is strictly positive for any log Fano manifold $(X,\Delta).$
 2.2. Preliminaries
@@
 where $s_{D}$ is a (multivalued) holomorphic section cutting out $D$ and $||\cdot||_{D}$, $||\cdot||_{\Delta}$ denotes any fixed smooth metrics on the line bundles corresponding to $D$ and $\Delta$, respectively.
-In order to calculate the log canonical thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ we will instead use the well-known valuative definition of $\mathrm{lct}(X,\Delta;D)$ (see, for example, [25, Thm B.5]):
-(2.5) $\mathrm{lct}(X,\Delta;D)=\inf_{v}\mathrm{lct}(X,\Delta;D)_{v},\,\,\,\,\mathrm{lct}_{v}(X,\Delta;D)\coloneqq\frac{A_{(X,\Delta)}(v)}{v(D)},$
+In order to calculate the log canonical thresholds $\gamma^{(N)}(X,\Delta)$ and $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$ we will instead use the well-known valuative definition of $\mathrm{lct}(X,\Delta;D)$ (see, for example, [27, Thm B.5]):
+(2.7) $\mathrm{lct}(X,\Delta;D)=\inf_{v}\mathrm{lct}(X,\Delta;D)_{v},\,\,\,\,\mathrm{lct}(X,\Delta;D)_{v}\coloneqq\frac{A_{(X,\Delta)}(v)}{v(D)},$
 where the inf ranges over all divisorial valuations $v$ over $X$ and $A_{(X,\Delta)}(v)$ denotes the log discrepancy of $v$ with respect to $(X,\Delta).$ We recall that a divisorial valuation $v$ over $X$ is an integer-valued function on the space $\mathbb{C}(X)$ of all rational functions $f$ on $X$ of the form
@@
 The log discrepancy $A_{(X,\Delta)}(v)$ of a divisorial valuation $v$ is defined as
-(2.6) $A_{(X,\Delta)}(v)=1+\mathrm{ord}_{E}(\pi^{*}K_{(X,D)}-K_{Y}),$
-where $\pi^{*}K_{(X,D)}-K_{Y}$ defines a divisor, supported on the exceptional divisors of $\pi.$
+(2.8) $A_{(X,\Delta)}(v)=1+\mathrm{ord}_{E}(K_{Y}-\pi^{*}K_{(X,D)}),$
+where $K_{Y}-\pi^{*}K_{(X,D)}$ defines a divisor, supported on the exceptional divisors of $\pi.$
 Example 2.4.
@@
 a birational morphism that is an isomorphism over the complement of $\text{supp}(D+\Delta)$ such that $\pi^{*}(D+\Delta)$ is a divisor on $Y$ with simple normal crossings.
-Denoting by $E_{1},\dots,E_{m}$ its components, the inf in 2.5 can be restricted to the divisorial valuations defined by $E_{1},\dots,E_{m}$.
-Using the existence of log resolutions it then follows readily that the two definitions 2.4 and 2.5 of the log canonical threshold agree.
+Denoting by $E_{1},\dots,E_{m}$ its components, the inf in 2.7 can be restricted to the divisorial valuations defined by $E_{1},\dots,E_{m}$.
+Using the existence of log resolutions it then follows readily that the two definitions 2.6 and 2.7 of the log canonical threshold agree.
 We will use the following basic product property: if $(X_{1},\Delta_{1})$ and $(X_{2},\Delta_{2})$ are log pairs, and $D_{1},D_{2}$ are effective $\mathbb{Q}$–divisors on $X_{1}$ and $X_{2}$ respectively, then
@@
 […]
-Next, if $U$ is an open subset in the Euclidean topology of $X$, then $\mathrm{lct}(U,\Delta|_{U},D|_{U})$ designates the inf 2.5 restricted to valuations whose center have non-zero intersection with $U$.
+Next, if $U$ is an open subset in the Euclidean topology of $X$, then $\mathrm{lct}(U,\Delta|_{U},D|_{U})$ designates the inf 2.7 restricted to valuations whose center have non-zero intersection with $U$.
 With this definition, analogous versions of the above properties hold, and moreover, if $X=\cup_{i}U_{i}$, then clearly,
 $\mathrm{lct}(X,\Delta;D)=\inf_{i}\mathrm{lct}(U_{i},\Delta|_{U_{i}},D|_{U_{i}}).$
-A divisorial valuation $v$ over $X$ is said to compute $\mathrm{lct}(X,\Delta;D)$ if the inf 2.5 is attained at $v.$ More generally, when considering a sequence $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ a sequence $v_{N}$ of valuations over $X_{N}$ will be said to asymptotically compute $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ if $\text{lct}_{v_{N}}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right)$ converges towards $\liminf_{N\rightarrow\infty}\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),$ as $N\rightarrow\infty.$ Finally, note that any $N$-tuple of divisorial valuations $v_{1},\dots,v_{N}$ over $X$ induces a valuation over $X^{N}$ by blowing up the variety $E_{1}\times\cdots\times E_{N}$ over $X^{N},$ where $E_{i}$ is the divisor over $X$ corresponding to $v_{i}.$ Such a divisorial valuation over $X^{N}$ will be called decomposable.
+A divisorial valuation $v$ over $X$ is said to compute $\mathrm{lct}(X,\Delta;D)$ if the inf 2.7 is attained at $v.$ More generally, when considering a sequence $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ a sequence $v_{N}$ of valuations over $X_{N}$ will be said to asymptotically compute $\mathrm{lct}(X_{N},\Delta_{N};D_{N})$ if $\text{lct}_{v_{N}}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right)$ converges towards $\liminf_{N\rightarrow\infty}\text{lct}\left(X^{N},\Delta_{N};\mathcal{D}^{(N)}\right),$ as $N\rightarrow\infty.$ Finally, note that any $N$-tuple of divisorial valuations $v_{1},\dots,v_{N}$ over $X$ induces a valuation over $X^{N}$ by blowing up the variety $E_{1}\times\cdots\times E_{N}$ over $X^{N},$ where $E_{i}$ is the divisor over $X$ corresponding to $v_{i}.$ Such a divisorial valuation over $X^{N}$ will be called decomposable.
 2.2.2. Semistability in GIT
 Let $L\rightarrow Y$ be an ample line bundle over a complex projective manifold $Y$ and $G$ a complex reductive group acting on $(Y,L).$ The corresponding semistable locus (in sense of Geometric Invariant Theory) is the Zariski open subset $Y_{\text{ss}}$ of $Y$ consisting of all $y\in Y$ such that there exists a GG–invariant section $s_{k}\in H^{0}(Y,kL)$ for some positive integer $k$ such that $s_{k}(y)\neq 0$ [83].
-When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathrm{PGL}(2,\mathbb{C}),$ acting diagonally, $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the the fraction $I/N$ of points $x_{i}$ coinciding satisfies $I/N\leq 1/2$, i.e.
+More generally, given an ample line bundle $L\rightarrow Y$ with an action of $G$ on $(X,mL)$ for some positive integer $m,$ the semistable locus $Y_{\text{ss }}$ is defined wrt $(X,mL).$ When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathrm{PGL}(2,\mathbb{C}),$ acting diagonally, $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the the fraction $I/N$ of points $x_{i}$ coinciding satisfies $I/N\leq 1/2$, i.e.
 (2.10) $\left(\frac{1}{N}\sum_{i=1}^{N}\delta_{x_{i}}\right)\{x\}\leq 1/2$
-for all $x\in\mathbb{P}^{1}$ (see [85, Section 3]).
-When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathbb{C}^{*}$ the semistable locus $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the inequality 2.8 holds for the two fixed points $x$ of the action of $\mathbb{C}^{*}$ on $\mathbb{P}^{1}.$ Just like the case $G=\mathrm{PGL}(2,\mathbb{C}$) this follow readily from the Hilbert–Mumford criterion for one-parameter subgroups of $\mathbb{C}^{*}\subset G.$
+for all $x\in\mathbb{P}^{1}$ (see [103, Section 3]).
+When $Y=(\mathbb{P}^{1})^{N}$ and $G=\mathbb{C}^{*}$ the semistable locus $(X^{N})_{\text{ss}}$ consists of all $(x_{1},\dots,x_{N})$ such that the inequality 2.10 holds for the two fixed points $x$ of the action of $\mathbb{C}^{*}$ on $\mathbb{P}^{1}.$ Just like the case $G=\mathrm{PGL}(2,\mathbb{C}$) this follow readily from the Hilbert–Mumford criterion for one-parameter subgroups of $\mathbb{C}^{*}\subset G.$
 2.2.3. The Fulton–MacPherson compactification of configuration spaces
@@
   1. (1)
-    $X^{[N]}$ is a non-singular variety, projective if $X$ is [58, Theorems 1 and 2].
+    $X^{[N]}$ is a non-singular variety, projective if $X$ is [74, Theorems 1 and 2].
   2. (2)
-    There is a canonical morphism $\pi_{\mathrm{FM}}\colon X^{[N]}\rightarrow X^{N}$, which restricts to an isomorphism $\pi_{\mathrm{FM}}\colon X^{[N]}\setminus Z\rightarrow X^{N}\setminus\cup_{i\neq j}\{x_{i}=x_{j}\},$ where $Z$ defines a simple normal crossings divisor [58, Theorem 3].
+    There is a canonical morphism $\pi_{\mathrm{FM}}\colon X^{[N]}\rightarrow X^{N}$, which restricts to an isomorphism $\pi_{\mathrm{FM}}\colon X^{[N]}\setminus Z\rightarrow X^{N}\setminus\cup_{i\neq j}\{x_{i}=x_{j}\},$ where $Z$ defines a simple normal crossings divisor [74, Theorem 3].
   3. (3)
@@
     $\pi_{FM}^{*}(\mathcal{I}(W))=\mathcal{I}\left(\sum_{W^{\prime}\subset W}E_{W^{\prime}}\right)$
-    where $\mathcal{I}(\cdot)$ denotes the corresponding ideal sheaves and the sum ranges over all $W^{\prime}$ in $\mathcal{W}$ contained in $E$ [58, Theorem 3].
+    where $\mathcal{I}(\cdot)$ denotes the corresponding ideal sheaves and the sum ranges over all $W^{\prime}$ in $\mathcal{W}$ contained in $E$ [74, Theorem 3].
 Note that when $\mathrm{dim(X)=1}$, the first two properties imply that the map $\pi_{FM}$ is a log resolution of the log pair $(X^{N},\sum_{i<j}\{x_{i}=x_{j}\})$.
@@
 (2.11) $f_{*}^{-1}(E_{\pi_{W}})=E_{W^{\prime}},\,\,\,W^{\prime}\subset W$
-Indeed, $\pi_{W}^{*}(\mathcal{I}(W))=\mathcal{I}(E_{\pi_{W}})$ and thus by the commutativity of the diagram, $f^{*}(\mathcal{I}(E_{\pi_{W}}))=f^{*}\circ\pi_{W}^{*}(\mathcal{I}(W))=\pi_{FM}(\mathcal{I}(E_{\pi_{W}})),$ which, by property 3 of the Fulton–MacPherson compactification above, coincides with the ideal defined by the divisor $\sum_{W^{\prime}\subset W}E_{W^{\prime}}.$ Since the $E_{W}$’s are prime this proves the identity 2.9.
-It follows from 2.9 that
+Indeed, $\pi_{W}^{*}(\mathcal{I}(W))=\mathcal{I}(E_{\pi_{W}})$ and thus by the commutativity of the diagram, $f^{*}(\mathcal{I}(E_{\pi_{W}}))=f^{*}\circ\pi_{W}^{*}(\mathcal{I}(W))=\pi_{FM}(\mathcal{I}(E_{\pi_{W}})),$ which, by property 3 of the Fulton–MacPherson compactification above, coincides with the ideal defined by the divisor $\sum_{W^{\prime}\subset W}E_{W^{\prime}}.$ Since the $E_{W}$’s are prime this proves the identity 2.11.
+It follows from 2.11 that
 $\pi_{FM}(E_{W^{\prime}})=W$
@@
 Furthermore, $\mathcal{D}^{(N)}|_{U_{q}}\eqqcolon\mathcal{D}_{k_{1}}\boxplus\cdots\boxplus\mathcal{D}_{k_{l}}\times\mathcal{D}_{r}$.
-By 2.7
+By 2.9
 $\mathrm{lct}(U_{q},\Delta_{N}|_{U_{q}};\mathcal{D}^{(N)}|_{U_{q}})=\min\left(\min_{1\leq j\leq l}\left(\mathrm{lct}\left(\left.V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}}\right|_{V_{j}};\left.\mathcal{D}_{k_{j}}\right|_{V_{j}}\right)\right),\mathrm{lct}\left((\mathbb{P}^{1})^{r};\mathcal{D}_{r}\right)\right).$
 But for any $1\leq j\leq l$, $\pi^{*}(\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}})\leq\Delta_{N}^{p_{i_{j}},w_{i_{j}}}$, and $\pi^{*}(\mathcal{D}_{k_{j}})\leq\mathcal{D}^{(N)}$ where $\pi\colon(\mathbb{P}^{1})^{N}\rightarrow(\mathbb{P}^{1})^{k_{j}}$ is the projection and similarly $\pi^{*}(\mathcal{D}_{r})\leq\mathcal{D}^{(N)}$.
-In view of the definition 2.5 and 2.7, $\mathrm{lct}(V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}_{k_{j}})\geq\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ with $q^{\prime}=(p_{i_{j}},\dots,p_{i_{j}})$ but the numbers $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ are contenders in the inf 2.13.
-Moreover $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})=\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}^{(N)})$ and any such term is lesser or equal to $\mathrm{lct}(U,\Delta_{N}|_{U};\mathcal{D}^{(N)}|_{U})$ in 2.14, and also $\mathrm{lct}((\mathbb{P}^{1})^{r};\mathcal{D}_{r})$.
-This proves formula 2.12.
+In view of the definition 2.7 and 2.9, $\mathrm{lct}(V_{j},\Delta_{k_{i}}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}_{k_{j}})\geq\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ with $q^{\prime}=(p_{i_{j}},\dots,p_{i_{j}})$ but the numbers $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})$ are contenders in the inf 2.15.
+Moreover $\mathrm{lct}(U_{q^{\prime}},\Delta_{N}|_{U_{q^{\prime}}};\mathcal{D}^{(N)}|_{U_{q^{\prime}}})=\mathrm{lct}((\mathbb{P}^{1})^{N},\Delta_{N}^{p_{i_{j}},w_{i_{j}}};\mathcal{D}^{(N)})$ and any such term is lesser or equal to $\mathrm{lct}(U,\Delta_{N}|_{U};\mathcal{D}^{(N)}|_{U})$ in 2.16, and also $\mathrm{lct}((\mathbb{P}^{1})^{r};\mathcal{D}_{r})$.
+This proves formula 2.14.
 Step 2: The following formula holds
@@
 (2.17) $\left(\prod_{i<j}(z_{i}-z_{j})\right)^{-1/k}\text{and\,\,}\left(\prod_{i}z_{i}\right)^{w}.$
-Now regard these as functions of $N+1$ variables $z_{1},\dots,z_{N}$ and $z_{N+1}$, which are independent of $z_{N+1}.$ In other words, we consider their pullbacks to $\mathbb{C}^{N+1}$ under the natural projection from $\mathbb{C}^{N+1}$ to $\mathbb{C}^{N}.$ This does not affect the computation of the corresponding log canonical threshold, by 2.7.
-After making the affine change of variables $z_{i}=u_{i}-u_{N+1}$ for $i=1,\dots,N$ and $z_{N+1}=u_{N+1}$ the two expressions 2.15 become
+Now regard these as functions of $N+1$ variables $z_{1},\dots,z_{N}$ and $z_{N+1}$, which are independent of $z_{N+1}.$ In other words, we consider their pullbacks to $\mathbb{C}^{N+1}$ under the natural projection from $\mathbb{C}^{N+1}$ to $\mathbb{C}^{N}.$ This does not affect the computation of the corresponding log canonical threshold, by 2.9.
+After making the affine change of variables $z_{i}=u_{i}-u_{N+1}$ for $i=1,\dots,N$ and $z_{N+1}=u_{N+1}$ the two expressions 2.17 become
 (2.18) $\left(\prod_{1\leq i<j\leq N}(u_{i}-u_{j})\right)^{-1/k}\text{and\,\,}\left(\prod_{1\leq i\leq N}(u_{i}-u_{N+1})\right)^{w}.$
@@
 which proves the claim when $\frac{2(N-1)}{VN}\leq\min_{l}\frac{2(1-w_{l})}{V}.$ The other case follows similarly by blowing up along a subvariety of the form $\{x_{1}=\dots=x_{N}=p_{l}\}$. ∎
-The formula for $\gamma^{(N)}(X,\Delta)$ in the previous proposition corrects the formula given in [18, Thm 4.5], which is only correct modulo an error term of order $O(1/N)$ (that comes from the discrepancy between the upper and the lower bound on $\gamma^{(N)}(X,\Delta)$ established by analytic means in [18]).
+The formula for $\gamma^{(N)}(X,\Delta)$ in the previous proposition corrects the formula given in [19, Thm 4.5], which is only correct modulo an error term of order $O(1/N)$ (that comes from the discrepancy between the upper and the lower bound on $\gamma^{(N)}(X,\Delta)$ established by analytic means in [19]).
 2.3.3. The case of non-trivial $\mathcal{G}$
@@
 It was previously known that Gibbs semistability was equivalent to K-semistability and that (uniform) Gibbs stability was equivalent to (uniform) K-stability (see [19]).
+2.5. Higher dimensional examples
+2.5.1. Gibbs stable orbifolds that are not exceptional
+Following [34, 33], a log Fano variety $(X,\Delta)$ is said to be exceptional if $\text{lct}(X,\Delta;D)>1$ for any $\mathbb{Q}$-divisor $D$ linearly equivalent to $K_{(X,\Delta)}.$ By [33, Thm 1.7], this is equivalent to $\alpha(X,\Delta)>1,$ where $\alpha(X,\Delta)$ denotes Tian’s alpha invariant, whose algebro-geometric definition appears in formula 2.5.
+If $(X,\Delta)$ is exceptional, then $(X,\Delta)$ is Gibbs stable at any level $k.$ In fact, $(X,\Delta)$ is even uniformly Gibbs stable, by the lower bound 2.5.
+Consider now the log pair $(X,\Delta)=(\mathbb{P}^{n},\Delta_{w}),$ where
+$\Delta_{w}:=w(H_{1}+\cdots+H_{n+2}),\,\,\,w\in[0,1[$
+for $(n+2)$ given hyperplanes $H_{i}$ in $\mathbb{P}^{n}$ with normal crossings.
+The log pair $(\mathbb{P}^{n},\Delta_{w})$ is Fano iff w<11/(n+2)w<1-1/(n+2) and exceptional iff w>11/n+1;w>1-1/n+1; a simple proof is provided in Remark 2.15 below.
+Example 2.12.
+When w=11/mw=1-1/m for $m\in\mathbb{Z}_{+}$ the corresponding orbifold $(\mathbb{P}^{n},\Delta_{w})$ is the global orbifold quotient of the Fermat hypersurface $X_{m}$ of degree $m$ in $\mathbb{P}^{n+1}$ under the standard action of the group $(\mathbb{Z}/\mathbb{Z}_{m})^{n}$ on $\mathbb{P}^{n+1},$ where $X_{m}$ is defined as the zero-locus of $x_{0}^{m}+\cdots+x_{n+1}^{m}$ in $\mathbb{P}^{n+1}.$ Indeed, $(\mathbb{Z}/\mathbb{Z}_{m})^{n}$ is the Galois group for the Galois branched cover $X_{m}\rightarrow\mathbb{P}^{n}$ defined by $(x_{0}:\cdots:x_{n+1})\mapsto(x_{0}^{m}:\cdots:x_{n}^{m}).$
+The following result provides the first examples of Gibbs stable log Fano varieties – beyond complex log Fano curves – that are not exceptional:
+Proposition 2.13.
+The log Fano variety $(\mathbb{P}^{n},\Delta_{1-1/(n+1)}),$ which may be identified with the Fano orbifold $X_{n+1}/(\mathbb{Z}/\mathbb{Z}_{n+1})^{n},$ is Gibbs stable at any level $k.$
+Although the group $\text{Aut}_{0}(\mathbb{P}^{n},\Delta)$ is trivial, the proof of the previous proposition exploits that $\text{Aut}_{0}(\mathbb{P}^{n})$ is non-trivial, by applying the following lemma to $X=\mathbb{P}^{n}$:
+Lemma 2.14.
+Let $(X,D)$ be a projective log pair and $G$ a connected algebraic group acting algebraically on $X.$ If $D$ is $G$-invariant, then so is the center $Z_{v}$ of any divisorial valuation $v$ over $X$ computing $\mathrm{lct}(X,D).$ In particular, given a Fano variety $X$ and a positive rational number $k$ such that $kK_{X}$ is defined as line bundle, any divisorial valuation computing the log canonical threshold of the $\mathbb{Q}$-divisor $\mathcal{D}^{(N_{k})}$ on $X_{N_{k}}$ is invariant under the diagonal action of $\text{Aut}_{0}(X).$
+Proof.
+Consider the action of $G$ on valuations $v$ over $X$ defined by
+$(g\cdot v)(f)=v(g^{-1})f)$
+for any $g\in G$ and rational function $f$ on $X.$ If $v$ is a divisorial valuation, then so is $g\cdot v.$ Indeed, if $v$ is represented by a prime divisor $E$ on a model $\pi:Y\rightarrow X,$ then $g\cdot v$ is represented by the same divisor $E$ wrt the model $g^{*}\pi:Y\rightarrow X.$ In particular, $A(v)=A(g\cdot v)$ and $Z_{gv}=g(Z_{v}).$ It follows that if $D$ is $G$-invariant and $v$ computes $\mathrm{lct}(X,D),$ then $g\cdot v$ also computes $\mathrm{lct}(X,D)$ and, as a consequence, $g(Z_{v})$ is the center of a valuation computing $\mathrm{lct}(X,D),$ for any $g\in G.$ But, for a general log pair $(X,D),$ the number of centers of valuations computing $\mathrm{lct}(X,D)$ is finite.
+Indeed, after scaling $D$ we may as well assume that $\mathrm{lct}(X,D)=1.$ Then, by definition, $v$ computes $\mathrm{lct}(X,D)$ iff $v$ is a log canonical place (in the terminology of the MMP) and it is well-known that there are only a finite number of centers in $X$ of log canonical places [1].
+Finally, we can identify the center $Z_{v}$ with a point in the Chow variety $\mathcal{C}$ induced by a fixed ample line bundle on $X.$ The group $G$ naturally acts algebraically on $\mathcal{C}$ and since $G$ is connected so is the $G$-orbit $GZ_{v},$ viewed as a subset of $\mathcal{C}.$ But since the orbit $GZ_{v}$ is finite this implies that it consists of a single point, i.e.
+$Z_{v}$ is $G$-invariant, as desired.
+Finally, the last statement of the lemma follows from the fact that $\mathcal{D}^{(N_{k})}$ is is invariant under the diagonal action of $\text{Aut}_{0}(X)$ on $X^{N_{k}}$ (see the proof of Lemma 6.1).
+∎
+To prove Proposition 2.13 we start by expressing $K_{(\mathbb{P}^{n},\Delta_{w})}=:d\mathcal{O}(1),$ where d=n+1w(n+2)d=n+1-w(n+2) and $\Delta_{w}=w\Delta_{1}.$ The assumption w=11/(n+1)w=1-1/(n+1) is equivalent to
+$d+w=1,$
+since d+w=(n+1)(1w).d+w=(n+1)(1-w). Likewise, rescaling
+$\mathcal{D}^{(N)}\eqqcolon dD_{N},\,\,\Delta_{N}=w\Delta_{1,N},$
+where $\Delta_{1,N}$ is the divisor on $(\mathbb{P}^{n})^{N}$ defined by $\pi_{1}^{*}\Delta_{1}+\cdots+\pi_{N}^{*}\Delta_{1},$ in terms of the $N$ different projections $\pi_{i}\colon(\mathbb{P}^{n})^{N}\rightarrow\mathbb{P}^{n},$ we note that
+(2.19) $\text{(i)\,}\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})=1,\,\,\,\text{(ii) }\mathrm{lct}((\mathbb{P}^{n})^{N},D_{N})\geq 1.$
+The first item follows directly from the fact that $\Delta_{1,N}$ has simple normal crossings.
+The prove the second one observe that, by definition, $D_{N}$ is linearly equivalent to $\pi_{1}^{*}\mathcal{O}(1)+\cdots+\pi_{N}^{*}\mathcal{O}(1)$ and since $\mathrm{lct}((\mathbb{P}^{n},D)\geq 1$ for any $\mathbb{Q}$-divisor $D$ on $\mathbb{P}^{n}$ [34, Ex1.3] item $(ii)$ thus follows from formula 2.9.
+Now fix a divisorial valuation $v$ over $(\mathbb{P}^{n})^{N}$ and note that that the inverse of $\mathrm{lct}(X,D)_{v},$ appearing in the valuative formula 2.7 for $\mathrm{lct}(X,D),$ is linear wrt $D.$ Hence,
+(2.20) $\frac{1}{\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)}+\Delta_{N}\right)_{v}}=d\frac{1}{\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right)_{v}}+w\frac{1}{\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})_{v}}\leq 1,$
+giving $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)};\Delta\right)_{v}\geq 1.$ As a consequence, $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)};\Delta\right)\geq 1.$ Now consider a divisorial valuation $v$ over $(\mathbb{P}^{n})^{N}$ computing $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)}+\Delta_{N}\right).$ If $\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right)_{v}>1,$ then $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)}+\Delta_{N}\right)>1,$ by inequality 2.20.
+Consider finally the case when $\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right)_{v}=1.$ This implies, by item $ii$ in formula 2.19, that $v$ computes $\mathrm{lct}\left((\mathbb{P}^{n})^{N},D_{N}\right).$ To prove that $\mathrm{lct}\left((\mathbb{P}^{n})^{N},\mathcal{D}^{(N)};\Delta\right)>1$ it will, by the inequality 2.20, be enough to show that $\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})_{v}>1.$ But, by the previous lemma, the center $Z_{v}$ of $v$ is invariant under group $\mathrm{PGL}(n+1,\mathbb{C})$ acting diagonally on $(\mathbb{P}^{n})^{N}.$ This, in fact, forces $v(\Delta_{1,N})=0$ and thus $\mathrm{lct}(\mathbb{P}^{n})^{N},\Delta_{1,N})_{v}=\infty.$ Indeed, assume, to get a contradiction, that $v(\Delta_{1,N})>0.$ Then the center $Z_{v}$ is contained in support of $\Delta_{1,N}.$ In particular, fixing any $(x_{1},...,x_{N})\in Z_{v}$ this implies, since $Z_{v}$ is $\mathrm{PGL}(n+1,\mathbb{C})$-invariant, that $(gx_{1},...,gx_{N})$ is contained in $\Delta_{1,N}$ for any $g\in\mathrm{PGL}(n+1,\mathbb{C}).$ But this implies, since $\mathrm{PGL}(n+1,\mathbb{C})$ acts transitively on $\mathbb{P}^{n},$ that the support of the divisor $\Delta_{1}$ is all of $\mathbb{P}^{n},$ which is a contradiction.
+Remark 2.15.
+Taking $N=1$ and replacing the divisor $\mathcal{D}^{(N)}$ in the previous proof with any $\mathbb{Q}$-divisor that is linearly equivalent to $K_{(\mathbb{P}^{n},\Delta_{w})}$ shows that $(\mathbb{P}^{n},\Delta_{1-1(n+1)})$ is exceptional iff $d+w<1.$ Indeed, when $d+w<1,$ i.e.
+when w>11/n+1,w>1-1/n+1, the inequality 2.20 is strict and when $d+w\leq 1$ its left hand side is greater than are equal to $1.$ This is seen by taking $v$ to be the divisorial valuation on $\mathbb{P}^{n}$ defined by any of the hyperplanes $H_{i}.$
+2.6. The minimal case on $\mathbb{P}^{n}$
+As a preparation, we recall the following result [67, Theorem 11.2] describing the semistable locus $(\mathbb{P}^{n})_{\mathrm{ss}}^{N}$ of $(\mathbb{P}^{n})^{N}$ for general $N$ with respect to the diagonal action of $\mathcal{G}=\mathrm{PGL}(2,\mathbb{C})$ and polarization $K_{(\mathbb{P}^{n})^{N}}$: A point $(x_{1},\dots,x_{N})\in(\mathbb{P}^{n})^{N}$ is semistable iff for any proper linear subspace $W\in\mathbb{P}^{n},$
+(2.21) $\{i:x_{i}\in W\}\leq\frac{\mathrm{dim}W+1}{n+1}N.$
+With this explicit description we can compute $\gamma^{(N)}(\mathbb{P}^{n})$ for the minimal value of $N$.
+Proposition 2.16.
+$\mathbb{P}^{n}$ is Gibbs polystable at the minimal level $k_{\mathrm{min}}\coloneqq 1/(n+1)$, more precisely
+$\gamma^{(N_{\mathrm{min}})}(\mathbb{P}^{n})^{\mathcal{G}}=\infty.$
+Proof.
+Consider $X=\mathbb{P}^{n}$ at the minimal level $k_{\min}\coloneqq 1/(n+1)$, i.e.
+$kK_{X}=\mathcal{O}(1)$ and $N_{\mathrm{min}}=n+1$.
+First recall the well-known fact that the quantized Futaki character of $\mathbb{P}^{n}$ vanishes at every level.
+Note that in homogeneous coordinates $X_{i,j}$, $1\leq i,j\leq N$ on $(\mathbb{P}^{n})^{n+1}$, $\mathrm{det}S^{(N_{\mathrm{min}})}=\mathrm{det}[X_{i,j}]_{i,j}$.
+Note that when $N=n+1$, condition 2.21 for $W$ of codimension 1 states that not all $x_{i}$ lie on $W$, or equivalently that $\mathrm{det}S^{(n+1)}(x_{1},\dots,x_{n+1})\neq 0$.
+Thus $(\mathbb{P}^{n})_{\mathrm{ss}}^{N_{\mathrm{min}}}\subset\{\mathbb{\mathrm{det}}S^{(N_{\mathrm{min}})}\neq 0\}$ so we get
+$\gamma^{(N_{\mathrm{min}})}(\mathbb{P}^{n})^{\mathcal{G}}=\mathrm{lct}((\mathbb{P}^{n})_{\mathrm{ss}}^{N_{\mathrm{min}}},\frac{1}{k_{\mathrm{min}}}\mathrm{det}S^{N_{\min}})=\infty.$
+In fact, when $N=n+1$ in 2.21, the conditions for arbitrary $W$ follows from the ones for $W$ of codimension 1 and thus $(\mathbb{P}^{n})_{\mathrm{ss}}^{N_{\mathrm{min}}}=\{\mathbb{\mathrm{det}}S^{(N_{\mathrm{min}})}\neq 0\}$ in this special case.
+By Theorem 3.13, $\gamma^{(N)}(X)^{\mathcal{G}}$ is in not increasing, in general.
+This phenomenon persists for $X=\mathbb{P}^{n}$ when $n\geq 2$ by the following observation.
+Consider $X=\mathbb{P}^{2}$ at level $k=\frac{2}{3}$ so that $kK_{\mathbb{P}^{2}}=\mathcal{O}(2)$.
+Then it does not hold that $(\mathbb{P}^{2})_{\mathrm{ss}}^{6}\subset\{\mathrm{det}S^{(6)}\neq 0\}$.
+To see this, pick $6$ distinct points $(y_{1},\dots,y_{6})$ on a smooth quadric in $\mathbb{P}^{2}$, then $\mathrm{det}S^{(6)}(y_{1},\dots,y_{6})=0$.
+Since they are distinct, the only condition to be checked for semistability is whether $5$ or more points are contained in a single line.
+But a line meets a quadric in $\mathbb{P}^{2}$ in at most $2$ points by Bezout, so at most two of the $y_{1},\dots,y_{6}$ are contained in a common line.
+By 2.21, $(y_{1},\dots,y_{6})$ is semistable, giving
+$\mathrm{lct}_{(y_{1},\dots,y_{6})}((\mathbb{P}^{2})^{6},\frac{3}{2}\mathrm{det}S^{(6)})\leq 2/3$
+and thus $\gamma^{(6)}(\mathbb{P}^{2})\leq 2/3$ and $\mathbb{P}^{2}$ is not Gibbs polystable at level $k=2/3$.
 3. Analytic stability thresholds and thermodynamics
@@
 (3.7) $\omega_{u_{\mu}}^{n}/V=\mu,$
-using the non-pluripolar product introduced in [32], to define $\omega_{u_{\mu}}^{n}$.
+using the non-pluripolar product introduced in [38] to define $\omega_{u_{\mu}}^{n}$.
 We will fix the additive constant by demanding that $\int u_{\mu}\omega_{0}^{n}=0$ and call $u_{\mu}$ the potential (function) of $\mu.$ Equivalently, the corresponding potential metric $\phi_{\mu}$ satisfies
@@
 where $\mathcal{I}$ and $\mathcal{J}$ are the standard energy functionals in Kähler geometry, defined by
-$\mathcal{J}(u)\coloneqq-\mathcal{E}(u)+\frac{1}{V}\int_{X}u\frac{\omega_{0}^{n}}{V},\,\,\,\,\,\mathcal{I}(u)\coloneqq\int_{X}u\frac{(\omega_{0}^{n}-\omega_{u}^{n})}{V}.$
+$\mathcal{J}(u)\coloneqq-\mathcal{E}(u)+\int_{X}u\frac{\omega_{0}^{n}}{V},\,\,\,\,\,\mathcal{I}(u)\coloneqq\int_{X}u\frac{(\omega_{0}^{n}-\omega_{u}^{n})}{V}.$
 Example 3.2.
@@
 if $E(\mu)<\infty$ or equivalently $u_{\mu}$ is in the Sobolev space $W^{2,1}(X)$ of all $u\in L^{2}(X)$ such that $du\in L^{2}(X).$ Moreover, when $n=1$ the sup in formula 3.6 can be taken over all of $C^{0}(X)$ or over $W^{2,1}(X).$ Indeed, by classical Hilbert space theory the latter sup is attained by $u_{\mu}\in W^{2,1}(X)$ solving the (Laplace) equation 3.7 and hence $u_{\mu}\in\text{PSH}(X,\omega_{0})$.
-3.2.5. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$
-The space $\mathcal{H}(X,\omega_{0})$ of Kähler potentials can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the squared length of a tangent vector $\dot{u}$ at $u$ is defined by $\int_{X}\dot{u}^{2}\omega_{u}^{n}.$ The corresponding notion of geodesics was extended to the space
+3.2.5. The actions of $\mathcal{G}$ on $\text{$\text{PSH}$(}L),$ $\text{PSH}(X,\omega_{0})$ and $\mathcal{P}(X)$
+The group $\mathcal{G}$ naturally acts from the right on $\text{$\text{PSH}$(}L)$ by pullback: $\phi\cdot g:=g^{*}\phi.$ Fixing a reference metric $\phi_{0}$ in $L$ and setting $\omega_{0}:=dd^{c}\phi_{0},$ the right action of $\mathcal{G}$ on $\text{$\text{PSH}$(}L)$ induces a right action on $\text{PSH(}(X,\omega_{0}),$ by identifying $\text{$\text{PSH}$(}L)$ with $\text{PSH}(X,\omega_{0}).$ In other words, $u\cdot g:=g^{*}(\phi_{0}+u)-\phi_{0}$ (which is thus different than the action of $\mathcal{G}$ on $\text{PSH(}(X,\omega_{0})$ by pullback).
+The group $\mathcal{G}$ also naturally acts on the right on $\mathcal{P}(X)$ by setting $\mu\cdot g:=(g_{-1})_{*}\mu:=g^{*}\mu.$ This action is compatible with the action of $\mathcal{G}$ on $\text{$\text{PSH}$(}L)$ in the following sense: if $\mu=(dd^{c}\phi)^{n}/V,$ then $g^{*}\mu=(dd^{c}g^{*}\phi)^{n}/V.$ Moreover, if $L=-K_{(X,\Delta)}$ then $g^{*}\mu_{\phi}=\mu_{g^{*}\phi}.$ Finally, there is also a natural left action of $\mathcal{G}$ on $\text{$\text{PSH}$(}L),$ defined by $g\cdot\phi:=(g^{-1})^{*}\phi,$ which is compatible with the natural left action of $\mathcal{G}$ on $\mathcal{P}(X),$ defined by $g\cdot\mu:=g_{*}\mu.$
+3.2.6. The space $\mathcal{R}_{1,1}(X,\omega)$ of geodesics rays in $\mathcal{H}_{1,1}(X,\omega)$ and the $L^{1}$-Finsler completion of $\mathcal{H}(X,\omega)$
+The space $\mathcal{H}(X,\omega_{0})$ of Kähler potentials can be endowed with a canonical Riemannian metric, first introduced by Mabuchi, where the length of a tangent vector $\dot{u}$ at $u$ is defined by
+(3.13) $\left\|\dot{u}\right\|_{u}:=\left(\int_{X}\dot{u}^{2}\omega_{u}^{n}\right)^{1/2}$
+The corresponding notion of geodesics was extended to the space
 $\mathcal{H}_{1,1}(X,\omega)\coloneqq\left\{u\colon\omega_{u}\in L_{\text{loc}}^{\infty},\,\,\omega_{u}\geq 0\right\}.$
-in [40]. We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$
+in [51].
+We will denote by $\mathcal{R}(X,\omega)_{1,1}$ the corresponding space of all geodesic rays $u_{t}$ in $\mathcal{H}_{1,1}(X,\omega),$ parametrized by $t\in[0,\infty[.$ More generally, following [59], fixing $p\in[1,\infty[$ and replacing the $L^{2}$-norm in formula 3.13 with the corresponding $L^{p}$-norm yiels a Finsler metric on $\mathcal{H}(X,\omega_{0}),$ which induces a metric $d_{p}$ on $\mathcal{H}(X,\omega_{0}).$ As shown in [59], for $p=1,$ the metric space completion $\overline{(\mathcal{H}(X,\omega_{0}),d_{1})}$ of $(\mathcal{H}(X,\omega_{0}),d_{1})$ may be identified with $\mathcal{E}^{1}(X,\omega)$ (defined in the previous section).
+Thus $\overline{(\mathcal{H}(X,\omega_{0}),d_{1})}$ contains $\mathcal{H}_{1,1}(X,\omega)$ and, as shown in [59], the geodesics in $\mathcal{H}_{1,1}(X,\omega),$ discussed above, yield constant speed geodesics in the metric space $\overline{(\mathcal{H}(X,\omega_{0}),d_{1})}$.
+Moreover, on the subset $\{\mathcal{E}(u)=0\}$ of $\mathcal{E}^{1}(X,\omega)$ the following inequalities hold:
+(3.14) $A^{-1}E\left(\frac{\omega_{u}^{n}}{V}\right)-B\leq d_{1}(0,u)\leq AE\left(\frac{\omega_{u}^{n}}{V}\right)+B$
+for some positive constants $A$ and $B.$ Indeed, by 3.11, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [11, Lemma 2.13] and by [61, Prop 5.5] the inequalities 3.14 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u)$ (see also [61, Remark 3.9]).
 3.2.7. The Mabuchi functional $\mathcal{M}$ versus the free energy functional $F_{\beta}$
@@
 (3.22) $\frac{d}{dt}\mathcal{M}\left(\exp(t\mathrm{Re}V^{1,0})^{*}\psi\right)_{|t=0}=\mathrm{Re}\text{$\mathrm{Fut}$}_{(X,\Delta)}(V^{1,0})$
-More generally, 3.19 holds if $\psi$ and $\phi$ have finite energy and 3.20 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy.
+More generally, 3.21 holds if $\psi$ and $\phi$ have finite energy and 3.22 holds if $(dd^{c}\psi)^{n}/V$ has finite entropy.
 In particular, if $(X,\Delta)$ admits a KE metric, then $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
 Proof.
-This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.19 first note that, for any given any $g\in\mathcal{G},$
+This is well-known, at least for smooth data, but for completeness we provide simple proofs. To prove 3.21 first note that, for any given any $g\in\mathcal{G},$
 $\mathcal{E}\left(g^{*}\psi\right)-\mathcal{E}\left(g^{*}\phi\right)=\mathcal{E}(\psi)-\mathcal{E}(\phi).$
-This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.5 defined by formula 3.5, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.19.
+This follows directly from the fact that the differential $d\mathcal{E},$ which is the one-form on 3.5 defined by formula 3.5, satisfies $g^{*}(d\mathcal{E})=d(g^{*}\mathcal{E}).$ Replacing $g$ with a one-parameter subgroup $g_{\tau}$ and differentiating with respect to $\tau$ then proves formula 3.21.
 Next, observe that $\text{Fut}_{(X,\Delta)}(V^{1,0})=0$ for all $V^{1,0}\in\mathfrak{g}$ iff $\int_{X}\mathcal{L}_{\text{Re}V}\phi_{0}(dd^{c}\phi_{0})^{n}=0$ for all $V^{1,0}\in\mathfrak{g}$ (by applying $\text{Fut}_{(X,\Delta)}$ to both $V^{1,0}$ and $iV^{1,0}).$ Now, using that $\mathfrak{g}$ is reductive, we can decompose $\text{Re}V=W_{1}+JW_{2}$ where $W_{i}\in\mathfrak{k}$, the Lie algebra of $\mathcal{K}$.
 Since $\phi_{0}$ is assumed $\mathcal{K}$–invariant $\text{Fut}_{(X,\Delta)}(V^{1,0})=\int h_{W_{2}}(\omega_{0})^{n}/V,$ which shows that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0$ iff $\int_{X}\boldsymbol{m}(\omega_{0})^{n}/V=0,$ i.e.
-iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.13, we can decompose
+iff $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ Finally, applying formula 3.15, we can decompose
 $\mathcal{M}(\psi)=-\mathcal{E}(\psi)+D(dd^{c}\psi)^{n}/V|\mu_{\psi}),$
@@
 $\mathcal{M}(g_{t}^{*}\psi)=-\mathcal{E}(g_{t}^{*}\psi)+D(dd^{c}\psi)^{n}/V|\mu_{\psi}).$
-Differentiating with respect to $t$ and using 3.4 thus proves 3.20.
-More generally, the proof of 3.19 only used that $\phi$ and $\psi$ have finite energy.
+Differentiating with respect to $t$ and using 3.4 thus proves 3.22.
+More generally, the proof of 3.21 only used that $\phi$ and $\psi$ have finite energy.
 Moreover, when $(dd^{c}\psi)^{n}$ has finite entropy it is shown in [25] that $\psi$ has finite energy, showing that $D(dd^{c}\psi)^{n}|\mu_{\psi})<\infty.$ Finally, if $(X,\Delta)$ admits a Kähler–Einstein metric $\psi,$ then, by [26], $(dd^{c}\psi)^{n}$ has finite entropy and $\psi$ minimizes the Mabuchi functional on the space of metrics with finite energy.
 Hence, $\frac{d}{dt}\mathcal{M}\left(\exp(t\text{Re}V^{1,0})^{*}\psi\right)_{|t=0}$, showing that $\text{Fut}_{(X,\Delta)}(\cdot)\equiv 0.$
-The independence of the choice of $\phi_{0}$ is also a consequence of the fact that $\text{Fut}_{(X,\Delta)}$ arises as the leading term of the large-$k$ asymptotics of the quantized Futaki character $\text{Fut}_{(X,\Delta),k}(\cdot)$ at level $k,$ defined by
+The independence of the choice of $\phi_{0}$ is also a consequence of the fact that $\text{Fut}_{(X,\Delta)}$ arises as the leading term of the large-$k$ asymptotics of the quantized Futaki character $\text{Fut}_{(X,\Delta),k}(\cdot)$ at level $k,$ defined as follows when $k$ is a positive integer such that $k\Delta$ is divisor:
 (3.23) $\text{Fut}_{(X,\Delta),k}(V^{1,0})\coloneqq-\frac{1}{kN}\text{Tr}(\mathcal{L}_{V^{1,0}}),$
-where $\mathcal{L}_{V^{1,0}}$ now denotes the endomorphism of $H^{0}(X,-kK_{(X,\Delta)})$ induced by the infinitesimal action of $V^{1,0}$ on $H^{0}(X,-kK_{(X,\Delta)})$ (see the proof of [26, Prop 4.7]).
+where $\mathcal{L}_{V^{1,0}}$ now denotes the endomorphism of $H^{0}(X,-kK_{(X,\Delta)})$ induced by the infinitesimal action of $V^{1,0}$ on $H^{0}(X,-kK_{(X,\Delta)})$ (see the proof of [28, Prop 4.7]).
+More generally, when it is only assumed that $kK_{(X,\Delta)}$ is a line bundle we define $\text{Fut}_{(X,\Delta),k}$ by formula 6.1 in Lemma 6.1.
 3.2.9. The moment maps $\boldsymbol{m}$ and $\boldsymbol{m}_{N}$
@@
 $\boldsymbol{m}\colon\,\,X\rightarrow\mathbb{R}^{r},\,\,\,x\mapsto(h_{e_{1}}(x),\dots,h_{e_{r}}(x)),$
-where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.18).
+where $h_{W}$ denotes the Hamiltonian of a real vector field $W$ on $X$ corresponding to an element $W$ in $\mathfrak{k},$ defined by $h_{W}\coloneqq\mathcal{L}_{JW}\phi$ (defined as in formula 3.20).
 We will denote by $\boldsymbol{m}_{N}$ the moment map of the diagonal action of the group $\mathcal{K}$ on $X^{N},$ endowed with the symplectic form
@@
 (3.24) $\mathcal{G}\left\{\boldsymbol{m}_{N}=0\right\}=(X^{N})_{\text{ps }}\subset X_{\text{ss}}$
-Moreover, 0 is a regular value of $\boldsymbol{m}_{N}$ iff all the stabilizers in $\mathcal{K}$ of the points in $\left\{\boldsymbol{m}_{N}=0\right\}$ are finite iff the semistable locus of $X^{N}$ coincides with the polystable locus [64, Cor 3.10].
+Moreover, 0 is a regular value of $\boldsymbol{m}_{N}$ iff all the stabilizers in $\mathcal{K}$ of the points in $\left\{\boldsymbol{m}_{N}=0\right\}$ are finite iff the semistable locus of $X^{N}$ coincides with the polystable locus [83, Cor 3.10].
 3.3. The structure of the space $\mathcal{P}(X)_{0}$ of probability measures with vanishing moment
@@
 Proposition 3.4.
-Consider the infimum of the pluricomplex energy $E$ on the orbit $\mathcal{G}\nu$ of any volume form in $\mathcal{P}(X)$ with divisorial singularities along $\Delta.$ It is attained at a unique $\mu$ in $(\mathcal{G}\nu)/\mathcal{K},$ which is the unique critical point of $E$ on $(\mathcal{G}\nu)/\mathcal{K}.$ Moreover, if the Futaki character of $(X,\Delta)$ vanishes, then $\mu$ is the unique element in $\left(\mathcal{G}\nu\cap\mathcal{P}(X)_{0}\right)/\mathcal{K}$ and, more generally, any $\mu\in\mathcal{P}(X)_{0}$ minimizes $E$ on $\mathcal{G}\mu.$
+Assume that the Futaki character of $(X,\Delta)$ vanishes.
+Then, given $\nu\in\mathcal{P}(X)$ such that $E(\nu)<\infty,$ there exists $\mu\in\mathcal{G}\nu$ minimizing $E$ on $\mathcal{G}\nu.$ Moreover, $\mu$ minimizes $E$ on $\mathcal{G}\nu$ iff $\mu\in\mathcal{G}\nu\cap\mathcal{P}(X)_{0}$ and such a measure $\mu$ is uniquely determined modulo the action of $\mathcal{K}.$
 Proof.
-First note that $E$ is $\mathcal{K}$–invariant (since $\omega_{0}$ is $\mathcal{K}$–invariant). We thus obtain a function
-$\mathcal{G}/\mathcal{K}\rightarrow\mathbb{R},\,\,\,g\mapsto E(g^{*}\nu),$
-The first part of the proposition to be proved says that this function has a unique critical point.
-To prove this first recall that $\mathcal{G}=\mathcal{K}\exp(J\mathfrak{k}),$ by polar decomposition of complex reductive Lie groups.
-Hence, by basic convex analysis it enough to show that $E(g^{*}\nu)$ is proper and strictly convex along curves $g_{t}$ of the form $g_{t}=\exp(tJW),$ where $W\in\mathfrak{k}-\{0\}.$ Expressing $\nu=(dd^{c}\phi_{\nu})^{n}/V,$ for some $\phi_{\nu}\in\text{PSH}(-K_{(X,\Delta)})_{b},$ gives $g_{t}^{*}\nu=dd^{c}(g_{t}^{*}\phi_{\nu}))^{n}/V.$ However, the real one-parameter curve $g_{t}^{*}\phi_{\nu}$ need not define a geodesic in the sense of [40], even when $\phi_{\nu}$ is smooth, unless $\phi_{\nu}$ is $\mathcal{K}$-invariant.
-To bypass this difficulty we will use a complexification argument, by setting $\tau=s+it$ for $s,t\in\mathbb{R}$ and $g_{\tau}\coloneqq\exp(sW+tJW).$ Since $E$ is $\mathcal{K}$-invariant we have $E(g_{{}_{\tau}}^{*}\nu)=E(g_{t}^{*}\nu).$ The corresponding properness on $\mathcal{G}/\mathcal{K}$ is well-known (see, for example, [21, Remark 9.2]).
-What remains is thus to show that
-(3.23) $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)>0.$
-[…]
-$\phi^{\tau}\coloneqq\exp(g_{\tau})^{*}\phi_{\nu},$
-we can express $g_{\tau}^{*}\nu=(dd^{c}\phi^{\tau})^{n}/V.$ We claim that there exists a positive constant $c$ such that
-(3.24) $\frac{\partial}{\partial\tau}\frac{\partial}{\partial\overline{\tau}}E\left(\frac{(dd^{c}\phi^{\tau})^{n}}{V}\right)\geq c\int_{X}i\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)\wedge\overline{\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)}\wedge dd^{c}\phi_{0}\wedge(dd^{c}\phi^{\tau})^{n-2}$
-(when $n=2$ and when $n=1$ the last factor above should be omitted).
-Accepting this for the moment, it follows that $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0.$ Moreover, if $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)=0,$ then $\partial(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}})=0$ on $X-\Delta,$ since both $dd^{c}\phi_{0}$ and $dd^{c}\phi^{\tau}$ are smooth and strictly positive on $X-\Delta.$ But this means that
-$\partial\bar{\partial}\phi^{\tau}(\cdot,\overline{W^{1,0}})=0\,\,\,\text{on $X-\Delta$}$
-and since $i\partial\bar{\partial}\phi^{\tau}$ is smooth and strictly positive on $X-\Delta$ it follows that $W=0$ on $X-\Delta.$ Thus, by continuity, $W=0$ on $X,$ which is a contradiction.
-Hence, 3.23 holds.
-Finally, the inequality 3.24 is shown in essentially the same way as in the proof of [21, Lemma 9.1].
-More precisely, in [21, Lemma 9.1] it is assumed that the corresponding metric $\phi^{\tau}$ is independent of the real part of $\tau$ (in our notation) and that $\phi^{\tau}$ satisfies the complex geodesic equation, that is the local function $\Phi(z,\tau)\coloneqq\phi^{\tau}(z)$ satisfies
-(3.25) $(dd^{c}\Phi)^{n+1}=0\,\,\,\text{on $X\times\mathbb{C}$}$
-But the proof only uses the latter equation and this equation is satisfied for our $\phi^{\tau}$ (as is well-known and can be seen by locally making the change of variable, $w\coloneqq g_{\tau}(z)).$
-Next, let us show that $\mu$ is a critical point of $E$ on the orbit $\mathcal{G}\nu$ iff $\mu\in\mathcal{P}(X)_{0}.$ Since $E$ is $\mathcal{K}$–invariant it will be enough to show that for any real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g},$
-(3.26) $\frac{d}{dt}E(\mu_{t})_{|t=0}=-\int h_{W}\mu,\,\,\,\>\mu_{t}\coloneqq\exp(tJW)^{*}\mu$
-where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ First recall that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0},$ (see Section 3.2.8).
-To prove formula 3.26 note that, by 3.10, denoting by $\phi$ the potential of $\mu,$
-$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int(\phi-\phi_{0})\frac{d\mu_{t}}{dt}|{}_{t=0}=-\int(\phi-\phi_{0})\mathcal{L}\mu,$
-Integrating by parts gives
-$\frac{d}{dt}E(\mu_{t})|_{t=0}=\int\mathcal{L}(\phi-\phi_{0})\mu=\int\mathcal{L}\phi\frac{(dd^{c}\phi)^{n}}{V}-\int h_{W}\mu,$
+First note that $E$ is $\mathcal{K}$–invariant (since $\omega_{0}$ is $\mathcal{K}$–invariant).
+Given $\nu\in\mathcal{P}(X)$ we thus obtain a function on the quotient $\mathcal{K}\backslash\mathcal{G}$ of $\mathcal{G}$ by the left action of $\mathcal{K}:$
+(3.25) $\mathcal{K}\backslash\mathcal{G}\rightarrow\mathbb{R},\,\,\,f(g):=E(g_{*}\nu),$
+Since $\mathcal{G}=\mathcal{K}\exp(J\mathfrak{k})$ (by polar decomposition of complex reductive Lie groups) the function $f$ can be identified with a function on $\exp(J\mathfrak{k}).$ Next, let us show that, given $\nu\in\mathcal{P}(X)$ such that $E(\nu)<\infty,$ $\mu$ is a critical point of $E$ on the orbit $\mathcal{G}\nu$ iff $\mu\in\mathcal{P}(X)_{0}.$ Since $E$ is $\mathcal{K}$–invariant it will be enough to show that for any real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g}$ the following formula holds for any given $\mu\in\mathcal{P}(X)$ such that $E(\mu)<\infty$:
+(3.26) $\left.\frac{df(\exp(tJW))}{dt}\right|_{t=0}\coloneqq\frac{d}{dt}E(\exp(tJW)_{*}\mu)_{|t=0}=\int h_{W}\mu,$
+where $h_{W}$ is the Hamiltonian function corresponding to $W$ and $\omega_{0}.$ We start by proving the formula when $\mu$ is a smooth form.
+Set $\mu_{t}\coloneqq\exp(tJW)_{*}\mu.$ By 3.10, denoting by $\phi$ the potential of $\mu,$
+$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int(\phi-\phi_{0})\frac{d\mu_{t}}{dt}|{}_{t=0}=\int(\phi-\phi_{0})\mathcal{L}_{JW}\mu,$
+Integrating by parts and using that $h_{W}\coloneqq\mathcal{L}_{JW}\phi_{0}$ (see Section 3.2.9) gives
+$\frac{d}{dt}E(\mu_{t})|_{t=0}=-\int\mathcal{L}_{JW}(\phi-\phi_{0})\mu=-\int\mathcal{L}_{JW}\phi\frac{(dd^{c}\phi)^{n}}{V}+\int h_{W}\mu,$
 which proves 3.26, using the previous lemma and the assumption that the Futaki character of $(X,\Delta)$ vanishes.
-Finally, take $\mu\in\mathcal{P}(X)_{0}.$ If $E(\mu)=\infty,$ then $E(g^{*}\mu)=\infty$ for all $g\in\mathcal{G}.$ Hence, we may assume that $E(\mu)<\infty,$ which means that $E(g^{*}\mu)<\infty$ for all $g\in\mathcal{G}.$ Consider, as before, the function $g\mapsto E(g^{*}\mu)$ on $\mathcal{G}/\mathcal{K}.$ It follows from the inequality $\partial_{\tau}\bar{\partial}_{\tau}E(g_{{}_{\tau}}^{*}\nu)\geq 0,$ shown above when $\nu$ is a volume form, and a standard approximation argument, that $E(g^{*}\mu)$ is convex on the symmetric space $\mathcal{G}/\mathcal{K}.$ Moreover, as before, $\frac{d}{dt}E(\mu_{t})_{|t=0}=0$ using the previous lemma, that applies since $E(\mu)<\infty$ (which equivalently means that $\mathcal{E}(\phi_{\mu})>-\infty).$ Since any critical point of a convex function $f$ on the symmetric space $\mathcal{G}/\mathcal{K}$ minimizes $f$ this shows that $\mu$ minimizes $E$ on $\mathcal{G}\mu,$ as desired.
-∎
-By the previous proposition, the following map is well-defined:
-$\pi_{0}\colon\,\mathcal{V}(X)\rightarrow\mathcal{V}(X)_{0}/\mathcal{K},\,\,\,E\left(\pi_{0}(\nu)\right)=\inf_{\mathcal{G}\nu}E$
-where $\mathcal{V}(X)$ denotes the subspace of $\mathcal{P}(X)$ consisting of volume forms, i.e.
-measures with smooth and strictly positive density and $\mathcal{V}(X)_{0}\coloneqq\mathcal{P}(X)_{0}\cap\mathcal{V}(X).$ More generally, $\pi_{0}$ is well-defined for any volume form $\mu$ in $\mathcal{P}(X)$ with divisorial singularities.
+Now assume only that $E(\mu)<\infty$ and take a sequence $u_{j}$ in $\mathcal{H}(X,\omega_{0})$ decreasing to the potential $u_{\mu}$ of $\mu.$ Since $u_{\mu}\in\mathcal{E}^{1}(X,\omega)$ it follows that $\mu_{j}:=\omega_{u_{j}}^{n}/V\rightarrow\mu$ and $E(\mu_{j})\rightarrow E(\mu),$ by [25, Prop 1.1].
+Likewise, given an automorphism $g$ of $X,$ the sequence $g\cdot u_{j}$ in $\mathcal{H}(X,\omega_{0})$ decreases to $g\cdot u_{\mu}$ and thus $g_{*}\mu_{j}\rightarrow g_{*}\mu$ and $E(g_{*}\mu_{j})\rightarrow E(g_{*}\mu)$.
+Now, applying the previous case to $\mu_{j}$ and $tW$ for all $t\in]0,\epsilon]$ gives
+$\epsilon^{-1}\left(E(\exp(\epsilon JW)_{*}\mu_{j})-E(\mu_{j})\right)=\epsilon^{-1}\int_{0}^{\epsilon}\left(\int_{X}h_{tW}dt\right)\mu_{j}dt.$
+Letting $j\rightarrow\infty$ and using that that $h_{tW}$ is continuous thus gives
+$\epsilon^{-1}\left(E(\exp(\epsilon JW)_{*}\mu)-E(\mu)\right)=\epsilon^{-1}\int_{X}\left(\int_{0}^{\epsilon}h_{tW}dt\right)\mu dt.$
+Finally, letting $\epsilon\rightarrow 0$ and using that $h_{tW}(x)$ is continuous wrt $(t,x)$ proves formula 3.26.
+Now assume that $\nu\in\mathcal{P}(X)$ and $E(\nu)<\infty.$ Then $E(g_{*}\nu)<\infty$ for all $g\in\mathcal{G}$ and we will next show that the function $f,$ defined by formula, is strictly convex, when identified with a function on the real vector space $\exp(J\mathfrak{k}).$ To this end fix a real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g}.$ Differentiating formula 3.26 gives
+$\frac{d^{2}}{d^{2}t}_{|t=0}E(\exp(tJW)_{*}\nu)=\int_{X}dH(JW)\nu=\int_{X}\omega_{0}(W,JW)\nu,$
+using that $h$ is a Hamiltonian function for $W.$ Since $\omega_{0}(W,JW)$ is non-negative, applying the previous formula to any element in $\exp(\mathbb{R}JW)_{*}\nu$ shows that $t\mapsto E(\exp(tJW)_{*}\nu)$ is a smooth convex function, which is, in fact, strictly convex.
+Indeed, $\omega(W,JW)$ only vanishes on the subvariety of $X$ defined by the zero-locus of $V$ and $\nu$ does not put mass on analytic subvarieties (since $E(\nu)<\infty,$ $\nu$ does not charge pluripolar subsets [25]).
+It follows directly from the strictly convexity of $f$ that its critical points are unique.
+All that remains is thus to prove that $f$ is proper on $\mathcal{K}/\mathcal{G}$ (and thus admits a critical point).
+First assume that $\nu=\omega_{0}^{n}/V,$ which minimizes $E$ on all of $\mathcal{P}(X).$ Then $f$ is a strictly convex function admitting a minimum (when $g$ is the identity), which implies that $f$ is proper.
+Next, for any given $\nu\in\mathcal{P}(X)$ such that $E(\nu)<\infty$ there exists a unique $u\in\mathcal{E}^{1}(X,\omega)$ such that $\omega_{u}^{n}/V=\nu$ and $\mathcal{E}(u)=0.$ Since we have assumed that the Futaki character of $(X,\Delta)$ vanishes, it follows that $\mathcal{E}(u\cdot g)=0$ for all $g\in\mathcal{G},$ using Lemma 3.3.
+Hence, using the inequality 3.14, it is equivalent to show that the function $g\mapsto d_{1}(u\cdot g,0)$ is proper.
+Having already shown that $g\mapsto E(g^{*}\omega_{0}^{n}/V)$ is proper – which is equivalent to $d_{1}(0\cdot g,0)$ being proper – it will thus be enough to show that $d_{1}(u\cdot g,0)-d_{1}(0\cdot g,0)$ is bounded by a constant independent of $g.$ But, applying the triangle inequality to $d_{1}$ yields
+$\left|d_{1}(u\cdot g,0)-d_{1}(0\cdot g,0)\right|\leq d_{1}(u\cdot g,0\cdot g)=d(u,0),$
+using, in the equality, that $\mathcal{G}$ preserves the metric $d_{1}$ [61, Lemma 5.9]. ∎
+By the previous proposition, the following map is well-defined, if the Futaki character of $(X,\Delta)$ vanishes:
+$\pi_{0}\colon\,\mathcal{P}(X)\cap\{E<\infty\}\rightarrow\mathcal{P}(X)_{0}\cap\{E<\infty\}/\mathcal{K},\,\,\,E\left(\pi_{0}(\nu)\right)=\inf_{\mathcal{G}\nu}E$
 Lemma 3.5.
@@
 Then $F_{-\gamma}\left(\pi_{0}(\nu)\right)\leq F_{-\gamma}\left(\nu\right)$ when $\gamma\leq 1$ and $F_{-\gamma}\left(\pi_{0}(\nu)\right)\geq F_{-\gamma}\left(\nu\right)$ when $\gamma\geq 1.$ As a consequence, for $\gamma\leq 1,$
-(3.27) $\inf_{\mathcal{V}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{\mathcal{V}}(X)}F_{-\gamma}=\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)}F_{-\gamma}$
+(3.27) $\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)}F_{-\gamma}$
 Proof.
@@
 Moreover, by definition, $F_{-\gamma}=(1-\gamma)E+F_{-1}.$ Since $F_{-1}$ is $\mathcal{G}$–invariant, under the assumption about the vanishing of the Futaki character, the first statement of the lemma thus follows and, as a consequence, also the first identity in formula 3.27.
-Thus, to prove the rest of the identities 3.27, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [27, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively.
+Thus, to prove the rest of the identities 3.27, it will be enough to show that for any given $\mu$ in $\mathcal{P}(X)$ such that $F_{-\gamma}(\mu)<\infty$, there exists $\nu_{j}\in\mathcal{V}(X)$ such that $F(\nu_{j})\rightarrow F(\mu).$ But this follows from [29, Lemma 3.1], implying that for any given $\mu$ such that $E(\mu)<\infty$ and $D(\mu|\mu_{0})<\infty,$ there exists a sequence $\mu_{j}$ converging towards $\mu$ in $\mathcal{P}(X)$ such that $E(\mu_{j})$ and $D(\mu_{j}|\mu_{0})$ converge towards $E(\mu)$ and $D(\mu|\mu_{0}),$ respectively.
 ∎
@@
 showing that $\left|E_{\phi_{0}}(\mu)-E_{\phi_{1}}(\mu)\right|\leq C+\left\|u\right\|_{L^{\infty}},$ as desired.
-Finally, the second statement follows from combining formulae 3.13 and 3.11 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.7).
+Finally, the second statement follows from combining formulae 3.15 and 3.11 with proposition 3.4 and using that $F_{-1}$ is $\mathcal{G}$–invariant (Section 3.2.8).
 ∎
 We are now ready for the proof of the following result, where the equivalence $1\iff 2$ follows directly from [26, 61].
-The equivalence $1\iff 4$ provides partial evidence for the Large Deviation Principle (LDP) in Conjecture A, stated in Section 6.2, as it establishes the lower semicontinuity of the rate functional – a property built into the definition of an LDP.
+The equivalence $1\iff 4$ provides partial evidence for the Large Deviation Principle (LDP) in Conjecture A, stated in Section 6.3, as it establishes the lower semicontinuity of the rate functional – a property built into the definition of a LDP.
 Theorem 3.7.
@@
 (3.29) $\mathcal{M}(u)\geq\epsilon\inf_{g\in\mathcal{G}}d_{1}(0,u\cdot g)-C\,\,\,\text{on $\mathcal{H}(X,\omega_{0}$})\cap\{\mathcal{E}(u)=0\},$
-where $d_{1}$ is the $L^{1}$–Finsler metric on $\mathcal{H}(X,\omega_{0})$, which, on $\{\mathcal{E}(u)=0\}$ satisfies
-(3.30) $A^{-1}E\left(\frac{\omega_{u}^{n}}{V}\right)-B\leq d_{1}(0,u)\leq AE\left(\frac{\omega_{u}^{n}}{V}\right)+B$
-for some positive constants $A$ and $B.$ Indeed, by 3.11, $E(\frac{\omega_{u}^{n}}{V})=(\mathcal{I}-\mathcal{J})(u)$, which is comparable to $\mathcal{J}(u)$ [10, Lemma 2.13] and by [48, Prop 5.5] the inequalities 3.30 do hold when $E(\frac{\omega_{u}^{n}}{V})$ is replaced by $\mathcal{J}(u).$ This proves $1\iff 2$.
+where $d_{1}$ is the $L^{1}$–Finsler metric on $\mathcal{H}(X,\omega_{0})$ (see Section 3.2.6).
+Using the inequality 3.14, this proves $1\iff 2$.
 Next, note that, $3\implies 2$ follows from the previous proposition.
-To prove that $2\implies 3$ we recall that it was pointed out in [48, Remark 3.9] that, by density, the inequality 3.29 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [46], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.30 this shows that (after perhaps decreasing $\epsilon$ and increasing $C$) there exist $\epsilon,C>0$ such that
+To prove that $1\implies 3$ we recall that it was pointed out in [61, Remark 3.9] that, by density, the inequality 3.29 is equivalent to the corresponding inequality on the metric completion of the metric space $\left(\mathcal{H}(X,\omega_{0}),d_{1}\right),$ which, by [59], may be identified with $\mathcal{E}^{1}(X,\omega_{0}).$ Combined with the inequalities 3.14, this shows that, assuming $1,$ there exist $\epsilon,C>0$ such that
 $F(\mu)\geq\epsilon\inf_{\mathcal{G}}E(\mu)-C\text{\,\,on $\mathcal{P}(X)\cap\{E<\infty\}.$ }$
-In particular, restricting this inequality to $\mathcal{P}(X)_{0},$ and using the last statement of Prop 3.4, yields $F(\mu)\geq\epsilon E(\mu)-C$ on $\mathcal{P}(X)\cap\{E<\infty\},$ i.e.
+In particular, restricting this inequality to $\mathcal{P}(X)_{0},$ and using Prop 3.4, yields $F(\mu)\geq\epsilon E(\mu)-C$ on $\mathcal{P}(X)\cap\{E<\infty\},$ i.e.
 $F_{-(1+\epsilon)}\geq-C$ on $\mathcal{P}(X)\cap\{E<\infty\}$ and hence, by definition, on $\mathcal{P}(X),$ as desired.
-To prove that $1\iff 4,$ first assume that $F_{-1}$ is lsc on $\mathcal{P}(X)_{0}.$ Since $\mathcal{P}(X)_{0}$ is compact it follows that $F_{-1}$ admits a minimizer in $\mathcal{P}(X)_{0}.$ Such a minimizer is KE, since the Futaki character vanishes (see below).
+To prove the remaining equivalence $1\iff 4,$ first assume that $F_{-1}$ is lsc on $\mathcal{P}(X)_{0}.$ Since $\mathcal{P}(X)_{0}$ is compact it follows that $F_{-1}$ admits a minimizer in $\mathcal{P}(X)_{0}.$ Such a minimizer is KE, since the Futaki character vanishes (see below).
 Conversely, assume that $(X,\Delta)$ admits a KE metric.
-Take a sequence $\mu_{j}$ converging towards $\mu_{\infty}$ in $\mathcal{P}(X)_{0}.$ To prove that $F_{-1}$ is lsc we may as well assume that $F_{-1}(\mu_{j})\leq C.$ Using $1\iff 3,$ it follows that $D(\mu_{j}|\mu_{\phi_{0}})\leq C.$ Thus, by [24, Thm 2.17], $F_{-1}(\mu_{\infty})\leq\liminf F_{-1}(\mu_{j}),$ as desired.
+Take a sequence $\mu_{j}$ converging towards $\mu_{\infty}$ in $\mathcal{P}(X)_{0}.$ To prove that $F_{-1}$ is lsc we may as well assume that $F_{-1}(\mu_{j})\leq C.$ Using $1\iff 3,$ it follows that $D(\mu_{j}|\mu_{\phi_{0}})\leq C.$ Thus, by [26, Thm 2.17], $F_{-1}(\mu_{\infty})\leq\liminf F_{-1}(\mu_{j}),$ as desired.
 To prove the last statement of the theorem, recall that any KE volume form $\mu$ in $\mathcal{P}(X)$ minimizes $F_{-1}$ on $\mathcal{P}(X)$ [26].
-Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [21, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e.
+Thus, by Lemma 3.5, there exists a KE volume form $\pi_{0}(\mu)$ minimizing $F_{-1}$ on $\mathcal{P}(X)_{0}$ (but, a priori, $\pi_{0}(\mu)$ is only determined modulo $\mathcal{K}).$ To prove that there exists a unique KE volume form $\mu_{\text{KE}}$ in $\mathcal{P}(X)_{0},$ we recall that, as shown in [22, Lemma 6.3], for any given $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)})$ there exists a unique KE volume form $\mu_{-1}$ in $\mathcal{P}(X)$ minimizing $E$ on the space of KE volume forms in $\mathcal{P}(X),$ i.e.
 on $\mathcal{G}\mu_{-1}.$ Moreover, $\mu_{-1}$ arises as the limit 3.28.
@@
 Consider the space $\mathcal{H}(X,\omega_{0})/\mathbb{R}$ of all Kähler potentials, modulo additive constants.
-The group $\mathcal{G}$ naturally acts from the right on $\mathcal{H}(X,\omega_{0})/\mathbb{R},$ by demanding that $g^{*}\omega_{u}=\omega_{0}+dd^{c}(u\cdot g).$ Now assume that $\omega_{0}$ is Kähler–Einstein and consider the following orthogonality condition on $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$:
+Assume that $\omega_{0}$ is Kähler–Einstein and consider the following orthogonality condition on $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$:
 (3.30) $\int_{X}u\psi\omega_{0}^{n}=0\,\,\,\forall\psi\in\text{$\ker$(}\Delta-\lambda_{1})$
 where $\Delta$ denotes the Laplacian with respect to $\omega$ and $\lambda_{1}$ denotes its smallest strictly positive eigenvalue.
-This condition appears in Tian’s conjecture [89, Conj 5.5], settled in [48], saying that if $\omega_{0}$ is Kähler–Einstein, then there exists positive constants $\epsilon$ and $C$ such that
+This condition appears in Tian’s conjecture [107, Conj 5.5], settled in [61], saying that if $\omega_{0}$ is Kähler–Einstein, then there exists positive constants $\epsilon$ and $C$ such that
 (3.31) $\mathcal{M}(u)\geq\epsilon(\mathcal{I}-\mathcal{J})(u)-C$
-for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.31.222Strictly speaking, in the original conjecture [89, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [48].
-It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [23, Section 2]).
-The inequality 3.32 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.32 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.31 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [95, Lemma A.2]:
+for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying the orthogonality condition 3.30.222Strictly speaking, in the original conjecture [107, Conj 5.5] the role of $\mathcal{M}$ is played by the Ding functional $\mathcal{D}$ (see Section 4.0.1); anyhow, both versions of the conjecture were established in [61].
+It should be emphasized that the functional $(\mathcal{I}-\mathcal{J})$ appearing in the right hand side above may as well be replaced by the functionals $\mathcal{I}$ or $\mathcal{J}$ (up to changing the constant $\epsilon),$ since all three functionals $\mathcal{I},(\mathcal{I}-\mathcal{J})$ and $\mathcal{J}$ are comparable (up to multiplicative constants [25, Section 2]).
+The inequality 3.31 should be compared to item 3 in Theorem 3.7, saying that the inequality 3.31 holds for all $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ such that $\omega_{u}^{n}\boldsymbol{m}$ integrates to zero on $X.$ Note that, while Tian’s orthogonality condition 3.30 is linear in $u$, the moment condition appearing in Theorem 3.7 is linear with respect to the Monge–Ampère measure $\omega_{u}^{n}.$ The relation between these to conditions is clarified in the following lemma, which gives a refinement of [113, Lemma A.2]:
 Lemma 3.9.
 Let $X$ be a Fano manifold and assume that the Futaki character of $X$ vanishes and that the group $\mathcal{G}$ is reductive.
-Fix a Kähler form $\omega_{0}$ that is invariant under a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}.$ Given $u\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ the following two conditions are equivalent for $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}:$
-$(i)\,\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)=\mathcal{J}(v),\,\,\,\,(ii)\int_{X}v\Delta\boldsymbol{m}\omega_{0}^{n}=0,$
-where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, there exists a unique $v\in u\cdot G\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions.
-When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.31 for $v.$
+Fix a Kähler form $\omega_{0}$ that is invariant under a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}.$ Given $u\in\mathcal{E}^{1}(X,\omega_{0})/\mathbb{R}$ the following conditions are equivalent for $v\in u\cdot G\subset\mathcal{E}^{1}(X,\omega_{0})/\mathbb{R}:$
+$(i)\,\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)=\mathcal{J}(v),\,\,\,(ii)\int_{X}v\Delta\boldsymbol{m}\omega_{0}^{n}=0,\,\,\,(iii)\int_{X}\boldsymbol{m}\omega_{v}\wedge\omega_{0}^{n-1}=0$
+where $\Delta$ denotes the Laplacian with respect to the Kähler metric $\omega_{0}.$ Moreover, modulo the action of $\mathcal{K},$ there exists a unique $v\in u\cdot G\subset\mathcal{E}^{1}(X,\omega_{0})/\mathbb{R}$ satisfying any of these two conditions.
+When $\omega_{0}$ is Kähler–Einstein the condition $(ii)$ is equivalent to Tian’s orthogonality condition 3.30 for $v.$
 Proof.
 First observe that $\mathcal{J}(u)$ is $\mathcal{K}$–invariant, since $\omega_{0}$ is assumed $\mathcal{K}$–invariant.
-It thus follows, precisely as in the proof of Prop 3.4, that the inf over $\mathcal{G}$ in $(i)$ may be replaced by the inf over $J\mathfrak{k}.$ The pullback of $\mathcal{J}$ to $J\mathfrak{k}$ is proper and strictly convex on $J\mathfrak{k}.$ Indeed, for any complex one-parameter curve $\phi^{\tau}$ in $\mathcal{H}(-K_{X})$ satisfying the complex geodesic equation 3.25 a direct calculation gives
-$V\frac{\partial^{2}\mathcal{J}(\phi^{\tau}-\phi_{0})}{\partial\tau\partial\overline{\tau}}=\int_{X}\frac{\partial^{2}\phi^{\tau}}{\partial\tau\partial\overline{\tau}}\omega_{0}^{n}=\int_{X}\left|\partial\left(\frac{\partial\phi^{\tau}}{\partial\overline{\tau}}\right)\right|_{dd^{c}\phi^{\tau}}^{2}\omega_{0}^{n}$
-and we can thus proceed precisely as in the proof of Prop 3.4.
-Hence, the inf in $(i)$ is attained for a unique $v\in u\cdot\mathcal{G}\subset\mathcal{H}(X,\omega_{0})/\mathbb{R}$ and this $v$ is the unique critical point of the the pullback of $\mathcal{J}$ to $J\mathfrak{k}.$ Now, fix $W\in\mathfrak{k}$ and $v\in\mathcal{H}(X,\omega_{0})/\mathbb{R}$ and set $v_{t}\coloneqq v\cdot\exp(tJW).$ Then
-(3.33) $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=\int_{X}\left.\frac{dv_{t}}{dt}\right|_{t=0}\omega_{0}^{n}=\int_{X}\left\langle dv,JW\right\rangle\omega_{0}^{n}=\int_{X}\nabla v\cdot JW\omega_{0}^{n}$
-using, in the first equality, that the Futaki character vanishes (formula 3.17).
-Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$.
-Using that $JW$ is proportional to $\nabla h_{W},$ where $h_{W}$ is the Hamiltonian function corresponding to $W,$ we deduce that
+It thus follows, precisely as in the proof of Prop 3.4, that the inf over $\mathcal{G}$ in $(i)$ may be replaced by the inf over $J\mathfrak{k},$ under the identification of $\mathcal{K}\backslash\mathcal{G}$ with $J\mathfrak{k}.$ Now fix a non-trivial real vector field $W$ on $X$ induced by an element in $\mathfrak{k}\subset\mathfrak{g}.$ Let us first show that
+(3.32) $V\frac{d\mathcal{J}(v\cdot\exp(-tJW))}{dt}_{|t=0}=c_{n}\int_{X}v\Delta h_{W}\omega_{0}^{n}$
+for a positive constant $c_{n}$ only depending on $n.$ It is enough to consider the case when $v\in\mathcal{H}(X,-K_{X}),$ using an approximation argument, precisely as in the proof of Prop 3.4.
+Setting $v_{t}\coloneqq v\cdot\exp(-tJW),$
+(3.33) $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=\int_{X}\left.\frac{dv_{t}}{dt}\right|_{t=0}\omega_{0}^{n}=-\int_{X}\left\langle dv,JW\right\rangle\omega_{0}^{n}=-\int_{X}\nabla v\cdot JW\omega_{0}^{n}$
+using, in the first equality, that the Futaki character vanishes (formula 3.19).
+Expressing $\left\langle dv,JW\right\rangle$ as the scalar product $\nabla v\cdot JW$ between the vector field $JW$ and the gradient $\nabla v$ of $v$ with respect to the Kähler metric $\omega_{0}$ and using that $JW=c_{n}\nabla h_{W}$ for a positive constant $c_{n}$ (where $h_{W}$ denotes the Hamiltonian function corresponding to $W)$ we deduce that
 $V\frac{d\mathcal{J}(v_{t})}{dt}_{|t=0}=-c_{n}\int_{X}\nabla v\cdot\nabla h_{W}\omega_{0}^{n}=c_{n}\int_{X}v\Delta h_{W}\omega_{0}^{n},$
-Hence, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ holds, as desired.
-Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [8, Thm 2.2].
-Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.31, when $\omega_{0}$ is Kähler–Einstein.
+In particular, $\frac{d}{dt}\mathcal{J}(v_{t})=0$ for all $W\in\mathfrak{k}$ iff condition $(ii)$ in the present lemma holds.
+To prove that $(i)$ is equivalent to $(ii),$ together with the uniqueness statement in the lemma, it is thus enough to show that $\mathcal{J}(v_{t})$ is strictly convex for any given $W\in\mathfrak{k}.$ To this end we first note that formula 3.32 can be reformulated as
+$V\frac{d\mathcal{J}(v\cdot\exp(-tJW))}{dt}_{|t=0}=c_{n}\int_{X}h_{W}(\Delta v+n)\omega_{0}^{n-1}=\frac{c_{n}}{(n-1)!}\int_{X}h_{W}\omega_{v}\wedge\omega_{0}^{n-1},$
+showing that condition $(ii)$ is also equivalent to condition $(iii).$ Indeed, integrating by parts, we can replace the integrand $v\Delta h_{W}$ in formula 3.32 with $\Delta vh_{W}$ and then use that the integral of $h_{W}\omega_{0}^{n}$ vanishes, by Lemma 3.3, since the Futaki character of $X$ is assumed to vanish.
+It now follows that the function $t\mapsto\mathcal{J}(v\cdot\exp(-tJW))$ is strictly convex, precisely as in the proof of Prop 3.4, using that the measure $\omega_{v}\wedge\omega_{0}^{n-1}$ does not charge pluripolar subsets.
+Indeed, since $v\in\mathcal{E}^{1}(X,\omega)$ the measure defined as the product $\omega_{v}\wedge\omega_{0}^{n-1}$ of the current $\omega_{v}$ with the smooth form $\omega_{0}^{n-1}$ is identical to the non-pluripolar product of $\omega_{v}$ with $\omega_{0}^{n-1},$ which, by definition, does not charge pluripolar subsets (this identity follows, for example, from writing $v$ as a decreasing limit of elements in $\mathcal{E}^{1}(X,\omega)$ and using that both types of products are continuos wrt decreasing limits, since $v$ has full Monge-Ampère mass [25, Prop 2.1]).
+Next note that the pullback of $\mathcal{J}$ to $J\mathfrak{k}$ is proper and thus admits a critical point.
+Indeed, since $\mathcal{J}$ is comparable to $\mathcal{I}-\mathcal{J},$ which can be identified with the functional $E$ on $\mathcal{P}(X),$ the properness of $\mathcal{J}$ follows directly from the properness of $E,$ established in the course of the proof of Prop 3.4.
+Finally, when $\omega_{0}$ is Kähler–Einstein $\Delta h_{W}=\lambda_{1}h_{W}$ for any $W\in\mathfrak{k},$ where $\lambda_{1}$ denotes the smallest strictly positive eigenvalue of $\Delta$ [9, Thm 2.2].
+Since $\boldsymbol{m}=(h_{W_{1}},\dots,h_{W_{r}})$ for a basis $W_{i}$ in $\mathfrak{k}$ it follows that $(ii)$ is equivalent to Tian’s condition 3.30, when $\omega_{0}$ is Kähler–Einstein.
 ∎
-In the case when $n=1,$ so that $X=\mathbb{P}^{1}$, the functionals $\mathcal{I}-\mathcal{J}$ and $\mathcal{J}$ coincide and thus the moment condition appearing in Theorem 3.7 then coincides with condition $(ii)$ in the previous lemma.
 3.4. The reduced analytic stability thresholds of K-polystable log Fano curves
@@
 Indeed, $\mathcal{P}(X)_{0}$ is a closed compact subset of the compact space $\mathcal{P}(X),$ hence compact and, as is well-known, the map $\mu\mapsto\phi_{\mu}$ is a homeomorphism.
-It will thus be enough to show that $\mathcal{I}_{t,\mu_{0}}$ is continuous on $K.$ But since $\mathcal{I}_{t+\delta,\mu_{0}}$ is pointwise finite on $K$ for $\delta$ a sufficiently small positive number, this is shown precisely as in the proof of [10, Cor A.2], using the corresponding local result of Demailly–Kollar.
+It will thus be enough to show that $\mathcal{I}_{t,\mu_{0}}$ is continuous on $K.$ But since $\mathcal{I}_{t+\delta,\mu_{0}}$ is pointwise finite on $K$ for $\delta$ a sufficiently small positive number, this is shown precisely as in the proof of [11, Cor A.2], using the corresponding local result of Demailly–Kollar.
 We next prove the lower bounds on $\alpha(X,\Delta_{w})_{0}.$ First recall that for any function $\psi$ in $L_{\text{loc}}^{1}(\mathbb{C})$ we have
@@
 $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\sup\left\{\gamma\in]0,\infty[\colon F(-\gamma)_{0}>-\infty\right\}=2\alpha(X,\Delta_{w})_{0}$
-where $F(-\gamma)_{0}$ is defined by formula 3.15.
+where $F(-\gamma)_{0}$ is defined by formula 3.17.
 Moreover, $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ and $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}=\min\left\{\frac{1}{1-w},2\right\}$ when $w\neq 0.$
@@
 Let us first show that for any log Fano curve $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$ We will adapt the argument in [11] (concerning the case $\mathcal{G}$ trivial) to the present setup.
-Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.16 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
+Fix $\gamma>0$ and take $\mu\in\mathcal{P}(X)_{0}.$ Using that $\gamma u_{\mu}$ is a candidate for the sup appearing in the Legendre duality formula 3.18 for $D(\mu)$ gives, for any $t<\alpha(X,\Delta)_{0}$
 $D(\mu)\geq t\int_{X}u_{\mu}\mu-\log\mathcal{I}_{t,\mu_{0}}(u_{\mu})\geq t\int_{X}u_{\mu}\mu-C_{t},$
 using Lemma 3.12. Next, setting $u_{\mu}\coloneqq\phi_{\mu}-\phi_{0}$ (formula 3.8) we deduce, using $E(\mu)=-\int_{X}u_{\mu}\mu/V2$ (Example 3.2),
-$F_{\gamma}(\mu)\geq(-\gamma+2t)E(\mu)-C_{t},$
+$F_{-\gamma}(\mu)\geq(-\gamma+2t)E(\mu)-C_{t},$
 which is bounded from below if $\gamma\leq 2t,$ proving $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}\geq 2\alpha(X,\Delta)_{0}.$
@@
 Hence, we can take $\phi_{0}$ to be induced by the Fubini–Study metric.
-It will be enough to show that there exists a curve $\mu^{t}$ in $\mathcal{P}(X)_{0}$ such that for any $\gamma>2,$ $F_{\gamma}(\mu^{t})\rightarrow$-$\infty$ as $t\rightarrow\infty.$ To this end first note that if $\mu$ is $S^{1}$–invariant under the standard $S^{1}$–action of $\mathbb{P}^{1}$ fixing 0 and $\infty$ and $\mathbb{Z}_{2}$–invariant under the $\mathbb{Z}_{2}$–action $(x_{1},x_{2},x_{3})\mapsto(-x_{1},-x_{2},-x_{3})$ then $\boldsymbol{m}(x_{1},x_{2},x_{3})=-\boldsymbol{m}(-x_{1},-x_{2},-x_{3})$ and it follows that $\mu\in\mathcal{P}(X)_{0}.$ We will construct $\mu^{t}$ as $\mu^{t}=dd^{c}\phi_{t}/V,$ where $\phi_{t}$ is an $S^{1}$–invariant geodesic in $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}},$ using the notation in [16] (recalled in the appendix), where such a curve $\phi_{t}$ is realized as the Legendre transform of a function $v_{t}(y)$ on the moment polytope P=[V/2,V/2]P=[-V/2,V/2] of $(X,-K_{(X,\Delta)})$ of the form
+It will be enough to show that there exists a curve $\mu^{t}$ in $\mathcal{P}(X)_{0}$ such that for any $\gamma>2,$ $F_{-\gamma}(\mu^{t})\rightarrow$-$\infty$ as $t\rightarrow\infty.$ To this end first note that if $\mu$ is $S^{1}$–invariant under the standard $S^{1}$–action of $\mathbb{P}^{1}$ fixing 0 and $\infty$ and $\mathbb{Z}_{2}$–invariant under the $\mathbb{Z}_{2}$–action $(x_{1},x_{2},x_{3})\mapsto(-x_{1},-x_{2},-x_{3})$ then $\boldsymbol{m}(x_{1},x_{2},x_{3})=-\boldsymbol{m}(-x_{1},-x_{2},-x_{3})$ and it follows that $\mu\in\mathcal{P}(X)_{0}.$ We will construct $\mu^{t}$ as $\mu^{t}=dd^{c}\phi_{t}/V,$ where $\phi_{t}$ is an $S^{1}$–invariant geodesic in $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}},$ using the notation in [17] (recalled in the appendix), where such a curve $\phi_{t}$ is realized as the Legendre transform of a function $v_{t}(y)$ on the moment polytope P=[V/2,V/2]P=[-V/2,V/2] of $(X,-K_{(X,\Delta)})$ of the form
 (3.37) $\phi_{t}=v_{t}^{*},\,\,\,\,v_{t}\coloneqq v_{0}+tv,$
-where $v$ is a piecewise affine convex function on [1,1][-1,1] and $v_{0}$ is the Legendre transform of $\phi_{0}.$ It follows from well-known general results that such a curve $\phi_{t}$ is a geodesic ray in $\mathcal{H}(-K_{X})_{1,1}^{S^{1}}$ (see [41, Section 7]).
+where $v$ is a piecewise affine convex function on [1,1][-1,1] and $v_{0}$ is the Legendre transform of $\phi_{0}.$ It follows from well-known general results that such a curve $\phi_{t}$ is a geodesic ray in $\mathcal{H}(-K_{X})_{1,1}^{S^{1}}$ (see [52, Section 7]).
 Here we will take
@@
 Moreover, as shown in the appendix, for any $\phi\in\mathcal{H}(-K_{X})_{1,1}^{S^{1}},$ setting $\mu=dd^{c}\phi/V$,
-$F_{\gamma}(\mu)=\gamma\frac{\mathcal{L}(v)}{V}+(1-\gamma)\frac{V}{2}\int_{P}\left|v^{\prime}\right|dy-\int_{P}v^{\prime\prime}dy+O(1),\,\,\,\frac{\mathcal{L}(v)}{V}\coloneqq\int_{\partial P}\frac{v}{2}-\int_{P}v\frac{dy}{V},$
-where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 9.1, combined with 9.3).
+$F_{-\gamma}(\mu)=\gamma\frac{\mathcal{L}(v)}{V}+(1-\gamma)\frac{V}{2}\int_{P}\left|v^{\prime}\right|dy-\int_{P}v^{\prime\prime}dy+O(1),\,\,\,\frac{\mathcal{L}(v)}{V}\coloneqq\int_{\partial P}\frac{v}{2}-\int_{P}v\frac{dy}{V},$
+where $v^{\prime\prime}(y)$ denotes the absolutely continuous part of the second distributional derivatives $\partial^{2}v$ and $v^{\prime}$ is the function in $L_{\text{loc}}^{\infty}$ defined by the first distributional derivative $\partial v$ (Prop 10.1, combined with 10.3).
 In particular,
@@
 as $t\rightarrow\infty.$ Since $v_{t}^{\prime\prime}$ is independent of $t,$ it follows that
-$F_{\gamma}(\mu^{t})=\gamma t\frac{V}{4}+(1-\gamma)t\frac{V}{2}+O(1)=tV\left(\gamma\left(\frac{1}{4}-\frac{1}{2}\right)+\frac{1}{2}\right)+O(1),$
+$F_{-\gamma}(\mu^{t})=\gamma t\frac{V}{4}+(1-\gamma)t\frac{V}{2}+O(1)=tV\left(\gamma\left(\frac{1}{4}-\frac{1}{2}\right)+\frac{1}{2}\right)+O(1),$
 which tends to -$\infty$ if $\gamma>2,$ proving the inequality 3.36. It remains to show that when $w\in]0,1[$
@@
 (3.39) $\phi_{t}(x)=A_{-t}^{*}\phi_{0}=\phi_{0}(x+t).$
-The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 9.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
+The corresponding Legendre transform $v_{t}$ is given by $v_{t}(y)=v_{0}(y)-ty.$ Hence, by formula 10.2 in the appendix, we get as before (but now using that $\mathcal{L}(v_{t})=0)$
 $E(\mu^{t})=0+\frac{V}{2}t+O(1).$
@@
 and using the dominated convergence theorem. ∎
-The proof of the previous theorem shows that when $\gamma>\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ the functional $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is destabilized by a geodesic ray, i.e.
-$F_{\gamma}(\mu_{t})\rightarrow-\infty,$ as $t\rightarrow\infty,$ for a curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ of the form $\mu_{t}=\omega_{u_{t}}/V,$ where $u_{t}$ is a geodesic ray in $\mathcal{H}_{1,1}(X,\omega_{0}).$ This is reminiscent of the general existence/properness principle in [48, Section 3].
-However, in the present setup the function $F_{\gamma}(\omega_{u_{t}}/V)$ need not be convex in $t$ (since $\gamma>1)$ which is required in [48].
+The proof of the previous theorem shows that when $\gamma>\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ the functional $F_{-\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is destabilized by a geodesic ray, i.e.
+$F_{-\gamma}(\mu_{t})\rightarrow-\infty,$ as $t\rightarrow\infty,$ for a curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ of the form $\mu_{t}=\omega_{u_{t}}/V,$ where $u_{t}$ is a geodesic ray in $\mathcal{H}_{1,1}(X,\omega_{0}).$ This is reminiscent of the general existence/properness principle in [61, Section 3].
+However, in the present setup the function $F_{-\gamma}(\omega_{u_{t}}/V)$ need not be convex in $t$ (since $\gamma>1)$ which is required in [61].
 Moreover, our geodesics are required to satisfy the constraint $\int_{X}\boldsymbol{m}\omega_{u_{t}}=0.$
@@
 $\overrightarrow{\delta}_{0}(X,\Delta)\coloneqq\inf_{\mathcal{R}_{1,1}(X,\omega_{0})_{0}}\frac{\boldsymbol{D}}{\boldsymbol{E}},$
-where $E(u_{t})\coloneqq E(\omega_{u_{t}}^{n}/V)\coloneqq(\mathcal{I}-\mathcal{J})(u_{t}),$ $D(u_{t})\coloneqq D(\omega_{u_{t}}^{n}/V|\mu_{0})$ and $(\boldsymbol{D}/\boldsymbol{E})(\boldsymbol{u})\coloneqq\infty$ if $\boldsymbol{E}(\boldsymbol{u})=0.$ We recall that, $E(u_{t})$ and $D(u_{t})$ are convex with respect to $t$ (by the proof of [21, Lemma 9.1] and [27, Thm 4.10], respectively) and thus $\boldsymbol{D}$ and $\boldsymbol{E}$ are well-defined.
+where $E(u_{t})\coloneqq E(\omega_{u_{t}}^{n}/V)\coloneqq(\mathcal{I}-\mathcal{J})(u_{t}),$ $D(u_{t})\coloneqq D(\omega_{u_{t}}^{n}/V|\mu_{0})$ and $(\boldsymbol{D}/\boldsymbol{E})(\boldsymbol{u})\coloneqq\infty$ if $\boldsymbol{E}(\boldsymbol{u})=0.$ We recall that, $E(u_{t})$ and $D(u_{t})$ are convex with respect to $t$ (by the proof of [22, Lemma 9.1] and [29, Thm 4.10], respectively) and thus $\boldsymbol{D}$ and $\boldsymbol{E}$ are well-defined.
 The following proposition shows, in particular, that $\overrightarrow{\delta}_{0}(X,\Delta_{w})=\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}.$
@@
 4. Applications to logarithmic HLS inequalities, Moser–Trudinger inequalities and spontaneous symmetry breaking
-We continue with the setup in the previous section applied to the case when $X$ is the Riemann sphere and $\Delta=\Delta_{w}$ for $w\in[0,1].$ The lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ resulting from Theorem 3.13 is equivalent to the inequality
+We continue with the setup in the previous section applied to the case when $X$ is the Riemann sphere and $\Delta=\Delta_{w}$ for $w\in[0,1[.$ The lower bound on $\delta^{A}(X,\Delta_{w})^{\mathcal{G}}$ resulting from Theorem 3.13 is equivalent to the inequality
 (4.1) $E(\mu)\leq\frac{1}{\gamma}D(\mu|\mu_{0})+C_{\gamma}\,\,\forall\mu\in\mathcal{P}(X):\int_{X}\boldsymbol{m}\mu=0.$
-When $w=0$ and $\gamma=1$ this inequality is known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $X$ [37, 38, 9].
-In this case it holds without the moment constraint (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant).
-Thus Theorem 3.13 yields an improvement of the logarithmic HLS inequality, allowing any $\gamma\in]0,2[,$ under the constraint that $\mu$ has vanishing moment.
-In contrast to [37, 38, 9], in this case it does not seem possible to use symmetric decreasing rearrangements, due to a phenomenon of spontaneous symmetry breaking (see Section 4.1.1).
+When $\gamma=1$ the inequality holds without the moment constraint, i.e.
+it holds on all of $\mathcal{P}(X)$ (since $F_{-1}(\mu)$ is $\mathcal{G}$–invariant).
+If moreover, $w=0,$ the corresponding inequality on $\mathcal{P}(X)$ is then known as the logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere $S^{2}$ [43, 44, 10], which is the critical (logarithmic) case of the HLS inequality (as recalled in Section 5.2).
+Thus Theorem 3.13 yields, in particular, an improvement of the logarithmic HLS inequality on $S^{2}$ - allowing any $\gamma\in]0,2[$ under the constraint that $\mu$ has vanishing moment - that will be further sharpened in Section 5.
+In contrast to [43, 44, 10], in this case, it does not seem possible to use symmetric decreasing rearrangements.
+This is related to a phenomenon of spontaneous symmetry breaking, discussed in Section 4.1.1.
 To facilitate the comparison between our complex-geometric notation and the PDE-literature [43, 48, 44, 10, 80] we provide the following basic lemma, proved in the last section of the appendix:
 Lemma 4.1.
-(normalizations) Let $\omega_{\text{$FS$}}$ denote the Fubini-Study Kähler form on $\mathbb{P}^{1}$ (which is $SU(2)$-invariant and thus coincides with the $SO(3)$-invariant probability measure on $S^{2})$ and denote by $\Delta_{\text{FS}}$ the Laplacian on $\mathbb{P}^{1}$ defined by
+(normalizations) Let $\omega_{\text{$FS$}}$ denote the Fubini-Study Kähler form on $\mathbb{P}^{1}$ (which is $SU(2)$-invariant and thus coincides with the $\mathrm{SO}(3)$-invariant probability measure $\mu_{0}$ on $S^{2})$ and denote by $\Delta_{\text{FS}}$ the Laplacian on $\mathbb{P}^{1}$ defined by
 (4.2) $dd^{c}u=\Delta_{\text{FS}}u\omega_{\text{FS}},$
-Then $\Delta_{\text{FS}}$ coincides with the Laplacian $\Delta_{g_{S^{2}}}$ wrt the metric $g_{S^{2}}$ on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}.$ Moreover, when $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta_{w})})$ is $SU(2)$-invariant, i.e.
-when $\omega_{0}=V\omega_{\text{FS}}$ for V=(22w),V=(2-2w),
-(4.3) $E(\mu)=(1-w)\int_{S^{2}\times S^{2}}G\mu\otimes\mu,$
-where $G(\boldsymbol{x},\boldsymbol{y})$ is the Green function for $\Delta_{g_{S^{2}}},$ i.e. $G(\boldsymbol{x},\boldsymbol{y})=\log\left(\frac{|\boldsymbol{x}-\boldsymbol{y}|^{2}}{4}\right)+1.$
-In Section 5 we will establish a sharp version of the constrained logarithmic HLS inequality 4.1 when $w=0.$ But first, in this section we show that Theorem 3.13 implies a generalization of Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [7, Cor 2] to $w\neq 0.$ To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
+Then $\Delta_{\text{FS}}$ coincides with the Laplacian $\Delta_{g}$ wrt the metric $g$ on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}.$ Moreover, when $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta_{w})})$ is $SU(2)$-invariant, i.e.
+when $\omega_{0}=V\mu_{0}$ for V=(22w)V=(2-2w)
+(4.3) $E(\mu)=(1-w)\int_{S^{2}\times S^{2}}G\mu\otimes\mu=\frac{V}{2}\left\|\mu-\mu_{0}\right\|_{H^{-1}}^{2}=\frac{V}{2}\left\|(-\Delta_{g})^{1/2}(\mu-\mu_{0})\right\|_{L^{2}(S^{2},dV_{g})}^{2}$
+where $G(\boldsymbol{x},\boldsymbol{y})$ is the Green function for $\Delta_{g_{S^{2}}},$ i.e.
+$G(\boldsymbol{x},\boldsymbol{y})=\log\left(\frac{|\boldsymbol{x}-\boldsymbol{y}|^{2}}{4}\right)+1$ and $\left\|\mu-\nu\right\|_{H^{-1}}^{2}$ is the Hilbert norm on the Hilbert space $H^{-1}(\mathbb{P}^{1})$ of signed measures on $\mathbb{P}^{1}$ with zero mean that is dual to the Hilbert space $H^{1}(\mathbb{P}^{1}):=L^{2,1}(\mathbb{P}^{1})/\mathbb{R},$ endowed with the Hilbert norm $\left\|v\right\|_{H^{1}}^{2}:=\int_{\mathbb{P}^{1}}dv\wedge d^{c}v.$
+In Section 5 we will establish a sharp version of the constrained logarithmic HLS inequality 4.1 when $w=0.$ But first, in this section we show that Theorem 3.13 implies a generalization of Aubin’s improvement of the Moser–Trudinger inequality on $S^{2}$ [8, Cor 2] to $w\neq 0.$ To this end, first recall that, given a log Fano manifold $(X,\Delta)$ and a reference metric $\phi_{0}\in\mathcal{H}(-K_{(X,\Delta)}),$ the twisted Ding functional on $\mathcal{H}(X,\omega_{0})$ is defined by
 (4.4) $\mathcal{D}_{\gamma}(u)\coloneqq-\mathcal{E}(u)-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0}.$
-It was introduced in [51] (when $\Delta=0)$ [^3] In the notation of [51], $\mathcal{D}_{\gamma}(u)=F_{\gamma}(u),$ but here the notation $F_{\gamma}$ is reserved for the free energy functional..
+It was introduced in [64] (when $\Delta=0)$ [^3] In the notation of [64], $\mathcal{D}_{\gamma}(u)=F_{-\gamma}(u),$ but here the notation $F_{-\gamma}$ is reserved for the free energy functional..
 Here we will consider the case of a log Fano curve and allow $u\in C^{1}(X)$ (thus, importantely, we do not assume $\omega_{u}\geq 0).$ In this case $\mathcal{E}(u)$ is given by the classical expression in Example 3.2.
@@
 Next, to destabilize $\mathcal{D}_{\gamma}(u),$ first recall [11] that for any $u\in\mathcal{H}(X,\omega_{0})$
-(4.7) $\mathcal{D}_{\gamma}(u)\leq\frac{1}{\gamma}F_{\gamma}\left(\frac{\omega_{u}}{V}\right).$
-In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed, $F_{\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$ using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e.
+(4.7) $\mathcal{D}_{\gamma}(u)\leq\frac{1}{\gamma}F_{-\gamma}\left(\frac{\omega_{u}}{V}\right).$
+In fact, this inequality holds also when $n>1$ by replacing $\omega_{u}$ with $\omega_{u}^{n}.$ Indeed,
+(4.8) $F_{-\gamma}(\frac{\omega_{u}}{V})-\mathcal{D}_{\gamma}(u)=D(\frac{\omega_{u}^{n}}{V}|\frac{\mu_{\phi_{0}+u}}{\int\mu_{\phi_{0}+u}})\geq 0,$
+using that $D(\cdot|\cdot)\geq$ on $\mathcal{P}(X)^{2}.$ All that remains is thus to show that the curve $\mu_{t}$ in $\mathcal{P}(X)_{0}$ destabilizing $F_{-\gamma},$ constructed in the proof of Theorem 3.13, has the property that it can be expressed as $\mu_{t}=\omega_{u_{t}}/V$ for $u_{t}$ in $\in C^{1,1}(X)_{0}\cap\text{PSH}(X,\omega_{0}).$ To this end first consider the case when $w<1/2,$ i.e.
 $\min\{\frac{1}{1-w},2\}<2.$ Then $u_{t}:=\phi_{t}-\phi_{0},$ where $\phi_{t}$ is defined by 3.39, has the desired properties.
@@
 […]
-is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint [^4] Strictly speaking, Aubin’s inequalities [7, Cor 2] are obtained by replacing $\varphi$ with $\varphi$:
+is equivalent to the following Moser–Trudinger inequality with Aubin’s constraint [^4] Strictly speaking, Aubin’s inequalities [8, Cor 2] are obtained by replacing $\varphi$ with $\varphi$:
 […]
@@
 Now specialize to the case when $\Delta=0$ and $\phi_{0}$ is $\mathrm{SU}(2)$–invariant, so that $\boldsymbol{m}=\boldsymbol{x}$ when $X$ is identified with the unit sphere in $\mathbb{R}^{3}$ (see Example 3.10).
-In this case Aubin [7, Cor 2] showed that the constrained Moser–Trudinger inequality 4.8 holds for any $\gamma<2.$ Moreover, in [55] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
+In this case Aubin [8, Cor 2] showed that the constrained Moser–Trudinger inequality 4.9 holds for any $\gamma<2.$ Moreover, in [71] it was shown that it fails for $\gamma>2,$ by constructing an explicit destabilizing curve $v_{t}.$ However, in contrast to the corollary above, the corresponding curve $u_{t}$ destabilizing $\mathcal{D}_{\gamma}$ does not appear to be contained in $\text{PSH}(X,\omega_{0}).$
 Building on Aubin’s result the following sharp form of the the Moser–Trudinger inequality with Aubin’s constraint was established in [80], that settles a conjecture in [48]:
 Theorem 4.5.
-[72] Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.8 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
-In Theorem 5.1 we will prove strengthening of this result, saying that the minimum of the corresponding free energy $F_{\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is attained at the $SO(3)$-invariant probability measure for any $\gamma\in]0,2]$ (by the inequality 4.6 this indeed implies Theorem 4.5).
-But we first give some applications of Theorem 4.5 in Kähler geometry and indicate why the constrained Moser-Trudinger inequality in Theorem 4.5 is, a priori, weaker than the sharp lower bound on $F_{\gamma}$ that will be established in Theorem 5.1.
+[80] Assume that $X$ is the two-sphere, $\Delta=0$ and $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then the constrained Moser–Trudinger inequality 4.9 holds for $\gamma\in]0,2]$ and $C_{\gamma}=0.$
+In Theorem 5.1 we will prove strengthening of this result, saying that the minimum of the corresponding free energy $F_{-\gamma}(\mu)$ on $\mathcal{P}(X)_{0}$ is attained at the $SO(3)$-invariant probability measure for any $\gamma\in]0,2]$ (by the inequality 4.6 this indeed implies Theorem 4.5).
+But we first give some applications of Theorem 4.5 in Kähler geometry and indicate why the constrained Moser-Trudinger inequality in Theorem 4.5 is, a priori, weaker than the sharp lower bound on $F_{-\gamma}$ that will be established in Theorem 5.1.
 4.0.1. Application to Moser–Trudinger inequalities on $S^{2}$ with Tian’s constraint
@@
 (4.10) $\mathcal{D}_{1}(u)\geq\epsilon(\mathcal{I}-\mathcal{J})(u)-C$
-for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.31.
-The conjecture was settled in [48] (with non-effective constants $\epsilon$ and $C)$ and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.8 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.31 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [87].
-Here we show, using Theorem 4.5, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.5 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.9):
+for any $u\in\mathcal{H}(X,\omega_{0})$ satisfying the orthogonality condition 3.30.
+The conjecture was settled in [61] (with non-effective constants $\epsilon$ and $C)$ and, as a consequence, a new Moser–Trudinger inequality was obtained when $X=S^{2}$, saying that the inequality 4.9 holds for some $\gamma>1$ if Aubin’s constraint on $v$ is replaced by Tian’s constraint, demanding that $v$ satisfies the orthogonality condition 3.30 and $v\in\mathcal{H}(X,\omega_{0}).$ It should be emphasized that – in contrast to Aubin’s constraint – for Tian’s constraint one needs to assume that $v\in\mathcal{H}(X,\omega_{0}),$ as shown in [105].
+Here we show, using Theorem 4.5, that the Moser–Trudinger inequality on $S^{2}$ in Theorem 4.5 also holds for $\gamma\leq 2$ and $C_{\gamma}=0$ when Aubin’s constraint is replaced by Tian’s constraint (this corresponds to $\epsilon=1/2$ and $C=0$ in the inequality 4.10):
 Proposition 4.6.
-Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X).$ For any $\varphi$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.31
+Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X).$ For any $\varphi$ in $\mathcal{H}(X,\omega_{0})$ satisfying Tian’s orthogonality condition 3.30
 $\frac{1}{2}\mathcal{J}(\varphi)-\int_{X}\varphi\frac{\omega_{0}}{2}-\log\int_{X}e^{-\varphi}\frac{\omega_{0}}{2}\geq 0,$
-i.e. the inequality is saturated when $\varphi$ is constant. Equivalently,
+Equivalently,
 (4.11) $\mathcal{D}_{1}(\varphi)\geq\frac{1}{2}\mathcal{J}(\varphi).$
@@
 $\mathcal{D}_{1}(u)=\mathcal{D}_{1}(u\cdot g_{0})\geq\frac{1}{2}\mathcal{J}(u\cdot g_{0})\geq\frac{1}{2}\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g)$
-by applying the inequality 4.10 to $\varphi\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.10 indeed holds if $u$ satisfies Tian’s constraint.
+by applying the inequality 4.11 to $\varphi\coloneqq u\cdot g_{0}.$ Since $\mathcal{J}(u)=E(\omega_{u}/V)$ it follows from Prop 3.4 that if $u\in\mathcal{H}(X,\omega_{0})$ and the integral of $\boldsymbol{x}\omega_{u}$ vanishes (which in the present case amounts to Tian’s constraint, by Lemma 3.9), then the inf above is attained when $g$ is the identity, showing that the inequality 4.11 indeed holds if $u$ satisfies Tian’s constraint.
 ∎
@@
 Let $X$ be the two-sphere and $\omega_{0}$ the $\mathrm{SU}(2)$–invariant Kähler form in $c_{1}(X)$ and assume $\gamma\leq 3/2.$ Then, for all measures $\mu$ in $\mathcal{P}(X)_{0}$
-$F_{\gamma}(\mu)\geq F_{\gamma}\left(\frac{\omega_{0}}{V}\right)(=0)$
+$F_{-\gamma}(\mu)\geq F_{-\gamma}\left(\frac{\omega_{0}}{V}\right)(=0)$
 If it is moreover assumed that $\mu\geq(1-1/\gamma)\omega_{0}/V,$ then the previous inequality holds when $\gamma\in[0,2].$
 Proof.
-Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.7 with the inequality 4.10 gives
-$F_{1}\left(\frac{\omega_{u}}{V}\right)\geq\frac{1}{2}\mathcal{J}(u)\,\,\,\,\text{if $\int\boldsymbol{x}\omega_{u}=0$. }$
-Since $\mathcal{J}(u)=E(\omega_{u}/V)$ this equivalently means that $F_{1}(\omega_{u}/V)-E(\omega_{u}/V)/2\geq 0.$ In other words, $F_{3/2}(\omega_{u}/V)\geq 0.$ Using that $F_{\gamma}(\mu)$ is decreasing with respect to $\gamma$ (since $E\geq 0)$ this proves the inequality in the corollary.
+Expressing $\mu=\omega_{u}/V$ and combining the inequality 4.7 with the inequality 4.11 gives
+$F_{-1}\left(\frac{\omega_{u}}{V}\right)\geq\frac{1}{2}\mathcal{J}(u)\,\,\,\,\text{if $\int\boldsymbol{x}\omega_{u}=0$. }$
+Since $\mathcal{J}(u)=E(\omega_{u}/V)$ this equivalently means that $F_{-1}(\omega_{u}/V)-E(\omega_{u}/V)/2\geq 0.$ In other words, $F_{-3/2}(\omega_{u}/V)\geq 0.$ Using that $F_{-\gamma}(\mu)$ is decreasing with respect to $\gamma$ (since $E\geq 0)$ this proves the inequality in the corollary.
 To prove the last statement first note that it follows from the previous proposition, combined with the scaling argument in the proof of Lemma 4.4, that
@@
 Furthermore, the condition that $\gamma\varphi\in\mathcal{H}(X,\omega_{0})$ is equivalent to $\omega_{\varphi}\geq(1-1/\gamma)\omega_{0},$ whose right hand side is non-negative, by assumption.
-We can thus apply the inequality 4.7 to $\varphi$ to conclude that $F_{\gamma}(\frac{\omega_{\varphi}}{V})\geq 0.$
+We can thus apply the inequality 4.7 to $\varphi$ to conclude that $F_{-\gamma}(\frac{\omega_{\varphi}}{V})\geq 0.$
 4.1. Spontaneous symmetry breaking
@@
 (4.12) $\infty=\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}<\inf_{\mathcal{P}(X)_{0}^{S^{1}}}F_{-\gamma},\,\,\,-\infty=\inf_{C^{1}(X)_{0}}\mathcal{D}_{\gamma}(u)<\inf_{C^{1}(X)_{0}^{S^{1}}}\mathcal{D}_{\gamma}(u)$
-although both $F_{\gamma}$ and $\mathcal{P}(X)_{0}$ are $S^{1}$–invariant (and likewise for $\mathcal{D}_{\gamma}$ and $C^{1}(X)_{0})$.
+although both $F_{-\gamma}$ and $\mathcal{P}(X)_{0}$ are $S^{1}$–invariant (and likewise for $\mathcal{D}_{\gamma}$ and $C^{1}(X)_{0})$.
 The point is that the destabilizing curve $\mu^{t}$, exhibited in the proof of Theorem 3.13, is not $S^{1}$–invariant (with respect to the $S^{1}$–action fixing 0 and $\infty).$ This is in contrast to the case when $\gamma=1,$ where the two infima above do coincide.
-Indeed, when $\gamma=-1$, the inf of $F_{\gamma}$ is attained at the unique Kähler–Einstein volume form in $\mathcal{P}(X)_{0}$ and this volume form is $S^{1}$–invariant (and likewise for $\mathcal{D}_{1}(u)).$
+Indeed, when $\gamma=-1$, the inf of $F_{-\gamma}$ is attained at the unique Kähler–Einstein volume form in $\mathcal{P}(X)_{0}$ and this volume form is $S^{1}$–invariant (and likewise for $\mathcal{D}_{1}(u)).$
 The previous phenomenon of spontaneous symmetry breaking is reinforced on the space $\mathcal{R}_{1,1}(X,\omega_{0})_{0}$ of geodesic rays with vanishing moment (defined in the beginning of Section 3.5).
@@
 The arrangement was defined with respect to Lebesgue measure on $\mathbb{R}^{2}$ in [43, 44] (using stereographic projection), while in [10] it was defined with respect to the $\mathrm{SO}(3)$–invariant measure on the two-sphere.
-Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{1}$ (and the arrangement in [9] decreases $F_{\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.11 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{\gamma}.$
+Both these two rearrangements define maps $\mathcal{P}(X)\rightarrow\mathcal{P}(X)^{S^{1}}$ decreasing the functional $F_{-1}$ (and the arrangement in [10] decreases $F_{-\gamma}$ for any $\gamma>0).$ However, neither of the rearrangements preserve the space $\mathcal{P}(X)_{0}$ (for any given weight $w).$ Moreover, the spontaneous symmetry breaking exhibited in formula 4.12 implies that when $w\in]0,1/2[$ and $\gamma\in]1/(1-w),2[$ there is no map $\mathcal{P}(X)_{0}\rightarrow\mathcal{P}(X)_{0}^{S^{1}}$ that decreases the functional $F_{-\gamma}.$
 4.2. Restricting to $\mathcal{K}$-invariant metrics
@@
 $\delta_{C(\mathcal{G})}^{A}(X,\Delta)\coloneqq 1+\sup\left\{\epsilon\in\mathbb{R}\colon\exists C_{\epsilon}\in\mathbb{R}\,\,\mathcal{M}\geq\epsilon\inf_{g\in C(g)}(\mathcal{I}-\mathcal{J})-C_{\epsilon}\,\,\text{on\,$\mathcal{H}(X,\omega_{0})^{\mathcal{K}}$}\right\}$
-By [77, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $(X,\Delta)$ admits a KE metric.
-Hence, by [48, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $\delta_{\mathcal{G}}^{A}(X,\Delta)>1.$ However, in general these two analytic invariants may differ:
+By [95, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $(X,\Delta)$ admits a KE metric.
+Hence, by [61, Thm 2.18], $\delta_{C(\mathcal{G})}^{A}(X,\Delta)>1$ iff $\delta_{\mathcal{G}}^{A}(X,\Delta)>1.$ However, in general these two analytic invariants may differ:
 Proposition 4.9.
@@
 ∎
-5. A sharp constrained log HLS inequality on $S^{2}$
-In this section we will prove the following result, which is equivalent to Theorem 1.4 stated in the introduction (by formula 3.13 and Prop 3.4):
+5. A sharp constrained log HLS inequality on $S^{2}$ and application to quantitative stability
+In this section we will prove the following result, which is equivalent to Theorem 1.6 stated in the introduction (by formulas 3.15, 3.11):
 Theorem 5.1.
-Let $X$ be the two-sphere, $\Delta=0$ and assume that $\phi_{0}$ is $\mathrm{SO}(3)$–invariant. Then for any given $\gamma\in]0,2]$
-$F_{\gamma}(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0},$
-[…]
-When $\gamma=1$ the inequality for $F_{1}(\mu)$ follows from the general fact that, when a Fano manifold $X$ admits a Kähler–Einstein metric $\omega_{\text{KE}}\in C^{1}(X),$ the Mabuchi functional $\mathcal{M}$ is minimized on the Kähler potential $u$ of $\omega_{\text{KE}}$ [8] and $\mathcal{M}$ is then $\mathcal{G}$-invariant – so that the moment condition on $\mu$ can be dispensed with.
-In the case of the two-sphere (or more generally, the unit sphere in $\mathbb{R}^{n+1})$ the inequality $F_{1}(\mu)\geq 0$ on $\mathcal{P}(X)$ is shown in [37, 38, 9] using symmetrization (the inequality $F_{\gamma}(\mu)\geq 0$ for $\gamma\in[0,1]$ then follows directly).
+Let $X$ be the two-sphere (viewed as the log pair $(X,0)$) and assume that $\omega_{0}\in c_{1}(-K_{X})$ is $\mathrm{SO}(3)$–invariant. Then for any given $\gamma\in]0,2]$
+(5.1) $F_{-\gamma}(\mu):=-\gamma E(\mu)+D(\mu)\geq 0\,\,\,\forall\mu\in\mathcal{P}(X)_{0},$
+where $\mathcal{P}(X)_{0}$ denotes the space of all probability measures on $X$ satisfying the moment condition $\int_{X}$$\boldsymbol{x}\mu=0$.
+Moreover, equality holds iff $\mu$ is the $\mathrm{SO}(3)$–invariant probability measure on $X.$ By Prop 3.4, this equivalently means that for any $\mathcal{\mu\in P}(X)$ such that $E(\mu)<\infty$
+(5.2) $F_{-1}(\mu)\geq 1\cdot\inf E_{g\in\mathcal{G}}(g^{*}\mu)$
+with equality iff $\mu$ is in the $\mathcal{G}$-orbit of $\mu_{0}.$
+We recall that, by Theorem 3.13, $F_{-\gamma}(\mu)$ is unbounded from below when $\gamma>2,$ which implies (by Prop 3.4 that the constant $1$ in the inequality 5.2 cannot be replaced by a strictly larger constant.
+When $\gamma=1$ the inequality 5.1 follows from the general fact that, when a Fano manifold $X$ admits a Kähler–Einstein metric $\omega_{\text{KE}}\in C^{1}(X),$ the Mabuchi functional $\mathcal{M}$ is minimized on the Kähler potential $u$ of $\omega_{\text{KE}}$ [9] and $\mathcal{M}$ is then $\mathcal{G}$-invariant – so that the moment condition on $\mu$ can be dispensed with.
+In the case of the two-sphere (or more generally, the unit sphere in $\mathbb{R}^{n+1})$ the inequality $F_{-1}(\mu)\geq 0$ on $\mathcal{P}(X)$ is shown in [43, 44, 10] using symmetrization (the inequality $F_{-\gamma}(\mu)\geq 0$ for $\gamma\in[0,1]$ then follows directly).
 Before embarking on the detailed proof of Theorem 5.1 it may be illuminating to compare its structure to the proof of Theorem 4.5 in [80].
@@
 Since this is an infinite dimensional Hilbert manifold it follows that $u_{\gamma}$ satisfies an Euler–Lagrange equation, involving the differential of $\mathcal{D}_{\gamma}$ and three Lagrange multipliers (incorporating Aubin’s constraint).
-As shown in [39, Lemma 3.1], the Lagrange multipliers vanish and, as a result, $u_{\gamma}$ is, in fact, a critical point of $\mathcal{D}_{\gamma}$ on $W^{2,1}(X)$ and thus satisfies a Liouville equation on $S^{2},$ that is shown to have only have constant solutions in [72] (thus confirming a conjecture in [39]).
-In order to prove Theorem 5.1 we replace $\mathcal{D}_{\gamma}$ with the free energy functional $F_{\gamma}$ on $\mathcal{P}(X)_{0}.$ It follows from the formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ in Theorem 3.13 that $F_{\gamma}$ admits a minimizer $\mu_{\gamma}$ in $\mathcal{P}(X)_{0}$ for any given $\in]0,2[.$ However, since the space $\mathcal{P}(X)$ has a boundary it seems hard to directly show that $\mu_{\gamma}$ satisfies an appropriate equation.
-Instead, we leverage a duality formula, expressing the infimum of $F_{\gamma}$ over $\mathcal{P}(X)_{0}$ as the infimum over $W^{2,1}(X)$ of a perturbation of the twisted Ding functional:
-$\inf_{\mathcal{P}(X)_{0}}F_{\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),$
-where $\delta_{\gamma}$ is an entropic distance, defining a function on $W^{2,1}(X)$ taking values in $0,\infty[,$ that vanishes precisely when $\gamma u$ satisfies Aubin’s constraint 4.8 (in particular, the unboundedness from above of $\delta_{\gamma}(u)$ compensates for the unboundedness from below of $\mathcal{D}_{\gamma}(u)).$ This bypasses the boundary problem and allows us to reduce the proof of Theorem 5.1 to the uniqueness result in [72].
+As shown in [48, Lemma 3.1], the Lagrange multipliers vanish and, as a result, $u_{\gamma}$ is, in fact, a critical point of $\mathcal{D}_{\gamma}$ on $W^{2,1}(X)$ and thus satisfies a Liouville equation on $S^{2},$ that is shown to have only have constant solutions in [80] (thus confirming a conjecture in [48]).
+In order to prove Theorem 5.1 we replace $\mathcal{D}_{\gamma}$ with the free energy functional $F_{-\gamma}$ on $\mathcal{P}(X)_{0}.$ It follows from the formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ in Theorem 3.13 that $F_{-\gamma}$ admits a minimizer $\mu_{\gamma}$ in $\mathcal{P}(X)_{0}$ for any given $\in]0,2[.$ However, since the space $\mathcal{P}(X)$ has a boundary it seems hard to directly show that $\mu_{\gamma}$ satisfies an appropriate equation.
+Instead, we leverage a duality formula, expressing the infimum of $F_{-\gamma}$ over $\mathcal{P}(X)_{0}$ as the infimum over $W^{2,1}(X)$ of a perturbation of the twisted Ding functional:
+$\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),$
+where $\delta_{\gamma}$ is an entropic distance, defining a function on $W^{2,1}(X)$ taking values in $0,\infty[,$ that vanishes precisely when $\gamma u$ satisfies Aubin’s constraint 4.9 (in particular, the unboundedness from above of $\delta_{\gamma}(u)$ compensates for the unboundedness from below of $\mathcal{D}_{\gamma}(u)).$ This bypasses the boundary problem and allows us to reduce the proof of Theorem 5.1 to the uniqueness result in [80].
 In fact, the proof will show that the inf in the right hand side is attained precisely when $u$ is constant and thus vanishes.
 Remark 5.2.
-A consequence of the proof is that the infimum of $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ over $W^{2,1}(X)$ coincides with the one obtained by replacing the functional $\delta_{\gamma}(u)$ with the indicator function $\chi(u)$ attached to Aubin’s constraint (i.e.
-$\chi(u)$ vanishes when $\gamma u$ satisfies Aubin’s constraint and $\chi(u)=\infty$ otherwise).
-From a general optimization perspective this is somewhat miraculous.
+A consequence of the proof is that
+(5.3) $\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right)=\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\chi(\gamma u)\right)$
+where $\chi$ is the indicator function attached to Aubin’s constraint, i.e.
+$\chi(v)$ vanishes when $v$ satisfies Aubin’s constraint and $\chi(v)=\infty$ otherwise.
+From a general optimization perspective - where the infimum in the left hand side in equation 5.3 can be viewed as a relaxation of the constained infimum in the right hand side - this is somewhat miraculous.
+Indeed, a priori, the left hand side of is merely smaller then or equal to the right hand side, since $\delta_{\gamma}(u)\leq\chi(\gamma u)$ and the infimum in the left hand side could even equal -$\infty$ when the right hand side is finite.
+This is a reflection of the fact that Theorem 1.6 is, a priori, stronger than Theorem 4.5.
 5.1. Proof of Theorem 5.1
@@
 which defines a probability measure on $X,$ by the previous lemma.
-It will be enough to prove the inequality in Theorem 5.1 for a fixed $\gamma\in]1,2[.$ By Theorem 3.13 the infimum of $F_{\gamma}$ on $\mathcal{P}(X)_{0}$ is then finite.
+It will be enough to prove the inequality in Theorem 5.1 for a fixed $\gamma\in]1,2[.$ By Theorem 3.13 the infimum of $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ is then finite.
 By the previous lemma we thus have
-$\inf_{\mathcal{P}(X)_{0}}F_{\gamma}=\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{\gamma}>-\infty.$
+$\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}=\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{-\gamma}>-\infty.$
 5.1.1. A duality formula
 We start by proving the following formula, which applies in the general setup of a log Fano curve $(X,\Delta):$
-(5.1) $\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),\,\,\,\delta_{\gamma}(u)\coloneqq\frac{1}{\gamma}\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\mu_{\gamma u}),$
+(5.4) $\inf_{\mathcal{P}(X)_{0}\cap W^{2,1*}}F_{-\gamma}=\gamma\inf_{u\in W^{2,1}(X)}\left(\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)\right),\,\,\,\delta_{\gamma}(u)\coloneqq\frac{1}{\gamma}\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\mu_{\gamma u}),$
 where $\mathcal{D}_{\gamma}(u)$ is defined by formula 4.4.
@@
 $D(\mu|\mu_{0})+\gamma\left\langle u,\mu\right\rangle=D(\mu|\mu_{u})-\log\int_{X}e^{-\gamma u}\mu_{0}$
-then proves 5.4.
-Next, note that the functional $\tilde{D}(\mu)$ on $W^{2,1}(X)^{*}$ is convex and lsc (since $\mathcal{P}(X)_{0}$ is compact and convex in $W^{2,1}(X)^{*})$ and $D(\mu|\mu_{0})$ is convex on $\mathcal{P}(X)_{0}).$ It thus follows from combining formula 5.4 with the duality 5.2 that
+then proves 5.7.
+Next, note that the functional $\tilde{D}(\mu)$ on $W^{2,1}(X)^{*}$ is convex and lsc (since $\mathcal{P}(X)_{0}$ is compact and convex in $W^{2,1}(X)^{*})$ and $D(\mu|\mu_{0})$ is convex on $\mathcal{P}(X)_{0}).$ It thus follows from combining formula 5.7 with the duality 5.5 that
 $\frac{1}{\gamma}\tilde{D}=\mathcal{A}^{*},$
-so that formula 5.3 translates into
-$\inf_{\mathcal{P}(X)_{0}\cap W^{2,1}(X)^{*}}F_{\gamma}=\inf_{W^{2,1}(X)^{*}}\left(-\mathcal{E}^{*}+\mathcal{A}^{*}\right).$
+so that formula 5.6 translates into
+$\inf_{\mathcal{P}(X)_{0}\cap W^{2,1}(X)^{*}}F_{-\gamma}=\inf_{W^{2,1}(X)^{*}}\left(-\mathcal{E}^{*}+\mathcal{A}^{*}\right).$
+We will use the following essentially well-known identity:
 Lemma 5.4.
@@
 $\mathcal{A}^{**}(u):=\inf_{\mu\in H^{*}}\left(\mathcal{A}^{*}(\mu)+\left\langle u,\mu\right\rangle\right)\geq\inf_{\mu\in H^{*}}\left(\mathcal{E}^{*}(\mu)+\left\langle u,\mu\right\rangle\right)+C:=\mathcal{E}^{**}(u)+C$
-By 5.2, this means that $\mathcal{A}(u)\geq\mathcal{E}(u)+C$ for any $u\in H.$ Conversely, if the latter inequality holds, then, by taking the sup over all $u\in H$ we directly deduce that $\mathcal{A}^{*}(\mu)\geq\mathcal{E^{*}}(\mu)+C$ for all $\mu\in H^{*}.$ Thus $(-\mathcal{E}^{*}+\mathcal{A}^{*})(\mu)\geq C$ for all $\mu\in H^{*}.$
-In the present case $\mathcal{A}$ is a Legendre–Fenchel transform (by formula 5.4) and thus usc and concave on $W^{2,1}(X).$ Furthermore, $\mathcal{E}$ is also usc and concave on $W^{2,1}(X)$ (using the formula in Example 3.2 and the fact that the Hilbert norm on a Hilbert space is convex and continuous on the Hilbert space.
-Thus applying formula 5.5 concludes the proof of the duality formula 5.1.
+By 5.5, this means that $\mathcal{A}(u)\geq\mathcal{E}(u)+C$ for any $u\in H.$ Conversely, if the latter inequality holds, then, by taking the sup over all $u\in H$ we directly deduce that $\mathcal{A}^{*}(\mu)\geq\mathcal{E^{*}}(\mu)+C$ for all $\mu\in H^{*}.$ Thus $(-\mathcal{E}^{*}+\mathcal{A}^{*})(\mu)\geq C$ for all $\mu\in H^{*}.$
+In the present case $\mathcal{A}$ is a Legendre–Fenchel transform (by formula 5.7) and thus usc and concave on $W^{2,1}(X).$ Furthermore, $\mathcal{E}$ is also usc and concave on $W^{2,1}(X)$ (using the formula in Example 3.2 and the fact that the Hilbert norm on a Hilbert space is convex and continuous on the Hilbert space.
+Thus applying formula 5.8 concludes the proof of the duality formula 5.4.
 5.1.2. Existence of a dual minimizer in $W^{2,1}(X)$
-We next specialize to the log Fano curve $(X,\Delta)=(X,0)$ and take $\phi_{0}$ to be SU(2)SU(2)–invariant, so that $\mu_{0}$ and $\omega_{0}/2$ both coincide with $\omega_{\text{FS }}$ and hence with the $SO(3)$-invariant probability measure on $S^{2}.$ We will show that the functional $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ admits a minimizer in $W^{2,1}(X)$ (the same argument applies to any log Fano curve).
+We next specialize to the log Fano curve $(X,\Delta)=(X,0)$ and take $\phi_{0}$ to be $\mathrm{SU}(2)$–invariant, so that $\mu_{0}$ and $\omega_{0}/2$ both coincide with $\omega_{\text{FS }}$ and hence with the $\mathrm{SO}(3)$-invariant probability measure on $S^{2}.$ We will show that the functional $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ admits a minimizer in $W^{2,1}(X)$ (the same argument applies to any log Fano curve).
 To prove this we make a change of variable from $u$ to $v$ on $W^{2,1}(X),$ defined by
@@
 $\gamma\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)=\mathcal{F}_{\gamma}(v)+\delta(v).$
-Now take $\gamma^{\prime}>\gamma$ such that $\gamma^{\prime}<2.$ Combining Theorem 3.13 for $(X,0)$ with the duality formula 5.1 shows that there exists a constant $C$ such that
+Now take $\gamma^{\prime}>\gamma$ such that $\gamma^{\prime}<2.$ Combining Theorem 3.13 for $(X,0)$ with the duality formula 5.4 shows that there exists a constant $C$ such that
 $\gamma^{\prime}\left(\mathcal{D}_{\gamma^{\prime}}(u)+\delta{}_{\gamma^{\prime}}(u)\right)\geq-C$
@@
 Lemma 5.5.
-The uniform $W^{2,1}$-bound 5.6 implies, that, after passing to a subsequence,
+The uniform $W^{2,1}$-bound 5.9 implies, that, after passing to a subsequence,
 (5.10) $e^{2v_{j}}\omega_{0}\rightarrow e^{2v_{\infty}}\omega_{0}\,\,\,\left(e^{pv_{\infty}}\in L^{1}(X)\right),$
@@
 This is standard, but for completeness we provide a proof.
-First, by the Rellich–Kondrachov theorem, the uniform bound 5.6 implies that, after passing to a subsequence, $v_{j}\rightarrow v_{\infty}$ in $L^{1}(X).$ As a consequence, after again passing to a subsequence, we may assume that $v_{j}(x)\rightarrow v_{\infty}(x)$ for almost any $x$ in $X.$ Next, by the Moser-Trudinger inequality on $X$ the uniform bound 5.6 also implies that, for any given $\delta>0,$ there exists $A_{\delta}>0$ such that
+First, by the Rellich–Kondrachov theorem, the uniform bound 5.9 implies that, after passing to a subsequence, $v_{j}\rightarrow v_{\infty}$ in $L^{1}(X).$ As a consequence, after again passing to a subsequence, we may assume that $v_{j}(x)\rightarrow v_{\infty}(x)$ for almost any $x$ in $X.$ Next, by the Moser-Trudinger inequality on $X$ the uniform bound 5.9 also implies that, for any given $\delta>0,$ there exists $A_{\delta}>0$ such that
 (5.11) $\int e^{(1+\delta)2v}\omega_{0}\leq A_{\delta}$
@@
 $\int e^{2v_{j}}g\omega_{0}=\int_{\left\{v_{j}\leq M\right\}}e^{2v_{j}}g\omega_{0}+\epsilon_{j,M},\,\,\,\,\epsilon_{j,M}:=\int_{\left\{v_{j}>M\right\}}e^{2v_{j}}g\omega_{0}\leq\frac{A_{\delta}\left\|g\right\|_{L^{\infty}}}{e^{2\delta M}},$
-using the uniform bound 5.8 for some $\delta>0.$ Finally, letting first $j\rightarrow\infty$ and using the dominated convergence theorem and then letting $M\rightarrow\infty$ and using the monotone convergence theorem, shows that $\int e^{2v_{j}}g\omega_{0}$ converges towards $\int e^{2v_{\infty}}g\omega_{0},$ as desired.
+using the uniform bound 5.11 for some $\delta>0.$ Finally, letting first $j\rightarrow\infty$ and using the dominated convergence theorem and then letting $M\rightarrow\infty$ and using the monotone convergence theorem, shows that $\int e^{2v_{j}}g\omega_{0}$ converges towards $\int e^{2v_{\infty}}g\omega_{0},$ as desired.
 ∎
-All that remains is to show that $\delta(v)$ is lsc along the sequence $v_{j}.$ To this end, take $\mu_{j}\in\mathcal{P}(X)$ such that $D(\mu_{j}|e^{2v_{j}}\omega_{\text{FS}}/\int e^{2v_{j}}\omega_{\text{FS}})\leq\delta(v_{j})+1/j.$ Since $\mathcal{P}(X)$ is compact, we may, after passing to a subsequence, assume that $\mu_{j}\rightarrow\mu_{\infty}$ in $\mathcal{P}(X)$ for some $\mu_{\infty}.$ Since $D(\cdot|\cdot)$ is lsc with respect to the product topology on $\mathcal{P}(X)^{2}$ it thus follows from the weak convergence 5.7 that
+All that remains is to show that $\delta(v)$ is lsc along the sequence $v_{j}.$ To this end, take $\mu_{j}\in\mathcal{P}(X)$ such that $D(\mu_{j}|e^{2v_{j}}\omega_{\text{FS}}/\int e^{2v_{j}}\omega_{\text{FS}})\leq\delta(v_{j})+1/j.$ Since $\mathcal{P}(X)$ is compact, we may, after passing to a subsequence, assume that $\mu_{j}\rightarrow\mu_{\infty}$ in $\mathcal{P}(X)$ for some $\mu_{\infty}.$ Since $D(\cdot|\cdot)$ is lsc with respect to the product topology on $\mathcal{P}(X)^{2}$ it thus follows from the weak convergence 5.10 that
 $\delta(v_{\infty})\leq D(\mu_{\infty}|e^{2v_{\infty}}\omega_{\text{FS}}/\int e^{2v_{\infty}}\omega_{\text{FS}})\leq\liminf_{j\rightarrow\infty}\delta(v_{j}),$
@@
 $\left\langle f\right\rangle:=\int_{X}fe^{2v}\omega_{\text{FS}},$
-assuming that the normalization condition 5.9 holds. The vector $\boldsymbol{\lambda}\in\mathbb{R}^{3},$ depending on $v,$ is uniquely determined by the condition that
+assuming that the normalization condition 5.12 holds. The vector $\boldsymbol{\lambda}\in\mathbb{R}^{3},$ depending on $v,$ is uniquely determined by the condition that
 (5.14) $\int_{X}\boldsymbol{x}e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}e^{2v}\omega_{\text{FS}}=0.$
-To see this, we express
-$\delta(v)\coloneqq\delta(\nu)\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\nu),\,\,\,\nu\coloneqq e^{2v}\omega_{\text{FS}}$
-It follows from classical convex analysis that $\delta(\nu)$ is differentiable along smooth curves $\nu_{t}$ in $\mathcal{P}(X).$ Indeed, $D(\cdot|\nu)$ is strictly convex and not identically $\infty$ on the non-empty convex and compact subspace $\mathcal{P}(X)_{0}$ of $\mathcal{P}(X),$ which ensures that the inf above is attained at unique minimizer $\mu\in\mathcal{P}(X)_{0}.$ Moreover, when $\nu_{0}:=\nu,$
-$\frac{d\delta(\nu_{t})}{dt}_{|t=0}=\left\langle\log(\frac{\mu}{\nu}),\frac{d\nu_{t}}{dt}\right\rangle.$
-Furthermore, the minimizer $\mu$ is given by
+Formula 5.13 follows directly from the following basic lemma applied to $\nu=e^{2v}\omega_{\text{FS}}$ and the curve $\nu_{t}$ corresponding to $v+t\dot{v}.$
+Lemma 5.6.
+Given $\nu\in\mathcal{P}(X)$ consider the “entropic distance” from $\nu$ to $\mathcal{P}(X)_{0}$ defined by
+$\delta(\nu)\coloneqq\inf_{\mu\in\mathcal{P}(X)_{0}}D(\mu|\nu)$
+The function $\delta(\nu)$ is differentiable along smooth curves $\nu_{t}$ in $\mathcal{P}(X).$ Morover, the inf defining $\delta(\nu)$ is attained unique $\mu\in\mathcal{P}(X)_{0}$ and $\mu$ is given by
 (5.15) $\mu=\frac{e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu}{\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu},$
-where $\boldsymbol{\lambda}\in\mathbb{R}^{3}$ is uniquely determined by the condition that $e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu\in\mathcal{P}(X)_{0}.$ Indeed, the existence and uniqueness of $\boldsymbol{\lambda}$ follows directly from the fact that the function $\boldsymbol{\lambda}\mapsto\log\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu$ on $\mathbb{R}^{3}$ is strictly convex and proper (using that $\nu$ and $\omega_{\text{FS }}$are mutually absolutely continuous) and thus admits a critical point $\boldsymbol{\lambda}.$ Moreover, a direct computation reveals that the measure $\mu$ in $\mathcal{P}(X)_{0}$ defined by formula 5.12 is a critical point of $D(\cdot|\nu)$ in $\mathcal{P}(X)_{0}$ and thus minimizers $D(\cdot|\nu).$ Finally, taking $\nu_{t}$ to be the curve corresponding to $v+t\dot{v}$ proves formula 5.10.
+where $\boldsymbol{\lambda}\in\mathbb{R}^{3}$ is uniquely determined by the condition that $e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu\in\mathcal{P}(X)_{0}.$
+Proof.
+This is, without doubt, essentially well-known, but we provide a simple proof.
+First, since $D(\cdot|\nu)$ is strictly convex and not identically $\infty$ on the non-empty convex and compact subspace $\mathcal{P}(X)_{0}$ of $\mathcal{P}(X),$ inf above is attained at unique minimizer $\mu\in\mathcal{P}(X)_{0}.$ As a consequence, it follows from general principles (see the appendix in [23]), that for any curve $\nu^{t}$ coinciding with $\nu$ at $t=0,$
+$\frac{d\delta(\nu^{t})}{dt}_{|t=0}=\left\langle\log(\frac{\mu}{\nu}),\frac{d\nu^{t}}{dt}\right\rangle.$
+Next, the existence and uniqueness of $\boldsymbol{\lambda}$ follows directly from the fact that the function $\boldsymbol{\lambda}\mapsto\log\int e^{\boldsymbol{\lambda}\cdot\boldsymbol{x}}\nu$ on $\mathbb{R}^{3}$ is strictly convex and proper (using that $\nu$ and $\mu$ are mutually absolutely continuous) and thus admits a critical point $\boldsymbol{\lambda}.$ Moreover, a direct computation reveals that the measure $\mu$ in $\mathcal{P}(X)_{0}$ defined by formula 5.15 is a critical point of $D(\cdot|\nu)$ in $\mathcal{P}(X)_{0}$ and thus minimizers $D(\cdot|\nu).$
 Let now $v$ be a minimizer of $\mathcal{F}_{\gamma}+\delta$ on $W^{2,1}(X).$ In particular, the Gateaux differential of $\mathcal{F}_{\gamma}+\delta$ vanishes at $v,$ forcing
@@
 To conclude the proof it will be enough to show that any minimizer $v$ of $\mathcal{F}_{\gamma}+\delta$ on $W^{2,1}(X)$ is constant.
-By the uniqueness result for the Liouville equation on $S^{2}$ for $\gamma\in]1,2]$ in [72] it is enough to show that $\boldsymbol{\lambda}=0.$ To this end first recall that, by a result of Kazdan-Warner [39, formula 2.11] (and the normalizations in Lemma 4.1) any solution $u$ of the equation
+By the uniqueness result for the Liouville equation on $S^{2}$ for $\gamma\in]1,2]$ in [80] it is enough to show that $\boldsymbol{\lambda}=0.$ To this end first recall that, by a result of Kazdan-Warner [48, formula 2.11] (and the normalizations in Lemma 4.1) any solution $u$ of the equation
 $\Delta_{\text{FS}}u+he^{u}=c$
@@
 $\boldsymbol{\lambda}=(0,0,\lambda)$
-Indeed, taking $F$ to be a rotation of $S^{2}$ having the property that $F^{*}\boldsymbol{\lambda}\cdot\boldsymbol{x}=\lambda x_{3}$ for some $\lambda\in\mathbb{R}$ we can replace $v$ with $F^{*}v.$ Now, applying the Kazdan-Warner condition 5.13 to $i=3$ gives, by multiplying both sides with $\lambda,$
+Indeed, taking $F$ to be a rotation of $S^{2}$ having the property that $F^{*}\boldsymbol{\lambda}\cdot\boldsymbol{x}=\lambda x_{3}$ for some $\lambda\in\mathbb{R}$ we can replace $v$ with $F^{*}v.$ Now, applying the Kazdan-Warner condition 5.16 to $i=3$ gives, by multiplying both sides with $\lambda,$
 (5.17) $\left\langle\nabla(\lambda x_{3})\cdot\nabla(\lambda x_{3})\right\rangle=(2-2\gamma)\left\langle-\lambda x_{3}+\lambda^{2}x_{3}-\left\langle\lambda x_{3}\right\rangle\lambda x_{3}\right\rangle.$
@@
 $\left\langle\lambda x_{3}\right\rangle\leq 0.$
-Indeed, setting $f:=\lambda x_{3}$ we have, by the defining property 5.11 of $\lambda,$ that $\left\langle fe^{f}\right\rangle=0.$ But, by Jensen’s inequality
+Indeed, setting $f:=\lambda x_{3}$ we have, by the defining property 5.14 of $\lambda,$ that $\left\langle fe^{f}\right\rangle=0.$ But, by Jensen’s inequality
 $\frac{\left\langle fe^{f}\right\rangle}{\left\langle e^{f}\right\rangle}\geq\log\left\langle e^{f}\right\rangle\geq\left\langle f\right\rangle.$
-All in all this means that the bracket in the rhs of formula 5.14 is non-negative and strictly positive unless $\lambda=0.$ Since the same thing goes for the lhs in formula 5.14 and $(2-2\gamma)<0$ we deduce that $\lambda=0,$ as desired.
+All in all this means that the bracket in the rhs of formula 5.17 is non-negative and strictly positive unless $\lambda=0.$ Since the same thing goes for the lhs in formula 5.17 and $(2-2\gamma)<0$ we deduce that $\lambda=0,$ as desired.
 5.1.5. Any minimizer of $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ equals $\omega_{\text{FS}}$ for any $\gamma\in]1,2]$
 Given $\mu\in\mathcal{P}(X)$ such that $D(\mu)<\infty$ the following formula follows directly from the definitions:
-$\gamma^{-1}F_{\gamma}(\mu)=\mathcal{D}_{\gamma}(u_{\mu})+\gamma^{-1}D(\mu|\mu_{\gamma u_{\mu}}),$
+$\gamma^{-1}F_{-\gamma}(\mu)=\mathcal{D}_{\gamma}(u_{\mu})+\gamma^{-1}D(\mu|\mu_{\gamma u_{\mu}}),$
 where $u_{\mu}$ denotes the potential of $\mu$ which, by Lemma 5.3, is in $W^{2,1}(X),$ since $D(\mu)<\infty.$ Now assume furthermore that $\mu\in\mathcal{P}(X)_{0}.$ Then it follows directly from the definition 5.4 of $\delta_{\gamma}(u)$ that
-$\gamma^{-1}F_{\gamma}(\mu)\geq\mathcal{D}_{\gamma}(u_{\mu})+\delta_{\gamma}(u_{\mu}).$
-Thus, if $\mu$ minimizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ then, by the first part of Theorem 5.1, proved above and the the duality formula 5.1, $u_{\mu}$ minimizes $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ on $W^{2,1}(X).$ But, as shown in Section 5.1.4, this implies that $u_{\mu}$ is constant, i.e.
+(5.18) $\gamma^{-1}F_{-\gamma}(\mu)\geq\mathcal{D}_{\gamma}(u_{\mu})+\delta_{\gamma}(u_{\mu}).$
+Thus, if $\mu$ minimizes $F_{-\gamma}$ on $\mathcal{P}(X)_{0}$ then, by the first part of Theorem 5.1, proved above and the the duality formula 5.4, $u_{\mu}$ minimizes $\mathcal{D}_{\gamma}(u)+\delta_{\gamma}(u)$ on $W^{2,1}(X).$ But, as shown in Section 5.1.4, this implies that $u_{\mu}$ is constant, i.e.
 $\mu=\omega_{\text{FS}},$ as desired.
+5.2. Application to optimal stability results for the sharp log HLS inequality on $S^{2}$
+Starting with the seminal stability result [32] for the sharp Sobolev inequality in $\mathbb{R}^{n},$ for $n\geq 3,$ a wide variety of quantitative stability results for various sharp functional inequalities $\text{LHS }(f)\leq\text{RHS }(f)$ have been established of the form
+(5.19) $\text{RHS }(f)-\text{LHS }(f)\geq\kappa d(f,M),\,\,\,\,d(f,M)\coloneqq\inf_{g\in M}d(f,g)$
+where $M$ denotes the set of optimizers, i.e.
+all $f$ for which $\text{LHS }(f)=\text{RHS }(f),$ $\kappa$ is a positive constant, that we shall call the stability constant and $d(f,g)$ is a given divergence, i.e.
+$d(f,g)\geq 0$ with equality iff $f=g.$ In the original case of the Sobolev inequality [32], the divergence was the squared Dirichlet-norm, but $\kappa$ was non-explicit.
+This stability result was extended in [50] to all sharp fractional Sobolev inequalities in $\mathbb{R}^{n}$ of order $\alpha\in]0,n/2[$ (coinciding with the ordinary Sobolev inequality when $n\geq 3$ and $\alpha=1),$ with a (non-explicit) constant $\kappa_{\mathcal{S}_{n,\alpha}}$ and then to the corresponding dual inequalities, the Hardy–Littlewood–Sobolev inequalities (HLS) of order $\alpha\in]0,n/2[$ in [46] using duality:
+(5.20) $C_{n,\alpha}^{*}\left\|f\right\|_{L^{\frac{2n}{n+2\alpha}}(\mathbb{R}^{n})}^{2}-\left\|(-\Delta)^{-\alpha/2}f\right\|_{(\mathbb{R}^{n})}^{2}\geq\kappa_{\mathcal{S}_{n,\alpha}^{*}}\left\|f-M_{S_{n,\alpha}^{*}}\right\|_{L^{\frac{2n}{n+2\alpha}}(\mathbb{R}^{n})}^{2},$
+where $C_{n,\alpha}^{*}$ denotes the sharp constant in the HLS inequality (which is explicitly known), $M_{S_{n,\alpha}^{*}}$ is the corresponding $(n+2)$-dimensional submanifold of optimizers and $\kappa_{\mathcal{S^{*}}_{n,\alpha}}$ is explicitly expressed in terms of $\kappa_{S_{n,\alpha}}.$ Since, as shown in [44], differentiating the HLS inequalities at the critical value $\alpha=n/2$ (in an appropriate one-sided sense) yields the log HLS inequality it is, as pointed out in [46, Section 3.2], tempting to try to extract a stability result for the log HLS inequality from the stability result 5.20.
+However, this would require detailed information about the dependence of $\kappa_{S_{n,\alpha}}$ on $\alpha.$ In the case $n=2$ (so that $\alpha=1$ is the critical case) a quantitative stability result for the log HLS inequality was first established in [45] and recently considerably improved in [47], showing that, after stereographic projection to the two-sphere $S^{2}$ and using our notation,
+(5.21) $D(\mu)-E(\mu)\geq\frac{1}{8}\left\|\mu-M\right\|_{\text{TV}}^{2},\,\,\,\mu\in\mathcal{P}(S^{2}),\,\,D(\mu)<\infty,$
+where $M$ denotes the 33-dimensional submanifold of $\mathcal{P}(S^{2})$ consisting of all optimizers $\mu$ and $\left\|\mu-\nu\right\|_{\text{TV}}^{2}$ is the total variation norm.
+Since the entropy $D(\mu)$ arises when differentiating the $L^{\frac{2n}{n+2\alpha}}$-norm at the critical case $\alpha=n/2,$ it seems natural to ask if the the norm $\left\|\mu-\nu\right\|_{\text{TV}}^{2}$ can be replaced by the relative entropy $D(\mu,\nu),$ i.e.
+by the Kullback–Leibler divergence.
+In fact, as next shown, this follows from Theorem 5.1 with an optimal stability constant $\kappa$:
+Corollary 5.7.
+Denote by $\Delta$ the Laplacian wrt any Riemannian metric on $S^{2}$ that is conformally equivalent to the one induced by the Euclidean metric on $\mathbb{R}^{3}$ and by $dV$ the corresponding volume form.
+The following stability inequalities hold for any $\mu\in\mathcal{P}(S^{2}),$ such that $D(\mu)<\infty,$
+(5.22) $D(\mu)-E(\mu)\geq 1\cdot\left\|(-\Delta)^{1/2}(\mu-M)\right\|_{L^{2}(S^{2},dV_{g})}^{2},$
+and
+(5.23) $D(\mu)-E(\mu)\geq\frac{1}{2}D\left(\mu|M\right),$
+Moreover, the constants $1$ and $1/2$ are optimal, i.e. the stability inequalities above fail when these constants are replaced by any strictly larger constants.
+Proof.
+By Theorem 5.1
+$D(\mu)-E(\mu)\geq 1\cdot\inf_{g\in\mathcal{G}}E(g^{*}\mu),$
+where $1$ is the optimal constant.
+Moreover, by formula 4.3, $E(g^{*}\mu)=\left\|(g^{*}\mu-\mu_{0})\right\|_{H^{-1}}^{2},$ which equals $\left\|(\mu-g_{*}\mu_{0})\right\|_{H^{-1}}^{2},$ since the $H^{-1}$-norm is $\mathcal{G}$-invariant.
+Equivalently, $E(g^{*}\mu)=\left\|(-\Delta)^{1/2}(\mu-g_{*}\mu_{0})\right\|_{L^{2}(S^{2},dV)}^{2}.$ Since $M$ coincides with the $\mathcal{G-}$orbit of $\mu_{0}$ this proves the first stability inequality 5.22 and shows that $1$ is the optimal stability constant.
+To prove the second stability inequality 5.23 first assume that $\mu\in\mathcal{P}(X)_{0}.$ Then Theorem 5.1 for $\gamma=2$ yields the first inequality in the following equivalence
+$F_{-2}(\mu)\coloneqq-2E(\mu)+D(\mu)\geq 0\iff F_{-1}(\mu)\coloneqq D(\mu)-E(\mu)\geq\frac{1}{2}D(\mu).$
+Now take any $\mu$ in $\mathcal{P}(X)$ such that $D(\mu)<\infty.$ In particular, $E(\mu)<\infty.$ Hence, by Prop 3.4, there exists $g\in\mathcal{G}$ such that $g_{*}\mu\in\mathcal{P}(X)_{0}.$ Moreover, by Prop 3.4 $g_{*}\mu$ minimizes $E(\mu)$ on the $\mathcal{G}$-orbit of $\mu.$ But this means, equivalently, that $g_{*}\mu$ minimizes $D(\mu)$ on the $\mathcal{G}$-orbit of $\mu,$ since $D(\mu)-E(\mu)(\coloneqq F_{-1}(\mu))$ is $\mathcal{G}$-invariant.
+Hence,
+$F_{-1}(\mu)=F_{-1}(g_{*}\mu)\geq\frac{1}{2}D(g_{*}\mu)=\frac{1}{2}\inf_{g\in\mathcal{G}}D(g_{*}\mu)$
+Next observe that
+(5.24) $\inf_{g\in\mathcal{G}}D(g_{*}\mu)=D(\mu|M).$
+Indeed, $D(g_{*}\mu)\coloneqq D(g_{*}\mu|\mu_{0})=D(\mu|g^{*}\mu_{0})$ and $F_{-1}(\mu)=0$ iff there exists $g\in\mathcal{G}$ such that $\mu=g^{*}\mu_{0}.$ Combining the previous two inequalities thus yields the inequality in the corollary.
+Finally, to prove that the optimality of the stability constant $1/2$ assume, to get a contradiction, that there exists a small positive number $\epsilon$ such that
+$F_{-1}(\mu)\geq\frac{1}{2-\epsilon}D(\mu|M).$
+By the previous argument this means, equivalently, that $D(\mu)-E(\mu)\geq\frac{1}{2-\epsilon}D(\mu)$ on $\mathcal{P}(X)_{0}$ i.e.
+that $D(\mu)-\gamma E(\mu)\geq 0$ on $\mathcal{P}(X)_{0}$ for $\gamma\coloneqq(2-\epsilon)/(1-\epsilon)>2,$ which contradicts Theorem 3.13 saying, in particular, that $F_{-\gamma}(\mu)$ is unbounded from below on $\mathcal{P}(X)_{0}$ when $\gamma>2.$
+5.2.1. Comparison with the stability inequality in [47] and stability of the sharp Moser-Trudinger inequality
+Combining the previous corollary with the sharp Pinsker inequality,
+$D(\mu|\nu)\geq\left\|\mu-\nu\right\|_{\text{TV}}^{2}/2,$
+yields
+(5.25) $D(\mu)-E(\mu)\geq\frac{1}{4}\left\|\mu-M\right\|_{\text{TV}}^{2},$
+which thus improves on the stability result 5.21 from [47].
+The main reason for this improvement is that the proof of the latter stability in [47] is based on the sharp improvement of Aubin’s constrained Moser–Trudinger inequality established in [80] (recalled in Theorem 4.5), while the stability result in Cor 5.7 is a consequence of the sharp constrained log HLS inequality in Theorem 5.1, which is, a priori, stronger than Theorem 4.5, as discussed in Section 4.0.1.
+For comparison, let us summarize the proof of 5.21 in [47].
+The starting point is Theorem 4.5 from [80] that, as shown in [47, Thm 4.3], implies an explicit stability inequality for the sharp Moser-Trudinger inequality, that in our notation may be expressed as
+(5.26) $\mathcal{D}_{1}(u)\geq\frac{1}{2}\inf_{g\in\mathcal{G}}\mathcal{J}(u\cdot g),\,\,\,\forall u\in W^{2,1}(S^{2})$
+(the proof is similar to the proof of Prop 4.6).
+The stability inequality 5.21 is then deduced, using the general duality theory for stability inequalities developed in [46], combined with a strong Young inequality for the relative entropy established in [47, Thm 3.2].
+But here we note that, starting from the stability inequality 5.26, there is, in fact, an alternative argument that yields a stronger stability inequality for the log HLS inequality than the inequality 5.21, but still weaker than the stability inequality 5.25.
+Indeed, on $W^{2,1}(S^{2})\cap\mathcal{H}(X,\omega_{0})$ the inequality 5.26 is equivalent to the inequality in Prop 4.6, which, as shown in the proof of Prop 4.7, implies that $F_{-3/2}(\mu)\geq 0$ on $\mathcal{P}(X)_{0}.$ Thus implies, precisely, as in the proof of Cor 5.7, that
+$D(\mu)-E(\mu)\geq\frac{1}{3}D\left(\mu|M\right),$
+giving $D(\mu)-E(\mu)\geq 6^{-1}\left\|\mu-M\right\|_{\text{TV}}^{2},$ by the sharp Pinsker inequality.
+Finally, we note that (as in the formulation of [47, Thm 4.3]), the inequality 5.26 is a stability inequality for the sharp Moser–Trudinger inequality of the general form 5.19 on $W^{2,1}(S^{2})/\mathbb{R}$ with explicit stability constant $\kappa=1/8$ wrt the divergence $d(u,v):=\left\|u-v\right\|_{H^{-1}}^{2}.$ It seems thus natural to ask what the value of the corresponding optimal stability $\kappa$ constant is?
+In this direction it follows from Theorem 1.3 together with the argument in the proof of Prop 4.7, that $\kappa\leq 1/4.$ If the inf over $\mathcal{J}(u\cdot g)$ over $\mathcal{G}$ were attained for $u_{g}:=u\cdot g$ satisfying Aubin’s constraint, then one could conclude from Theorem 4.5 that $\kappa=1/8$ is optimal, using that $\mathcal{D}_{1}$ is $\mathcal{G}$-invariant.
+However, this is not the case, in general.
+For example, when $u\in\mathcal{H}(S^{2},\omega_{0})$ the inf in question is, as explained in Section 4.0.1, attained for $u_{g}$ satisfying Tian’s orthogonality constraint, which, in this case is equivalent to $\omega_{u_{g}}\in\mathcal{P}(S^{2})_{0}.$ In contrast, Aubin’s constraint says, in our notation, that $e^{-u_{g}}\omega_{0}\in\mathcal{P}(S^{2})_{0}.$
+5.2.2. Comparison with other recent stability results
+To the best of our knowledge Cor 5.7 provides the first instance of stability results of the form 5.19, where the optimal stability constant $\kappa$ is explicitly computed, in the context of fractional Sobolev and HLS inequalities or their logarithmic versions.
+The original compactness argument employed in [32], generalized to fractional Sobolev inequalities (with $\alpha\in]0,n/2[$) in [50], did not yield any explicit value for the corresponding optimal stability constants $\kappa_{S_{n,\alpha}},$ but it did yield an upper bound, showing that $\kappa_{S_{n,\alpha}}\leq 4\alpha/(n+2+2\alpha).$ Only recently an explicit lower bound on the optimal constant $\kappa_{\mathcal{S}_{n,1}}$ (for $n\geq 3)$ was established in [65] of the form $\kappa_{S_{n,1}}\geq\beta/n$ for an explicit constant $\beta,$ whose dimension dependence is thus of the same asymptotic order as the upper bound.
+By letting $n\rightarrow\infty$ a stability result for the sharp Gaussian log Sobolev inequality in $\mathbb{R}^{n}$ was deduced in [65], with an explicit (but not optimal) constant $\kappa,$ where the corresponding divergence $d$ is the squared Gaussian $L^{2}$-norm.
+Very recently the explicit lower bound on $\kappa_{\mathcal{S}_{n,1}}$ in [65] (for $n\geq 3)$ was extended to $\kappa_{\mathcal{S}_{n,\alpha}}$ and $\kappa_{\mathcal{S^{*}}_{n,\alpha}}$ in [49] for any $\alpha\in]0,n/2[,$ replacing the constant $\beta$ with a constant $\beta_{\alpha}$ depending on $\alpha.$
+The relative entropy $D(\cdot|\cdot)$ (or more precisely, functions of $D(\cdot|\cdot)$) has previously appeared as a divergence in stability results of the form 5.19.
+For example, under a moment constraint, a stability results for the sharp Gaussian log Sobolev inequality in $\mathbb{R}^{n}$ was established in [66], where the divergence is an explicit function $\Phi(D(\cdot|\cdot))$ of the relative entropy $D(\cdot|\cdot)$ (see also the recent survey [37]).
 6. The $\mathcal{K}$–reduced probability measures on $X^{N}$
 We continue with the setup introduced in Section 3.
-We thus fix a compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and, given a log Fano manifold $(X,\Delta),$ we fix a $\mathcal{K}$–invariant positively curved metric on $K_{(X,\Delta)}.$ Recall that $\boldsymbol{m}_{N}$ denotes the corresponding moment map for the diagonal action of $\mathcal{K}$ on $X^{N}$, defined $X^{N}$ with values in the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}$ (Section 3.2.8).
-Before introducing the probabilistic setup we start with some preparations.
-First assume that 0 is a regular value of the moment map $\boldsymbol{m}_{N}.$ Given a measure $\nu$ on $X^{N}$ with a smooth density and a basis $e_{1},\dots,e_{r}$ of $\mathfrak{k}$, we will denote by $\nu_{\boldsymbol{m}_{N}}$ the corresponding Gelfand–Leray form on the submanifold $\left\{\boldsymbol{m}_{N}=0\right\}$ of $X^{N},$ uniquely determined by the following property:
+Given a log Fano manifold $(X,\Delta),$we thus fix a compact subgroup $\mathcal{K}$ of $\mathcal{G}$ and a $\mathcal{K}$–invariant positively curved metric on $K_{(X,\Delta)}.$ Recall that $\boldsymbol{m}_{N}$ denotes the corresponding moment map for the diagonal action of $\mathcal{K}$ on $X^{N}$, taking values in the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $\mathcal{K}$ (Section 3.2.9).
+Given a positive rational number $k$ such that $kK_{(X,\Delta)}$ defines a line bundle we denote by $N$ the dimension of $H^{0}(X,-kK_{(X,\Delta)})$ and by $\nu^{(N)}$ the measure on $X^{N}$ defined by
+$\nu^{(N)}=\frac{i^{(Nn)^{2}}}{2^{Nn}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
+where $\det S^{(N)}$ denotes the section of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ obtained by taking $s_{1},\dots,s_{N}$ in formula 1.3 to be a basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$
+Lemma 6.1.
+Given a log Fano manifold $(X,\Delta),$ there exists a real character $\chi_{N}$ on $\mathcal{G}$ such that $g^{*}\nu_{N}=\chi_{N}(g)\nu_{N}$ for any $g\in\mathcal{G}.$ More precisely, given $W$ in the Lie algebra of $\mathcal{G}$,
+(6.1) $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)},$
+for $t\in\mathbb{C},$ where $\text{Fut}_{(X,\Delta),k}(W)$ denotes the quantized Futaki invariant at level $k.$ Moreover, if $k$ is a positive integer and $k\Delta$ is a divisor, then $\text{Fut}_{(X,\Delta),k}=0$ iff $\det S^{(N)}$ is $\mathcal{G}$-invariant.
+Proof.
+First assume that $k$ is a positive integer and $k\Delta$ is a divisor.
+Then $\mathcal{G}$ acts naturally on $(X,-kK_{(X,\Delta)})$ and, as a consequence, also on $H^{0}(X,-kK_{(X,\Delta)}).$ By basic linear algebra, $g^{*}\nu_{N}=\left|\det A(g)\right|^{-2/k}\nu_{N}$ where $A(\cdot)$ denotes the representation of $\mathcal{G}$ on $H^{0}(X,-kK_{(X,\Delta)})$ and $\det A(g)$ defines a $\mathbb{C}^{*}$-valued algebraic character on $\mathcal{G}.$ Hence, $\chi_{N}(g)\coloneqq\left|\det A(g)\right|^{-2/k}$ is a real-valued character on $\mathcal{G}$ and $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)}$ by the very definition of $\text{Fut}_{(X,\Delta),k}(W).$ In particular, if $\text{Fut}_{(X,\Delta),k}=0$ then $\det A(g)$ is a holomorphic function on $\mathcal{G}$ taking values in $S^{1}.$ But thus implies that $\det A(g)$ is identically equal to $1,$ showing that $\det S^{(N)}$ is is $\mathcal{G}$-invariant.
+Finally, consider the case when it is only assumed that $kK_{(X,\Delta)}$ defines a line bundle.
+Then $\mathcal{G}$ still acts naturally on the real line bundle over $X^{N}$ defined as the the top exterior power of the real cotangent bundle of $X^{N},$ having $\left|\det S^{(N)}\right|^{-2/k}$ as a smooth section.
+We can thus proceed essentially as before.
+∎
+Define the energy per NN particles as the following lsc function on $X^{N}$
+$E^{(N)}(x_{1},\dots,x_{N})\coloneqq-(kN)^{-1}\log\left(\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}\right)$
+By 2.1,
+(6.2) $E^{(N)}(x_{1},\dots,x_{N})=\infty\iff\exists s\in H^{0}(X,-kK_{(X,\Delta)})-\{0\}\text{ such that }s(x_{1})=\cdots=s(x_{N})$
+As shown in [13], as $N\rightarrow\infty,$ $E^{(N)}$ converges towards the pluricomplex energy $E$ on $\mathcal{P}(X),$ in the sense of Gamma-convergence.
+The first part of the following result can thus be viewed as an analog of Prop 3.4 for finite $N$ and the second part is an analog of the fact that the unique minimizer $\omega_{0}^{n}/V$ of the functional $E$ on $\mathcal{P}(X)$ is in $\mathcal{P}(X)_{0},$ if the Futaki character of $(X,\Delta)$ vanishes (by Lemma 3.3).
+Proposition 6.2.
+Let $(X,\Delta)$ be a log Fano manifold. Assume that the quantized Futaki character of $(X,\Delta)$ at level $k$ vanishes. Then, under the diagonal action of $\mathcal{G}$ of $X^{N},$
+$E^{(N)}(x_{1},\dots,x_{N})<\infty\implies(x_{1},...,x_{N})\,\text{is semi-stable }$
+As a consequence, the closure of the $\mathcal{G}$-orbit of $(x_{1},...,x_{N})$ intersects $\left\{\boldsymbol{m}_{N}=0\right\}.$ Morover, any configuration $(x_{1},\dots,x_{N})_{*}$ minimizing $E^{(N)}$ satisfies $\boldsymbol{m}_{N}((x_{1},\dots,x_{N})_{*})=0.$ Additionally, if $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}$, then $E^{(N)}(x_{1},\dots,x_{N})<\infty$ implies that $(x_{1},...,x_{N})$ has a finite stabilizer.
+In particular, then $\boldsymbol{m}_{N}$ is a submersion at $(x_{1},\dots,x_{N})_{*}.$
+Proof.
+First assume that $k$ is a positive integer such that $k\Delta$ is a divisor.
+Then $\mathcal{G}$ naturally acts on $kK_{(X,\Delta)}$ and the section $\det S^{(N)}$ is $G$-invariant, by Lemma 6.1.
+Hence, if $\det S^{(N)}(x_{1},\dots,x_{N})\neq 0,$ then $(x_{1},\dots,x_{N})$ is semi-stable, by the very definition of semi-stability.
+Next, if $\boldsymbol{x}:=(x_{1},\dots,x_{N})_{*}$ minimizes $E^{(N)}$ then it follows, in particular, that $\boldsymbol{x}$ minimizes the restriction of $E^{(N)}$ to the orbit $\mathcal{G}\boldsymbol{x}.$ Note that since $E^{(N)}$ is $G$-invariant its restriction $f$ to $\mathcal{G}\boldsymbol{x}$ defines a Kempf-Ness, i.e.
+up to scaling,
+$f(g\boldsymbol{x})=-\log\left\|gS_{\boldsymbol{x}}\right\|(g\boldsymbol{x})$
+for a non-zero vector $S_{\boldsymbol{x}}$ in the complex line over $\boldsymbol{x},$ which in this case is $(-kK_{X^{N_{k}}})_{|\boldsymbol{x}}.$ But in general a point $\boldsymbol{x}$ is a minimum of the Kempf-Ness function on $\mathcal{G}\boldsymbol{x}$ iff the corresponding moment map vanishes at $\boldsymbol{x}$ [83].
+In the general case when it is only assumed that $kK_{(X,\Delta)}$ is defined as a line bundle, one can replace $\det S^{(N)}$ with a tensor power $(\det S^{(N)})^{\otimes m}$ for $m$ a sufficently divisible positive integer and proceed essentially as before.
+Finally, assume that $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}.$ Assume that $E^{(N)}(x_{1},\dots,x_{N})\neq\infty$ and - in order to get a contradiction - that the stabilizer of $(x_{1},\dots,x_{N})$ is infinite.
+Then there exists a non-trivial holomorphic vector field $V$ on $X$ vanishing at all points $x_{1},...,x_{N}.$ Take an element $S_{k}$ in $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)-\{0\}.$ Contracting $S_{k}$ with $V$ then gives a non-trivial holomorphic section of $kK_{(X,\Delta)}$ vanishing at all points $x_{1},...,x_{N},$ which contradicts $E^{(N)}(x_{1},\dots,x_{N})\neq\infty.$ Finally, we recall that, in general, the moment map is a submersion at $\boldsymbol{x}$ iff the stabilizier is finite iff $\boldsymbol{x}$ is stable.
+∎
+Note that the assumption $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}$ is satisfied when $k$ is sufficently large, since $K_{(X,\Delta)}$ is ample.
+In general, this assumption is needed to get finite stabilizers, as illustrated in the following example.
+Example 6.3.
+Let $(X,\Delta)=(\mathbb{P}^{1},0).$ In this case $k\in\mathbb{Z}_{+}/2,$ $N=2k+1=2,3,$...
+and $H^{0}(X,-kK_{(X,\Delta)}\otimes T^{*}X)\neq\{0\}$ holds iff $k\geq 1,$ i.e.
+$N\geq 3.$ For any $N\geq 2$, $E^{(N)}(x_{1},\dots,x_{N})\neq 0$ iff all $x_{i}$ are distinct.
+Hence, $\left\{E^{(N)}(x_{1},\dots,x_{N})\neq\infty\right\}$ is a subset of $(X^{N})_{ss}$ , which is a proper subset when $N>3$ (by the condition 2.10).
+Consider now the case when $k=1/2$ (i.e.
+$N=2)$ and take $\phi_{0}$ to be the Fubini-Study metric i.e.
+$\phi_{0}$ is an $SO(3)$-invariant metric.
+Identifying $\mathbb{P}^{1}$ with the unit-sphere in $\mathbb{R}^{3},$ $\left\|\det S^{(2)}(x_{1},x_{2})\right\|_{\phi_{0}}$ is proportional to $|x_{1}-x_{2}|$ and $\boldsymbol{m}_{2}=(x_{1}+x_{2})/2.$ Hence, $(x_{1},x_{2})$ minimizes $E^{(2)}(x_{1},x_{2})=0$ iff $x_{1}$and $x_{2}$ are anti-podal points, which is equivalent to $\boldsymbol{m}_{2}(x_{1},x_{2})=0.$ This is also a consequence of the previous proposition.
+However, in this case the stabilizer is infinite (and, as consequence, $\boldsymbol{m}_{2}$ is not a submersion at $(x_{1},x_{2}).$).
+Indeed, any two distinct points on $\mathbb{P}^{1}$ are fixed by $\mathbb{C}^{*}-a$ction.
+Note also that $N=1$ on $\mathbb{P}^{1}$ is not covered by the previous proposition, which is consistent with the fact that there are no semi-stable points in this case (since the condition 2.10 is never satisfied).
+In particular, the moment map has no zeroes.
+6.1. Gelfand–Leray measures and contracted measures
+Before introducing the probabilistic setup we start with some general definitions.
+Assume that 0 is a regular value of the moment map $\boldsymbol{m}_{N}.$ Given a measure $\nu$ on $X^{N}$ with a smooth density and a basis $e_{1},\dots,e_{r}$ of $\mathfrak{k}$, we will denote by $\nu_{\boldsymbol{m}_{N}}$ the corresponding Gelfand–Leray form on the submanifold $\left\{\boldsymbol{m}_{N}=0\right\}$ of $X^{N},$ uniquely determined by the following property:
 (6.3) $\nu_{\boldsymbol{m}_{N}}\wedge dh_{1}\wedge\cdots\wedge dh_{r}=\nu$
@@
 6.2. Probabilistic and statistical mechanical setup
-6.1.1.
-The Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ and the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ and partition function $Z_{\boldsymbol{m}_{N}}.$
-Given a log Fano manifold $(X,\Delta),$ denote by $\nu^{(N)}$ the measure on $X^{N}$ defined by
-$\nu^{(N)}=\frac{i^{(Nn)^{2}}}{2^{Nn}}\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)^{-1/k}\wedge\overline{\left(\det S^{(N)}\otimes S_{\Delta_{N}}\right)}^{-1/k},$
-where $\det S^{(N)}$ denotes the section of $kK_{(X^{N},\Delta_{N})}\rightarrow X^{N}$ obtained by taking $s_{1},\dots,s_{N}$ in formula 1.3 to be a basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$ When 0 is a regular value of $\boldsymbol{m}_{N}$ we will denote by $\nu_{\boldsymbol{m}_{N}}^{(N)}$ the corresponding Gelfand–Leray measures, supported on $X^{N}.$ While this measure depends on the choice of basis in $\mathfrak{k},$ changing the basis only has the effect of multiplying $\nu_{\boldsymbol{m}_{N}}$ with a positive constant $C$ (the determinant of the corresponding change of bases matrix).
+6.2.1. The Gelfand–Leray measure $\nu_{0}^{(N)}$ and the $\mathcal{K}$–reduced canonical probability measure $\mu_{0}^{(N)}$ and partition function $Z_{N,0}.$
+Assume that 0 is a regular value of $\boldsymbol{m}_{N}.$ We then denote by $\nu_{0}^{(N)}$ the Gelfand–Leray measure supported on $\left\{\boldsymbol{m}_{N}=0\right\}\subset X^{N},$ corresponding to the measure $\nu^{(N)}$ on $X^{N}.$ While this measure depends on the choice of basis in $\mathfrak{k},$ changing the basis only has the effect of multiplying $\nu_{0}^{(N)}$ with a positive constant $C$ (the determinant of the corresponding change of bases matrix).
 Since $C$ is independent of $N$ this does not effect the large-$N$ limits in Conjecture 1.8 and Theorem 1.9.
-When 0 is not a regular value of $\boldsymbol{m}_{N}$ we propose, tentatively, to define $\nu_{\boldsymbol{m}_{N}}$ as
+Set
+(6.4) $\mu_{0}^{(N)}\coloneqq\nu_{0}^{(N)}/Z_{N,0},\,\,\,Z_{N,0}\coloneqq\int\nu_{0}^{(N)}.$
+Assuming that $Z_{N,0}<\infty$ the measure $\mu_{0}^{(N)}$ defines a probability measure on $X^{N},$ that we shall call the $\mathcal{K}$–reduced canonical probability measure and the corresponding normalizing constant $Z_{N,0}$ will be called the $\mathcal{K}$–reduced partition function.
+6.2.2. The thickened setup
+When 0 is not a regular value of $\boldsymbol{m}_{N}$ we replace the measure $\nu_{0}^{(N)}$ with
 […]
-[…]
+for an appropriate sequence of positive numbers $\epsilon_{N},$ called the thickening sequence.
+The corresponding normalization constants $Z_{N,\epsilon_{N}}$ and probability measurse $\mu_{\epsilon_{N}}^{(N)}$ will be called the thickened partition functions and probability measures.
 Remark 6.4.
-When 0 is a regular value of $\boldsymbol{m}_{N}$ the previous definition coincides with the limit of the rhs of formula 6.2 as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed).
-It would be interesting to know if, in general, this limit exists and if it coincides with the measure supported on the regular locus, defined by the Gelfand–Leray type measure on the regular locus.
-Set
-(6.3) $\mu_{\boldsymbol{m}_{N}}^{(N)}\coloneqq\nu_{\boldsymbol{m}_{N}}^{(N)}/Z_{\boldsymbol{m}_{N}},\,\,\,Z_{\boldsymbol{m}_{N}}\coloneqq\int\nu_{\boldsymbol{m}_{N}}^{(N)}.$
-Assuming that $Z_{\boldsymbol{m}_{N}}<\infty$ the measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$ defines a probability measure on $X^{N},$ that we shall call the $\mathcal{K}$–reduced canonical probability measure and the corresponding normalizing constant $Z_{\boldsymbol{m}_{N}}$ will be called the $\mathcal{K}$–reduced partition function.
+When 0 is a regular value of $\boldsymbol{m}_{N}$, $\nu_{0}^{(N)}$ coincides with the limit of $\nu_{\epsilon_{N}}^{(N)}$ as $\epsilon_{N}\rightarrow 0$ (for $N$ fixed).
+It would be interesting to know if, when $\{\boldsymbol{m}_{N}=0\}$ is non-empty, this limit exists and if it coincides with the measure supported on the regular locus of $\{\boldsymbol{m}_{N}=0\},$ defined by the Gelfand–Leray type measure on the regular locus.
 6.2.3. The contracted setup
-One could also define an alternative probability measures $(\mu^{(N)})^{\mathrm{c}}$ by replacing the Gelfand–Leray measure $\nu_{\boldsymbol{m}_{N}}^{(N)}$ with the contracted measure $(\nu^{(N)})^{\mathrm{c}}$ and the $\mathcal{K}$–reduced partition function $Z_{\boldsymbol{m}_{N}}$ with the contracted partition function, defined by
+One could also define an alternative probability measures $(\mu^{(N)})^{\mathrm{c}}$ by replacing the Gelfand–Leray measure $\nu_{0}^{(N)}$ with the contracted measure $(\nu^{(N)})^{\mathrm{c}}$ and the $\mathcal{K}$–reduced partition function $Z_{N,0}$ with the contracted partition function, defined by
 (6.6) $Z_{N}^{\mathrm{c}}\coloneqq\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}$
@@
 Given a number $\beta\in\mathbb{R},$ set
-$\nu_{\beta}^{(N)}\coloneqq e^{-\beta NE^{(N)}}\mu_{0}^{\otimes N},\,\,\,\,\,E^{(N)}(x_{1},\dots,x_{N})\coloneqq-(kN)^{-1}\log\left(\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}\right).$
-When $\beta=-1$ it coincides with the measure $\nu^{(N)}$ on $X^{N}$ (but $\nu_{\beta}^{(N)}$ depends on the fixed metric $\phi_{0}$ when $\beta\neq-1).$ The measures $\nu_{\beta}^{(N)}$ on $X^{N}$ induce a family of Gelfand–Leray measures $\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}.$ Assuming that the $\mathcal{K}$–reduced partition function $Z_{\boldsymbol{m}_{N}}(\beta)$ at inverse temperature $\beta$ is finite,
-$Z_{\boldsymbol{m}_{N}}(\beta)\coloneqq\int\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}<\infty,$
+$\nu_{\beta}^{(N)}\coloneqq e^{-\beta NE^{(N)}}\mu_{0}^{\otimes N}.$
+When $\beta=-1$ it coincides with the measure $\nu^{(N)}$ on $X^{N}$ (but $\nu_{\beta}^{(N)}$ depends on the fixed metric $\phi_{0}$ when $\beta\neq-1).$ The measures $\nu_{\beta}^{(N)}$ on $X^{N}$ induce a family of Gelfand–Leray measures $\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}.$ Assuming that the $\mathcal{K}$–reduced partition function $Z_{N,0}(\beta)$ at inverse temperature $\beta$ is finite,
+$Z_{N,0}(\beta)\coloneqq\int\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}<\infty,$
 we obtain a measure on $X^{N}$ that we shall call the $\mathcal{K}$–reduced Gibbs measures at inverse temperature $\beta$ by setting
-(6.5) $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}\coloneqq\frac{\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}}{Z_{\boldsymbol{m}_{N}}(\beta)}=\frac{e^{-\beta NE^{(N)}}\left(\mu_{0}^{\otimes N}\right)_{\boldsymbol{m}_{N}}}{Z_{\boldsymbol{m}_{N}}(\beta)}.$
-When $\beta=-1$ this measure coincides with the $\mathcal{K}$–reduced canonical probability measure $\mu_{\boldsymbol{m}_{N}}^{(N)}$, defined in the previous section, and when $\beta=0$ it coincides with the classical microcanonical measure associated to the Hamiltonians $h_{1},\dots,h_{r},$ that is ubiquitous in the statistical mechanics literature.
+(6.7) $\mu_{0,\beta}^{(N)}\coloneqq\frac{\left(\nu_{\beta}^{(N)}\right)_{\boldsymbol{m}_{N}}}{Z_{N,0}(\beta)}=\frac{e^{-\beta NE^{(N)}}\left(\mu_{0}^{\otimes N}\right)_{\boldsymbol{m}_{N}}}{Z_{N,0}(\beta)}.$
+When $\beta=-1$ this measure coincides with the $\mathcal{K}$–reduced canonical probability measure $\mu_{0}^{(N)}$, defined in the previous section, and when $\beta=0$ it coincides with the classical microcanonical measure associated to the Hamiltonians $h_{1},\dots,h_{r},$ that is ubiquitous in the statistical mechanics literature (especially when $r=1)$.
+In general, $Z_{N,0}(\beta)$ is finite when $\beta>-\alpha(X,\Delta)$ (by the inequality 2.5).
+But if Conjecture 1.2 is correct, then this bound can be considerably improved.
 Remark 6.5.
-From a statistical mechanics perspective, the function $E^{(N)}$ plays the role of the $N$-particle energy (or Hamiltonian) and the corresponding Gibbs measure $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ describes the corresponding equilibrium state of $N$ interacting particles confined to $\left\{\boldsymbol{m}_{N}=0\right\},$ at inverse temperature $\beta.$ The corresponding probability space $(X^{N},\mu_{\boldsymbol{m}_{N},\beta}^{(N)})$ may be viewed as a singular incarnation of the notion of mixed ensemble introduced in [54].
+From a statistical mechanics perspective, the function $E^{(N)}$ plays the role of the $N$-particle energy (or Hamiltonian) and the corresponding Gibbs measure $\mu_{0,\beta}^{(N)}$ describes the corresponding equilibrium state of $N$ interacting particles confined to $\left\{\boldsymbol{m}_{N}=0\right\},$ at inverse temperature $\beta.$ The corresponding probability space $(X^{N},\mu_{0,\beta}^{(N)})$ may be viewed as a singular incarnation of the notion of mixed ensemble introduced in [70].
 In particular, when $X=\mathbb{P}^{1},$ identified with the two-sphere and $\omega_{0}$ is the standard $\mathrm{SO}(3)$–invariant symplectic form on $X,$ $E^{(N)}$ coincides with the energy particle in Onsager’s statistical mechanical description of the 2D incompressible Euler equations, adapted to the two-sphere [87, 88].
-The corresponding Gibbs measures $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ are invariant under the Hamiltonian flow of $E^{(N)}$ on $X^{N}$ and, building on Theorem 1.7, we show in the companion paper [4] that the empirical measure $\delta_{N},$ viewed as a random variable on the probability space $(X^{N},\mu_{\boldsymbol{m}_{N},\beta}^{(N)})$ converges in law towards the $\mathrm{SO}(3)$–invariant probability measure on the two-sphere, for any $\beta>-1.$
+The corresponding Gibbs measures $\mu_{0,\beta}^{(N)}$ are invariant under the Hamiltonian flow of $E^{(N)}$ on $X^{N}$ and, building on Theorem 1.9, we show in the companion paper [5] that the empirical measure $\delta_{N},$ viewed as a random variable on the probability space $(X^{N},\mu_{0,\beta}^{(N)})$ converges in law towards the $\mathrm{SO}(3)$–invariant probability measure on the two-sphere, for any $\beta>-1.$
 Proposition 6.6.
-For any $\beta\in\mathbb{R}$ the probability measure $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ is $\mathcal{K}$–invariant, when it is well-defined.
+For any $\beta\in\mathbb{R}$ the probability measure $\mu_{0,\beta}^{(N)}$ is $\mathcal{K}$–invariant, when it is well-defined.
 Proof.
 First observe that for any $g\in\mathcal{G},$ $g^{*}\det S^{(N)}=\chi_{N}(g)\det S^{(N)}$ for a homomorphism $\chi_{N}\colon\mathcal{G}\rightarrow\mathbb{C}^{*}$ (see Lemma 6.1 below).
-It follows that the restriction of $\chi_{N}$ to $\mathcal{K}$ takes values in $S^{1}\subset\mathbb{C}^{*}.$ Indeed, since $\mathcal{K}$ is compact the image $\chi_{N}(\mathcal{G})$ is a compact subgroup of $\mathbb{C}^{*}$ and thus contained in the unique maximally compact subgroup $S^{1}$ of $\mathbb{C}^{*}.$ Since $\phi_{0}$ is $\mathcal{K}$–invariant, it follows that $\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}$ is $\mathcal{K}$–invariant on $X^{N}.$ Finally, since the measure $\mu_{0}$ and $|\boldsymbol{m}|$ are also $\mathcal{K}$–invariant we deduce that $\mu_{\boldsymbol{m}_{N},\beta}^{(N)}$ is $\mathcal{K}$–invariant.
+It follows that the restriction of $\chi_{N}$ to $\mathcal{K}$ takes values in $S^{1}\subset\mathbb{C}^{*}.$ Indeed, since $\mathcal{K}$ is compact the image $\chi_{N}(\mathcal{G})$ is a compact subgroup of $\mathbb{C}^{*}$ and thus contained in the unique maximally compact subgroup $S^{1}$ of $\mathbb{C}^{*}.$ Since $\phi_{0}$ is $\mathcal{K}$–invariant, it follows that $\left\|\det S^{(N)}(x_{1},\dots,x_{N})\right\|_{\phi_{0}}^{2}$ is $\mathcal{K}$–invariant on $X^{N}.$ Finally, since the measure $\mu_{0}$ and $|\boldsymbol{m}|$ are also $\mathcal{K}$–invariant we deduce that $\mu_{0,\beta}^{(N)}$ is $\mathcal{K}$–invariant.
 ∎
@@
 We begin by formulating a conjecture, which serves as motivation for the conjectures stated in the introduction of the paper.
-Recall that a sequence of measures $\Gamma_{N}$ on a topological space $\mathcal{M}$ satisfies a Large Deviation Principle (LDP) with speed $N$ and rate function $F\colon\mathcal{M}\rightarrow$$\infty$$\infty$] if $F$ is lsc and for any (Borel) measurable subset $\mathcal{B}$ of $\mathcal{M}$
+Recall that a sequence of measures $\Gamma_{N}$ on a topological space $\mathcal{X}$ satisfies a Large Deviation Principle (LDP) with speed $N$ and rate function $F\colon\mathcal{X}\rightarrow]-\infty,\infty]$ if $F$ is lsc and for any (Borel) measurable subset $\mathcal{B}$ of $\mathcal{X}$
 […]
@@
 Conjecture 6.7.
-The following conjectures hold:
+The following conjectures hold when $N$ is a sequence of regular values:
   • A(i) If $F(\beta-\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ then the measures $(\delta_{N})_{*}\left((\nu_{\beta}^{(N)})_{\boldsymbol{m}_{N}}\right)$ on $\mathcal{P}(X)_{0}$ satisfy an LDP with speed $N$ and rate functional $F_{\beta}(\mu)$.
@@
 (ii) if $\nu_{\beta}^{(N)}/\int\nu_{\beta}^{(N)}$ are well-defined probability measures for $N\gg 1,$ then the probability measures $(\delta_{N})_{*}\text{$\left(\nu_{\beta}^{(N)}/\int\nu_{\beta}^{(N)}\right)$}$ on $\mathcal{P}(X)$ satisfy an LDP with speed $N$ and rate functional $F_{\beta}(\mu)-C$ for some $C\in\mathbb{R}.$
+Without assuming that $N$ is a sequence of regular values the corresponding LDPs also hold on $\mathcal{P}(X)$ when $\left\{\boldsymbol{m}_{N}=0\right\}$ is replaced by $\left\{|\boldsymbol{m}_{N}|\leq\epsilon_{N}\right\}$ for any appropriate sequence $\epsilon_{N},$ with rate functional $F_{\beta}+\chi_{0},$ where $\chi_{0}=0$ on $\mathcal{P}(X)_{0}$ and $\chi_{0}=\infty$ away from $\mathcal{P}(X)_{0}.$
 Conjecture B follows from results in [19], when $n=1$ and when $\beta>0$ it is established in [13, 14] for any $n.$ In the general case, Conjecture A is motivated by the following heuristic argument.
@@
 in the sense of $\Gamma$-convergence. This suggests the formal approximation
-$(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}\sim e^{-N\left(\beta E+\chi\right)}(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right),$
-as probability measures on $\mathcal{P}(X),$ where $\chi=0$ on $\mathcal{P}(X)_{0}$ and $\chi=\infty$ away from $\mathcal{P}(X)_{0}.$ Next, according to Sanov’s classical LDP [50]
+$(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}\sim e^{-N\left(\beta E+\chi_{0}\right)}(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right),$
+as probability measures on $\mathcal{P}(X),$ where $\chi_{0}=0$ on $\mathcal{P}(X)_{0}$ and $\chi_{0}=\infty$ away from $\mathcal{P}(X)_{0}.$ Next, according to Sanov’s classical LDP [63]
 $(\delta_{N})_{*}\left(\mu_{0}^{\otimes N}\right)\sim e^{-ND(\mu|\mu_{0})}.$
-Hence, formally, $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ satisfies a LDP with rate functional $\beta F+\chi.$ Finally, since $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ is supported on $\mathcal{P}(X)_{0}$ such an LDP is equivalent to the LDP for $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ in Conjecture 6.4.
+Hence, formally, $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ satisfies a LDP with rate functional $\beta F+\chi_{0}.$ Finally, since $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ is supported on $\mathcal{P}(X)_{0}$ such an LDP is equivalent to the LDP for $(\delta_{N})_{*}\text{$\nu_{\beta}^{(N)}$}$ in Conjecture 6.7.
 The argument for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ is similar.
@@
 (6.9) $\gamma(X,\Delta)^{\mathcal{G}}\geq\sup_{\gamma>0}\left\{F(-\gamma)_{0}>-\infty\right\}.$
-Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 6.11.
+Conversely, take $\gamma>\gamma(X,\Delta)^{\mathcal{G}}.$ Then $\mu_{-\gamma}^{(N)}$ is a well-defined probability measure for $N\gg 1,$ as follows from the argument in the proof of Prop 6.13.
 Hence, the assumed LDP for the probability measures $(\delta_{N})_{*}\text{$\mu_{\beta}^{(N)}$}$ yields
 $0=-\lim_{N\rightarrow\infty}N^{-1}\log\int_{X^{N}}\mu_{\beta}^{(N)}=\inf_{\mathcal{P}(X)_{0}}F_{\beta}(\mu)-C,$
-showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 6.7 is an equality.
+showing that $F(\beta)_{0}=C>-\infty.$ Thus, the inequality 6.9 is an equality.
 As a consequence, $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff there exists $\gamma>1$ such that $F(-\gamma)_{0}>-\infty.$ By Theorem 3.7, assuming that the Futaki character of $(X,\Delta)$ vanishes, this means that $\gamma(X,\Delta)^{\mathcal{G}}>1$ iff $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}>1$ iff $(X,\Delta)$ admits a KE metric, which by the solution of the YTD conjecture equivalently means that $(X,\Delta)$ is K-polystable (see [53] for the case $\Delta=0$ with $X$ smooth and [95, 97] for the case of a general log Fano variety $(X,\Delta)).$ Finally, the last statement of the proposition is shown as above.
 Next, to prove that Conjecture 6.7 A implies Conjecture 1.8, assume that $(X,\Delta)$ admits a KE metric.
-It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 6.6 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.14 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric).
+It follows, from Theorem 3.7 , that $F(\beta+\epsilon)_{0}>-\infty,$ for some $\epsilon>0,$ Hence, the convergence 6.8 for $\beta=-1$ implies, by Lemma 3.5 , the convergence 1.16 towards $\inf\mathcal{M}$ which is finite (since the inf is attained at any KE metric).
 As a consequence, $\int(\nu^{(N)})_{\boldsymbol{m}_{N}}$ is finite for $N$ sufficiently large.
-This means that $\mu_{-1}^{(N)}(=\mu_{\boldsymbol{m}_{N}}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional.
+This means that $\mu_{-1}^{(N)}(=\mu_{0}^{(N)})$ is a well-defined probability measure for $N\gg 1.$ Thus, by assumption, $(\delta_{N})_{*}\text{$\mu_{-1}^{(N)}$ }$ satisfies a LDP with rate functional $F_{-1}(\mu)-C$ on $\mathcal{P}(X)_{0}.$ By standard results for LDPs, to prove the convergence towards $\mu_{\text{KE}}$ in 1.1 it is thus enough to show that $\mu_{\text{KE}}$ is the unique minimizer of the rate functional.
 But this is shown in Theorem 3.7.
@@
 But this implication fails, in general.
-Indeed, by [15, Thm 5.11], $F_{-1}$ is lsc on $\mathcal{P}(X)$ iff $X$ admits a KE metric and $\mathcal{G}$ is trivial (which, in turn, is equivalent to $X$ being K-stable [42]).
+Indeed, by [16, Thm 5.11], $F_{-1}$ is lsc on $\mathcal{P}(X)$ iff $X$ admits a KE metric and $\mathcal{G}$ is trivial (which, in turn, is equivalent to $X$ being K-stable [53]).
 Hence, already $X=\mathbb{P}^{1}$ is a counterexample.
 But there are also counterexamples when $\mathcal{G}$ is trivial.
-Indeed, $F_{-1}$ is bounded from below on $\mathcal{P}(X)$ iff $(X,\Delta)$ is K-semistable (see [75] for the case when $X$ is smooth and $\Delta$ is trivial and the proof of [1, Thm 2.5] for the general case).
-However, there are examples of K-semistable log Fano manifolds $(X,\Delta)$ with $\mathcal{G}$ is trivial, such that $(X,\Delta)$ is not K-stable (for example, this can be achieved when $X=\mathbb{P}^{1}$ and $\Delta$ consists of three points).
+Indeed, $F_{-1}$ is bounded from below on $\mathcal{P}(X)$ iff $(X,\Delta)$ is K-semistable (see [93] for the case when $X$ is smooth and $\Delta$ is trivial and the proof of [2, Thm 2.5] for the general case).
+However, there are examples of K-semistable log Fano manifolds $(X,\Delta)$ with $\mathcal{G}$ trivial, such that $(X,\Delta)$ is not K-stable (for example, this can be achieved when $X=\mathbb{P}^{1}$ and $\Delta$ consists of three points).
 It is natural to ask about the relation between the assumptions in Conjectures A(i) and A(ii) (and likewise for B(i) and B(ii)).
-If inequality 6.9 can be strengthened as follows, then the second assumption implies the first:
+If inequality 6.11 can be strengthened as follows, then the second assumption implies the first:
 Lemma 6.10.
@@
 The first statement follows directly from the observation that $\int_{X^{N}}\nu_{-\gamma}^{(N)}<\infty\implies$ $\gamma<\gamma^{(N)}(X,\Delta)$ (by the openness property of local integrability thresholds).
-The second statement is shown in a similar way, using Prop 6.11 below.
+The second statement is shown in a similar way, using Prop 6.13 below.
 ∎
@@
 As shown in [73] (see also [20] for an analytic proof) $\delta_{k}(X)\geq\gamma^{(N_{k})}(X),$ where $\delta_{k}(X)$ is the invariant introduced in [73], defined in terms of log canonical thresholds of basis divisors on $X.$ Next, in the case of a toric Fano variety which is Gorenstein (i.e.
-the corresponding moment polytope is reflexive) it follows from [65, Cor 2.13] that $\delta_{k}(X)<\delta(X)$ for $k$ sufficiently large, when $\delta(X)<1.$ Finally, when $n=2$ it is shown in [65, Cor 2.13] that
+the corresponding moment polytope is reflexive) it follows from [85, Cor 2.13] that $\delta_{k}(X)<\delta(X)$ for $k$ sufficiently large, when $\delta(X)<1.$ Finally, when $n=2$ it is shown in [85, Cor 2.13] that
 $\delta_{k}(X)^{-1}=1+c_{k}\left(\delta(X)^{-1}-1\right),$
@@
 ∎
-6.3. Gibbs polystability vs finiteness of the restricted partition function $Z_{\boldsymbol{m}_{N}}$
-Lemma 6.9.
-Given a log Fano manifold $(X,\Delta),$ there exists a real character $\chi_{N}$ on $\mathcal{G}$ such that $g^{*}\nu_{N}=\chi_{N}(g)\nu_{N}$ for any $g\in\mathcal{G}.$ More precisely, given $W$ in the Lie algebra of $\mathcal{G}$, $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)},$ for $t\in\mathbb{C},$ where $\text{Fut}_{(X,\Delta),k}(W)$ denotes the quantized Futaki invariant at level $k.$
-Proof.
-By basic linear algebra, $g^{*}\nu_{N}=\left|\det A(g)\right|^{-2/k}\nu_{N}$ where $A(\cdot)$ denotes the representation of $\mathcal{G}$ on $H^{0}(X,-kK_{(X,\Delta)}).$ Hence, $\chi_{N}(g)\coloneqq\left|\det A(g)\right|^{-2/k}$ is a real character on $\mathcal{G}$ and $\chi_{N}(e^{tW})=e^{tN2\text{Re}\left(\text{Fut}_{(X,\Delta),k}(W)\right)}$ by the very definition of $\text{Fut}_{(X,\Delta),k}(W).$
+6.4. Gibbs polystability vs finiteness of the restricted partition function $Z_{N,0}$
 Lemma 6.12.
@@
 Proof.
-First, by the analytic definition of the lct (formula 2.4), $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}).$ Next, note that there exists $\epsilon_{N}\in]0,\infty[$ such that the closure of $\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ is contained in $(X^{N})_{\text{ss}}$ (since $\left\{\boldsymbol{m}_{N}=0\right\}$ is compact and contained in the open set $(X^{N})_{\text{ss}}).$ Hence, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ implies that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right).$ To prove the converse, assume that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right)$ and consider the integral of $\nu^{(N)}$ over a neighbourhood $U$ of a given point $(x_{1},\dots,x_{N})$ in $(X^{N})_{\text{ss }}.$ By the Kempf–Ness theorem [64, Thm 3.3], a point $(x_{1},\dots,x_{N})\in(X^{N})_{\text{ss}}$ iff $\overline{\mathcal{G}(x_{1},\dots,x_{N})}$ intersects $\left\{\left\|\boldsymbol{m}_{N}\right\|=0\right\}.$ This means that there exists $g\in\mathcal{G}$ such that $g(x_{1},\dots,x_{N})\in\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ and, as a consequence, there exists a neighborhood $U$ of $(x_{1},\dots,x_{N})$ such that $gU\subset\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}.$ Thus, by assumption, $\int_{gU}\nu^{(N)}<\infty$ and since the integrability of $\nu^{(N)}$ is $\mathcal{G}$–invariant (by the previous lemma) we deduce that $\int_{U}\nu^{(N)}<\infty,$ showing that $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}),$ as desired.
+First, by the analytic definition of the lct (formula 2.6), $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}).$ Next, note that there exists $\epsilon_{N}\in]0,\infty[$ such that the closure of $\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ is contained in $(X^{N})_{\text{ss}}$ (since $\left\{\boldsymbol{m}_{N}=0\right\}$ is compact and contained in the open set $(X^{N})_{\text{ss}}).$ Hence, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ implies that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right).$ To prove the converse, assume that $\nu^{(N)}\in L_{\text{}}^{1}\left(\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}\right)$ and consider the integral of $\nu^{(N)}$ over a neighbourhood $U$ of a given point $(x_{1},\dots,x_{N})$ in $(X^{N})_{\text{ss }}.$ By the Kempf–Ness theorem [83, Thm 3.3], a point $(x_{1},\dots,x_{N})\in(X^{N})_{\text{ss}}$ iff $\overline{\mathcal{G}(x_{1},\dots,x_{N})}$ intersects $\left\{\left\|\boldsymbol{m}_{N}\right\|=0\right\}.$ This means that there exists $g\in\mathcal{G}$ such that $g(x_{1},\dots,x_{N})\in\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}$ and, as a consequence, there exists a neighborhood $U$ of $(x_{1},\dots,x_{N})$ such that $gU\subset\left\{\left\|\boldsymbol{m}_{N}\right\|<\epsilon_{N}\right\}.$ Thus, by assumption, $\int_{gU}\nu^{(N)}<\infty$ and since the integrability of $\nu^{(N)}$ is $\mathcal{G}$–invariant (by the previous lemma) we deduce that $\int_{U}\nu^{(N)}<\infty,$ showing that $\nu^{(N)}\in L_{\text{loc}}^{1}((X^{N})_{\text{ss }}),$ as desired.
 ∎
 Proposition 6.13.
-$\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{\boldsymbol{m}_{N}}<\infty.$ In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{\boldsymbol{m}_{N}}<\infty$ for all sufficiently large $N.$
+First assume that 0 is a regular value of $\boldsymbol{m}_{N}.$ Then $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{N,0}<\infty.$ In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{N,0}<\infty$ for all sufficiently large $N.$ In general, $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1$ iff $Z_{N,\epsilon_{N}}<\infty$ when $\epsilon_{N}$ is sufficently small.
+In particular, $(X,\Delta)$ is Gibbs polystable iff it is Gibbs semistable and $Z_{N,\epsilon_{N}}<\infty$ when $N$ is sufficiently large and $\epsilon_{N}$ is sufficently small.
 Proof.
@@
 This follows from the observation that, along $\left\{\boldsymbol{m}_{N}=0\right\},$ the tangent bundle of $X$ is the direct sum of the tangent bundle of $\left\{\boldsymbol{m}_{N}=0\right\}$ and the bundle spanned by the vector fields induced by $Je_{1},\dots,Je_{r}.$ Indeed, the vector fields induced by $e_{1},\dots,e_{r}$ are tangent to $\left\{\boldsymbol{m}_{N}=0\right\}$ (since $\mathcal{K}$ preserves $\left\{\boldsymbol{m}_{N}=0\right\})$ and the vector fields corresponding to $Je_{1},\dots,Je_{r}$ are linearly independent (since $Je_{i}$ is the gradient of $\left\langle\text{$\boldsymbol{m}_{N}$},e_{i}\right\rangle$ with respect to the Kähler metric $\omega_{0}$ and 0 is a regular value of $\boldsymbol{m}_{N}).$ Hence, by the previous lemma, it will be enough to show that there exists a constant $C$ (depending on $N)$ such that
-(6.11) $C^{-1}\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}$
-But this follows from combining the inequalities 6.12 below with the first identity in Proposition 6.14 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
-6.4. Relating Gibbs polystability and the restricted partition function $Z_{\boldsymbol{m}_{N}}$ to integrals over $(X^{N})_{\mathrm{s}}/\mathcal{G}$
-The main virtue of the following proposition, as compared with Prop 6.11, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
+(6.13) $C^{-1}\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{0}^{(N)}\leq\int_{U_{\epsilon}}\nu^{(N)}\leq C\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{0}^{(N)}$
+But this follows from combining the inequalities 6.14 below with the first identity in Proposition 6.16 below and using that $f_{N}$ is a strictly positive function, when 0 is a regular value of $\boldsymbol{m}_{N}.$
+6.5. Relating Gibbs polystability and the restricted partition function $Z_{N,0}$ to integrals over $(X^{N})_{\mathrm{s}}/\mathcal{G}$
+The main virtue of the following proposition, as compared with Prop 6.13, is that it also applies when 0 is not a regular value of the moment map $\boldsymbol{m}_{N}.$
 Proposition 6.14.
@@
 Proof.
-Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 6.10, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small.
-Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 6.10.
+Denote by $U_{\epsilon}$ the image in $X^{N}$ of $\left\{\boldsymbol{m}_{N}=0\right\}\times]-\epsilon,\epsilon[^{r}$ under the smooth map $\Phi,$ defined by formula 6.12, and note that its closure is a compact subset of $(X^{N})_{\text{ss}},$ for $\epsilon$ sufficiently small.
+Indeed, since $(X^{N})_{\text{ss}},$ is open and contains the compact set $\left\{\boldsymbol{m}_{N}=0\right\},$ this follows from the continuity of the map 6.12.
 As a consequence, if $\gamma^{(N)}(X,\Delta)^{\mathcal{G}}>1,$ then the integral of $\nu^{(N)}$ over $U_{\epsilon}$ is finite.
@@
 $\rho_{N}(\boldsymbol{y},0)d\boldsymbol{y}=\left.\Phi^{*}((\nu^{(N)})^{\mathrm{c}})\right|_{(\mathbf{y},0)}.$
-Hence, the desired formula 6.13 follows from formula 6.14. ∎
+Hence, the desired formula 6.15 follows from formula 6.16. ∎
 Remark 6.15.
-Formula 6.14 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [20], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
+Formula 6.16 yields a new proof of the result that $\int_{X^{N}}\nu^{(N)}=\infty$ when $\mathcal{G}$ is nontrivial [21], by integrating $\nu^{(N)}$ over $\Phi(\left\{\boldsymbol{m}_{N}=0\right\}\times\mathbb{R}^{r}).$
 Consider now the following function on $X^{N},$
@@
 Then the contracted form $(\nu^{(N)})^{c}$ on $X^{N}$ is $\mathcal{G}$–invariant and if 0 is a regular value of $\boldsymbol{m}_{N},$ then
-$\int_{\left\{\boldsymbol{m}_{N}=0\right\}}f_{N}\nu_{\boldsymbol{m}_{N}}^{(N)}=\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}$
+$\int_{\left\{\boldsymbol{m}_{N}=0\right\}}f_{N}\nu_{0}^{(N)}=\int_{\left\{\boldsymbol{m}_{N}=0\right\}}(\nu^{(N)})^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}$
 for a positive constant $C,$ depending only on the choice of $e_{1},\dots,e_{r}$ (and thus independent of $N)$.
@@
 Moreover, in the general case when 0 is potentially not a regular value of $\mathbf{m}_{N}$, these identities hold if $\left\{\boldsymbol{m}_{N}=0\right\}$ is replaced by its intersection with the stable locus $(X^{N})_{\text{s}}.$
-We recall that 0 is a regular value of $\boldsymbol{m}_{N}$ iff the stable locus $(X^{N})_{\text{s}}$ coincides with both the semistable and the polystable locus [64, Cor 3.10].
+We recall that 0 is a regular value of $\boldsymbol{m}_{N}$ iff the stable locus $(X^{N})_{\text{s}}$ coincides with both the semistable and the polystable locus [83, Cor 3.10].
 In general, $(X^{N})_{\text{s}}$ is a Zariski open subset of the semistable locus and $(X^{N})_{\text{s}}/\mathcal{G}$ is a Zariski open subset (which as an orbifold) of the projective variety defined by the GIT quotient of $X^{N}$ under the action of $\mathcal{G}.$ Before turning to the proof of this proposition we note that, combining the previous two propositions gives the following
@@
 $\int_{(X^{N})_{\text{s}}/\mathcal{G}}\delta_{e_{1}}\cdots\delta_{e_{m}}(\nu^{(N)})^{c}<\infty$
-6.4.1. Proof of Prop 6.14
+6.5.1. Proof of Prop 6.16
 We will consider a general setup of a complex reductive Lie group $G$ acting on a polarized complex projective manifold $(X,L)$ (in the previous proposition the role of $X$ is played by $X^{N})$.
@@
 $\int_{\left\{m=0\right\}/K}\beta=C\int\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu.$
-To this end it is enough to show that the projection $p$ in 6.16 satisfies
+To this end it is enough to show that the projection $p$ in 6.18 satisfies
 $p_{*}\delta_{Je_{1}}\cdots\delta_{Je_{m}}\nu=C\beta,$
@@
 where the dots indicates a sum of forms that contain some factor of $dh_{i}$ for some $i.$ In particular, such forms vanish when restricted to $\{m=0\}.$
-Finally, Prop 6.14 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
+Finally, Prop 6.16 follows directly by combining the previous lemma and the previous proposition, applied to $X^{N}$ when 0 is a regular value the moment map $\boldsymbol{m}_{N}.$ In the general case we can apply the previous argument on $\boldsymbol{m}_{N}\cap(X^{N})_{\mathrm{s}}.$
 6.6. Examples where the quantized Futaki character vanishes and reductivity holds
@@
 As consequence, $\mathcal{G}$ is then reductive, if $\Delta$ is assumed to have simple normal crossings.
-For example, $\boldsymbol{b}_{k}=0$ for all $k$ when $X$ is $\mathbb{P}^{m}\times\mathbb{P}^{l}$ (since $X$ is homogenous) and when $X$ is $\mathbb{P}^{2}$ blown up in three non-collinear points (as follows from [65, Thm 6.1]).
+For example, $\boldsymbol{b}_{k}=0$ for all $k$ when $X$ is $\mathbb{P}^{m}\times\mathbb{P}^{l}$ (since $X$ is homogenous) and when $X$ is $\mathbb{P}^{2}$ blown up in three non-collinear points (as follows from [85, Thm 6.1]).
 In general, given a toric Fano manifold $X,$ $\boldsymbol{b}_{k}=0$ for all $k$ iff $\boldsymbol{b}_{k}=0$ for at least $n+1$ distinct values of $k,$ as a shown in shown in [85]; see also the survey [54].
+7. Strong uniform Gibbs polystability and thickened partition functions
+7.1. Strong uniform Gibbs polystability
+We start by recalling some general notions from GIT.
+Let $G$ be a complex reductive group acting on a polarized projective manifold $(X,L)$ and fix a maximally compact subgroup $K$ of $G$ and an $\text{ad}K$-invariant scalar product on the dual $\mathfrak{k}^{*}$ of the Lie algebra $\mathfrak{k}$ of $K.$ The corresponding norms on $\mathfrak{k}^{*}$ and $\mathfrak{k}$ will both be denoted by $\left|\cdot\right|.$ Denote by $w_{\lambda}(x)$ the weight of a point $x\in X$ wr a non-trivial subgroup $\lambda$ of $G$ isomorphic to $\mathbb{C}^{*},$ which takes values in $\mathbb{Z}$ (also known as the Hilbert-Mumford weight) and denote by $\left|\lambda\right|$ the norm of generator of the compact subgroup of $\lambda,$ identified with an element in the the Lie algebra of $K.$ Given a “thickening parameter” $\epsilon\in[0,\infty[$ set
+$X_{\text{ss,$\epsilon$ }}:=\left\{x\in X:\,M(x)\geq-\epsilon\right\},\,\,\,M(x):=\inf_{\lambda}\frac{w_{\lambda}(x)}{\left|\lambda\right|}.$
+$(X^{N})_{\text{ss},\epsilon}$ coincides with the semi-stable locus $X_{\text{ss }}$ when $\epsilon=0$ and contains $X_{\text{ss }}$ for any $\epsilon.$ The sets $X_{\text{ss,$\epsilon$ }}$ are Zariski open for a generic $\epsilon$ (avoiding the finite set $M(X)\cap]-\infty[$).
+This follows from the properties of the Hesselink stratification of $X-X_{\text{ss }}$ (where the function M(x)-M(x) defines the order of the different of strata [84, Section 5.6]).
+Let now $(X,\Delta)$ be a log Fano manifold and consider the action of the group $\mathcal{G}$ on $X^{N}$ and set, for $N\epsilon$ generic,
+$(X^{N})_{\text{ss},\epsilon}:=\left\{X^{N}:\,\frac{1}{N}M(x_{1},...,x_{N})\geq-\epsilon\right\},\,\,\,\,\gamma_{\epsilon}^{(N)}(X,\Delta)^{\mathcal{G}}\coloneqq\text{lct}\left((X^{N})_{\text{ss},\epsilon},\Delta_{N};\mathcal{D}^{(N)}\right)$
+Definition 7.1.
+A log Fano manifold $(X,\Delta)$ is said to be strongly uniformly Gibbs polystable if $(X,\Delta)$ is Gibbs semi-stable and there exists a sequence $\epsilon_{N}\in[0,\infty[,$ called the thickening sequence, such that $\epsilon_{N}N^{1/2}\rightarrow\infty$ and
+$\liminf_{N\rightarrow\infty}\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}}>1$
+In particular, when $\epsilon_{N}=0$ this means that $X$ is uniformly Gibbs polystable.
+The condition that $(X,\Delta)$ be strongly uniformly Gibbs polystable is independent of the choice of scalar product on $\mathfrak{k}^{*}.$ Indeed, since any two norms are comparable, strong uniform Gibbs polystability is preserved when changing the norm, by rescaling the thickening sequence.
+We expect that a log Fano manifold is uniformly Gibbs polystable iff it is strongly uniformly Gibbs polystable, as illustrated by the following
+Example 7.2.
+A log Fano curve $(\mathbb{P}^{1},\Delta)$ is uniformly Gibbs polystable iff it is is strongly uniformly Gibbs polystable. Indeed, for any thickening sequence $\epsilon_{N}$ tending to zero
+$\lim_{N\rightarrow\infty}\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}}=\lim_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)^{\mathcal{G}}$
+as follows from a slight generalization of the proof of Theorem 1.3.
+Indeed, when, for example $\Delta=0,$ so that $\mathcal{G}=SL(2,\mathbb{C}),$we get, for an appropriate choice of norm, $M_{N}(x_{1},...,x_{N})/N=1-2I/N$ where $I$ denotes the numbers of $x_{i}$ coinciding (as follows from the formula for the weights $w_{\lambda}(x_{1},...,x_{N})$ in [84, Example 6.6]).
+Hence, $M_{N}(x_{1},...,x_{N})/N\geq-\epsilon_{N}$ iff $I/N\leq 1/2+\epsilon_{N}/2$ and one can then proceed precisely as in the proof of Theorem 1.3.
+7.2. An effective large deviation bound
+Fix a $\mathcal{K}$–invariant metric $\phi_{0}$ on $K_{(X,\Delta)}$ with positive curvature form $\omega_{0}$ and denote, as before, the corresponding measure on $X$ by $\mu_{0},$ by $\boldsymbol{m}$ the corresponding moment map $X\rightarrow\mathfrak{k}^{*}$ and by $\boldsymbol{m}_{N}$ the avarage of the moment map $X^{N}\rightarrow\mathfrak{k}^{*}.$ Fixing an $\text{ad}K$-invariant $L^{2}$-norm $\left|\cdot\right|$ on $\mathfrak{k}^{*},$ we can identify $\mathfrak{k}^{*}$ with $\mathbb{R}^{r}$ and write $\boldsymbol{m}=(h_{1},...,h_{r}).$ Given $\epsilon>0$ consider the thickened partition function $Z_{N,\epsilon}(-\gamma),$ defined by
+$Z_{N,\epsilon}(-\gamma):=\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\left\|\det S^{(N)}\right\|_{\phi_{0}}^{-2\gamma/k}\mu_{0}^{\otimes N}$
+Lemma 7.3.
+Assume that $(X,\Delta)$ is strongly uniformly Gibbs polystable and denote by $\epsilon_{N}$ the corresponding thickening sequence. There exists $\gamma>1$ such that for any sufficently large NN
+$Z_{N,\epsilon^{\prime}_{N}}(-\gamma)<\infty,$
+for any sequence $\epsilon^{\prime}_{N}$ such that $\epsilon^{\prime}_{N}<\epsilon_{N}.$
+Proof.
+By the moment-weight inequality in GIT [77]
+$\frac{1}{N}\inf_{\lambda}\frac{w_{\lambda}(x_{1},..,x_{N})}{\left|\lambda\right|}\leq\inf_{g\in G}\left|\boldsymbol{m}_{N}(g\cdot(x_{1},..,x_{N}))\right|$
+In particular, if $\boldsymbol{m}_{N}((x_{1},..,x_{N}))<\epsilon_{N}$ then $(x_{1},..,x_{N})\in(X^{N})_{\text{ss},\epsilon_{N}}.$ This shows that the closure of $U_{\epsilon^{\prime}_{N}}$ is contained in $(X^{N})_{\text{ss},\epsilon_{N}}.$ Hence, the integrability in the lemma follows directly from the analytic definition of the log canonical threshold.
+∎
+The following result, which builds on [20], yields an effective large deviation bound, motivated, by the conjectural LDP in Conjecture 6.7:
+Theorem 7.4.
+Let $(X,\Delta)$ be a log Fano manifold and fix $\gamma>0.$ There exists a constant $C$ only depending on $\phi_{0}$ and $\gamma$ such that
+$\frac{1}{N\gamma}\log Z_{N,\epsilon}(-\gamma)\leq\left(1+Ck^{-1}\right)\inf_{\mathcal{P}(X)_{0}}F_{-\gamma\left(1-Ck^{-1}\right)}+Ck^{-1}+C\epsilon-\log\left(1-Ce^{-2\epsilon^{2}N/C}\right)$
+if $\epsilon^{2}N$ is sufficiently large so that $Ce^{-2\epsilon^{2}N/C}<1.$ Moreover, if the Futaki character of $(X,\Delta)$ vanishes, then it also follows that $\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}$ is non-empty and, as a consequence, so is $(X^{N})_{\text{ss},\epsilon}.$
+A direct consequence of the previous theorem is the following result, containing Theorem 1.4, stated in the introduction (using that Gibbs-semistability implies that the Futaki character vanishes):
+Corollary 7.5.
+Assume that the Futaki character of $(X,\Delta)$ vanishes and $\epsilon_{N}N^{1/2}\rightarrow\infty.$ Then
+$\limsup_{N\rightarrow\infty}\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}}\leq\delta_{A}(X,\Delta)^{\mathcal{G}}.$
+As a consequence, if $(X,\Delta)$ is strongly uniformly Gibbs polystable, then $(X,\Delta)$ admits a Kähler-Einstein metric. More precisely, there exists a constant $C$ such that
+$(1-Ck^{-1})\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}}\leq\delta_{A}(X,\Delta)^{\mathcal{G}}.$
+Letting $N\rightarrow\infty$ we also deduce the following result, which yields the upper bound in Conjecture 1.8, for the “thickened” partition function $Z_{N,\epsilon_{N}}:$
+Corollary 7.6.
+Let $\epsilon_{N}$ be a sequence of positive numbers tending to zero sufficiently slowly so that $\epsilon_{N}N^{1/2}\rightarrow\infty.$ Then
+$\limsup_{N\rightarrow\infty}\left(-\frac{1}{N\gamma}\log Z_{N,\epsilon_{N}}(-\gamma)\right)\leq\inf_{\mathcal{P}(X)_{0}}F_{-\gamma}$
+In particular, if the Futaki character of $(X,\Delta)$ vanishes, then
+$\limsup_{N\rightarrow\infty}\left(-\frac{1}{N}\log Z_{N,\epsilon_{N}}(-1)\right)\leq\inf_{\mathcal{H}(X,\omega_{0})}\mathcal{M}$
+7.3. The proof of Theorem 7.4
+Given a function $u$ on $X$ such that $e^{-\gamma u}\in L^{1}(X)$ we will denote by $\mu_{\gamma u}$ the probability measure $\mu_{\gamma u}:=e^{-\gamma u}\mu_{0}/\int_{X}e^{-\gamma u}\mu_{0}$ on $X.$ We also set
+$U_{\epsilon}:=\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}\subset X^{N}$
+7.3.1. Inequality in terms of a quantized Ding functional and a mean entropy penalty for a finite NN
+We start by establishing the following inequality for any $u$ such that $e^{-\gamma u}\in L^{1}(X):$
+(7.1) $\frac{1}{N}\log Z_{N,\epsilon}(-\gamma)\leq\mathcal{D}_{k,\gamma}(u)-\frac{1}{\gamma}\frac{1}{N}\log\mu_{\gamma u}^{\otimes N}\left(\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}\right)+\frac{\log N!}{N(k-\gamma)}$
+where $\mathcal{D}_{k,-\gamma}$ denotes the functional on $C^{0}(X)$ defined by
+(7.2) $\mathcal{D}_{k,\gamma}(u):=-\mathcal{L}_{k}(u)-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0},$
+where $\mathcal{L}_{k}(u)$ is the following variant of Donaldson’s $\mathcal{L}_{k}$-functional:
+$\mathcal{L}_{k}(u):=-\frac{1}{N(k+\gamma)})\log\det\left(\int_{X}\left\langle s_{i},s_{j}\right\rangle_{\phi_{0}+u}e^{-\gamma u}\mu_{0}\right),$
+with $\left\langle\cdot,\cdot\right\rangle_{\phi_{0}+u}$ denoting the Hermian metric on $kK_{(X,\Delta)}$ induced by the metric $\phi_{0}+u$ on $K_{(X,\Delta)}$ and $s_{1},...,s_{N}$ any fixed basis of $H^{0}(X,-kK_{(X,\Delta)})$ which is orthonormal with respect to the Hermitian product on $H^{0}(X,-kK_{(X,\Delta)})$ induced by the metric $\phi_{0}$ on $K_{(X,\Delta)}$ and the measure $\mu_{\phi_{0}}$ on $X.$ The normalization have been chosen to ensure that $\mathcal{D}_{k,-\gamma}(u)$ is invariant under $u\mapsto u+c,$ for any $x\in\mathbb{R}.$
+To prove the inequality 7.1, we first rewrite
+$\left\|\det S^{(N)}\right\|_{\phi_{0}}^{-2\gamma/k}\mu_{0}^{\otimes N}=\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{-2\gamma/k}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}=$
+$\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{-2\gamma/k}\left(\frac{\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}}{\int_{U_{\epsilon}}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}}\right)\int_{U_{\epsilon}}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}$
+Applying Hölder’s inequality with negative exponent $\gamma/k$ (or applying Jensen’s inequality to the convex function $t\mapsto t^{-\gamma/k}$ on $\infty$$\infty$[) thus yields
+$Z_{N,\epsilon}(-\gamma)^{-1}\leq\left(\int_{U_{\epsilon}}\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{2}\frac{\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}}{\int_{U_{\epsilon}}e^{-\gamma u}\mu_{0}}\right)^{\gamma/k}\left(\int_{U_{\epsilon}}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}\right)^{-1}=$
+$\left(\int_{U_{\epsilon}}\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{2}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}\right)^{\gamma/k}\left(\int_{U_{\epsilon}}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}\right)^{-(1-\gamma/k)}$
+Bounding the integral over $U_{\epsilon}$ in the first bracket above by an integral over all of $X^{N},$ taking logarithms and dividing by $\gamma(1-\gamma/k),$ we deduce that
+$\frac{1}{\gamma N(1-\gamma/k)}\log Z_{N,\epsilon}(-\gamma)\leq\frac{\log\left(\int_{X^{N}}\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{2}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}\right)}{\gamma N(k-\gamma)}-\frac{\log\int_{U_{\epsilon}}\left(e^{-\gamma u}\mu_{0}\right)^{\otimes N}}{\gamma N}$
+Next, applying the previous inequality to $u-\log\int_{X}e^{-\gamma u}\mu_{0}$ the right hand side becomes
+$\frac{1}{N(k-\gamma)}\log\left(\int_{X^{N}}\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{2}\mu_{\gamma u}^{\otimes N}\right)-\frac{1}{\gamma}\log\int_{X}e^{-\gamma u}\mu_{0}-\frac{1}{\gamma}\frac{1}{N}\log\int_{U_{\epsilon}}\mu_{\gamma u}^{\otimes N},$
+using that the latter right hand side is invariant under $u\mapsto u-c$ for any $c\in\mathbb{R}.$ Finally, the desired inequality 7.1 follows from the following basic identity [23, Lemma 5.3]:
+$\int_{X^{N}}\left\|\det S^{(N)}\right\|_{\phi_{0}+u}^{2}\mu_{\gamma u}^{\otimes N}=N!\det\left(\int_{X}\left\langle s_{i},s_{j}\right\rangle_{\phi_{0}+u}e^{-\gamma u}\mu_{0}\right).$
+7.3.2. An effective upper bound on the quantized Ding functional
+Now assume that $u\in\mathcal{H}(X,\omega_{0}).$ Then it follows from [20, Lemma 2.4] (together with the scaling argument in the beginning of [20, Section 2.3]) that
+$\mathcal{D}_{k,\gamma}(u)\leq\mathcal{D}_{\gamma_{k}}(u)+(C_{0}+1)k^{-1}E(\frac{\omega_{u}^{n}}{V}),\,\,\,\gamma_{k}:=(1-\frac{\gamma-1}{k+\gamma})\gamma$
+(we recall that the proof of [20, Lemma 2.4] leverages Berndtsson’s positivity result for direct image bundles and weak geodesics in the space $\mathcal{H}(X,\omega_{0})).$ Since any $u\in\text{PSH }(X,\omega_{0})$ is a decreasing limit of functions $u_{j}\in\mathcal{H}(X,\omega_{0})$ (by Demailly’s approximation result), it follows that the inequality in fact holds for any $u\in\mathcal{E}^{1}(X,\omega),$ using also that $e^{-\gamma u}\in L^{1}(X)$ for any $\gamma>0$ if $u\in\mathcal{E}^{1}(X,\omega)$ [26].
+7.3.3. An effective upper bound on the mean entropy
+Next, we establish the following effective upper bound, expressed in terms of the entropy distance $\delta(\nu)$ to $\mathcal{P}(X)_{0}$ defined in Lemma 5.6, where $C$ is a constant only depending on $\boldsymbol{m}$ (and hence only on $\phi_{0}):$
+(7.3) $\frac{1}{N}\log\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\mu_{\gamma u}^{\otimes N}\leq\delta(\mu_{\gamma u})+C\epsilon-\log\left(1-2re^{-\frac{1}{R^{2}}2\epsilon^{2}N/r^{2}}\right),\,\,R:=\max_{i\leq r}\left(\sup_{X}h_{i}-\inf_{X}h_{i}\right)$
+when $\epsilon^{2}N$ is sufficently large so that $2re^{-\frac{1}{R^{2}}2\epsilon^{2}N/r^{2}}<1.$ To prove this we will use some standard arguments in the context of Cramér’s large deviation principle for iid random variables.
+Set $\nu:=\mu_{\gamma u}$ and $\nu_{0}:=\nu e^{\left\langle\lambda,\boldsymbol{m}\right\rangle}/\int_{X}\nu e^{\left\langle\lambda,\boldsymbol{m}\right\rangle},$ where $\lambda$ is the unique vector in $\mathbb{R}^{r}$ such that $\int_{X}\boldsymbol{m}\nu_{0}=0$ (see Lemma 5.6).
+Then
+$\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\nu^{\otimes N}=\left(\int_{X}\nu e^{\left\langle\lambda,\boldsymbol{m}\right\rangle}\right)^{N}\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}e^{-N\left\langle\lambda,\boldsymbol{m}_{N}\right\rangle}\nu_{0}^{\otimes N}$
+Hence,
+$\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\nu^{\otimes N}\geq\left(\int_{X}\nu e^{\left\langle\lambda,\boldsymbol{m}\right\rangle}\right)^{N}e^{-N|\lambda|\epsilon}\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\nu_{0}^{\otimes N},$
+giving
+$\frac{1}{N}\log\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\nu^{\otimes N}\leq-\log\left(\int_{X}\nu e^{\left\langle\lambda,\boldsymbol{m}\right\rangle}\right)+|\lambda|\epsilon-\frac{1}{N}\log\int_{\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}}\nu_{0}^{\otimes N}.$
+By the argument in the proof of Lemma 5.6, $\log\left(\int_{X}\nu e^{\left\langle\lambda,\boldsymbol{m}\right\rangle}\right)=\delta(\nu).$ Since $\left|\boldsymbol{m}_{N}(x_{1},..,x_{N})\right|_{L^{2}}\leq\left|\boldsymbol{m}_{N}(x_{1},..,x_{N})\right|r$ it will thus be enough to prove the following effective lower bound for any $\nu\in\mathcal{P}(X)$ satisfying $\int_{X}\boldsymbol{m}\nu_{0}=0:$
+(7.4) $\nu_{0}^{\otimes N}\left(\left\{\left|\boldsymbol{m}_{N}\right|_{L^{\infty}}<\epsilon\right\}\right)\geq 1-2re^{-\frac{1}{R^{2}}2N\epsilon^{2}},\,\,\,$,
+where $\left|\boldsymbol{m}_{N}(x_{1},...,x_{N})\right|_{L^{\infty}}$ denotes the maximum of the absolute values of the $r$ components of $\boldsymbol{m}_{N}(x).$ To this end consider first the case when $r=1.$ Then $\boldsymbol{m}_{N}=N^{-1}\sum_{i}h(x_{i})$ is the empirical sum $H_{N}$ of $N$ random real-valued indendent random variables on the probability space $(X^{N},\nu_{0}^{\otimes N})$ with identical distribution $h_{*}\nu_{0}$ that, by assumption, has mean zero, i.e.
+$\int_{t\in\mathbb{R}}th_{*}\nu_{0}=0.$ Hence, by Hoeffding’s inequality [82]:
+(7.5) $\nu_{0}^{\otimes N}\left(\left\{\left|H_{N}\right|\geq\epsilon\right\}\right)\leq 2e^{-\frac{1}{R^{2}}2N\epsilon^{2}},\,R:=\sup h-\inf h(X)$
+which is equivalent to the inequality 7.4.
+To prove the general case, denote by $H_{N}^{(1)},...,H_{N}^{(r)},$ the empirical sums of the real-valued random variables on the probability space $(X^{N},\nu_{0}^{\otimes N})$ defined by the coordinates $h_{1},...,h_{r}$ of $\boldsymbol{m}_{N}.$ Since
+$\left\{\left|\boldsymbol{m}_{N}\right|_{L^{\infty}}\geq\epsilon\right\}\subset\left\{\left|H_{N}^{(1)}\right|\geq\epsilon\right\}\cup\cdots\cup\left\{\left|H_{N}^{(r)}\right|\geq\epsilon\right\},$
+it follows from Hoeffding’s inequality 7.5 applied to $h_{N}^{(1)},...,h_{N}^{(r)}$ that
+$\nu_{0}^{\otimes N}\left(\left\{\left|\boldsymbol{m}_{N}\right|_{L^{\infty}}\geq\epsilon\right\}\right)\leq 2re^{-\frac{1}{R^{2}}2N\epsilon^{2}},\,R:=\max_{i\leq r}\left(\sup_{X}h_{i}-\inf_{X}h_{i}\right),$
+which concludes the proof of the lower bound 7.4.
+7.3.4. Conclusion of the proof of Theorem 1.4
+By the previous steps
+$\frac{1}{N\gamma}\log Z_{N,\epsilon}(-\gamma)\leq\mathcal{D}_{\gamma_{k}}(u)+\delta_{\gamma}(u)+O(1/k)E(\frac{\omega_{u}^{n}}{V})+O(1/k),\,\,\,\gamma_{k}=\gamma(1-O(k^{-1}))$
+Hence, the proof of the inequality in Theorem 1.4 is concluded by invoking inequality 5.18 and then taking the inf over all $u\in\mathcal{E}^{1}(X,\omega_{0}).$ Finally, assume that the Futaki character of $(X,\Delta)$ vanishes.
+Then, by Lemma 3.3, $\omega_{0}^{n}/V\in\mathcal{P}(X)_{0}.$ It thus follows from the lower bound 7.4, applied to $\nu_{0}:=\omega_{0}^{n}/V,$ that $\left\{\left|\boldsymbol{m}_{N}\right|<\epsilon\right\}$ is non-empty.
 8. Asymptotics of the $\mathcal{K}$-reduced partition function on log Fano curves
@@
 To this end it will be convenient to replace the basis of $H^{0}(X,-kK_{(X,\Delta)}),$ appearing in the definition of $\nu^{(N)},$ by the basis of monomials $1,z,z^{2},\dots$ induced by standard identification of $kK_{(X,\Delta)}$ with $kV\mathcal{O}(1)\rightarrow\mathbb{P}^{1}.$ This amounts to adding an appropriate constant to the corresponding Mabuchi functional $\mathcal{M}$ of $(X,\Delta)$ to get a new functional that we denote by $\mathcal{M}_{\mathbb{Z}}$ (the subscript indicates the arithmetic interpretation of this setup, discussed in Section 9).
-The inf of $\mathcal{M}_{\mathbb{Z}}$ was explicitly computed in [3, Thm 6.1], when $\Delta$ is any divisor on $\mathbb{P}^{1}$ with support $\{0,1,\infty\},$ assuming that $(X,\Delta)$ is K-stable (in particular, $\mathcal{G}$ is then trivial).
-In a nutshell, we will reduce the proof of Theorem 1.7, where $\mathcal{G}$ is non-trivial to the computation in [3, Thm 6.1], by showing that breaking the $\mathcal{G}$-symmetry corresponds, effectively, to perturbing the given K-polystable log Fano curve $(X,\Delta)$ to a K-stable log Fano curve $(X,\Delta_{N})$.
+The inf of $\mathcal{M}_{\mathbb{Z}}$ was explicitly computed in [4, Thm 6.1], when $\Delta$ is any divisor on $\mathbb{P}^{1}$ with support $\{0,1,\infty\},$ assuming that $(X,\Delta)$ is K-stable (in particular, $\mathcal{G}$ is then trivial).
+In a nutshell, we will reduce the proof of Theorem 1.9, where $\mathcal{G}$ is non-trivial to the computation in [4, Thm 6.1], by showing that breaking the $\mathcal{G}$-symmetry corresponds, effectively, to perturbing the given K-polystable log Fano curve $(X,\Delta)$ to a K-stable log Fano curve $(X,\Delta_{N})$.
 Lemma 8.1.
-On any log Fano curve $(X,\Delta)$ with vanishing Futaki character, $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$
+On any log Fano curve $(X,\Delta)$ with vanishing Futaki character, $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$ Moreover, for any log Fano curve 0 is a regular value of the moment map $\boldsymbol{m}_{N}$ if $N$ is odd and $N\geq 3.$
 Proof.
 We can express $(X,\Delta)=(\mathbb{P}^{1},\Delta_{w})$ for $w\in[0,1[$ (see the discussion proceeding Theorem 2.11).
-Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 6.20, which implies the lemma.
+Fix a $\mathbb{C}^{*}$–action preserving $\Delta_{w}.$ Then the corresponding moment polytope $P_{w}$ is symmetric (P=[1+w,1w])(P=[-1+w,1-w]) and thus $\boldsymbol{b}_{k}=0$ in Prop 6.22, which implies the first part of the lemma.
+Alternatively, one can show by direct computation that $g^{*}\nu_{N}=\nu_{N}$ for any $g\in\mathcal{G}.$ To prove the last statement, first note that for any $N\geq 2,$ $\left\{\boldsymbol{m}_{N}=0\right\}$ is non-empty, since it intersects the $\mathcal{G}$-orbit of any configuration $(x_{1},...,x_{N})$ of mutually distinct points.
+Indeed, by [103, Section 3], a configuration $(x_{1},...,x_{N})$ is stable iff the inequality 2.10 is strict and by the general Kempf-Ness isomorphism [83], the $\mathcal{G}$-orbit of any polystable point intersects the zero-locus of the moment map (and stability implies polystability, in general).
+Finally, assume that $N$ is odd and $N\geq 3.$ Then $(x_{1},...,x_{N})$ is semi-stable iff it is stable.
+Indeed, when $N$ is odd the semi-stability inequality 2.10 holds iff it is a strict inequality.
+Hence, the semi-stable locus coincides with the stable locus, which is equivalent to 0 beeing a regular value of the moment map [83, Cor 3.10].
 ∎
@@
 8.1.1. Expressing the contracted partition function $Z_{N}^{\mathrm{c}}$ as a complex Selberg integral $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
-By Prop 6.14, $Z_{N}^{\mathrm{c}}$ (defined by formula 6.4) may be expressed as
+By Prop 6.16, $Z_{N}^{\mathrm{c}}$ (defined by formula 6.6) may be expressed as
 $Z_{N}^{\mathrm{c}}=C\int_{(X^{N})_{\text{s}}/\mathcal{G}}(\nu^{(N)})^{\mathrm{cc}}$
@@
 (8.3) $\nu^{(N)}=\eta\wedge\rho_{N-3}^{\boldsymbol{w}}(y_{4},\dots,y_{N}),\,\,\,\eta\coloneqq\frac{(\frac{i}{2})^{3}dz_{1}\wedge d\overline{z}_{1}\wedge\cdots\wedge dz_{3}\wedge d\overline{z}_{3}}{\left|(z_{1}-z_{2})(z_{1}-z_{3})(z_{2}-z_{3})\right|^{2}}$
-Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 7.1, applied to $N=3).$ Thus formula 7.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 6.9.
+Note that $\eta$ defines a $\mathcal{G}$–invariant measure on $X^{3}$ (this also follows from Lemma 8.1, applied to $N=3).$ Thus formula 8.3 recovers the $\mathcal{G}$–invariance of $\nu^{(N)}$ in Lemma 6.1.
 Moreover, it follows from the $\mathcal{G}$–invariance of $\eta$ that the contracted form $\eta^{\mathrm{c}}$ is a constant.
@@
 $\eta^{\mathrm{cc}}=1.$
-Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 6.4.1. Thus, accepting formula 7.3, we conclude that
+Moreover, using that $y_{i}$ is $\mathcal{G}$–invariant we have $\delta_{e_{i}}dy_{j}=\delta_{Je_{i}}dy_{j}=0,$ in the notation of Section 6.5.1. Thus, accepting formula 8.3, we conclude that
 $(\nu^{(N)})^{\mathrm{cc}}=\rho_{N-3}^{\boldsymbol{w}}(y_{4},\dots,y_{N}),$
-which implies the desired formula 7.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
-Finally, as next described, the deferred proof of formula 7.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 7.2).
-First note that since $V=2$ and $N=2k+1$ [18, Lemma 4.3] yields
+which implies the desired formula 8.2 (with $C=1),$ since $y_{i}$ is $\mathcal{G}$–invariant.
+Finally, as next described, the deferred proof of formula 8.3 follows from a direct, but somewhat unwieldy calculations (the argument is guided by the simpler case when $\Delta$ is nontrivial and $\mathcal{G}=\mathbb{C}^{*},$ described in Section 8.2).
+First note that since $V=2$ and $N=2k+1$ [19, Lemma 4.3] yields
 $\nu^{(N)}=\left(\prod_{i\neq j}\left|z_{i}-z_{j}\right|\right)^{-\frac{1}{k}}\left(\frac{i}{2}dz\wedge d\overline{z}\right)^{\otimes N}.$
@@
 $=\prod_{1\leq i<j\leq N}|z_{i}-z_{j}|^{-2/k}\bigwedge_{\ell=1}^{N}\mathrm{d}z_{\ell}\wedge\mathrm{d}\bar{z}_{\ell}=\nu^{(N)}(z),$
-which concludes the proof of formula 7.3.
-7.1.2. Estimating the distortion $f_{N}$ between $Z_{\boldsymbol{m}_{N}}$ and $Z_{N}^{\mathrm{c}}$
-We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 6.15 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
+which concludes the proof of formula 8.3.
+8.1.2. Estimating the distortion $f_{N}$ between $Z_{N,0}$ and $Z_{N}^{\mathrm{c}}$
+We next show that there exists a positive constant $C$ such that function $f_{N}$ in formula 6.17 satisfies $N^{-3}C\leq f_{N}\leq C$ when $\boldsymbol{m}_{N}=0.$ By symmetry, it will be enough to show that
 (8.4) $N^{-1}C\leq\omega_{N}(e_{3},Je_{3})\leq C\,\,\text{on $\{\left\langle\boldsymbol{m}_{N},e_{3}\right\rangle=0\}\subset X^{N}$.}$
@@
 Denote by $P$ the convex polytope in $\mathbb{R}^{N}$ defined by the constraints on $p_{i}$ above.
-Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 7.4.
+Since the function to be maximized is convex its maximum is attained at the extremal points $(p_{1},\dots,p_{N})$ of $P.$ We claim that these satisfy $p_{i}\in\{0,\pm 1\}.$ It then readily follows that the maximum in question is attained when $p_{i}=0$ for one index $i$ and the remaining N1N-1 indices are split into two groups of (N1)/2(N-1)/2 members satisfying $p_{i}=-1$ and $p_{i}=1,$ respectively, which proves the lower bound in 8.4.
 Finally, to prove the previous claim, consider the hypersimplex $\Delta_{N/2}$ obtained by replacing $p_{i}$ with $t_{i}\coloneqq(p_{i}+1)/2$:
@@
 8.1.3. Computing the large-$N$ limit of $\mathcal{Z}_{N-3}^{\boldsymbol{w}}$
-As explained in [3], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 7.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial).
-By [3, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
+As explained in [4], the general integral $\mathcal{Z}_{N}^{\boldsymbol{w}},$ defined in formula 8.1, can be viewed as an arithmetic partition function on $(\mathbb{P}^{1})^{N}$ corresponding to the log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}}),$ where $\Delta^{\boldsymbol{w}}$ denotes the divisor supported on the three points $0,1$ and $\infty$ with weights $w_{1},w_{2}$ and $w_{3},$ respectively (in this case the corresponding automorphism group $\mathcal{G}$ is trivial).
+By [4, Thm 5.3], given any log Fano curve $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ that is K-stable (or, equivalently, Gibbs stable), the corresponding partition function on $(\mathbb{P}^{1})^{N}$ satisfies
 (8.5) $\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathcal{Z}_{N}^{\boldsymbol{w}}=\inf\mathcal{M}_{\mathbb{Z}}$
@@
 (8.6) $\boldsymbol{w}_{N}\coloneqq\frac{2}{N+2}(1,1,1)\implies-\lim_{N\rightarrow\infty}\frac{1}{N}\log\mathcal{Z}_{N}^{\boldsymbol{w}_{N}}=\inf\mathcal{M}_{\mathbb{Z}},$
-where $\mathcal{M}_{\mathbb{Z}}$ denote the arithmetic Mabuchi functional attached to $(\mathbb{P}^{1},0).$ As pointed out in [3, Section 6.2], it follows from [53, Formula B.9] (and [60, formula 3.1] and [92, Section 2.3]) that for any weight vector $\boldsymbol{w}\in]0,1[^{3}$ such that $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ is a log Fano curve that is K-stable, i.e.
+where $\mathcal{M}_{\mathbb{Z}}$ denote the arithmetic Mabuchi functional attached to $(\mathbb{P}^{1},0).$ As pointed out in [4, Section 6.2], it follows from [69, Formula B.9] (and [76, formula 3.1] and [110, Section 2.3]) that for any weight vector $\boldsymbol{w}\in]0,1[^{3}$ such that $(\mathbb{P}^{1},\Delta^{\boldsymbol{w}})$ is a log Fano curve that is K-stable, i.e.
 such that
@@
 (8.9) $\mathcal{Z}_{1}^{\boldsymbol{w}}\coloneqq\int_{\mathbb{C}}\left|z\right|^{-2w_{1}}\left|z-1\right|^{-2w_{2}}\frac{i}{2}dz\wedge d\bar{z}=\pi\frac{\Gamma(1-w_{1})}{\Gamma(w_{1})}\frac{\Gamma(1-w_{1})}{\Gamma(w_{1})}\frac{\Gamma(-1+w_{1}+w_{2})}{\Gamma(2-w_{1}-w_{2})}$
-(as follows from formula 7.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [3], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
+(as follows from formula 8.8 applied to $V=0$ and any $N\geq 2).$ When $\boldsymbol{w}$ is independent of $N$ it follows from elementary arguments, as observed in [4], that $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}}$ converges, as $N\rightarrow\infty,$ towards
 $\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right)$
-Combining this convergence with the convergence 7.5 gives the following explicit formula
+Combining this convergence with the convergence 8.5 gives the following explicit formula
 (8.10) $\inf\mathcal{M}_{\mathbb{Z}}=\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right),$
-which is shown in [3] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 7.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 7.6.
+which is shown in [4] to be valid also when $(X,\Delta)$ is merely K-polystable, using a continuity argument with respect to $\boldsymbol{w}.$ In particular, the formula applies to $\Delta=0.$ Now apply the explicit formula 8.10 to the case of the weight vector $\boldsymbol{w}_{N}$ in formula 8.6.
 Note that
@@
 where the error terms are positive.
-Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 7.8:
+Hence, Stirling’s approximation and the fact that $\Gamma(x)$ has a simple pole with residue 1 at $x=0$ leads to the following logarithmic asymptotics of the first factor in formula 8.8:
 $\frac{1}{N}\left(\log N!-N\log\left(-l\left(-\frac{1}{2}\frac{\left|V_{N}\right|}{N-1}\right)\right)\right)=\log\frac{|V|}{2}-1+\frac{1}{2}\frac{\log N}{N}+O\left(N^{-1}\right).$
@@
 $\int_{0}^{1}\log\Gamma\left(w_{N}+t\frac{V_{N}}{2}\right)dt=\int_{0}^{1}\log\Gamma\left(w+t\frac{V}{2}\right)dt+O(N^{-1}\log N),$
-using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 7.11.
-The nominator in the second factor in formula 7.8 is handled in essentially the same manner, giving
+using that the Gamma function has a simple pole at $x=0$ and that $x\log x-x$ is a primitive of $\log x.$ This proves the asymptotics 8.11.
+The nominator in the second factor in formula 8.8 is handled in essentially the same manner, giving
 $N^{-1}\log\mathcal{Z}_{N}^{\boldsymbol{w}_{N}}=\log\frac{\left|V\right|}{2}+\log\pi-1+\frac{2}{\left|V\right|}\left(\int_{-\frac{\left|V\right|}{2}}^{0}\log(-l(x))\mathrm{d}x-\sum_{k=1}^{3}\int_{w_{k}}^{w_{k}+\frac{\left|V\right|}{2}}\log l(x)\mathrm{d}x\right)+O(N^{-1}\log N)$
-for $w_{i}=0$ and $V=2.$ The proof of formula 7.6 is thus concluded by invoking formula 7.10.
+for $w_{i}=0$ and $V=2.$ The proof of formula 8.6 is thus concluded by invoking formula 8.10.
 8.2. Proof of Theorem 1.9 when $\Delta$ is nontrivial
@@
 (8.13) $\nu^{(N)}=\eta\wedge\rho_{N-1}^{\boldsymbol{w}}(y_{2},\dots,y_{N-2})\left(\frac{i}{2}dy\wedge d\overline{y}\right)^{\otimes N-1},\,\,\,\eta=\frac{i}{2}\frac{dz_{1}}{z_{1}}\wedge\overline{(\frac{dz_{1}}{z_{1}})}$
-To prove this formula recall that, by [18, Lemma 4.3], we may express
+To prove this formula recall that, by [19, Lemma 4.3], we may express
 $\nu^{(N)}=\left(\prod_{i\neq j}\left|z_{i}-z_{j}\right|\right)^{-\frac{V}{N-1}}\prod_{i\leq N}\left|z_{i}\right|^{-2w}\left(\frac{i}{2}dz\wedge d\overline{z}\right)^{\otimes N},\,\,\,V\coloneqq 2-2w.$
@@
 for some real number $r$ and a direct computation reveals that $r=0.$ Alternatively, using that both $\nu^{(N)}$ and the second factor in the right hand side of the previous equation is $\mathcal{G}$–invariant (i.e.
-$\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 7.13.
+$\mathbb{C}^{*}$–invariant) forces $r=0.$ This proves formula 8.13.
 9. Relations to arithmetic geometry
@@
 Then
-(8.1) $\lim_{N\rightarrow\infty}N^{-1}\log\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{\boldsymbol{m}_{N}}^{(N)}=\inf\mathcal{M}_{\mathbb{Z}}.$
-In fact, this conjecture is equivalent to the convergence in Conjecture 1.6, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [3, Thm 5.3]).
-One drawback of Conjecture 1.6 – from the arithmetic point of view – is that left hand side in formula 8.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).).
+(9.1) $\lim_{N\rightarrow\infty}N^{-1}\log\int_{\left\{\boldsymbol{m}_{N}=0\right\}}\nu_{0}^{(N)}=\inf\mathcal{M}_{\mathbb{Z}}.$
+In fact, this conjecture is equivalent to the convergence in Conjecture 1.8, as follows readily from the arithmetic Hilbert–Samuel theorem (by the argument in the proof of [4, Thm 5.3]).
+One drawback of Conjecture 1.8 – from the arithmetic point of view – is that left hand side in formula 9.1 depends on non-arithmetic data (the choice of the group $\mathcal{K}$ and the metric $\phi_{0}$).).
 This can be remedied if the quantized Futaki character vanishes for $k$ sufficiently large (Section 6.6) and the Lie algebra $\mathfrak{g}$ of the $\mathcal{G}$ is the complexification of a natural $\mathbb{Z}$-module $\mathfrak{g_{\mathbb{Z}}}$ attached to $(\mathcal{X},\mathcal{D})$:
@@
 $\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})\coloneqq\int_{(X^{N})_{\text{s}}/\mathcal{G}}(\nu^{(N)})^{\mathrm{cc}}$
-realized as periods. The following conjecture generalizes [18, Conjecture 5.2] (see also [3, Section 5.2]) to the case when $\mathcal{G}$ is nontrivial:
+realized as periods. The following conjecture generalizes [19, Conjecture 5.2] (see also [4, Section 5.2]) to the case when $\mathcal{G}$ is nontrivial:
 Conjecture 9.2.
@@
 $\lim_{N\rightarrow\infty}N^{-1}\log\mathcal{Z}_{N}(\mathcal{X},\mathcal{D})=\inf\mathcal{M}_{\mathbb{Z}}$
-By Prop 6.14, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 6.15 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.7).
+By Prop 6.16, the two previous conjectures are equivalent if the $f_{N}$ defined in formula 6.17 satisfies $\log f_{N}=o(N^{-1}).$ We expect this to be the case, in general (the case when $X=\mathbb{P}^{1}$ was established in the proof of Theorem 1.9).
 9.1. Examples
-The condition 8.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e.
-$\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [66, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [81, Section 2].
-By Prop 6.20, the assumptions of Conjecture 8.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish.
+The condition 9.2 is, for example, satisfied when $\mathcal{X}$ is smooth over $\mathbb{Z},$ if $\mathfrak{g_{\mathbb{Z}}}$ is taken as the relative tangent bundle $T_{\mathcal{X}/\mathbb{Z}}$ of $\mathcal{X}$ over $\mathbb{Z}.$ This is the case if $\mathcal{X}$ is homogeneous over $\mathbb{Z},$ i.e.
+$\mathcal{X}=G/P,$ where $G$ is a semisimple Chevalley group over $\mathbb{Z}$ and $P$ a parabolic subgroup of $G$ [86, Section 2], or if $\mathcal{X}$ is the canonical model over $\mathbb{Z}$ of a toric Fano manifold $X$ [99, Section 2].
+By Prop 6.22, the assumptions of Conjecture 9.2 are then satisfied (for all $k)$ if, in the toric case, we also assume that the quantized barycenters vanish.
 It should be stressed that in these cases, the arithmetic data that go into the definition of the periods $Z_{N}(\mathcal{X},\mathcal{D})$ is, essentially, combinatorial.
 One is thus led to ask if $Z_{N}(\mathcal{X},\mathcal{D})$ can be computed explicitly in these cases.
-For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 7.1 and formula 7.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
+For example, when $\mathcal{X}=\mathbb{P}_{\mathbb{Z}}^{1}$ and $\mathcal{D}$ is trivial, combining formula 8.1 and formula 8.8 shows that $Z_{N}(\mathcal{X},\mathcal{D})$ is a product of Gamma-factors at rational numbers (up to factors of $\pi)$ with rational exponents, when $N\geq 5$:
 $Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1})=\pi^{N-3}\frac{(N-3)!}{-l(-\frac{1}{N-4})^{N}}\prod_{j=0}^{N-4}\frac{-l(-(j+1)\frac{1}{N-4})}{l(\frac{2}{N-1}+j\frac{1}{N-4})^{3}},\,\,\,\,\,l(x)\coloneqq\frac{\Gamma(x)}{\Gamma(1-x)}$
-(and a similar formula holds when $N=4,$ by formula 7.9).
-Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 7.12 and formula 7.8 gives, when $N\geq 3,$
+(and a similar formula holds when $N=4,$ by formula 8.9).
+Likewise, when $\mathcal{D}$ is the Zariski closure $\mathcal{D}_{w}$ in $\mathbb{P}_{\mathbb{Z}}^{1}$ of the divisor $\Delta$ over $\mathbb{Q}$ supported at $\{0,\infty\}$ with equal weights $w\in]0,\infty[,$ combining formula 8.12 and formula 8.8 gives, when $N\geq 3,$
 $Z_{N}(\mathbb{P}_{\mathbb{Z}}^{1},\mathcal{D}_{w})=\pi^{N-1}\left(\frac{(N-1)!}{-l(-\frac{1-w}{N-2})^{N}}\right)\prod_{j=0}^{N-2}\frac{-l(-(j+1)\frac{1-w}{N-2})}{l(w+j\frac{1-w}{N-2})^{2}l(\frac{1-w}{N-1}+j\frac{1-w}{N-2})}.$
@@
 Proposition 10.1.
-Denote by $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}}$ the space of all $S^{1}$–invariant psh metrics $\phi$ on $K_{(X,\Delta)}$ such that $\phi(x)$ is in $C^{1,1}(\mathbb{R}).$ Given $\phi_{0}\in\text{PSH(}-K_{(X,\Delta})_{\mathrm{b}}$ which is assumed $S^{1}$–invariant, the corresponding free energy $F_{\gamma}(\mu)$ satisfies the following formula, when $\mu=dd^{c}\phi/V$ for $\phi\in\mathcal{H}(-K_{(X,\Delta)})_{1,1}^{S^{1}}$:
+Denote by $\mathcal{H}_{1,1}(-K_{(X,\Delta)})^{S^{1}}$ the space of all $S^{1}$–invariant psh metrics $\phi$ on $K_{(X,\Delta)}$ such that $\phi(x)$ is in $C^{1,1}(\mathbb{R}).$ Given $\phi_{0}\in\text{PSH(}-K_{(X,\Delta})_{\mathrm{b}}$ which is assumed $S^{1}$–invariant, the corresponding free energy $F_{-\gamma}(\mu)$ satisfies the following formula, when $\mu=dd^{c}\phi/V$ for $\phi\in\mathcal{H}(-K_{(X,\Delta)})_{1,1}^{S^{1}}$:
 $VF_{-\gamma}(\mu)=\gamma\mathcal{L}(v)+(1-\gamma)\int_{P}\phi_{0}(v^{\prime})dy-\int_{P}v^{\prime\prime}dy,\,\,\,\mathcal{L}(v)\coloneqq\frac{V}{2}\int_{\partial P}v-\int_{P}vdy$
@@
 Proof.
-When $\gamma=1$ it is shown in the proof of [16, Lemma 5, step 1] that
+When $\gamma=1$ it is shown in the proof of [17, Lemma 5, step 1] that
 $V\mathcal{M}(\phi)=\mathcal{L}(v)+\int_{\mathbb{R}}(\log\phi^{\prime\prime})\phi^{\prime\prime}dx$
-(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [16]). By formula 3.13, this means that
+(in our normalization $V\mathcal{M}$ corresponds to the functional $\mathcal{M}$ in [17]). By formula 3.15, this means that
 $F_{-1}(\mu)=\gamma E(\mu)+D(\mu|\mu_{\phi})=\mathcal{L}(v)-\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx+\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$
-Note that, by formula 9.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
+Note that, by formula 10.1, $D(\mu|\mu_{\phi})=\int_{\mathbb{R}}\log\frac{\phi^{\prime\prime}}{e^{-\phi_{0}}}\phi^{\prime\prime}dx.$ Hence,
 (10.2) $E(\mu)=-\mathcal{L}(v)+\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx,$
@@
 $F_{-\gamma}(\mu)=\gamma\mathcal{L}(v)+(1-\gamma)\int_{\mathbb{R}}\phi_{0}\phi^{\prime\prime}dx+\int_{\mathbb{R}}(\log\phi^{\prime\prime})\phi^{\prime\prime}dx.$
-As shown in proof of [16, Lemma 5, step 2], the last term above equals $\int_{P}v^{\prime\prime}dy$ (as shown by making the change of variables $y=\phi^{\prime}(x)).$ Likewise,
+As shown in proof of [17, Lemma 5, step 2], the last term above equals $\int_{P}v^{\prime\prime}dy$ (as shown by making the change of variables $y=\phi^{\prime}(x)).$ Likewise,
 $\int_{\mathbb{R}}\phi_{0}(x)\phi^{\prime\prime}(x)dx=\int_{P}\phi_{0}(v^{\prime})dy,$
@@
 $\mathcal{D}_{\gamma}(u)=-\frac{\mathcal{E}(u)}{V}--\frac{1}{\gamma}\log\int e^{-\left(\gamma\phi+(1-\gamma)\phi_{0}\right)}dx,\,\,\,\,-\frac{\mathcal{E}(u)}{V}=\int_{P}v\frac{dy}{V}$
-(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [22, Prop 2.9] in the second equality).
+(using in the first equality that $\text{(Log)}_{*}\left(|z|^{-2}\frac{i}{2\pi}dz\wedge d\bar{z}\right)=dx$ and [24, Prop 2.9] in the second equality).
 For example, when $w=0,$ the destabilizing curve for $\mathcal{D}_{\gamma},$ appearing in Cor 4.2 may, if one drops the requirement that it be in $C^{1}(X),$ be expressed explicitly as follows, by taking $v_{0}=0$ in formula 3.37 (which amounts to taking $\phi_{0}(x)\coloneqq|x|)$:
@@
 10.3. Proof of Lemma 4.1 (normalizations)
-For any metric $g$ on $S^{2}$ compatible with the complex structure on $\mathbb{P}^{1}$
+For any Riemannian metric $g$ on $S^{2}$ compatible with the complex structure on $\mathbb{P}^{1}$
 $dd^{c}u:=\frac{1}{\pi}\frac{i}{2}\partial\bar{\partial}u=\Delta_{g}u\frac{dV_{g}}{4\pi}$
-In particular, if $g=g_{S^{2}}$ then $dV_{g}/4\pi$ has unit-mass and thus coincides with $\omega_{\text{FS}}.$ Hence, the defining formula 4.2 for $\Delta_{\text{FS}}$ forces $\Delta_{\text{FS}}=\Delta_{g_{S^{2}}}.$ Next, by formula 3.12,
-$E(\mu)=-\frac{1}{2}\int_{X}u_{\mu}\mu,\,\,\text{where}\,dd^{c}(\frac{u_{\mu}}{V})+\omega_{\text{FS}}=\mu,\,\int u_{\mu}\omega_{\text{FS}}=0.$
+In particular, if $g$ is the metric on $S^{2}$ induced by the Euclidean metric on $\mathbb{R}^{3}$ then $dV_{g}/4\pi$ has unit-mass and thus coincides with $\omega_{\text{FS}}.$ Hence, the defining formula 4.2 for $\Delta_{\text{FS}}$ forces $\Delta_{\text{FS}}=\Delta_{g}.$ Next, by formula 3.12,
+(10.4) $E(\mu)=-\frac{1}{2}\int_{X}u_{\mu}\mu,\,\,\text{where}\,dd^{c}(\frac{u_{\mu}}{V})=\mu-\omega_{\text{FS}}\,\int u_{\mu}\omega_{\text{FS}}=0.$
 Let $G(x,y)$ be the symmetric lsc function on $X^{2}$ defined by $dd^{c}G(,\cdot y)=\omega_{\text{FS}}-\delta_{y},$ where $\delta_{y}$ denotes the Dirac measure at $y$ and $\int_{X}G(\cdot,y)\omega_{\text{FS }}=0.$ Then the previous equation yields
 $E(\mu)=\frac{V}{2}\int_{X^{2}}G(x,y)\mu\otimes\mu=(1-w)\int_{X^{2}}G(x,y)\mu\otimes\mu.$
-Finally, in the holomorphic coordinate $z$ on $\mathbb{C}\subset\mathbb{P}^{1}$ we have $dd^{c}(\log(1+|z|^{2})=\omega_{\text{FS}}$ and $dd^{c}(\log|z|^{2})=\delta_{0}.$ Thus
+Moreover, by formula 3.12 we also have
+$E(\mu)=\frac{V}{2}\int_{X}d(\frac{u_{\mu}}{V})\wedge d^{c}(\frac{u_{\mu}}{V})=\frac{V}{2}\left\|\mu-\omega_{\text{FS}}\right\|_{H^{-1}}^{2},$
+using the equation for $\frac{u_{\mu}}{V}$ in formula 10.4. Equivalently, integrating by parts,
+$E(\mu)=\frac{V}{2}\int_{X}(-dd^{c})(\frac{u_{\mu}}{V})(\frac{u_{\mu}}{V})=\int_{X}(-\Delta_{g})(\frac{u_{\mu}}{V})(\frac{u_{\mu}}{V})dV_{g}=\int_{X}(-\Delta_{g})^{1/2}(\frac{u_{\mu}}{V})(\frac{u_{\mu}}{V})dV_{g}$
+which proves the last equality in formula 4.3, using again the equation for $\frac{u_{\mu}}{V}$ in formula 10.4.
+Finally, to prove the explicit formula for $G(x,y)$ first note that in the holomorphic coordinate $z$ on $\mathbb{C}\subset\mathbb{P}^{1}$ we have $dd^{c}(\log(1+|z|^{2})=\omega_{\text{FS}}$ and $dd^{c}(\log|z|^{2})=\delta_{0}.$ Thus
 $G(z,w)=\log\frac{\left|z-w\right|^{2}}{\left(1+\left|z\right|^{2}\right)\left(1+\left|w\right|^{2}\right)}-C,$
 $C:=\int_{\mathbb{C}}\log\frac{\left|z\right|^{2}}{1+\left|z\right|^{2}}\omega_{\text{FS}}=\int_{\mathbb{C}}\log\frac{\left|z\right|^{2}}{1+\left|z\right|^{2}}\frac{1}{\pi}\frac{i}{2}\frac{dz\wedge d\bar{z}}{(1+\left|z\right|^{2})^{2}}=\int_{0}^{\infty}\log\frac{r^{2}}{1+r^{2}}\frac{d(r^{2})}{(1+r^{2})^{2}}=-1$
-Finally, using stereographic projection gives the explicit formula for $G(\boldsymbol{x},\boldsymbol{y}).$ Indeed, by $SO(3)$-symmetry, the argument of log above equals $(c\left|\boldsymbol{x}-\boldsymbol{y}\right|)^{2}$ for some $c>0$ and using that the diameter of $S^{2}$ equals $2$ forces $c=1/2.$
+By stereographic projection this proves the explicit formula for $G(\boldsymbol{x},\boldsymbol{y}).$ Indeed, by $\mathrm{SO}(3)$-symmetry, the argument of log above equals $(c\left|\boldsymbol{x}-\boldsymbol{y}\right|)^{2}$ for some $c>0$ and using that the diameter of $S^{2}$ equals $2$ forces $c=1/2.$